GNU LibreDWG

for version 0.13.3, 16 February 2024

GNU LibreDWG Developers and Thien-Thi Nguyen

This manual is for GNU LibreDWG (version 0.13.3, 16 February 2024). Copyright \begin{matrix}c \\ ◯ \end{matrix} 2010-2024 Free Software Foundation, Inc.

Permission is granted to copy, distribute and/or modify this document under the terms of the GNU Free Documentation License, Version 1.3 or any later version published by the Free Software Foundation; with no Invariant Sections, with no Front-Cover Texts, and with no Back-Cover Texts. A copy of the license is included in the section entitled “GNU Free Documentation License”.

All other trademarks, trade names or company names referenced herein are used for identification purposes only and are the property of their respective owners. DWG is the native and proprietary file format for AutoCAD◯^{R} and a trademark of Autodesk, Inc. LibreDWG is neither associated with Autodesk, nor with the Open Design Alliance.

Table of Contents

1 Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

1.1 API/ABI version . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1.2 Coverage . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1.3 Related projects . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

2 Usage . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

3 Types. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

4 Objects . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

4.1 HEADER . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 4.2 ENTITIES . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 4.3 OBJECTS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96

5 Sections. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 264

5.1 HEADER Section . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 264 5.2 OBJECTS Section . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 264 5.3 CLASSES Section . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 264 5.4 HANDLES Section . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 265 5.5 R2004 Header . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 265 5.6 UNKNOWN Section . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 265 5.7 SummaryInfo. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 265 5.8 Preview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 266 5.9 VBAProject. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 266 5.10 AppInfo. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 266 5.11 AppInfoHistory . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 266 5.12 FileDepList . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 266 5.13 AcDS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 266 5.14 RevHistory. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 266 5.15 Security. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 266 5.16 ObjFreeSpace . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 266 5.17 Template. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 266 5.18 AuxHeader. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 266 5.19 Signature . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 266 5.20 INFO . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 267 5.21 SYSTEM MAP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 267 5.22 Tables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 267 5.23 EXTRAS entities section . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 267

6 Structures . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 269

6.1 EED . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 269 6.2 XDATA . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 270

7 Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 271

7.1 Decoding. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 271 7.2 Encoding. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 272 7.3 add api . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 273 7.4 dynapi . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 274 7.5 strings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 276 7.6 Other Formats . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 276

7.6.1 DXF . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 276 7.6.2 DXFB. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 277 7.6.3 JSON . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 277 7.6.4 GeoJSON . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 277

8 Errors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 278

9 Programs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 279

10 Bindings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 282

11 Reference API . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 283

12 Reporting bugs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 284

Appendix A GNU Free Documentation License . . 285

13 Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 293

13.1 General Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 293 13.2 Object and Field Index. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 297

1 Overview

LibreDWG is a free C library to read and write DWG files. The DWG file format is proprietary and was created in the 1970s for one then-emerging CAD application.

This library is part of the GNU project, released under the aegis of GNU. It is made available under GPLv3+, i.e., under the terms of the GNU General Public License version 3, or (at your option) any later version.

It came out of code from the QCad Community Edition product from Ribbonsoft.

1.1 API/ABI version

LibreDWG 0.13.3 provides the API/ABI version 1. We hope that this doesn’t need to change much in the future.

   See include/dwg.h for LIBREDWG_VERSION_MAJOR, LIBREDWG_VERSION_MINOR and
LIBREDWG_SO_VERSION.

1.2 Coverage

Because the DWG file format is not open, its specification had to be reverse-engineered. The specification is almost complete. The LibreDWG implementation of the specification is an ongoing process; as of version 0.13.3, coverage is approximately 99%.

It can read all DWG formats from r1.2 to r2018 for 99%. The internally supported DWG revisions are identified as r1.1, r1.2, r1.3, r1.4, r2.0b, r2.0, r2.10, r2.21, r2.22, r2.4, r2.5, r2.6, r9, r9c1, r10, r11b1, r11b2, r11, r13b1, r13b2, r13, r13c3, r14, r2000b, r2000, r2000i, r2002, r2004a, r2004b, r2004c, r2004, r2007a, r2007b, r2007, r2010b, r2010, r2013b, r2013, r2018b and r2018.

Here is a list of features that are still missing:

Writing r2004+ DWG Writing DWG formats for r2004 and later: r2007, r2010, r2013, r2018 is an ongoing effort. You need to patch the code to enable writing to the r2004+ format. See the work/2004 branch. Writing to the specific r2007 format is not implemented. We write as default in the r2000 or older formats only.

Reading binary DXF DXF support is now mostly implemented. ASCII DXF’s are generated fully, with much more elements and fields and other free libraries, but AutoCAD ◯^{R} fails to import some of them. See the TODO file for a detailed coverage report. Reading binary DXF should be complete but is undertested.

Reading and writing pre-R13 DXF Reading and writing pre-R13 DXF’s is work in progress. There is no support planned for the very old pre-R2.0 DXF format.

Enabled entities and objects, but unstable, undertested. Field names may change: For a detailed overview see the file src/classes.inc or src/classes.c. ACSH BREP CLASS ACSH CHAMFER CLASS ACSH CONE CLASS

ACSH PYRAMID CLASS  ARC DIMENSION  ASSOCACTION

ASSOCBLENDSURFACEACTIONBODY ASSOCEXTENDSUR-FACEACTIONBODY ASSOCEXTRUDEDSURFACEACTIONBODY ASSOCFILLETSURFACEACTIONBODY ASSOCGEOMDEPEN-DENCY ASSOCLOFTEDSURFACEACTIONBODY ASSOCNETWORK ASSOCDEPENDENCY ASSOCVALUEDEPENDENCY ASSOCNETWORK-SURFACEACTIONBODY ASSOCOFFSETSURFACEACTIONBODY ASSOCPATCHSURFACEACTIONBODY ASSOCPLANESURFACEAC-TIONBODY ASSOCREVOLVEDSURFACEACTIONBODY ASSOCTRIM-SURFACEACTIONBODY BACKGROUND BLOCKLINEARPARAMETER BLOCKROTATIONPARAMETER BLOCKXYPARAMETER BLOCKVISI-BILITYGRIP BLOCKVISIBILITYPARAMETER EVALUATION GRAPH HELIX LARGE RADIAL DIMENSION LIGHTLIST MATERIAL MENTALRAYRENDERSETTINGS OBJECT PTR RAPIDRTRENDER-SETTINGS RENDERSETTINGS SECTION SETTINGS SPATIAL INDEX TABLESTYLE (works only pre-2010)

Unhandled (fields spec’ed, but broken/undertested):

ACMECOMMANDHISTORY ACMESCOPE ACMESTATEMGR ACSH EXTRUSION CLASS ACSH LOFT CLASS ACSH REVOLVE CLASS ACSH SWEEP CLASS ALDIMOBJECTCONTEXTDATA ANNOTSCA-LEOBJECTCONTEXTDATA ASSOC2DCONSTRAINTGROUP ASSOCACTION ASSOCALIGNEDDIMACTIONBODY ASSOCEX-TRUDEDSURFACEACTIONBODY ASSOCGEOMDEPENDENCY ASSOCLOFTEDSURFACEACTIONBODY ASSOCNETWORK ASSOCOSNAPPOINTREFACTIONPARAM ASSOCOSNAPPOINTRE-FACTIONPARAM ASSOCPERSSUBENTMANAGER ASSOCREVOLVED-SURFACEACTIONBODY ASSOCVERTEXACTIONPARAM ATEXT BLKREFOBJECTCONTEXTDATA CONTEXTDATAMANAGER CSACDOCUMENTOPTIONS CURVEPATH DATALINK DATATABLE DIMASSOC DYNAMICBLOCKPROXYNODE EXTRUDEDSURFACE FCFOBJECTCONTEXTDATA GEOMAPIMAGE GEOPOSITION-MARKER LAYOUTPRINTCONFIG LEADEROBJECTCONTEXTDATA LOFTEDSURFACE MLEADEROBJECTCONTEXTDATA MOTIONPATH MTEXTATTRIBUTEOBJECTCONTEXTDATA MTEXTOBJECT-CONTEXTDATA NAVISWORKSMODEL NAVISWORKSMODELDEF NURBSURFACE PERSUBENTMGR PLANESURFACE POINTPATH RENDERENTRY RENDERGLOBAL REVOLVEDSURFACE RTEXT SUN-STUDY SWEPTSURFACE TABLE (works only pre-2010) TABLECONTENT TEXTOBJECTCONTEXTDATA TVDEVICEPROPERTIES

ASSOCDIMDEPENDENCYBODY BLOCKPARAMDEPENDENCYBODY ALIGNMENTPARAMETERENTITY BASEPOINTPARAMETEREN-TITY FLIPPARAMETERENTITY LINEARPARAMETERENTITY POINTPARAMETERENTITY ROTATIONPARAMETERENTITY VISIBILITYPARAMETERENTITY VISIBILITYGRIPENTITY XYPA-RAMETERENTITY BLOCKALIGNEDCONSTRAINTPARAMETER BLOCKANGULARCONSTRAINTPARAMETER BLOCKARRAYACTION BLOCKDIAMETRICCONSTRAINTPARAMETER BLOCKHORIZONTAL-CONSTRAINTPARAMETER BLOCKLINEARCONSTRAINTPARAME-TER BLOCKLOOKUPACTION BLOCKLOOKUPPARAMETER BLOCK-POINTPARAMETER BLOCKPOLARGRIP BLOCKPOLARPARAMETER BLOCKPOLARSTRETCHACTION BLOCKPROPERTIESTABLE BLOCK-PROPERTIESTABLEGRIP BLOCKRADIALCONSTRAINTPARAMETER BLOCKREPRESENTATION BLOCKSTRETCHACTION BLOCKUSER-PARAMETER BLOCKVERTICALCONSTRAINTPARAMETER BLOCKXYGRIP POINTCLOUD POINTCLOUDEX POINTCLOUDDEF POINTCLOUDDEFEX POINTCLOUDDEF REACTOR POINTCLOUD-DEF REACTOR EX POINTCLOUDCOLORMAP

See src/classes.inc.

Missing:

* PROXY subentities, PROXY ENTITY Halfway:

SUNSTUDY  VBA PROJECT  ASSOCACTION  ASSOCNET-
WORK  ASSOCALIGNEDDIMACTIONBODY  ASSOCOSNAP-

POINTREFACTIONPARAM ASSOCPERSSUBENTMANAGER PERSUBENTMGR ASSOC2DCONSTRAINTGROUP EVAL-UATION GRAPH ASSOCOSNAPPOINTREFACTIONPARAM ACSH BOX CLASS ACSH EXTRUSION CLASS ACSH HISTORY CLASS ACSH SWEEP CLASS NAVISWORKSMODEL (i.e. COORDINATION MODEL) NAVISWORKSMODELDEF DATATABLE TABLESTYLE ASSOCGEOMDEPENDENCY LAYOUTPRINTCONFIG RENDERENVI-RONMENT RENDERGLOBAL LIGHTLIST SECTION SETTINGS

Unhandled (i.e. passed through, no DXF and fields):

ACDSRECORD ACDSSCHEMA NPOCOLLECTION RAPIDRTREN-DERENVIRONMENT XREFPANELOBJECT

no test coverage for entities:

I.e. we need an extended example_2018.dwg with all types, with the following missing entities: ARCALIGNEDTEXT BODY CAMERA DIMENSION ANG3PT DIMEN-SION DIAMETER DIMENSION RADIUS DGNUNDERLAY DWFUNDER-LAY GEOPOSITIONMARKER IMAGE LEADER LONG TRANSACTION MESH MINSERT OLE2FRAME OLEFRAME POLYLINE 2D POLY-LINE MESH PROXY ENTITY PROXY LWPOLYLINE SHAPE TOLERANCE VERTEX 2D VERTEX MESH and objects: CSACDOCUMENTOPTIONS XREFPANELOBJECT IDBUFFER IMAGEDEF IMAGEDEF REACTOR LAYER INDEX LIGHTLIST NPOCOLLECTION OBJECT PTR PLOTSETTINGS PROXY OBJECT RASTERVARIABLES SPATIAL INDEX UCS VBA PROJECT

1.3 Related projects

Some projects that use DWG (and specifically LibreDWG) are:

FreeCAD https://freecadweb.org/

GRASS GIS

          http://grass.osgeo.org/

Plans are to add support for SolveSpace, LibreCAD, FreeCAD, OpenSCAD and Python-CAD.

Related libraries:

libdwg The old version (documented in Esperanto) which was forked to LibreDWG in 2009. But in the meantime it got a DXF reader.

libdxfrw Read the DWG format for all versions r13+ but with much less elements, only those needed for DXF. Written in C++, under the GPLv2 license.

libopencad Read the r2000 DWG format. Written in C++, under the GPLv2 license.

ACadSharp Read and write the basic objects in most DWG and DXF formats. Written in C#, under the MIT license.

OpenDWG The OpenDWG’s license does not allow the usage in free software projects.

Compared to libdwg, libdxfrw and libopencad, LibreDWG can read and write much more DWG versions and details. Which is especially important for attached links and data from third party applications: BIM, MAP, GIS, AEC, MECH, ..., for 3D solids and dynamic parametric constraints.

2 Usage

This chapter describes how to compile and link a program against LibreDWG. To access LibreDWG interface elements (see Chapter 3 [Types], page 6, see Chapter 7 [Functions], page 271), include its header file in the C code.

     #include <dwg.h>

Optionally you can also use the

     #include <dwg_api.h>

API. Make sure you specify ‘-lredwg’ when linking, such as in this Makefile.am fragment:

     AM_LDFLAGS += -lredwg

Note that the shared object library is named libredwg (with some system-specific extension, e.g., .so), so you do not want to specify -llibredwg, as that would (try to) link against

liblibredwg and fail.

3 Types

LibreDWG types map closely to the type system of the DWG file format. This chapter describes the enums and structs used to define the single DWG structure, which is passed around the functions (see Chapter 7 [Functions], page 271).

[define] BITCODE_RC char

1 raw unsigned char, uint8 t

[define] BITCODE_RS short

1 raw unsigned short int, uint16 t

[define] BITCODE_RL long

1 raw unsigned long int, uint32 t

[define] BITCODE_RD double

1 raw IEEE-754 double

[define] BITCODE_B byte

1 bit

[define] BITCODE_BB byte

2 bits

[define] BITCODE_3B byte

1-3 bits

[define] BITCODE_4BITS byte

4 bits (for VIEW view mode)

[define] BITCODE_BS short

1 bit-encoded unsigned short

[define] BITCODE_BL long

1 bit-encoded unsigned long (max 32bit)

[define] BITCODE_BLL uint64_t

1 bit-encoded unsigned 64bit long

[define] BITCODE_BD double

1 bit-encoded double

[define] BITCODE_DD double

1 bit-encoded double with default

[define] BITCODE_MC long int

1-4 modular chars

[define] BITCODE_UMC long unsigned int

1-4 unsigned modular chars

[define] BITCODE_MS long unsigned int

1 modular short, max 2 words

[define] BITCODE_BE double[3]

1 bitencoded extrusion vector. Note that this specifies an OCS (Object Coordinate System) for each entity, with the default (0, 0, 1). An extrusion of (0, 0, -1) is typically caused by exploding a block inserted with a negative x scale, i.e. the sign of each X point needs to be flipped. For more see the vendor DXF documentation on OCS and programs/geom.c.

[define] BITCODE_BT double

1 bitencoded thickness value

[define] BITCODE_TV char*

length + ASCIIZ string The default text type until r2004.

[define] BITCODE_TU wchar*

length + windows 2-byte wchar string (UCS-2). The default text type since r2007.

[define] BITCODE_TF char*

Fixed length buffer, which can include NUL characters.

[define] BITCODE_TFF char*

Embedded fixed length string, which can include NUL characters.

[define] BITCODE_H void*

handle-references

[define] BITCODE_CMC struct Dwg Color

Dwg Color struct with index or rgb, alpha and optional DBCOLOR handle, name, book name.

[and some more] Two types that do not derive from the type system of the DWG file format are the

enums for return codes and error codes.

On non-C99 systems ensure that stdint.h and inttypes.h are available to use the proper C99 int32_t,... types, and not just the native fallback types int/long, which are different across platforms.

4 Objects

4.1 HEADER

All header variables.

ACADMAINTVER

RC, DXF 90

ANGBASE BD, DXF 50 ANGDIR B, DXF 70

APPID_CONTROL_OBJECT

H

ATTDIA B, DXF 70 ATTMODE BS, DXF 70 ATTREQ B, DXF 70 AUNITS BS, DXF 70 AUPREC BS, DXF 70

AXISMODE BS, DXF 70

AXISUNIT 2RD, DXF 20

BACKZ  BD
BLIPMODE B, DXF 70
BLOCK_CONTROL_OBJECT

H

BLOCK_RECORD_MSPACE

H

BLOCK_RECORD_PSPACE

H

CAMERADISPLAY

B, DXF 290

CAMERAHEIGHT

BD, DXF 40

CECOLOR CMC, DXF 62

CELTSCALE

BD, DXF 40

CELTYPE H, DXF 6
CELWEIGHT

BSd, DXF 370

CEPSNTYPE

BS, DXF 380

CHAMFERA BD, DXF 40
CHAMFERB BD, DXF 40
CHAMFERC BD, DXF 40
CHAMFERD BD, DXF 40
CLAYER  H, DXF 8
CMATERIAL

H, DXF 347

CMLJUST BS, DXF 70

CMLSCALE BD, DXF 40
CMLSTYLE H, DXF 2

COORDS BS, DXF 70 CPSNID H, DXF 390 CSHADOW RC, DXF 280 DELOBJ B, DXF 70 DGNFRAME RC, DXF 280

DICTIONARY_ACAD_GROUP

H

DICTIONARY_ACAD_MLINESTYLE

H

DICTIONARY_COLOR

H

DICTIONARY_LAYOUT

H

DICTIONARY_LIGHTLIST

H

DICTIONARY_MATERIAL

H

DICTIONARY_NAMED_OBJECT

H

DICTIONARY_PLOTSETTINGS

H

DICTIONARY_PLOTSTYLENAME

H

DICTIONARY_VISUALSTYLE

H DIMADEC BS, DXF 70 DIMALT B, DXF 70 DIMALTD BS, DXF 70 DIMALTF BD, DXF 40

DIMALTMZF

BD

DIMALTMZS

T, DXF 1

DIMALTRND

BD, DXF 40

DIMALTTD BS, DXF 70
DIMALTTZ BS, DXF 70

DIMALTU BS, DXF 70 DIMALTZ BS, DXF 70

DIMAPOST TV, DXF 1
DIMAPOST_T

T

DIMARCSYM

BS, DXF 70

DIMARROW BD

DIMASO B, DXF 70 DIMASSOC RC, DXF 280 DIMASZ BD, DXF 40

DIMATFIT BS, DXF 70
DIMAUNIT BS, DXF 70

DIMAZIN BS, DXF 70

DIMBLK  H, DXF 1
DIMBLK1 H, DXF 1
DIMBLK1_T

T

DIMBLK2 H, DXF 1
DIMBLK2_T

T

DIMBLK_T T

DIMCEN BD, DXF 40 DIMCLRD CMC, DXF 70

DIMCLRD_C

RS

DIMCLRE CMC, DXF 70

DIMCLRE_C

RS

DIMCLRT CMC, DXF 70

DIMCLRT_C

RS

DIMDEC BS, DXF 70 DIMDLE BD, DXF 40 DIMDLI BD, DXF 40 DIMDSEP BS, DXF 70 DIMEXE BD, DXF 40 DIMEXO BD, DXF 40 DIMFIT BS, DXF 70 DIMFRAC BS, DXF 70 DIMFXL BD, DXF 40

DIMFXLON B, DXF 70

DIMGAP BD, DXF 40

DIMJOGANG

BD, DXF 40

DIMJUST BS, DXF 70

DIMLDRBLK

H, DXF 1

DIMLFAC BD, DXF 40 DIMLIM B, DXF 70

DIMLTEX1 H, DXF 6
DIMLTEX2 H, DXF 6
DIMLTYPE H, DXF 6
DIMLUNIT BS, DXF 70

DIMLWD BSd, DXF 70 DIMLWE BSd, DXF 70

DIMMZF  BD

DIMMZS  T, DXF 1
DIMPOST TV, DXF 1
DIMPOST_T

T

DIMRND BD, DXF 40 DIMSAH B, DXF 70

DIMSAV  B
DIMSCALE BD, DXF 40

DIMSD1 B, DXF 70 DIMSD2 B, DXF 70 DIMSE1 B, DXF 70 DIMSE2 B, DXF 70 DIMSHO B, DXF 70

DIMSOXD B, DXF 70
DIMSTYLE H, DXF 2
DIMSTYLE_CONTROL_OBJECT

H

DIMTAD BS, DXF 70 DIMTDEC BS, DXF 70 DIMTFAC BD, DXF 40

DIMTFILL BS, DXF 70
DIMTFILLCLR

CMC, DXF 70

DIMTIH B, DXF 70 DIMTIX B, DXF 70 DIMTM BD, DXF 40

DIMTMOVE BS, DXF 70
DIMTOFL B, DXF 70

DIMTOH B, DXF 70 DIMTOL B, DXF 70 DIMTOLJ BS, DXF 70 DIMTP BD, DXF 40 DIMTSZ BD, DXF 40 DIMTVP BD, DXF 40

DIMTXSTY H, DXF 7

DIMTXT BD, DXF 40

DIMTXTDIRECTION

B, DXF 70

DIMTZIN BS, DXF 70 DIMUNIT BS, DXF 70 DIMUPT B, DXF 70 DIMZIN BS, DXF 70

DISPSILH B, DXF 70
DRAGMODE BS, DXF 70

DRAGVS H, DXF 349 DWFFRAME RC, DXF 280

DWGCODEPAGE

TV, DXF 3

ELEVATION

BD, DXF 40

ENDCAPS B, DXF 280 EXTMAX 3BD, DXF 30 EXTMIN 3BD, DXF 30

EXTNAMES B, DXF 290
FACETRES BD
FASTZOOM B, DXF 70
FILLETRAD

BD, DXF 40

FILLMODE B, DXF 70
FINGERPRINTGUID

TV, DXF 2

FLAGS  BL
FLATLAND B, DXF 70
FRONTZ  BD
GRIDMODE RS, DXF 70

GRIDUNIT 2RD, DXF 20 HALOGAP RC, DXF 280

HANDLING BS, DXF 70

HANDSEED H, DXF 5

HIDETEXT RC, DXF 280

HYPERLINKBASE

T, DXF 1

INDEXCTL RC, DXF 280 INSBASE 3BD, DXF 30

INSUNITS BS, DXF 70
INTERFERECOLOR

CMC, DXF 62

INTERFEREOBJVS

H, DXF 345

INTERFEREVPVS

H, DXF 346

INTERSECTIONCOLOR

BS, DXF 70

INTERSECTIONDISPLAY

RC, DXF 280

ISOLINES BS
JOINSTYLE

B, DXF 280

LATITUDE BD, DXF 40
LAYER_CONTROL_OBJECT

H

LENSLENGTH

BD, DXF 40

LIGHTGLYPHDISPLAY

RC, DXF 280

LIMCHECK B, DXF 70

LIMMAX 2DPOINT, DXF 20 LIMMIN 2DPOINT, DXF 20

LOFTANG1 BD, DXF 40
LOFTANG2 BD, DXF 40
LOFTMAG1 BD, DXF 40
LOFTMAG2 BD, DXF 40
LOFTNORMALS

RC, DXF 280

LOFTPARAM

BS, DXF 70

LONGITUDE

BD, DXF 40

LTSCALE BD, DXF 40

LTYPE_BYBLOCK

H

LTYPE_BYLAYER

H

LTYPE_CONTINUOUS

H

LTYPE_CONTROL_OBJECT

H

LUNITS BS, DXF 70 LUPREC BS, DXF 70

LWDISPLAY

B, DXF 290

MAXACTVP BS, DXF 70
MEASUREMENT

BS, DXF 70

MENU TV, DXF 1

MENUEXT RC
MIRRTEXT B, DXF 70
NORTHDIRECTION

BD, DXF 40

OBSCOLOR BS, DXF 70

OBSLTYPE RC, DXF 280

OLESTARTUP

B, DXF 290

ORTHOMODE

B, DXF 70

OSMODE BS, DXF 70 PDMODE BS, DXF 70 PDSIZE BD, DXF 40

PELEVATION

BD, DXF 40

PELLIPSE B

PEXTMAX 3BD, DXF 30 PEXTMIN 3BD, DXF 30

PICKSTYLE

BS, DXF 70

PINSBASE 3BD, DXF 30

PLIMCHECK

B, DXF 70

PLIMMAX 2DPOINT, DXF 20 PLIMMIN 2DPOINT, DXF 20

PLINEGEN B, DXF 70
PLINEWID BD, DXF 40
PROJECTNAME

TV, DXF 1

PROXYGRAPHICS

BS, DXF 70

PSLTSCALE

B, DXF 70

PSOLHEIGHT

BD, DXF 40

PSOLWIDTH

BD, DXF 40

PSTYLEMODE

B, DXF 290

PSVPSCALE

BD, DXF 40

PUCSBASE H, DXF 2
PUCSNAME H, DXF 2

PUCSORG 3BD, DXF 30

PUCSORGBACK

3BD, DXF 30

PUCSORGBOTTOM

3BD, DXF 30

PUCSORGFRONT

3BD, DXF 30

PUCSORGLEFT

3BD, DXF 30

PUCSORGRIGHT

3BD, DXF 30

PUCSORGTOP

3BD, DXF 30

PUCSORTHOREF

H, DXF 2

PUCSORTHOVIEW

BS, DXF 70

PUCSXDIR 3BD, DXF 30 PUCSYDIR 3BD, DXF 30

QTEXTMODE

B, DXF 70

REALWORLDSCALE

B, DXF 290

REGENMODE

B, DXF 70

REQUIREDVERSIONS

BLL, DXF 160

SAVEIMAGES

BS

SHADEDGE BS, DXF 70
SHADEDIF BS, DXF 70
SHADOWPLANELOCATION

BD, DXF 40

SHOWHIST RC, DXF 280

SKETCHINC

BD, DXF 40

SKPOLY B, DXF 70 SNAPANG RD, DXF 50 SNAPBASE 2RD, DXF 20

SNAPISOPAIR

RS, DXF 70

SNAPMODE RS, DXF 70
SNAPSTYLE

RS, DXF 70

SNAPUNIT 2RD, DXF 20

SOLIDHIST

RC, DXF 280

SORTENTS RC, DXF 280

SPLFRAME B, DXF 70
SPLINESEGS

BS, DXF 70

SPLINETYPE

BS, DXF 70

STEPSIZE BD, DXF 40
STEPSPERSEC

BD, DXF 40

STYLESHEET

TV, DXF 1

STYLE_CONTROL_OBJECT

H

SURFTAB1 BS, DXF 70
SURFTAB2 BS, DXF 70
SURFTYPE BS, DXF 70

SURFU BS, DXF 70 SURFV BS, DXF 70

TARGET  3BD

TDCREATE TIMEBLL, DXF 40 TDINDWG TIMEBLL, DXF 40

TDUCREATE

TIMEBLL, DXF 40

TDUPDATE TIMEBLL, DXF 40

TDUSRTIMER

TIMEBLL, DXF 40

TDUUPDATE

TIMEBLL, DXF 40

TEXTQLTY BS
TEXTSIZE BD, DXF 40
TEXTSTYLE

H, DXF 7

THICKNESS

BD, DXF 40

TILEMODE B, DXF 70
TILEMODELIGHTSYNCH

RC, DXF 280

TIMEZONE BL, DXF 70
TRACEWID BD, DXF 40
TREEDEPTH

BSd, DXF 70

TSTACKALIGN

BS

TSTACKSIZE

BS

UCSBASE H, DXF 2
UCSICON RC
UCSNAME H, DXF 2

UCSORG 3BD, DXF 30

UCSORGBACK

3BD, DXF 30

UCSORGBOTTOM

3BD, DXF 30

UCSORGFRONT

3BD, DXF 30

UCSORGLEFT

3BD, DXF 30

UCSORGRIGHT

3BD, DXF 30

UCSORGTOP

3BD, DXF 30

UCSORTHOREF

H, DXF 2

UCSORTHOVIEW

BS, DXF 70

UCSXDIR 3BD, DXF 30 UCSYDIR 3BD, DXF 30

UCS_CONTROL_OBJECT

H

UNITMODE BS, DXF 70

USERI1 BSd, DXF 70 USERI2 BSd, DXF 70 USERI3 BSd, DXF 70 USERI4 BSd, DXF 70 USERI5 BSd, DXF 70 USERR1 BD, DXF 40 USERR2 BD, DXF 40 USERR3 BD, DXF 40 USERR4 BD, DXF 40 USERR5 BD, DXF 40

USRTIMER B, DXF 70
VERSIONGUID

TV, DXF 2

VIEWCTR 3RD, DXF 20 VIEWDIR 3BD, DXF 30

VIEWMODE B
VIEWSIZE RD, DXF 40
VIEWTWIST

BD

VIEW_CONTROL_OBJECT

H

VISRETAIN

B, DXF 70

VPOINTX 3RD
VPOINTXALT

3RD

VPOINTY 3RD
VPOINTYALT

3RD

VPOINTZ 3RD
VPOINTZALT

3RD

VPORT_CONTROL_OBJECT

H

VX_CONTROL_OBJECT

H

VX_TABLE_RECORD

H

WIREFRAME

B

WORLDVIEW

B, DXF 70

XCLIPFRAME

RC, DXF 290

XEDIT B, DXF 290

_3DDWFPREC

BD, DXF 40

aspect_ratio

RD

bitsize RL
bitsize_hi

RL

circle_zoom_percent

RS

codepage RS
dwg_size RL
flag_3d RS
layer_colors

RS

numentities

RS

oldCECOLOR_hi

RL

oldCECOLOR_lo

RL

size  RL
unit1_name

TV

unit1_ratio

BD

unit2_name

TV

unit2_ratio

BD

unit3_name

TV

unit3_ratio

BD

unit4_name

TV

unit4_ratio

BD

unknown_10

BS

unknown_11

B

unknown_12

BL

unknown_13

BL

unknown_14

BL

unknown_14b

BL

unknown_15

BL

unknown_16

BL

unknown_17

BL

unknown_20

H

unknown_21

BL

unknown_22

BL

unknown_23

BD

unknown_4f2

RL

unknown_5

RS

unknown_51e

RS

unknown_520

RS

unknown_52c

RSd

unknown_52e

RS

unknown_530

RC

unknown_54

BS

unknown_55

BS

unknown_56

BS

unknown_57

BS

unknown_59

RS

unknown_6

RS

unknown_6a

RS

unknown_6b

RS

unknown_6c

RS

unknown_8

BL

unknown_9

BL

unknown_day

RS

unknown_hour

RS

unknown_min

RS

unknown_mon

RS

unknown_ms

RS

unknown_sec

RS

unknown_string

T

unknown_year

RS

4.2 ENTITIES

All graphical objects with its fields. See [Common Entity fields], page 260,

3DFACE parent struct dwg object entity*

has_no_flags

B

z_is_zero

B

corner1 3BD, DXF 10 corner2 3BD, DXF 11 corner3 3BD, DXF 12 corner4 3BD, DXF 13

invis_flags

BS, DXF 70

3DLINE parent struct dwg object entity* start 3RD, DXF 10 end 3RD, DXF 11

extrusion

3RD, DXF 210

thickness

RD, DXF 39

3DSOLID parent struct dwg object entity*

acis_empty

B, DXF 290

unknown B

version BS, DXF 70

num_blocks

BL

block_size

BL*

encr_sat_data

char **, DXF 1

sab_size BL
acis_data

RC*

wireframe_data_present

B

point_present

B

point  3BD
isolines BL
isoline_present

B

num_wires

BL

wires Dwg 3DSOLID wire*

num_silhouettes

BL

silhouettes

Dwg 3DSOLID silhouette*

_dxf_sab_converted

B

acis_empty2

B

extra_acis_data

struct dwg entity 3DSOLID*

num_materials

BL

materials

Dwg 3DSOLID material*

revision_guid

RC, DXF 2

revision_major

BL

revision_minor1

BS

revision_minor2

BS

revision_bytes

RC

end_marker

BL

history_id

H, DXF 350

has_revision_guid

B

acis_empty_bit

B

ALIGNMENTPARAMETERENTITY

parent struct dwg object entity* ARC

parent struct dwg object entity* center 3BD, DXF 10 radius BD, DXF 40

thickness

BT, DXF 39

extrusion

BE, DXF 210

start_angle

BD, DXF 50

end_angle

BD, DXF 51

ARCALIGNEDTEXT parent struct dwg object entity*

text_size

D2T, DXF 42

xscale D2T, DXF 41

char_spacing

D2T, DXF 43 style T, DXF 7 t2 T, DXF 2 t3 T, DXF 3

text_value

T, DXF 1

offset_from_arc

D2T, DXF 44

right_offset

D2T, DXF 45

left_offset

D2T, DXF 46

center 3BD, DXF 10 radius BD, DXF 40

start_angle

BD, DXF 50

end_angle

BD, DXF 51

extrusion

3BD, DXF 210

color BL, DXF 90

is_reverse

BS, DXF 70

text_direction

BS, DXF 71

alignment

BS, DXF 72

text_position

BS, DXF 73

font_19 BS, DXF 74 bs2 BS, DXF 75

is_underlined

BS, DXF 76

bs1 BS, DXF 77 font BS, DXF 78 is_shx BS, DXF 79

wizard_flag

BS, DXF 280

arc_handle

H, DXF 330

ARC DIMENSION parent struct dwg object entity*

class_version

RC, DXF 280

extrusion

BE, DXF 210

def_pt 3BD, DXF 10

text_midpt

2RD, DXF 11

elevation

BD, DXF 31

flag RC, DXF 70

flag1  RC
user_text

T, DXF 1

text_rotation

BD, DXF 53

horiz_dir

BD, DXF 51

ins_scale

3BD 1

ins_rotation

BD, DXF 54

attachment

BS, DXF 71

lspace_style

BS, DXF 72

lspace_factor

BD, DXF 41

act_measurement

BD, DXF 42

unknown B, DXF 73
flip_arrow1

B, DXF 74

flip_arrow2

B, DXF 75

clone_ins_pt

2RD, DXF 12

dimstyle H, DXF 3

block H, DXF 2

xline1_pt

3BD, DXF 13

xline2_pt

3BD, DXF 14

center_pt

3BD, DXF 15

is_partial

B, DXF 70

arc_start_param

BD, DXF 41

arc_end_param

BD, DXF 42

has_leader

B, DXF 71

leader1_pt

3BD, DXF 16

leader2_pt

3BD, DXF 17

ATTDEF parent struct dwg object entity*

elevation

BD, DXF 30

ins_pt 2DPOINT, DXF 10

alignment_pt

2DPOINT, DXF 11

extrusion

BE, DXF 210

thickness

RD, DXF 39

oblique_angle

RD, DXF 51

rotation RD, DXF 50

height RD, DXF 40

width_factor

RD, DXF 41

default_value

T, DXF 1

generation

BS, DXF 71

horiz_alignment

BS, DXF 72

vert_alignment

BS, DXF 74

dataflags

RC

class_version

RC, DXF 280

type RC, DXF 70 tag T, DXF 2

field_length

BS

flags RC, DXF 70

lock_position_flag

B

style H, DXF 7

mtext_style

H, DXF 340

annotative_data_size

BS, DXF 70

annotative_data_bytes

RC

annotative_app

H

annotative_short

BS

attdef_class_version

RC

prompt  T, DXF 3

ATTRIB parent struct dwg object entity*

elevation

BD, DXF 30

ins_pt 2DPOINT, DXF 10

alignment_pt

2DPOINT, DXF 11

extrusion

BE, DXF 210

thickness

RD, DXF 39

oblique_angle

RD, DXF 51

rotation RD, DXF 50

height RD, DXF 40

width_factor

RD, DXF 41

text_value

T, DXF 1

generation

BS, DXF 71

horiz_alignment

BS, DXF 72

vert_alignment

BS, DXF 74

dataflags

RC

class_version

RC, DXF 280

type RC, DXF 70 tag T, DXF 2

field_length

BS

flags RC, DXF 70

lock_position_flag

B

style H, DXF 7

mtext_style

H, DXF 340

annotative_data_size

BS, DXF 70

annotative_data_bytes

RC

annotative_app

H

annotative_short

BS

BASEPOINTPARAMETERENTITY

parent struct dwg object entity* BLOCK

parent struct dwg object entity* name TV, DXF 2

xref_pname

TV, DXF 1

base_pt 2RD, DXF 10 BODY

See [3DSOLID], page 24, CAMERA

parent struct dwg object entity*

view  H

CIRCLE parent struct dwg object entity* center 3BD, DXF 10 radius BD, DXF 40

thickness

BT, DXF 39

extrusion

BE, DXF 210

DGNUNDERLAY parent struct dwg object entity*

extrusion

BE, DXF 210

ins_pt 3BD, DXF 10 scale 3BD 1, DXF 41 angle BD, DXF 50 flag RC, DXF 280 contrast RC, DXF 281 fade RC, DXF 282

num_clip_verts

BL

clip_verts

2RD*, DXF 11

num_clip_inverts

BS, DXF 170

clip_inverts

2RD*, DXF 12

definition_id

H, DXF 340

DIMENSION ALIGNED parent struct dwg object entity*

class_version

RC, DXF 280

extrusion

BE, DXF 210

def_pt 3BD, DXF 10

text_midpt

2RD, DXF 11

elevation

BD, DXF 31

flag RC, DXF 70

flag1  RC
user_text

T, DXF 1

text_rotation

BD, DXF 53

horiz_dir

BD, DXF 51

ins_scale

3BD 1

ins_rotation

BD, DXF 54

attachment

BS, DXF 71

lspace_style

BS, DXF 72

lspace_factor

BD, DXF 41

act_measurement

BD, DXF 42

unknown B, DXF 73
flip_arrow1

B, DXF 74

flip_arrow2

B, DXF 75

clone_ins_pt

2RD, DXF 12

dimstyle H, DXF 3

block H, DXF 2

xline1_pt

3BD, DXF 13

xline2_pt

3BD, DXF 14

oblique_angle

BD, DXF 50

DIMENSION ANG2LN parent struct dwg object entity*

class_version

RC, DXF 280

extrusion

BE, DXF 210

def_pt 3BD, DXF 10

text_midpt

2RD, DXF 11

elevation

BD, DXF 31

flag RC, DXF 70

flag1  RC
user_text

T, DXF 1

text_rotation

BD, DXF 53

horiz_dir

BD, DXF 51

ins_scale

3BD 1

ins_rotation

BD, DXF 54

attachment

BS, DXF 71

lspace_style

BS, DXF 72

lspace_factor

BD, DXF 41

act_measurement

BD, DXF 42

unknown B, DXF 73
flip_arrow1

B, DXF 74

flip_arrow2

B, DXF 75

clone_ins_pt

2RD, DXF 12

dimstyle H, DXF 3

block H, DXF 2

xline1start_pt

3BD, DXF 13

xline1end_pt

3BD, DXF 14

xline2start_pt

3BD, DXF 15

xline2end_pt

3BD, DXF 16

DIMENSION ANG3PT parent struct dwg object entity*

class_version

RC, DXF 280

extrusion

BE, DXF 210

def_pt 3BD, DXF 10

text_midpt

2RD, DXF 11

elevation

BD, DXF 31

flag RC, DXF 70

flag1  RC
user_text

T, DXF 1

text_rotation

BD, DXF 53

horiz_dir

BD, DXF 51

ins_scale

3BD 1

ins_rotation

BD, DXF 54

attachment

BS, DXF 71

lspace_style

BS, DXF 72

lspace_factor

BD, DXF 41

act_measurement

BD, DXF 42

unknown B, DXF 73
flip_arrow1

B, DXF 74

flip_arrow2

B, DXF 75

clone_ins_pt

2RD, DXF 12

dimstyle H, DXF 3

block H, DXF 2

xline1_pt

3BD, DXF 13

xline2_pt

3BD, DXF 14

center_pt

3BD, DXF 15

xline2end_pt

3RD

DIMENSION DIAMETER parent struct dwg object entity*

class_version

RC, DXF 280

extrusion

BE, DXF 210

def_pt 3BD, DXF 10

text_midpt

2RD, DXF 11

elevation

BD, DXF 31

flag RC, DXF 70

flag1  RC
user_text

T, DXF 1

text_rotation

BD, DXF 53

horiz_dir

BD, DXF 51

ins_scale

3BD 1

ins_rotation

BD, DXF 54

attachment

BS, DXF 71

lspace_style

BS, DXF 72

lspace_factor

BD, DXF 41

act_measurement

BD, DXF 42

unknown B, DXF 73

flip_arrow1

B, DXF 74

flip_arrow2

B, DXF 75

clone_ins_pt

2RD, DXF 12

dimstyle H, DXF 3

block H, DXF 2

first_arc_pt

3BD, DXF 15

leader_len

BD, DXF 40

DIMENSION LINEAR parent struct dwg object entity*

class_version

RC, DXF 280

extrusion

BE, DXF 210

def_pt 3BD, DXF 10

text_midpt

2RD, DXF 11

elevation

BD, DXF 31

flag RC, DXF 70

flag1  RC
user_text

T, DXF 1

text_rotation

BD, DXF 53

horiz_dir

BD, DXF 51

ins_scale

3BD 1

ins_rotation

BD, DXF 54

attachment

BS, DXF 71

lspace_style

BS, DXF 72

lspace_factor

BD, DXF 41

act_measurement

BD, DXF 42

unknown B, DXF 73
flip_arrow1

B, DXF 74

flip_arrow2

B, DXF 75

clone_ins_pt

2RD, DXF 12

dimstyle H, DXF 3

block H, DXF 2

xline1_pt

3BD, DXF 13

xline2_pt

3BD, DXF 14

oblique_angle

BD, DXF 52

dim_rotation

BD, DXF 50

DIMENSION ORDINATE parent struct dwg object entity*

class_version

RC, DXF 280

extrusion

BE, DXF 210

def_pt 3BD, DXF 10

text_midpt

2RD, DXF 11

elevation

BD, DXF 31

flag RC, DXF 70

flag1  RC
user_text

T, DXF 1

text_rotation

BD, DXF 53

horiz_dir

BD, DXF 51

ins_scale

3BD 1

ins_rotation

BD, DXF 54

attachment

BS, DXF 71

lspace_style

BS, DXF 72

lspace_factor

BD, DXF 41

act_measurement

BD, DXF 42

unknown B, DXF 73
flip_arrow1

B, DXF 74

flip_arrow2

B, DXF 75

clone_ins_pt

2RD, DXF 12

dimstyle H, DXF 3

block H, DXF 2

feature_location_pt

3BD, DXF 13

leader_endpt

3BD, DXF 14

flag2  RC

DIMENSION RADIUS parent struct dwg object entity*

class_version

RC, DXF 280

extrusion

BE, DXF 210

def_pt 3BD, DXF 10

text_midpt

2RD, DXF 11

elevation

BD, DXF 31

flag RC, DXF 70

flag1  RC
user_text

T, DXF 1

text_rotation

BD, DXF 53

horiz_dir

BD, DXF 51

ins_scale

3BD 1

ins_rotation

BD, DXF 54

attachment

BS, DXF 71

lspace_style

BS, DXF 72

lspace_factor

BD, DXF 41

act_measurement

BD, DXF 42

unknown B, DXF 73
flip_arrow1

B, DXF 74

flip_arrow2

B, DXF 75

clone_ins_pt

2RD, DXF 12

dimstyle H, DXF 3

block H, DXF 2

first_arc_pt

3BD, DXF 15

leader_len

BD, DXF 40 DWFUNDERLAY

See [UNDERLAY], page 32, ELLIPSE

parent struct dwg object entity* center 3BD, DXF 10 sm_axis 3BD, DXF 11

extrusion

BE, DXF 210

axis_ratio

BD, DXF 40

start_angle

BD, DXF 41

end_angle

BD, DXF 42

ENDBLK

parent struct dwg object entity* ENDREP

parent struct dwg object entity* numcols RS, DXF 70 numrows RS, DXF 71

colspacing

RD, DXF 40

rowspacing

RD, DXF 41

EXTRUDEDSURFACE parent struct dwg object entity*

acis_empty

B, DXF 290

unknown B

version BS, DXF 70

num_blocks

BL

block_size

BL*

encr_sat_data

char **, DXF 1

sab_size BL
acis_data

RC*

wireframe_data_present

B

point_present

B

point  3BD
isolines BL
isoline_present

B

num_wires

BL

wires Dwg 3DSOLID wire*

num_silhouettes

BL

silhouettes

Dwg 3DSOLID silhouette*

_dxf_sab_converted

B

acis_empty2

B

extra_acis_data

struct dwg entity 3DSOLID*

num_materials

BL

materials

Dwg 3DSOLID material*

revision_guid

RC, DXF 2

revision_major

BL

revision_minor1

BS

revision_minor2

BS

revision_bytes

RC

end_marker

BL

history_id

H, DXF 350

has_revision_guid

B

acis_empty_bit

B

modeler_format_version

BS

bindata_size

BL

bindata TF
u_isolines

BS, DXF 71

v_isolines

BS, DXF 72

class_version

BL

draft_angle

BD, DXF 42

draft_start_distance

BD, DXF 43

draft_end_distance

BD, DXF 44

twist_angle

BD, DXF 45

scale_factor

BD, DXF 48

align_angle

BD, DXF 49

sweep_entity_transmatrix

BD*, DXF 46

path_entity_transmatrix

BD*, DXF 47

is_solid B, DXF 290
sweep_alignment_flags

BS, DXF 70

path_flags

BS, DXF 71

align_start

B, DXF 292

bank B, DXF 293

base_point_set

B, DXF 294

sweep_entity_transform_computed

B, DXF 295

path_entity_transform_computed

B, DXF 296

reference_vector_for_controlling_twist

3BD, DXF 11

sweep_entity

H

path_entity

H

sweep_vector

3BD, DXF 10

sweep_transmatrix

BD*, DXF 40

FLIPGRIPENTITY

parent struct dwg object entity* FLIPPARAMETERENTITY

parent struct dwg object entity* GEOPOSITIONMARKER

parent struct dwg object entity*

class_version

BS, DXF 90

position 3BD, DXF 10 radius BD, DXF 40

landing_gap

BD, DXF 40

notes T, DXF 1

text_alignment

RC, DXF 280

mtext_visible

B, DXF 290

enable_frame_text

B, DXF 290

mtext struct dwg object* HATCH

parent struct dwg object entity*

is_gradient_fill

BL, DXF 450

reserved BL, DXF 451

gradient_angle

BD, DXF 460

gradient_shift

BD, DXF 461

single_color_gradient

BL, DXF 452

gradient_tint

BD, DXF 462

num_colors

BL, DXF 453

colors Dwg HATCH Color*

gradient_name

T, DXF 470

elevation

BD, DXF 30

extrusion

BE, DXF 210

name T, DXF 2

is_solid_fill

B, DXF 70

is_associative

B, DXF 71

num_paths

BL, DXF 91

paths Dwg HATCH Path* style BS, DXF 75

pattern_type

BS, DXF 76 angle BD, DXF 52

scale_spacing

BD, DXF 41

double_flag

B, DXF 77

num_deflines

BS, DXF 78

deflines Dwg HATCH DefLine*

has_derived

B

pixel_size

BD, DXF 47

num_seeds

BL, DXF 98

seeds 2RD*, DXF 10 HELIX

parent struct dwg object entity* flag BS, DXF 70

scenario BS

degree BS, DXF 71

splineflags

BL

knotparam

BL

fit_tol BD, DXF 44

beg_tan_vec

3BD, DXF 12

end_tan_vec

3BD, DXF 13

rational B
closed_b B
periodic B
weighted B
knot_tol BD, DXF 42
ctrl_tol BD, DXF 43

num_fit_pts

BS, DXF 74

fit_pts 3DPOINT*, DXF 11

num_knots

BL, DXF 72

knots BD*, DXF 40

num_ctrl_pts

BL, DXF 73

ctrl_pts Dwg SPLINE control point*

major_version

BL, DXF 90

maint_version

BL, DXF 91

axis_base_pt

3BD, DXF 10

start_pt 3BD, DXF 11

axis_vector

3BD, DXF 12

radius BD, DXF 40 turns BD, DXF 41

turn_height

BD, DXF 42

handedness

B, DXF 290

constraint_type

RC, DXF 280

IMAGE parent struct dwg object entity*

class_version

BL, DXF 90

pt0 3BD, DXF 10 uvec 3BD, DXF 11 vvec 3BD, DXF 12 size 2RD, DXF 13

display_props

BS, DXF 70

clipping B, DXF 280
brightness

RC, DXF 281

contrast RC, DXF 282 fade RC, DXF 283

clip_mode

B, DXF 290

clip_boundary_type

BS, DXF 71

num_clip_verts

BL, DXF 91

clip_verts

2RD*, DXF 14

imagedef H, DXF 340
imagedefreactor

H, DXF 360

INSERT parent struct dwg object entity* ins_pt 3DPOINT, DXF 10

scale_flag

BB

scale 3BD 1, DXF 41

rotation BD, DXF 50
extrusion

BE, DXF 210

has_attribs

B, DXF 66

num_owned

BL

block_header

H, DXF 2

first_attrib

H

last_attrib

H

attribs H*
seqend  H

num_cols RS, DXF 70
num_rows RS, DXF 71
col_spacing

RD, DXF 44

row_spacing

RD, DXF 45

block_name

TV, DXF 2

JUMP parent struct dwg object entity*

jump_address_raw

RL

jump_address

RL

jump_entity_section

Dwg Entity Sections

LARGE RADIAL DIMENSION parent struct dwg object entity*

class_version

RC, DXF 280

extrusion

BE, DXF 210

def_pt 3BD, DXF 10

text_midpt

2RD, DXF 11

elevation

BD, DXF 31

flag RC, DXF 70

flag1  RC
user_text

T, DXF 1

text_rotation

BD, DXF 53

horiz_dir

BD, DXF 51

ins_scale

3BD 1

ins_rotation

BD, DXF 54

attachment

BS, DXF 71

lspace_style

BS, DXF 72

lspace_factor

BD, DXF 41

act_measurement

BD, DXF 42

unknown B, DXF 73
flip_arrow1

B, DXF 74

flip_arrow2

B, DXF 75

clone_ins_pt

2RD, DXF 12

dimstyle H, DXF 3

block H, DXF 2

first_arc_pt

3BD, DXF 15

leader_len

BD, DXF 40

ovr_center

3BD, DXF 12

jog_point

3BD, DXF 13

LAYOUTPRINTCONFIG parent struct dwg object entity*

class_version

BS

flag BS, DXF 93 LEADER

parent struct dwg object entity*

unknown_bit_1

B

path_type

BS, DXF 72

annot_type

BS, DXF 73

num_points

BL, DXF 76

points 3DPOINT*, DXF 10 origin 3DPOINT

extrusion

BE, DXF 210

x_direction

3DPOINT, DXF 211

inspt_offset

3DPOINT, DXF 212

endptproj

3DPOINT, DXF 213

dimgap  BD
box_height

BD, DXF 40

box_width

BD, DXF 41

hookline_dir

B, DXF 74

hookline_on

B, DXF 75

arrowhead_on

B, DXF 71

arrowhead_type

BS

dimasz  BD
unknown_bit_2

B

unknown_bit_3

B

unknown_short_1

BS

byblock_color

BS, DXF 77

unknown_bit_4

B

unknown_bit_5

B

associated_annotation

H, DXF 340

dimstyle H, DXF 3

LIGHT parent struct dwg object entity*

class_version

BL, DXF 90

name T, DXF 1 type BL, DXF 70 status B, DXF 290

light_color

CMC, DXF 63

plot_glyph

B, DXF 291

intensity

BD, DXF 40

position 3BD, DXF 10 target 3BD, DXF 11

attenuation_type

BL, DXF 72

use_attenuation_limits

B, DXF 292

attenuation_start_limit

BD, DXF 41

attenuation_end_limit

BD, DXF 42

hotspot_angle

BD, DXF 50

falloff_angle

BD, DXF 51

cast_shadows

B, DXF 293

shadow_type

BL, DXF 73

shadow_map_size

BS, DXF 91

shadow_map_softness

RC, DXF 280

is_photometric

B

has_photometric_data

B, DXF 1

has_webfile

B, DXF 290

webfile T, DXF 300

physical_intensity_method

BS, DXF 70

physical_intensity

BD, DXF 40

illuminance_dist

BD, DXF 41

lamp_color_type

BS, DXF 71

lamp_color_temp

BD, DXF 42

lamp_color_preset

BS, DXF 72

lamp_color_rgb

BL

web_rotation

3BD 1, DXF 43

extlight_shape

BS, DXF 73

extlight_length

BD, DXF 46

extlight_width

BD, DXF 47

extlight_radius

BD, DXF 48

webfile_type

BS, DXF 74

web_symetry

BS, DXF 75

has_target_grip

BS, DXF 76

web_flux BD, DXF 49
web_angle1

BD, DXF 50

web_angle2

BD, DXF 51

web_angle3

BD, DXF 52

web_angle4

BD, DXF 53

web_angle5

BD, DXF 54

glyph_display_type

BS, DXF 77

LINE parent struct dwg object entity*

z_is_zero

RC

start 3BD, DXF 10 end 3BD, DXF 11

thickness

BT, DXF 39

extrusion

BE, DXF 210

unknown_r11

2RD

LINEARGRIPENTITY

parent struct dwg object entity* LINEARPARAMETERENTITY

parent struct dwg object entity* LOAD

parent struct dwg object entity*

file_name

TV, DXF 1

LOFTEDSURFACE parent struct dwg object entity*

acis_empty

B, DXF 290

unknown B

version BS, DXF 70

num_blocks

BL

block_size

BL*

encr_sat_data

char **, DXF 1

sab_size BL
acis_data

RC*

wireframe_data_present

B

point_present

B

point  3BD
isolines BL
isoline_present

B

num_wires

BL

wires Dwg 3DSOLID wire*

num_silhouettes

BL

silhouettes

Dwg 3DSOLID silhouette*

_dxf_sab_converted

B

acis_empty2

B

extra_acis_data

struct dwg entity 3DSOLID*

num_materials

BL

materials

Dwg 3DSOLID material*

revision_guid

RC, DXF 2

revision_major

BL

revision_minor1

BS

revision_minor2

BS

revision_bytes

RC

end_marker

BL

history_id

H, DXF 350

has_revision_guid

B

acis_empty_bit

B

modeler_format_version

BS, DXF 70

u_isolines

BS, DXF 71

v_isolines

BS, DXF 72

loft_entity_transmatrix

BD*, DXF 40

plane_normal_lofting_type

BL, DXF 70

start_draft_angle

BD, DXF 41

end_draft_angle

BD, DXF 42

start_draft_magnitude

BD, DXF 43

end_draft_magnitude

BD, DXF 44

arc_length_parameterization

B, DXF 290

no_twist B, DXF 291
align_direction

B, DXF 292

simple_surfaces

B, DXF 293

closed_surfaces

B, DXF 294

solid B, DXF 295

ruled_surface

B, DXF 296

virtual_guide

B, DXF 297

num_cross_sections

BS

num_guide_curves

BS

cross_sections

H*, DXF 310

guide_curves

H*, DXF 310

path_curve

H

LWPOLYLINE parent struct dwg object entity* flag BS, DXF 70

const_width

BD, DXF 43

elevation

BD, DXF 38

thickness

BD, DXF 39

extrusion

BE, DXF 210

num_points

BL, DXF 90

points 2RD*, DXF 10

num_bulges

BL

bulges BD*, DXF 42

num_vertexids

BL

vertexids

BL*, DXF 91

num_widths

BL

widths Dwg LWPOLYLINE width* MESH

parent struct dwg object entity* dlevel BS, DXF 71

is_watertight

B, DXF 72

num_subdiv_vertex

BL, DXF 91

subdiv_vertex

3DPOINT*, DXF 10

num_vertex

BL, DXF 92

vertex 3DPOINT*, DXF 10

num_faces

BL, DXF 93

faces BL*, DXF 90

num_edges

BL, DXF 94

edges Dwg MESH edge*

num_crease

BL, DXF 95

crease BD*, DXF 140

unknown_b1

B

unknown_b2

B

MINSERT parent struct dwg object entity* ins_pt 3DPOINT, DXF 10

scale_flag

BB

scale 3BD 1, DXF 41

rotation BD, DXF 50
extrusion

BE, DXF 210

has_attribs

B, DXF 66

num_owned

BL

num_cols BS, DXF 70
num_rows BS, DXF 71
col_spacing

BD, DXF 44

row_spacing

BD, DXF 45

block_header

H, DXF 2

first_attrib

H

last_attrib

H

attribs H*
seqend  H

MLINE parent struct dwg object entity* scale BD, DXF 40

justification

RC, DXF 70

base_point

3BD, DXF 10

extrusion

BE, DXF 210

flags BS, DXF 71

num_lines

RC, DXF 73

num_verts

BS, DXF 72

verts Dwg MLINE vertex*

mlinestyle

H, DXF 340

MPOLYGON parent struct dwg object entity*

is_gradient_fill

BL

reserved BL
gradient_angle

BD

gradient_shift

BD

single_color_gradient

BL

gradient_tint

BD

num_colors

BL

colors Dwg HATCH Color*

gradient_name

T

elevation

BD, DXF 30

extrusion

BE, DXF 210

name T, DXF 2

is_solid_fill

B, DXF 70

is_associative

B, DXF 71

num_paths

BL, DXF 91

paths Dwg HATCH Path* style BS, DXF 75

pattern_type

BS, DXF 76

angle BD, DXF 52

scale_spacing

BD, DXF 41

double_flag

B, DXF 77

num_deflines

BS, DXF 78

deflines Dwg HATCH DefLine*

hatch_color

CMC, DXF 62

x_dir 2RD, DXF 11

num_boundary_handles

BL, DXF 99

MTEXT parent struct dwg object entity* ins_pt 3BD, DXF 10

extrusion

BE, DXF 210

x_axis_dir

3BD, DXF 11

rect_height

BD, DXF 41

rect_width

BD, DXF 40

text_height

BD, DXF 40

attachment

BS, DXF 71

flow_dir BS, DXF 72
extents_width

BD, DXF 42

extents_height

BD, DXF 43

text T, DXF 1 style H, DXF 7

linespace_style

BS, DXF 73

linespace_factor

BD, DXF 44

unknown_b0

B

bg_fill_flag

BL, DXF 90

bg_fill_scale

BL, DXF 45

bg_fill_color

CMC, DXF 63

bg_fill_trans

BL, DXF 441

is_not_annotative

B

class_version

BS

default_flag

B, DXF 70

appid  H
ignore_attachment

BL

column_type

BS, DXF 71

numfragments

BL, DXF 72

column_width

BD, DXF 44

gutter BD, DXF 45

auto_height

B, DXF 73

flow_reversed

B, DXF 74

num_column_heights

BL, DXF 72

column_heights

BD*, DXF 46

MULTILEADER parent struct dwg object entity*

class_version

BS, DXF 270

ctx Dwg MLEADER AnnotContext

mleaderstyle

H, DXF 340

flags BL, DXF 90 type BS, DXF 170

line_color

CMC, DXF 91

line_ltype

H, DXF 341

line_linewt

BLd, DXF 171

has_landing

B, DXF 290

has_dogleg

B, DXF 291

landing_dist

BD, DXF 41

arrow_handle

H, DXF 342

arrow_size

BD, DXF 42

style_content

BS, DXF 172

text_style

H, DXF 343

text_left

BS, DXF 173

text_right

BS, DXF 95

text_angletype

BS, DXF 174

text_alignment

BS, DXF 175

text_color

CMC, DXF 92

has_text_frame

B, DXF 292

block_style

H, DXF 344

block_color

CMC, DXF 93

block_scale

3BD, DXF 10

block_rotation

BD, DXF 43

style_attachment

BS, DXF 176

is_annotative

B, DXF 293

num_arrowheads

BL

arrowheads

Dwg LEADER ArrowHead*

num_blocklabels

BL

blocklabels

Dwg LEADER BlockLabel*

is_neg_textdir

B, DXF 294

ipe_alignment

BS, DXF 178

justification

BS, DXF 179

scale_factor

BD, DXF 45

attach_dir

BS, DXF 271

attach_top

BS, DXF 273

attach_bottom

BS, DXF 272

is_text_extended

B, DXF 295

NAVISWORKSMODEL parent struct dwg object entity* flags BS, DXF 70

definition

H, DXF 340

transmatrix

BD*, DXF 40

unitfactor

BD, DXF 40

NURBSURFACE parent struct dwg object entity*

acis_empty

B, DXF 290

unknown B

version BS, DXF 70

num_blocks

BL

block_size

BL*

encr_sat_data

char **, DXF 1

sab_size BL
acis_data

RC*

wireframe_data_present

B

point_present

B

point  3BD
isolines BL
isoline_present

B

num_wires

BL

wires Dwg 3DSOLID wire*

num_silhouettes

BL

silhouettes

Dwg 3DSOLID silhouette*

_dxf_sab_converted

B

acis_empty2

B

extra_acis_data

struct dwg entity 3DSOLID*

num_materials

BL

materials

Dwg 3DSOLID material*

revision_guid

RC, DXF 2

revision_major

BL

revision_minor1

BS

revision_minor2

BS

revision_bytes

RC

end_marker

BL

history_id

H, DXF 350

has_revision_guid

B

acis_empty_bit

B

u_isolines

BS, DXF 71

v_isolines

BS, DXF 72

short170 BS, DXF 170

cv_hull_display

B, DXF 290

uvec1 3BD, DXF 10 vvec1 3BD, DXF 11 uvec2 3BD, DXF 12 vvec2 3BD, DXF 13

OLE2FRAME parent struct dwg object entity* type BS, DXF 71 mode BS, DXF 72

lock_aspect

RC, DXF 73

data_size

BL, DXF 90

data TF, DXF 310

oleversion

BS, DXF 70

oleclient

TF, DXF 3

pt1 3BD, DXF 10 pt2 3BD, DXF 11

OLEFRAME parent struct dwg object entity* flag BS, DXF 70

mode  BS
data_size

BL, DXF 90

data TF, DXF 310 PDFUNDERLAY

See [UNDERLAY], page 32, PLANESURFACE

parent struct dwg object entity*

acis_empty

B, DXF 290

unknown B

version BS, DXF 70

num_blocks

BL

block_size

BL*

encr_sat_data

char **, DXF 1

sab_size BL

acis_data

RC*

wireframe_data_present

B

point_present

B

point  3BD
isolines BL
isoline_present

B

num_wires

BL

wires Dwg 3DSOLID wire*

num_silhouettes

BL

silhouettes

Dwg 3DSOLID silhouette*

_dxf_sab_converted

B

acis_empty2

B

extra_acis_data

struct dwg entity 3DSOLID*

num_materials

BL

materials

Dwg 3DSOLID material*

revision_guid

RC, DXF 2

revision_major

BL

revision_minor1

BS

revision_minor2

BS

revision_bytes

RC

end_marker

BL

history_id

H, DXF 350

has_revision_guid

B

acis_empty_bit

B

modeler_format_version

BS, DXF 70

u_isolines

BS, DXF 71

v_isolines

BS, DXF 72

class_version

BL

POINT parent struct dwg object entity* x BD, DXF 10 y BD, DXF 20 z BD, DXF 30

thickness

BT, DXF 39

extrusion

BE, DXF 210

x_ang BD, DXF 50 POINTCLOUD

parent struct dwg object entity*

class_version

BS, DXF 70

origin 3BD, DXF 10

saved_filename

T, DXF 1

num_source_files

BL, DXF 90

source_files

TV*, DXF 2

extents_min

3BD, DXF 11

extents_max

3BD, DXF 12

numpoints

RLL, DXF 92

ucs_name T, DXF 3
ucs_origin

3BD, DXF 13

ucs_x_dir

3BD, DXF 210

ucs_y_dir

3BD, DXF 211

ucs_z_dir

3BD, DXF 212

pointclouddef

H, DXF 330

reactor H, DXF 360

show_intensity

B

intensity_scheme

BS, DXF 71

intensity_style

Dwg POINTCLOUD IntensityStyle

show_clipping

B

num_clippings

BL

clippings

Dwg POINTCLOUD Clippings*

POINTCLOUDEX parent struct dwg object entity*

class_version

BS, DXF 70

extents_min

3BD, DXF 10

extents_max

3BD, DXF 11

ucs_origin

3BD, DXF 12

ucs_x_dir

3BD, DXF 210

ucs_y_dir

3BD, DXF 211

ucs_z_dir

3BD, DXF 212

is_locked

B, DXF 290

pointclouddefex

H, DXF 330

reactor H, DXF 360 name T, DXF 1

show_intensity

B, DXF 291

stylization_type

BS, DXF 71

intensity_colorscheme

T, DXF 1

cur_colorscheme

T, DXF 1

classification_colorscheme

T, DXF 1

elevation_min

BD, DXF 40

elevation_max

BD, DXF 41

intensity_min

BL, DXF 90

intensity_max

BL, DXF 91

intensity_out_of_range_behavior

BS, DXF 71

elevation_out_of_range_behavior

BS, DXF 72

elevation_apply_to_fixed_range

B, DXF 292

intensity_as_gradient

B, DXF 293

elevation_as_gradient

B, DXF 294

show_cropping

B, DXF 295

unknown_bl0

BL, DXF 93

unknown_bl1

BL, DXF 93

num_croppings

BL, DXF 92

croppings

Dwg POINTCLOUDEX Croppings*

POINTPARAMETERENTITY

parent struct dwg object entity* POLARGRIPENTITY

parent struct dwg object entity* POLYLINE 2D

parent struct dwg object entity*

has_vertex

B, DXF 66

num_owned

BL

first_vertex

H

last_vertex

H

vertex  H*
seqend  H

flag BS, DXF 70

curve_type

BS, DXF 75

start_width

BD, DXF 40

end_width

BD, DXF 41

thickness

BT, DXF 39

elevation

BD

extrusion

BE, DXF 210

extra_r11_size

BL

extra_r11_text

TV

num_m_verts

BS, DXF 71

num_n_verts

BS, DXF 72

POLYLINE 3D parent struct dwg object entity*

has_vertex

B, DXF 66

num_owned

BL

first_vertex

H

last_vertex

H

vertex  H*
seqend  H
curve_type

RC, DXF 75

start_width

BD, DXF 40

end_width

BD, DXF 41

flag RC, DXF 70

extrusion

BE, DXF 210

POLYLINE MESH parent struct dwg object entity*

has_vertex

B, DXF 66

num_owned

BL

first_vertex

H

last_vertex

H

vertex  H*
seqend  H

flag BS, DXF 70

curve_type

BS, DXF 75

num_m_verts

BS, DXF 71

num_n_verts

BS, DXF 72

m_density

BS, DXF 73

n_density

BS, DXF 74

POLYLINE PFACE parent struct dwg object entity*

has_vertex

B, DXF 66

num_owned

BL

first_vertex

H

last_vertex

H

vertex  H*
seqend  H

flag BS, DXF 70

numverts BS, DXF 71
numfaces BS, DXF 72

PROXY ENTITY parent struct dwg object entity*

class_id BL, DXF 90

version BL, DXF 95

maint_version

BL, DXF 97

from_dxf B, DXF 70
data_numbits

BL

data_size

BL, DXF 160

data TF, DXF 310

num_objids

BL

objids H*, DXF 340 RAY

parent struct dwg object entity* point 3BD, DXF 10 vector 3BD, DXF 11

REGION

See [3DSOLID], page 24, REPEAT

parent struct dwg object entity* REVOLVEDSURFACE

parent struct dwg object entity*

acis_empty

B, DXF 290

unknown B

version BS, DXF 70

num_blocks

BL

block_size

BL*

encr_sat_data

char **, DXF 1

sab_size BL
acis_data

RC*

wireframe_data_present

B

point_present

B

point  3BD
isolines BL
isoline_present

B

num_wires

BL

wires Dwg 3DSOLID wire*

num_silhouettes

BL

silhouettes

Dwg 3DSOLID silhouette*

_dxf_sab_converted

B

acis_empty2

B

extra_acis_data

struct dwg entity 3DSOLID*

num_materials

BL

materials

Dwg 3DSOLID material*

revision_guid

RC, DXF 2

revision_major

BL

revision_minor1

BS

revision_minor2

BS

revision_bytes

RC

end_marker

BL

history_id

H, DXF 350

has_revision_guid

B

acis_empty_bit

B

modeler_format_version

BS, DXF 70

u_isolines

BS, DXF 71

v_isolines

BS, DXF 72

class_version

BL, DXF 90

id BL, DXF 90

axis_point

3BD, DXF 10

axis_vector

3BD, DXF 11

revolve_angle

BD, DXF 40

start_angle

BD, DXF 41

revolved_entity_transmatrix

BD*, DXF 42

draft_angle

BD, DXF 43

draft_start_distance

BD, DXF 44

draft_end_distance

BD, DXF 45

twist_angle

BD, DXF 46

solid B, DXF 290

close_to_axis

B, DXF 291

ROTATIONGRIPENTITY

parent struct dwg object entity* ROTATIONPARAMETERENTITY

parent struct dwg object entity* RTEXT

parent struct dwg object entity* pt 3BD, DXF 10

extrusion

BE, DXF 210

rotation BD, DXF 50

height BD, DXF 50 flags BS, DXF 70

text_value

T, DXF 1

style H, DXF 7 SECTIONOBJECT

parent struct dwg object entity* state BL, DXF 90 flags BL, DXF 91 name T, DXF 1 vert_dir 3BD, DXF 10

top_height

BD, DXF 40

bottom_height

BD, DXF 41

indicator_alpha

BS, DXF 70

indicator_color

CMC, DXF 62

num_verts

BL, DXF 92

verts 3BD*, DXF 11

num_blverts

BL, DXF 93

blverts 3BD*, DXF 12

section_settings

H, DXF 360

SEQEND parent struct dwg object entity*

begin_addr_r11

RL

SHAPE parent struct dwg object entity* ins_pt 3BD, DXF 10 scale BD, DXF 40

rotation BD, DXF 50
width_factor

BD, DXF 41

oblique_angle

BD, DXF 51

thickness

BD, DXF 39

style_id BS, DXF 2
extrusion

BE, DXF 210

style H, DXF 7 SOLID

parent struct dwg object entity*

thickness

BT, DXF 39

elevation

BD, DXF 38

corner1 2RD, DXF 10 corner2 2RD, DXF 11 corner3 2RD, DXF 12 corner4 2RD, DXF 13

extrusion

BE, DXF 210

SPLINE parent struct dwg object entity*

flag  RS
scenario BS

degree BS, DXF 71

splineflags

BL

knotparam

BL

fit_tol BD, DXF 44

beg_tan_vec

3BD, DXF 12

end_tan_vec

3BD, DXF 13

closed_b B
periodic B
rational B
weighted B
knot_tol BD, DXF 42
ctrl_tol BD, DXF 43
num_fit_pts

BS, DXF 74

fit_pts 3DPOINT*, DXF 11

num_knots

BL, DXF 72

knots BD*, DXF 40

num_ctrl_pts

BL, DXF 73

ctrl_pts Dwg SPLINE control point* SWEPTSURFACE

parent struct dwg object entity*

acis_empty

B, DXF 290

unknown B

version BS, DXF 70

num_blocks

BL

block_size

BL*

encr_sat_data

char **, DXF 1

sab_size BL

acis_data

RC*

wireframe_data_present

B

point_present

B

point  3BD
isolines BL
isoline_present

B

num_wires

BL

wires Dwg 3DSOLID wire*

num_silhouettes

BL

silhouettes

Dwg 3DSOLID silhouette*

_dxf_sab_converted

B

acis_empty2

B

extra_acis_data

struct dwg entity 3DSOLID*

num_materials

BL

materials

Dwg 3DSOLID material*

revision_guid

RC, DXF 2

revision_major

BL

revision_minor1

BS

revision_minor2

BS

revision_bytes

RC

end_marker

BL

history_id

H, DXF 350

has_revision_guid

B

acis_empty_bit

B

modeler_format_version

BS, DXF 70

u_isolines

BS, DXF 71

v_isolines

BS, DXF 72

class_version

BL, DXF 90

sweep_entity_id

BL, DXF 90

sweepdata_size

BL, DXF 90

sweepdata

TF, DXF 310

path_entity_id

BL, DXF 90

pathdata_size

BL, DXF 90

pathdata TF, DXF 310

draft_angle

BD, DXF 42

draft_start_distance

BD, DXF 43

draft_end_distance

BD, DXF 44

twist_angle

BD, DXF 45

scale_factor

BD, DXF 48

align_angle

BD, DXF 49

sweep_entity_transmatrix

BD*, DXF 46

path_entity_transmatrix

BD*, DXF 47

is_solid B, DXF 290
sweep_alignment_flags

BS, DXF 70

path_flags

BS, DXF 71

align_start

B, DXF 292

bank B, DXF 293

base_point_set

B, DXF 294

sweep_entity_transform_computed

B, DXF 295

path_entity_transform_computed

B, DXF 296

reference_vector_for_controlling_twist

3BD, DXF 11

sweep_entity

H

path_entity

H

TABLE parent struct dwg object entity* ldata Dwg LinkedData tdata Dwg LinkedTableData fdata Dwg FormattedTableData

tablestyle

H, DXF 342

unknown_rc

RC

unknown_h

H

unknown_bl

BL

unknown_b

B

unknown_bl1

BL

ins_pt 3BD, DXF 10 scale 3BD 1, DXF 41

scale_flag

BB

rotation BD, DXF 50
extrusion

BE, DXF 210

has_attribs

B, DXF 66

num_owned

BL

flag_for_table_value

BS, DXF 90

horiz_direction

3BD, DXF 11

num_cols BL, DXF 92
num_rows BL, DXF 91
num_cells

unsigned long

col_widths

BD*, DXF 142

row_heights

BD*, DXF 141

cells Dwg TABLE Cell*

has_table_overrides

B

table_flag_override

BL, DXF 93

title_suppressed

B, DXF 280

header_suppressed

B, DXF 281

flow_direction

BS, DXF 70

horiz_cell_margin

BD, DXF 40

vert_cell_margin

BD, DXF 41

title_row_color

CMC, DXF 64

header_row_color

CMC, DXF 64

data_row_color

CMC, DXF 64

title_row_fill_none

B, DXF 283

header_row_fill_none

B, DXF 283

data_row_fill_none

B, DXF 283

title_row_fill_color

CMC, DXF 63

header_row_fill_color

CMC, DXF 63

data_row_fill_color

CMC, DXF 63

title_row_alignment

BS, DXF 170

header_row_alignment

BS, DXF 170

data_row_alignment

BS, DXF 170

title_text_style

H, DXF 7

header_text_style

H, DXF 7

data_text_style

H, DXF 7

title_row_height

BD, DXF 140

header_row_height

BD, DXF 140

data_row_height

BD, DXF 140

has_border_color_overrides

B

border_color_overrides_flag

BL, DXF 94

title_horiz_top_color

CMC, DXF 64

title_horiz_ins_color

CMC, DXF 65

title_horiz_bottom_color

CMC, DXF 66

title_vert_left_color

CMC, DXF 63

title_vert_ins_color

CMC, DXF 68

title_vert_right_color

CMC, DXF 69

header_horiz_top_color

CMC, DXF 64

header_horiz_ins_color

CMC, DXF 65

header_horiz_bottom_color

CMC, DXF 66

header_vert_left_color

CMC, DXF 63

header_vert_ins_color

CMC, DXF 68

header_vert_right_color

CMC, DXF 69

data_horiz_top_color

CMC, DXF 64

data_horiz_ins_color

CMC, DXF 65

data_horiz_bottom_color

CMC, DXF 66

data_vert_left_color

CMC, DXF 63

data_vert_ins_color

CMC, DXF 68

data_vert_right_color

CMC, DXF 69

has_border_lineweight_overrides

B

border_lineweight_overrides_flag

BL, DXF 95

title_horiz_top_linewt

BS

title_horiz_ins_linewt

BS

title_horiz_bottom_linewt

BS

title_vert_left_linewt

BS

title_vert_ins_linewt

BS

title_vert_right_linewt

BS

header_horiz_top_linewt

BS

header_horiz_ins_linewt

BS

header_horiz_bottom_linewt

BS

header_vert_left_linewt

BS

header_vert_ins_linewt

BS

header_vert_right_linewt

BS

data_horiz_top_linewt

BS

data_horiz_ins_linewt

BS

data_horiz_bottom_linewt

BS

data_vert_left_linewt

BS

data_vert_ins_linewt

BS

data_vert_right_linewt

BS

has_border_visibility_overrides

B

border_visibility_overrides_flag

BL, DXF 96

title_horiz_top_visibility

BS

title_horiz_ins_visibility

BS

title_horiz_bottom_visibility

BS

title_vert_left_visibility

BS

title_vert_ins_visibility

BS

title_vert_right_visibility

BS

header_horiz_top_visibility

BS

header_horiz_ins_visibility

BS

header_horiz_bottom_visibility

BS

header_vert_left_visibility

BS

header_vert_ins_visibility

BS

header_vert_right_visibility

BS

data_horiz_top_visibility

BS

data_horiz_ins_visibility

BS

data_horiz_bottom_visibility

BS

data_vert_left_visibility

BS

data_vert_ins_visibility

BS

data_vert_right_visibility

BS

block_header

H, DXF 2

first_attrib

H

last_attrib

H

attribs H*
seqend  H
title_row_style_override

H, DXF 7

header_row_style_override

H

data_row_style_override

H

unknown_bs

BS

hor_dir 3BD, DXF 11

has_break_data

BL

break_flag

BL

break_flow_direction

BL

break_spacing

BD

break_unknown1

BL

break_unknown2

BL

num_break_heights

BL

break_heights

Dwg TABLE BreakHeight*

num_break_rows

BL

break_rows

Dwg TABLE BreakRow*

TEXT parent struct dwg object entity*

dataflags

RC

elevation

RD, DXF 30

ins_pt 2DPOINT, DXF 10

alignment_pt

2DPOINT, DXF 11

extrusion

BE, DXF 210

thickness

RD, DXF 39

oblique_angle

RD, DXF 51

rotation RD, DXF 50

height RD, DXF 40

width_factor

RD, DXF 41

text_value

T, DXF 1

generation

BS, DXF 71

horiz_alignment

BS, DXF 72

vert_alignment

BS, DXF 73

style H, DXF 7 TOLERANCE

parent struct dwg object entity*

unknown_short

BS

height  BD

dimgap  BD

ins_pt 3BD, DXF 10

x_direction

3BD, DXF 11

extrusion

BE

text_value

T, DXF 1

dimstyle H, DXF 3

TRACE parent struct dwg object entity*

thickness

BT, DXF 39

elevation

BD, DXF 38

corner1 2RD, DXF 10 corner2 2RD, DXF 11 corner3 2RD, DXF 12 corner4 2RD, DXF 13

extrusion

BE, DXF 210

UNKNOWN ENT

parent struct dwg object entity* VERTEX 2D

parent struct dwg object entity* flag RC, DXF 70 point 3BD, DXF 10

start_width

BD, DXF 40

end_width

BD, DXF 41

id BL, DXF 91 bulge BD, DXF 42

tangent_dir

BD, DXF 50 VERTEX 3D

parent struct dwg object entity* flag RC, DXF 70 point 3BD, DXF 10

VERTEX MESH

See [VERTEX 3D], page 93, VERTEX PFACE

See [VERTEX 3D], page 93, VERTEX PFACE FACE

parent struct dwg object entity* flag RC, DXF 70 vertind BS, DXF 71

VIEWPORT parent struct dwg object entity* center 3BD, DXF 10 width BD, DXF 40 height BD, DXF 41 on_off RS, DXF 68 id RS, DXF 69

view_target

3BD, DXF 17

VIEWDIR 3BD, DXF 16

twist_angle

BD, DXF 51

VIEWSIZE BD, DXF 45
lens_length

BD, DXF 42

front_clip_z

BD, DXF 43

back_clip_z

BD, DXF 44

SNAPANG BD, DXF 50 VIEWCTR 2RD, DXF 12 SNAPBASE 2RD, DXF 13 SNAPUNIT 2RD, DXF 14 GRIDUNIT 2RD, DXF 15

circle_zoom

BS, DXF 72

grid_major

BS, DXF 61

num_frozen_layers

BL

status_flag

BL, DXF 90

style_sheet

T, DXF 1

render_mode

RC, DXF 281

ucs_at_origin

B, DXF 74

UCSVP B, DXF 71 ucsorg 3BD, DXF 110 ucsxdir 3BD, DXF 111 ucsydir 3BD, DXF 112

ucs_elevation

BD, DXF 146

UCSORTHOVIEW

BS, DXF 79

shadeplot_mode

BS, DXF 170

use_default_lights

B, DXF 292

default_lighting_type

RC, DXF 282

brightness

BD, DXF 141

contrast BD, DXF 142

ambient_color

CMC, DXF 63

vport_entity_header

H

frozen_layers

H*, DXF 341

clip_boundary

H, DXF 340

named_ucs

H, DXF 345

base_ucs H, DXF 346
background

H, DXF 332

visualstyle

H, DXF 348

shadeplot

H, DXF 333

sun H, DXF 361 VISIBILITYGRIPENTITY

parent struct dwg object entity* VISIBILITYPARAMETERENTITY

parent struct dwg object entity* WIPEOUT

parent struct dwg object entity*

class_version

BL, DXF 90

pt0 3BD, DXF 10 uvec 3BD, DXF 11 vvec 3BD, DXF 12 size 2RD, DXF 13

display_props

BS, DXF 70

clipping B, DXF 280
brightness

RC, DXF 281

contrast RC, DXF 282 fade RC, DXF 283

clip_mode

B, DXF 290

clip_boundary_type

BS, DXF 71

num_clip_verts

BL, DXF 91

clip_verts

2RD*, DXF 14

imagedef H, DXF 340
imagedefreactor

H, DXF 360

XLINE

See [RAY], page 76, XYGRIPENTITY

parent struct dwg object entity* XYPARAMETERENTITY

parent struct dwg object entity*

4.3 OBJECTS

All non-graphical objects with its fields. See [Common Object fields], page 262,

ACMECOMMANDHISTORY parent struct dwg object object*

class_version

BS

ACMESCOPE parent struct dwg object object*

class_version

BS

ACMESTATEMGR parent struct dwg object object*

class_version

BS

ACSH BOOLEAN CLASS parent struct dwg object object* evalexpr Dwg EvalExpr

history_node

Dwg ACSH HistoryNode

major BL, DXF 90 minor BL, DXF 91

operation

RCd, DXF 280

operand1 BL, DXF 92
operand2 BL, DXF 93

ACSH BOX CLASS parent struct dwg object object* evalexpr Dwg EvalExpr

history_node

Dwg ACSH HistoryNode

major BL, DXF 90 minor BL, DXF 91 length BD, DXF 40 width BD, DXF 41 height BD, DXF 42

ACSH BREP CLASS parent struct dwg object object*

acis_empty

B, DXF 290

unknown B

version BS, DXF 70

num_blocks

BL

block_size

BL*

encr_sat_data

char **, DXF 1

sab_size BL
acis_data

RC*

wireframe_data_present

B

point_present

B

point  3BD
isolines BL

isoline_present

B

num_wires

BL

wires Dwg 3DSOLID wire*

num_silhouettes

BL

silhouettes

Dwg 3DSOLID silhouette*

_dxf_sab_converted

B

acis_empty2

B

extra_acis_data

struct dwg entity 3DSOLID*

num_materials

BL

materials

Dwg 3DSOLID material*

revision_guid

RC, DXF 2

revision_major

BL

revision_minor1

BS

revision_minor2

BS

revision_bytes

RC

end_marker

BL

history_id

H, DXF 350

has_revision_guid

B

acis_empty_bit

B

evalexpr Dwg EvalExpr

history_node

Dwg ACSH HistoryNode

major BL, DXF 90 minor BL, DXF 91

ACSH CHAMFER CLASS parent struct dwg object object* evalexpr Dwg EvalExpr

history_node

Dwg ACSH HistoryNode

major BL, DXF 90 minor BL, DXF 91 bl92 BL, DXF 92

base_dist

BD, DXF 41

other_dist

BD, DXF 42

num_edges

BL, DXF 93

edges BL*, DXF 94 bl95 BL, DXF 95

ACSH CONE CLASS parent struct dwg object object* evalexpr Dwg EvalExpr

history_node

Dwg ACSH HistoryNode

major BL, DXF 90 minor BL, DXF 91 height BD, DXF 40

major_radius

BD, DXF 41

minor_radius

BD, DXF 42

x_radius BD, DXF 43

ACSH CYLINDER CLASS parent struct dwg object object* evalexpr Dwg EvalExpr

history_node

Dwg ACSH HistoryNode

major BL, DXF 90 minor BL, DXF 91 height BD, DXF 40

major_radius

BD, DXF 41

minor_radius

BD, DXF 42

x_radius BD, DXF 43

ACSH EXTRUSION CLASS parent struct dwg object object* evalexpr Dwg EvalExpr

history_node

Dwg ACSH HistoryNode

major BL, DXF 90 minor BL, DXF 91

direction

3BD, DXF 10

bl92 BL, DXF 92

shsw_text_size

BL, DXF 90

shsw_text

TF, DXF 310

shsw_bl93

BL, DXF 93

shsw_text2_size

BL, DXF 90

shsw_text2

TF, DXF 310

draft_angle

BD, DXF 42

start_draft_dist

BD, DXF 43

end_draft_dist

BD, DXF 44

scale_factor

BD, DXF 45

twist_angle

BD, DXF 48

align_angle

BD, DXF 49

sweepentity_transform

BD*, DXF 46

pathentity_transform

BD*, DXF 47

align_option

RC, DXF 70

miter_option

RC, DXF 71

has_align_start

B, DXF 290

bank B, DXF 292

check_intersections

B, DXF 293

shsw_b294

B, DXF 294

shsw_b295

B, DXF 295

shsw_b296

B, DXF 296

pt2 3BD, DXF 11 ACSH FILLET CLASS

parent struct dwg object object* evalexpr Dwg EvalExpr

history_node

Dwg ACSH HistoryNode

major BL, DXF 90 minor BL, DXF 91 bl92 BL, DXF 92

num_edges

BL, DXF 93

edges BL*, DXF 94

num_radiuses

BL, DXF 95

num_startsetbacks

BL, DXF 96

num_endsetbacks

BL, DXF 97

radiuses BD*, DXF 41

startsetbacks

BD*, DXF 42

endsetbacks

BD*, DXF 43

ACSH HISTORY CLASS parent struct dwg object object* major BL, DXF 90 minor BL, DXF 91 owner H, DXF 360

h_nodeid BL, DXF 92
show_history

B, DXF 280

record_history

B, DXF 281

ACSH LOFT CLASS parent struct dwg object object* evalexpr Dwg EvalExpr

history_node

Dwg ACSH HistoryNode

major BL, DXF 90 minor BL, DXF 91

num_crosssects

BL, DXF 92

crosssects

H*

num_guides

BL, DXF 95

guides  H*

ACSH PYRAMID CLASS parent struct dwg object object* evalexpr Dwg EvalExpr

history_node

Dwg ACSH HistoryNode

major BL, DXF 90 minor BL, DXF 91 height BD, DXF 40 sides BL, DXF 92 radius BD, DXF 41

topradius

BD, DXF 42

ACSH REVOLVE CLASS parent struct dwg object object* evalexpr Dwg EvalExpr

history_node

Dwg ACSH HistoryNode

major BL, DXF 90 minor BL, DXF 91 axis_pt 3BD, DXF 10

direction

2RD, DXF 11

revolve_angle

BD, DXF 40

start_angle

BD, DXF 41

draft_angle

BD, DXF 43

bd44 BD, DXF 44 bd45 BD, DXF 45

twist_angle

BD, DXF 46

b290 B, DXF 290

is_close_to_axis

B, DXF 291

sweep_entity

H ACSH SPHERE CLASS

parent struct dwg object object* evalexpr Dwg EvalExpr

history_node

Dwg ACSH HistoryNode

major BL, DXF 90 minor BL, DXF 91 radius BD, DXF 40

ACSH SWEEP CLASS parent struct dwg object object* evalexpr Dwg EvalExpr

history_node

Dwg ACSH HistoryNode

major BL, DXF 90 minor BL, DXF 91

direction

3BD, DXF 10

bl92 BL, DXF 92

shsw_text_size

BL, DXF 90

shsw_text

TF, DXF 310

shsw_bl93

BL, DXF 93

shsw_text2_size

BL, DXF 90

shsw_text2

TF, DXF 310

draft_angle

BD, DXF 42

start_draft_dist

BD, DXF 43

end_draft_dist

BD, DXF 44

scale_factor

BD, DXF 45

twist_angle

BD, DXF 48

align_angle

BD, DXF 49

sweepentity_transform

BD*, DXF 46

pathentity_transform

BD*, DXF 47

align_option

RC, DXF 70

miter_option

RC, DXF 71

has_align_start

B, DXF 290

bank B, DXF 292

check_intersections

B, DXF 293

shsw_b294

B, DXF 294

shsw_b295

B, DXF 295

shsw_b296

B, DXF 296

pt2 3BD, DXF 11 ACSH TORUS CLASS

parent struct dwg object object* evalexpr Dwg EvalExpr

history_node

Dwg ACSH HistoryNode

major BL, DXF 90 minor BL, DXF 91

major_radius

BD, DXF 40

minor_radius

BD, DXF 41

ACSH WEDGE CLASS parent struct dwg object object* evalexpr Dwg EvalExpr

history_node

Dwg ACSH HistoryNode

major BL, DXF 90 minor BL, DXF 91 length BD, DXF 40 width BD, DXF 41 height BD, DXF 42

ALDIMOBJECTCONTEXTDATA parent struct dwg object object*

class_version

BS, DXF 70

is_default

B, DXF 290

scale H, DXF 340

dimension

Dwg OCD Dimension

dimline_pt

3BD, DXF 11

ANGDIMOBJECTCONTEXTDATA parent struct dwg object object*

class_version

BS, DXF 70

is_default

B, DXF 290

scale H, DXF 340

dimension

Dwg OCD Dimension

arc_pt 3BD, DXF 11 ANNOTSCALEOBJECTCONTEXTDATA

parent struct dwg object object*

class_version

BS, DXF 70

is_default

B, DXF 290

scale H, DXF 340 APPID APPID is a table object.

parent struct dwg object object*

flag  RC
name  TV
used  RSd
is_xref_ref

B

is_xref_resolved

BS

is_xref_dep

B

xref  H

unknown RC, DXF 71 APPID CONTROL APPID CONTROL is a table control object.

parent struct dwg object object*

num_entries

BS, DXF 70

entries H*
flags_r11

RS

ASSOC2DCONSTRAINTGROUP parent struct dwg object object*

class_version

BS, DXF 90

geometry_status

BL, DXF 90

owningnetwork

H, DXF 330

actionbody

H, DXF 360

action_index

BL, DXF 90

max_assoc_dep_index

BL, DXF 90

num_deps BL, DXF 90

deps Dwg ASSOCACTION Deps*

num_owned_params

BL

owned_params

H*

num_values

BL

values Dwg VALUEPARAM* version BL, DXF 90 b1 B, DXF 70

workplane

3BD

h1 H, DXF 360

num_actions

BL, DXF 90

actions H*, DXF 360

num_nodes

BL, DXF 90

nodes Dwg CONSTRAINTGROUPNODE* ASSOC3POINTANGULARDIMACTIONBODY

parent struct dwg object object*

aaab_version

BS, DXF 90

assoc_dep

H, DXF 330

aab_version

BS, DXF 90

actionbody

H, DXF 360

pab Dwg ASSOCPARAMBASEDACTIONBODY

class_version

BS, DXF 90

r_node H, DXF 330 d_node H, DXF 330

assocdep H, DXF 330

ASSOCACTION parent struct dwg object object*

class_version

BS, DXF 90

geometry_status

BL, DXF 90

owningnetwork

H, DXF 330

actionbody

H, DXF 360

action_index

BL, DXF 90

max_assoc_dep_index

BL, DXF 90

num_deps BL, DXF 90

deps Dwg ASSOCACTION Deps*

num_owned_params

BL

owned_params

H*

num_values

BL

values Dwg VALUEPARAM* ASSOCACTIONPARAM

parent struct dwg object object*

is_r2013 BS, DXF 90
aap_version

BL, DXF 90

name T, DXF 1 ASSOCALIGNEDDIMACTIONBODY

parent struct dwg object object*

aaab_version

BS, DXF 90

assoc_dep

H, DXF 330

aab_version

BS, DXF 90

actionbody

H, DXF 360

pab Dwg ASSOCPARAMBASEDACTIONBODY

class_version

BL, DXF 90

r_node H, DXF 330 d_node H, DXF 330

ASSOCARRAYACTIONBODY parent struct dwg object object*

aab_version

BL, DXF 90

pab Dwg ASSOCPARAMBASEDACTIONBODY

aaab_version

BL, DXF 90

paramblock

T, DXF 1

transmatrix

BD*, DXF 40

ASSOCARRAYMODIFYACTIONBODY parent struct dwg object object*

aab_version

BL, DXF 90

pab Dwg ASSOCPARAMBASEDACTIONBODY

aaab_version

BL, DXF 90

paramblock

T, DXF 1

transmatrix

BD*, DXF 40

status BS, DXF 70

num_items

BL, DXF 90

items Dwg ARRAYITEMLOCATOR* ASSOCASMBODYACTIONPARAM

parent struct dwg object object*

is_r2013 BS, DXF 90

aap_version

BL, DXF 90

name T, DXF 1

asdap_class_version

BL, DXF 90

dep H, DXF 330

class_version

BL, DXF 90

acis_empty

B

unknown B
version BS
num_blocks

BL

block_size

BL*

encr_sat_data

char **

sab_size BL
acis_data

RC*

wireframe_data_present

B

point_present

B

point  3BD
isolines BL
isoline_present

B

num_wires

BL

wires Dwg 3DSOLID wire*

num_silhouettes

BL

silhouettes

Dwg 3DSOLID silhouette*

_dxf_sab_converted

B

acis_empty2

B

extra_acis_data

struct dwg entity 3DSOLID*

num_materials

BL

materials

Dwg 3DSOLID material*

revision_guid

RC

revision_major

BL

revision_minor1

BS

revision_minor2

BS

revision_bytes

RC

end_marker

BL

history_id

H

has_revision_guid

B

acis_empty_bit

B

ASSOCBLENDSURFACEACTIONBODY parent struct dwg object object*

aab_version

BL, DXF 90

pab Dwg ASSOCPARAMBASEDACTIONBODY sab Dwg ASSOCSURFACEACTIONBODY

pbsab_status

BL, DXF 90

class_version

BL, DXF 90

b1 B, DXF 290 b2 B, DXF 291 b3 B, DXF 292 b4 B, DXF 293 b5 B, DXF 294

blend_options

BS, DXF 72

bs2 BS, DXF 73 ASSOCCOMPOUNDACTIONPARAM

parent struct dwg object object*

is_r2013 BS, DXF 90
aap_version

BL, DXF 90

name T, DXF 1

class_version

BS, DXF 90

bs1 BS, DXF 90

num_params

BL, DXF 90

params H*, DXF 360

has_child_param

B

child_status

BS, DXF 90

child_id BL, DXF 90
child_param

H, DXF 330

h330_2 H, DXF 330 bl2 BL, DXF 90 h330_3 H, DXF 330

ASSOCDEPENDENCY parent struct dwg object object*

class_version

BS, DXF 90

status BL, DXF 90

is_read_dep

B, DXF 290

is_write_dep

B, DXF 290

is_attached_to_object

B, DXF 290

is_delegating_to_owning_action

B, DXF 290

order BLd, DXF 90 dep_on H, DXF 330

has_name B, DXF 290

name T, DXF 1

depbodyid

BLd, DXF 90

readdep H, DXF 330

dep_body H, DXF 360

node H, DXF 330 ASSOCDIMDEPENDENCYBODY

parent struct dwg object object*

adb_version

BS, DXF 90

dimbase_version

BS, DXF 90

name T, DXF 1

class_version

BS, DXF 90

ASSOCEDGEACTIONPARAM parent struct dwg object object*

is_r2013 BS, DXF 90
aap_version

BL, DXF 90

name T, DXF 1

asdap_class_version

BL, DXF 90

dep H, DXF 330

class_version

BL, DXF 90

param H, DXF 330

has_action

B, DXF 290

action_type

BL, DXF 90

subent  H

ASSOCEDGECHAMFERACTIONBODY parent struct dwg object object*

aab_version

BL, DXF 90

pab Dwg ASSOCPARAMBASEDACTIONBODY sab Dwg ASSOCSURFACEACTIONBODY

pbsab_status

BL, DXF 90

ASSOCEDGEFILLETACTIONBODY parent struct dwg object object*

aab_version

BL, DXF 90

pab Dwg ASSOCPARAMBASEDACTIONBODY sab Dwg ASSOCSURFACEACTIONBODY

pbsab_status

BL, DXF 90

ASSOCEXTENDSURFACEACTIONBODY parent struct dwg object object*

aab_version

BL, DXF 90

pab Dwg ASSOCPARAMBASEDACTIONBODY sab Dwg ASSOCSURFACEACTIONBODY

pbsab_status

BL, DXF 90

class_version

BL, DXF 90

option RC, DXF 280 ASSOCEXTRUDEDSURFACEACTIONBODY

parent struct dwg object object*

aab_version

BL, DXF 90 pab Dwg ASSOCPARAMBASEDACTIONBODY sab Dwg ASSOCSURFACEACTIONBODY

pbsab_status

BL, DXF 90

class_version

BL, DXF 90

ASSOCFACEACTIONPARAM parent struct dwg object object*

is_r2013 BS, DXF 90
aap_version

BL, DXF 90

name T, DXF 1

asdap_class_version

BL, DXF 90

dep H, DXF 330

class_version

BL, DXF 90

index BL, DXF 90 ASSOCFILLETSURFACEACTIONBODY

parent struct dwg object object*

aab_version

BL, DXF 90

pab Dwg ASSOCPARAMBASEDACTIONBODY sab Dwg ASSOCSURFACEACTIONBODY

pbsab_status

BL, DXF 90

class_version

BL, DXF 90

status BS, DXF 70 pt1 2RD, DXF 10 pt2 2RD, DXF 10

ASSOCGEOMDEPENDENCY parent struct dwg object object* assocdep Dwg Object ASSOCDEPENDENCY

class_version

BS, DXF 90 enabled B, DXF 290

classname

T, DXF 1

dependent_on_compound_object

B, DXF 290

ASSOCLOFTEDSURFACEACTIONBODY parent struct dwg object object*

aab_version

BL, DXF 90

pab Dwg ASSOCPARAMBASEDACTIONBODY sab Dwg ASSOCSURFACEACTIONBODY

pbsab_status

BL, DXF 90

class_version

BL, DXF 90

ASSOCMLEADERACTIONBODY parent struct dwg object object*

aaab_version

BS, DXF 90

assoc_dep

H, DXF 330

aab_version

BS, DXF 90

actionbody

H, DXF 360

pab Dwg ASSOCPARAMBASEDACTIONBODY

class_version

BL, DXF 90

num_actions

BL, DXF 90

actions Dwg ASSOCACTIONBODY action* ASSOCNETWORK

parent struct dwg object object*

class_version

BS, DXF 90

geometry_status

BL, DXF 90

owningnetwork

H, DXF 330

actionbody

H, DXF 360

action_index

BL, DXF 90

max_assoc_dep_index

BL, DXF 90

num_deps BL, DXF 90

deps Dwg ASSOCACTION Deps*

num_owned_params

BL

owned_params

H*

num_values

BL

values Dwg VALUEPARAM*

network_version

BS, DXF 90

network_action_index

BL, DXF 90

num_actions

BL, DXF 90

actions Dwg ASSOCACTION Deps*

num_owned_actions

BL, DXF 90

owned_actions

H*, DXF 330

ASSOCNETWORKSURFACEACTIONBODY parent struct dwg object object*

aab_version

BL, DXF 90

pab Dwg ASSOCPARAMBASEDACTIONBODY sab Dwg ASSOCSURFACEACTIONBODY

pbsab_status

BL, DXF 90

class_version

BL, DXF 90 ASSOCOBJECTACTIONPARAM

parent struct dwg object object*

is_r2013 BS, DXF 90
aap_version

BL, DXF 90

name T, DXF 1

asdap_class_version

BL, DXF 90

dep H, DXF 330

class_version

BS, DXF 90

ASSOCOFFSETSURFACEACTIONBODY parent struct dwg object object*

aab_version

BL, DXF 90

pab Dwg ASSOCPARAMBASEDACTIONBODY sab Dwg ASSOCSURFACEACTIONBODY

pbsab_status

BL, DXF 90

class_version

BL, DXF 90

b1 B, DXF 290 ASSOCORDINATEDIMACTIONBODY

parent struct dwg object object*

aaab_version

BS, DXF 90

assoc_dep

H, DXF 330

aab_version

BS, DXF 90

actionbody

H, DXF 360

pab Dwg ASSOCPARAMBASEDACTIONBODY

class_version

BL, DXF 90

r_node H, DXF 330

d_node H, DXF 330 ASSOCOSNAPPOINTREFACTIONPARAM

parent struct dwg object object*

is_r2013 BS, DXF 90
aap_version

BL, DXF 90

name T, DXF 1

class_version

BS, DXF 90

bs1 BS, DXF 90

num_params

BL, DXF 90

params H*, DXF 360

has_child_param

B

child_status

BS, DXF 90

child_id BL, DXF 90
child_param

H, DXF 330

h330_2 H, DXF 330 bl2 BL, DXF 90 h330_3 H, DXF 330 status BS, DXF 90

osnap_mode

RC, DXF 90

param BD, DXF 40 ASSOCPATCHSURFACEACTIONBODY

parent struct dwg object object*

aab_version

BL, DXF 90

pab Dwg ASSOCPARAMBASEDACTIONBODY sab Dwg ASSOCSURFACEACTIONBODY

pbsab_status

BL, DXF 90

class_version

BL, DXF 90

ASSOCPATHACTIONPARAM parent struct dwg object object*

is_r2013 BS, DXF 90
aap_version

BL, DXF 90

name T, DXF 1

class_version

BS, DXF 90

bs1 BS, DXF 90

num_params

BL, DXF 90

params H*, DXF 360

has_child_param

B

child_status

BS, DXF 90

child_id BL, DXF 90
child_param

H, DXF 330

h330_2 H, DXF 330 bl2 BL, DXF 90 h330_3 H, DXF 330 version BL, DXF 90

ASSOCPERSSUBENTMANAGER parent struct dwg object object*

class_version

BL, DXF 90

unknown_3

BL, DXF 90

unknown_0

BL, DXF 90

unknown_2

BL, DXF 90

num_steps

BL, DXF 90

num_subents

BL, DXF 90

steps BL*, DXF 90

subents BL*
unknown_bl6

BL, DXF 90

unknown_bl6a

BL, DXF 90

unknown_bl7a

BL, DXF 90

unknown_bl7

BL, DXF 90

unknown_bl8

BL, DXF 90

unknown_bl9

BL, DXF 90

unknown_bl10

BL, DXF 90

unknown_bl11

BL, DXF 90

unknown_bl12

BL, DXF 90

unknown_bl13

BL, DXF 90

unknown_bl14

BL, DXF 90

unknown_bl15

BL, DXF 90

unknown_bl16

BL, DXF 90

unknown_bl17

BL, DXF 90

unknown_bl18

BL, DXF 90

unknown_bl19

BL, DXF 90

unknown_bl20

BL, DXF 90

unknown_bl21

BL, DXF 90

unknown_bl22

BL, DXF 90

unknown_bl23

BL, DXF 90

unknown_bl24

BL, DXF 90

unknown_bl25

BL, DXF 90

unknown_bl26

BL, DXF 90

unknown_bl27

BL, DXF 90

unknown_bl28

BL, DXF 90

unknown_bl29

BL, DXF 90

unknown_bl30

BL, DXF 90

unknown_bl31

BL, DXF 90

unknown_bl32

BL, DXF 90

unknown_bl33

BL, DXF 90

unknown_bl34

BL, DXF 90

unknown_bl35

BL, DXF 90

unknown_bl36

BL, DXF 90

unknown_b37

B, DXF 290

ASSOCPLANESURFACEACTIONBODY parent struct dwg object object*

aab_version

BL, DXF 90 pab Dwg ASSOCPARAMBASEDACTIONBODY sab Dwg ASSOCSURFACEACTIONBODY

pbsab_status

BL, DXF 90

class_version

BL, DXF 90

ASSOCPOINTREFACTIONPARAM parent struct dwg object object*

is_r2013 BS, DXF 90
aap_version

BL, DXF 90

name T, DXF 1

class_version

BS, DXF 90

bs1 BS, DXF 90

num_params

BL, DXF 90

params H*, DXF 360

has_child_param

B

child_status

BS, DXF 90

child_id BL, DXF 90
child_param

H, DXF 330

h330_2 H, DXF 330 bl2 BL, DXF 90 h330_3 H, DXF 330

ASSOCRESTOREENTITYSTATEACTIONBODY parent struct dwg object object*

aab_version

BL, DXF 90

class_version

BL, DXF 90

entity H, DXF 330 ASSOCREVOLVEDSURFACEACTIONBODY

parent struct dwg object object*

aab_version

BL, DXF 90

pab Dwg ASSOCPARAMBASEDACTIONBODY sab Dwg ASSOCSURFACEACTIONBODY

pbsab_status

BL, DXF 90

class_version

BL, DXF 90

ASSOCROTATEDDIMACTIONBODY parent struct dwg object object*

aaab_version

BS, DXF 90

assoc_dep

H, DXF 330

aab_version

BS, DXF 90

actionbody

H, DXF 360

pab Dwg ASSOCPARAMBASEDACTIONBODY

class_version

BS, DXF 90

r_node H, DXF 330 d_node H, DXF 330

ASSOCSWEPTSURFACEACTIONBODY parent struct dwg object object*

aab_version

BL, DXF 90

pab Dwg ASSOCPARAMBASEDACTIONBODY sab Dwg ASSOCSURFACEACTIONBODY

pbsab_status

BL, DXF 90

class_version

BL, DXF 90

ASSOCTRIMSURFACEACTIONBODY parent struct dwg object object*

aab_version

BL, DXF 90

pab Dwg ASSOCPARAMBASEDACTIONBODY sab Dwg ASSOCSURFACEACTIONBODY

pbsab_status

BL, DXF 90

class_version

BL, DXF 90

b1 B, DXF 290 b2 B, DXF 290

distance BD, DXF 40

ASSOCVALUEDEPENDENCY parent struct dwg object object* assocdep Dwg Object ASSOCDEPENDENCY

ASSOCVARIABLE parent struct dwg object object*

av_class_version

BS, DXF 90

class_version

BS, DXF 90

geometry_status

BL, DXF 90

owningnetwork

H, DXF 330

actionbody

H, DXF 360

action_index

BL, DXF 90

max_assoc_dep_index

BL, DXF 90

num_deps BL, DXF 90

deps Dwg ASSOCACTION Deps*

num_owned_params

BL

owned_params

H*

num_values

BL

values Dwg VALUEPARAM* name T, DXF 1 t58 T, DXF 1

evaluator

T, DXF 1

desc T, DXF 1 value Dwg EvalVariant has_t78 B, DXF 290 t78 T, DXF 1 b290 B, DXF 290

ASSOCVERTEXACTIONPARAM parent struct dwg object object*

is_r2013 BS, DXF 90
aap_version

BL, DXF 90

name T, DXF 1

asdap_class_version

BL, DXF 90

dep H, DXF 330

class_version

BL, DXF 90

pt 3BD, DXF 10 BLKREFOBJECTCONTEXTDATA

parent struct dwg object object*

class_version

BS, DXF 70

is_default

B, DXF 290

scale H, DXF 340

rotation BD, DXF 50

ins_pt 3BD, DXF 10

scale_factor

3BD 1, DXF 42 BLOCKALIGNEDCONSTRAINTPARAMETER

parent struct dwg object object* evalexpr Dwg EvalExpr name T, DXF 300

be_major BL, DXF 98
be_minor BL, DXF 99

eed1071 BL, DXF 1071

show_properties

B, DXF 280

chain_actions

B, DXF 281

def_basept

3BD, DXF 1010

def_endpt

3BD, DXF 1011

prop1 Dwg BLOCKPARAMETER PropInfo prop2 Dwg BLOCKPARAMETER PropInfo prop3 Dwg BLOCKPARAMETER PropInfo prop4 Dwg BLOCKPARAMETER PropInfo

prop_states

BL*, DXF 91

parameter_base_location

BS, DXF 177

upd_basept

3BD

basept  3BD
upd_endpt

3BD

endpt  3BD
dependency

H, DXF 330

expr_name

T, DXF 305

expr_description

T, DXF 306

value BD, DXF 140

value_set

Dwg BLOCKPARAMVALUESET

BLOCKALIGNMENTGRIP parent struct dwg object object* evalexpr Dwg EvalExpr name T, DXF 300

be_major BL, DXF 98
be_minor BL, DXF 99

eed1071 BL, DXF 1071 bg_bl91 BL, DXF 91 bg_bl92 BL, DXF 92

bg_location

3BD, DXF 1010

bg_insert_cycling

B, DXF 280

bg_insert_cycling_weight

BLd, DXF 93

orientation

3BD 1, DXF 140

BLOCKALIGNMENTPARAMETER parent struct dwg object object* evalexpr Dwg EvalExpr name T, DXF 300

be_major BL, DXF 98
be_minor BL, DXF 99

eed1071 BL, DXF 1071

show_properties

B, DXF 280

chain_actions

B, DXF 281

def_basept

3BD, DXF 1010

def_endpt

3BD, DXF 1011

prop1 Dwg BLOCKPARAMETER PropInfo prop2 Dwg BLOCKPARAMETER PropInfo prop3 Dwg BLOCKPARAMETER PropInfo prop4 Dwg BLOCKPARAMETER PropInfo

prop_states

BL*, DXF 91

parameter_base_location

BS, DXF 177

upd_basept

3BD

basept  3BD
upd_endpt

3BD

endpt  3BD
align_perpendicular

B, DXF 280

BLOCKANGULARCONSTRAINTPARAMETER parent struct dwg object object* evalexpr Dwg EvalExpr name T, DXF 300

be_major BL, DXF 98
be_minor BL, DXF 99

eed1071 BL, DXF 1071

show_properties

B, DXF 280

chain_actions

B, DXF 281

def_basept

3BD, DXF 1010

def_endpt

3BD, DXF 1011

prop1 Dwg BLOCKPARAMETER PropInfo prop2 Dwg BLOCKPARAMETER PropInfo prop3 Dwg BLOCKPARAMETER PropInfo prop4 Dwg BLOCKPARAMETER PropInfo

prop_states

BL*, DXF 91

parameter_base_location

BS, DXF 177

upd_basept

3BD

basept  3BD
upd_endpt

3BD

endpt  3BD
dependency

H, DXF 330

center_pt

3BD, DXF 1011

end_pt 3BD, DXF 1012

expr_name

T, DXF 305

expr_description

T, DXF 306

angle BD, DXF 140

orientation_on_both_grips

B, DXF 280

value_set

Dwg BLOCKPARAMVALUESET

BLOCKARRAYACTION parent struct dwg object object* evalexpr Dwg EvalExpr name T, DXF 300

be_major BL, DXF 98
be_minor BL, DXF 99

eed1071 BL, DXF 1071

display_location

3BD, DXF 1010

num_actions

BL, DXF 70

actions BL*, DXF 91

num_deps BL, DXF 71

deps H*, DXF 330 conn_pts Dwg BLOCKACTION connectionpts

column_offset

BD, DXF 140

row_offset

BD, DXF 141

BLOCKBASEPOINTPARAMETER parent struct dwg object object* evalexpr Dwg EvalExpr name T, DXF 300

be_major BL, DXF 98
be_minor BL, DXF 99

eed1071 BL, DXF 1071

show_properties

B, DXF 280

chain_actions

B, DXF 281

def_pt 3BD, DXF 1010

num_propinfos

BL, DXF 93

prop1 Dwg BLOCKPARAMETER PropInfo prop2 Dwg BLOCKPARAMETER PropInfo pt 3BD, DXF 1011 base_pt 3BD, DXF 1012

BLOCKDIAMETRICCONSTRAINTPARAMETER parent struct dwg object object* evalexpr Dwg EvalExpr name T, DXF 300

be_major BL, DXF 98
be_minor BL, DXF 99

eed1071 BL, DXF 1071

show_properties

B, DXF 280

chain_actions

B, DXF 281

def_basept

3BD, DXF 1010

def_endpt

3BD, DXF 1011

prop1 Dwg BLOCKPARAMETER PropInfo prop2 Dwg BLOCKPARAMETER PropInfo prop3 Dwg BLOCKPARAMETER PropInfo prop4 Dwg BLOCKPARAMETER PropInfo

prop_states

BL*, DXF 91

parameter_base_location

BS, DXF 177

upd_basept

3BD

basept  3BD
upd_endpt

3BD

endpt  3BD
dependency

H, DXF 330

expr_name

T, DXF 305

expr_description

T, DXF 306

distance BD, DXF 140

orientation_on_both_grips

B

value_set

Dwg BLOCKPARAMVALUESET

BLOCKFLIPACTION parent struct dwg object object* evalexpr Dwg EvalExpr name T, DXF 300

be_major BL, DXF 98
be_minor BL, DXF 99

eed1071 BL, DXF 1071

display_location

3BD, DXF 1010

num_actions

BL, DXF 70

actions BL*, DXF 91

num_deps BL, DXF 71

deps H*, DXF 330 conn_pts Dwg BLOCKACTION connectionpts

action_offset_x

BD

action_offset_y

BD

angle_offset

BD

BLOCKFLIPGRIP parent struct dwg object object* evalexpr Dwg EvalExpr name T, DXF 300

be_major BL, DXF 98
be_minor BL, DXF 99

eed1071 BL, DXF 1071 bg_bl91 BL, DXF 91 bg_bl92 BL, DXF 92

bg_location

3BD, DXF 1010

bg_insert_cycling

B, DXF 280

bg_insert_cycling_weight

BLd, DXF 93

combined_state

BL, DXF 93

orientation

3BD 1, DXF 140

upd_state

BS

state  BS

BLOCKFLIPPARAMETER parent struct dwg object object* evalexpr Dwg EvalExpr name T, DXF 300

be_major BL, DXF 98
be_minor BL, DXF 99

eed1071 BL, DXF 1071

show_properties

B, DXF 280

chain_actions

B, DXF 281

def_basept

3BD, DXF 1010

def_endpt

3BD, DXF 1011

prop1 Dwg BLOCKPARAMETER PropInfo prop2 Dwg BLOCKPARAMETER PropInfo prop3 Dwg BLOCKPARAMETER PropInfo prop4 Dwg BLOCKPARAMETER PropInfo

prop_states

BL*, DXF 91

parameter_base_location

BS, DXF 177

upd_basept

3BD

basept  3BD
upd_endpt

3BD

endpt  3BD
flip_label

T, DXF 305

flip_label_desc

T, DXF 306

base_state_label

T, DXF 307

flipped_state_label

T, DXF 308

def_label_pt

3BD, DXF 1012

bl96 BL, DXF 96 tooltip T, DXF 309

BLOCKGRIPLOCATIONCOMPONENT parent struct dwg object object* evalexpr Dwg EvalExpr

grip_type

BL, DXF 91

grip_expr

T, DXF 300

BLOCKHORIZONTALCONSTRAINTPARAMETER parent struct dwg object object* evalexpr Dwg EvalExpr name T, DXF 300

be_major BL, DXF 98
be_minor BL, DXF 99

eed1071 BL, DXF 1071

show_properties

B, DXF 280

chain_actions

B, DXF 281

def_basept

3BD, DXF 1010

def_endpt

3BD, DXF 1011

prop1 Dwg BLOCKPARAMETER PropInfo prop2 Dwg BLOCKPARAMETER PropInfo prop3 Dwg BLOCKPARAMETER PropInfo prop4 Dwg BLOCKPARAMETER PropInfo

prop_states

BL*, DXF 91

parameter_base_location

BS, DXF 177

upd_basept

3BD

basept  3BD
upd_endpt

3BD

endpt  3BD
dependency

H, DXF 330

expr_name

T, DXF 305

expr_description

T, DXF 306

value BD, DXF 140

value_set

Dwg BLOCKPARAMVALUESET

BLOCKLINEARCONSTRAINTPARAMETER parent struct dwg object object* evalexpr Dwg EvalExpr name T, DXF 300

be_major BL, DXF 98
be_minor BL, DXF 99

eed1071 BL, DXF 1071

show_properties

B, DXF 280

chain_actions

B, DXF 281

def_basept

3BD, DXF 1010

def_endpt

3BD, DXF 1011

prop1 Dwg BLOCKPARAMETER PropInfo prop2 Dwg BLOCKPARAMETER PropInfo prop3 Dwg BLOCKPARAMETER PropInfo prop4 Dwg BLOCKPARAMETER PropInfo

prop_states

BL*, DXF 91

parameter_base_location

BS, DXF 177

upd_basept

3BD

basept  3BD
upd_endpt

3BD

endpt  3BD
dependency

H, DXF 330

expr_name

T, DXF 305

expr_description

T, DXF 306

value BD, DXF 140

value_set

Dwg BLOCKPARAMVALUESET

BLOCKLINEARGRIP parent struct dwg object object* evalexpr Dwg EvalExpr name T, DXF 300

be_major BL, DXF 98
be_minor BL, DXF 99

eed1071 BL, DXF 1071 bg_bl91 BL, DXF 91 bg_bl92 BL, DXF 92

bg_location

3BD, DXF 1010

bg_insert_cycling

B, DXF 280

bg_insert_cycling_weight

BLd, DXF 93

orientation

3BD 1, DXF 140

BLOCKLINEARPARAMETER parent struct dwg object object* evalexpr Dwg EvalExpr name T, DXF 300

be_major BL, DXF 98
be_minor BL, DXF 99

eed1071 BL, DXF 1071

show_properties

B, DXF 280

chain_actions

B, DXF 281

def_basept

3BD, DXF 1010

def_endpt

3BD, DXF 1011

prop1 Dwg BLOCKPARAMETER PropInfo prop2 Dwg BLOCKPARAMETER PropInfo prop3 Dwg BLOCKPARAMETER PropInfo prop4 Dwg BLOCKPARAMETER PropInfo

prop_states

BL*, DXF 91

parameter_base_location

BS, DXF 177

upd_basept

3BD

basept  3BD
upd_endpt

3BD

endpt  3BD
distance_name

T, DXF 305

distance_desc

T, DXF 306

distance BD, DXF 140

value_set

Dwg BLOCKPARAMVALUESET

BLOCKLOOKUPACTION parent struct dwg object object* evalexpr Dwg EvalExpr name T, DXF 300

be_major BL, DXF 98
be_minor BL, DXF 99

eed1071 BL, DXF 1071

display_location

3BD, DXF 1010

num_actions

BL, DXF 70

actions BL*, DXF 91

num_deps BL, DXF 71

deps H*, DXF 330

numelems BL

numrows BL, DXF 92 numcols BL, DXF 93 lut Dwg BLOCKLOOKUPACTION lut* exprs TV*, DXF 302 b280 B, DXF 280

BLOCKLOOKUPGRIP parent struct dwg object object* evalexpr Dwg EvalExpr name T, DXF 300

be_major BL, DXF 98
be_minor BL, DXF 99

eed1071 BL, DXF 1071 bg_bl91 BL, DXF 91 bg_bl92 BL, DXF 92

bg_location

3BD, DXF 1010

bg_insert_cycling

B, DXF 280

bg_insert_cycling_weight

BLd, DXF 93

BLOCKLOOKUPPARAMETER parent struct dwg object object* evalexpr Dwg EvalExpr name T, DXF 300

be_major BL, DXF 98
be_minor BL, DXF 99

eed1071 BL, DXF 1071

show_properties

B, DXF 280

chain_actions

B, DXF 281

def_pt 3BD, DXF 1010

num_propinfos

BL, DXF 93

prop1 Dwg BLOCKPARAMETER PropInfo prop2 Dwg BLOCKPARAMETER PropInfo

lookup_name

T, DXF 303

lookup_desc

T, DXF 304

index BL, DXF 94

unknown_t

T

BLOCKMOVEACTION parent struct dwg object object* evalexpr Dwg EvalExpr name T, DXF 300

be_major BL, DXF 98
be_minor BL, DXF 99

eed1071 BL, DXF 1071

display_location

3BD, DXF 1010

num_actions

BL, DXF 70

actions BL*, DXF 91

num_deps BL, DXF 71

deps H*, DXF 330 conn_pts Dwg BLOCKACTION connectionpts

action_offset_x

BD, DXF 140

action_offset_y

BD, DXF 141

angle_offset

BD

BLOCKPARAMDEPENDENCYBODY parent struct dwg object object*

adb_version

BS, DXF 90

dimbase_version

BS, DXF 90

name T, DXF 1

class_version

BS, DXF 90

BLOCKPOINTPARAMETER parent struct dwg object object* evalexpr Dwg EvalExpr name T, DXF 300

be_major BL, DXF 98
be_minor BL, DXF 99

eed1071 BL, DXF 1071

show_properties

B, DXF 280

chain_actions

B, DXF 281

def_pt 3BD, DXF 1010

num_propinfos

BL, DXF 93

prop1 Dwg BLOCKPARAMETER PropInfo prop2 Dwg BLOCKPARAMETER PropInfo

position_name

T, DXF 303

position_desc

T, DXF 304

def_label_pt

3BD, DXF 1011 BLOCKPOLARGRIP

parent struct dwg object object* evalexpr Dwg EvalExpr name T, DXF 300

be_major BL, DXF 98
be_minor BL, DXF 99

eed1071 BL, DXF 1071 bg_bl91 BL, DXF 91 bg_bl92 BL, DXF 92

bg_location

3BD, DXF 1010

bg_insert_cycling

B, DXF 280

bg_insert_cycling_weight

BLd, DXF 93

BLOCKPOLARPARAMETER parent struct dwg object object* evalexpr Dwg EvalExpr name T, DXF 300

be_major BL, DXF 98
be_minor BL, DXF 99

eed1071 BL, DXF 1071

show_properties

B, DXF 280

chain_actions

B, DXF 281

def_basept

3BD, DXF 1010

def_endpt

3BD, DXF 1011

prop1 Dwg BLOCKPARAMETER PropInfo prop2 Dwg BLOCKPARAMETER PropInfo prop3 Dwg BLOCKPARAMETER PropInfo prop4 Dwg BLOCKPARAMETER PropInfo

prop_states

BL*, DXF 91

parameter_base_location

BS, DXF 177

upd_basept

3BD

basept  3BD
upd_endpt

3BD

endpt  3BD
angle_name

T, DXF 305

angle_desc

T, DXF 306

distance_name

T, DXF 305

distance_desc

T, DXF 306

offset BD, DXF 140

angle_value_set

Dwg BLOCKPARAMVALUESET

distance_value_set

Dwg BLOCKPARAMVALUESET

BLOCKPOLARSTRETCHACTION parent struct dwg object object* evalexpr Dwg EvalExpr name T, DXF 300

be_major BL, DXF 98
be_minor BL, DXF 99

eed1071 BL, DXF 1071

display_location

3BD, DXF 1010

num_actions

BL, DXF 70

actions BL*, DXF 91

num_deps BL, DXF 71

deps H*, DXF 330 conn_pts Dwg BLOCKACTION connectionpts num_pts BL, DXF 72 pts 2RD*, DXF 10

num_hdls BL, DXF 73

hdls H*, DXF 331 shorts BS*, DXF 74

num_codes

BL, DXF 75

codes BL*, DXF 76 BLOCKPROPERTIESTABLE

parent struct dwg object object* BLOCKPROPERTIESTABLEGRIP

parent struct dwg object object* evalexpr Dwg EvalExpr name T, DXF 300

be_major BL, DXF 98
be_minor BL, DXF 99

eed1071 BL, DXF 1071 bg_bl91 BL, DXF 91 bg_bl92 BL, DXF 92

bg_location

3BD, DXF 1010

bg_insert_cycling

B, DXF 280

bg_insert_cycling_weight

BLd, DXF 93

BLOCKRADIALCONSTRAINTPARAMETER parent struct dwg object object* evalexpr Dwg EvalExpr name T, DXF 300

be_major BL, DXF 98
be_minor BL, DXF 99

eed1071 BL, DXF 1071

show_properties

B, DXF 280

chain_actions

B, DXF 281

def_basept

3BD, DXF 1010

def_endpt

3BD, DXF 1011

prop1 Dwg BLOCKPARAMETER PropInfo prop2 Dwg BLOCKPARAMETER PropInfo prop3 Dwg BLOCKPARAMETER PropInfo prop4 Dwg BLOCKPARAMETER PropInfo

prop_states

BL*, DXF 91

parameter_base_location

BS, DXF 177

upd_basept

3BD

basept  3BD
upd_endpt

3BD

endpt  3BD
dependency

H, DXF 330

expr_name

T, DXF 305

expr_description

T, DXF 306

distance BD, DXF 140

value_set

Dwg BLOCKPARAMVALUESET

BLOCKREPRESENTATION parent struct dwg object object* flag BS, DXF 70 block H, DXF 340

BLOCKROTATEACTION parent struct dwg object object* evalexpr Dwg EvalExpr name T, DXF 300

be_major BL, DXF 98
be_minor BL, DXF 99

eed1071 BL, DXF 1071

display_location

3BD, DXF 1010

num_actions

BL, DXF 70

actions BL*, DXF 91

num_deps BL, DXF 71

deps H*, DXF 330 offset 3BD, DXF 1011 conn_pts Dwg BLOCKACTION connectionpts

dependent

B, DXF 280

base_pt 3BD, DXF 1012 BLOCKROTATIONGRIP

parent struct dwg object object* evalexpr Dwg EvalExpr name T, DXF 300

be_major BL, DXF 98
be_minor BL, DXF 99

eed1071 BL, DXF 1071 bg_bl91 BL, DXF 91 bg_bl92 BL, DXF 92

bg_location

3BD, DXF 1010

bg_insert_cycling

B, DXF 280

bg_insert_cycling_weight

BLd, DXF 93

BLOCKROTATIONPARAMETER parent struct dwg object object* evalexpr Dwg EvalExpr name T, DXF 300

be_major BL, DXF 98
be_minor BL, DXF 99

eed1071 BL, DXF 1071

show_properties

B, DXF 280

chain_actions

B, DXF 281

def_basept

3BD, DXF 1010

def_endpt

3BD, DXF 1011

prop1 Dwg BLOCKPARAMETER PropInfo prop2 Dwg BLOCKPARAMETER PropInfo prop3 Dwg BLOCKPARAMETER PropInfo prop4 Dwg BLOCKPARAMETER PropInfo

prop_states

BL*, DXF 91

parameter_base_location

BS, DXF 177

upd_basept

3BD

basept  3BD
upd_endpt

3BD

endpt  3BD
def_base_angle_pt

3BD, DXF 1011

angle_name

T, DXF 305

angle_desc

T, DXF 306

angle BD, DXF 140

angle_value_set

Dwg BLOCKPARAMVALUESET

BLOCKSCALEACTION parent struct dwg object object* evalexpr Dwg EvalExpr name T, DXF 300

be_major BL, DXF 98
be_minor BL, DXF 99

eed1071 BL, DXF 1071

display_location

3BD, DXF 1010

num_actions

BL, DXF 70

actions BL*, DXF 91

num_deps BL, DXF 71

deps H*, DXF 330 offset 3BD, DXF 1011 conn_pts Dwg BLOCKACTION connectionpts

dependent

B, DXF 280

base_pt 3BD, DXF 1012 BLOCKSTRETCHACTION

parent struct dwg object object* evalexpr Dwg EvalExpr name T, DXF 300

be_major BL, DXF 98
be_minor BL, DXF 99

eed1071 BL, DXF 1071

display_location

3BD, DXF 1010

num_actions

BL, DXF 70

actions BL*, DXF 91

num_deps BL, DXF 71

deps H*, DXF 330 conn_pts Dwg BLOCKACTION connectionpts num_pts BL, DXF 72 pts 2RD*, DXF 1011

num_hdls BL, DXF 73

hdls Dwg BLOCKSTRETCHACTION handles*

num_codes

BL, DXF 75

codes Dwg BLOCKSTRETCHACTION codes*

action_offset_x

BD, DXF 140

action_offset_y

BD, DXF 141

angle_offset

BD

BLOCKUSERPARAMETER parent struct dwg object object* evalexpr Dwg EvalExpr name T, DXF 300

be_major BL, DXF 98
be_minor BL, DXF 99

eed1071 BL, DXF 1071

show_properties

B, DXF 280

chain_actions

B, DXF 281

def_pt 3BD, DXF 1010

num_propinfos

BL, DXF 93

prop1 Dwg BLOCKPARAMETER PropInfo prop2 Dwg BLOCKPARAMETER PropInfo flag BS, DXF 90

assocvariable

H, DXF 330

expr T, DXF 301 value Dwg EvalVariant type BS, DXF 170

BLOCKVERTICALCONSTRAINTPARAMETER parent struct dwg object object* evalexpr Dwg EvalExpr name T, DXF 300

be_major BL, DXF 98
be_minor BL, DXF 99

eed1071 BL, DXF 1071

show_properties

B, DXF 280

chain_actions

B, DXF 281

def_basept

3BD, DXF 1010

def_endpt

3BD, DXF 1011

prop1 Dwg BLOCKPARAMETER PropInfo prop2 Dwg BLOCKPARAMETER PropInfo prop3 Dwg BLOCKPARAMETER PropInfo prop4 Dwg BLOCKPARAMETER PropInfo

prop_states

BL*, DXF 91

parameter_base_location

BS, DXF 177

upd_basept

3BD

basept  3BD
upd_endpt

3BD

endpt  3BD
dependency

H, DXF 330

expr_name

T, DXF 305

expr_description

T, DXF 306

value BD, DXF 140

value_set

Dwg BLOCKPARAMVALUESET BLOCKVISIBILITYGRIP

parent struct dwg object object* evalexpr Dwg EvalExpr name T, DXF 300

be_major BL, DXF 98
be_minor BL, DXF 99

eed1071 BL, DXF 1071 bg_bl91 BL, DXF 91 bg_bl92 BL, DXF 92

bg_location

3BD, DXF 1010

bg_insert_cycling

B, DXF 280

bg_insert_cycling_weight

BLd, DXF 93

BLOCKVISIBILITYPARAMETER parent struct dwg object object* evalexpr Dwg EvalExpr name T, DXF 300

be_major BL, DXF 98
be_minor BL, DXF 99

eed1071 BL, DXF 1071

show_properties

B, DXF 280

chain_actions

B, DXF 281

def_pt 3BD, DXF 1010

num_propinfos

BL, DXF 93

prop1 Dwg BLOCKPARAMETER PropInfo prop2 Dwg BLOCKPARAMETER PropInfo

is_initialized

B, DXF 281

unknown_bool

B, DXF 91

blockvisi_name

T, DXF 301

blockvisi_desc

T, DXF 302

num_blocks

BL, DXF 93

blocks H*, DXF 331

num_states

BL, DXF 92

states Dwg BLOCKVISIBILITYPARAMETER state* BLOCKXYGRIP

parent struct dwg object object* evalexpr Dwg EvalExpr name T, DXF 300

be_major BL, DXF 98
be_minor BL, DXF 99

eed1071 BL, DXF 1071 bg_bl91 BL, DXF 91 bg_bl92 BL, DXF 92

bg_location

3BD, DXF 1010

bg_insert_cycling

B, DXF 280

bg_insert_cycling_weight

BLd, DXF 93

BLOCKXYPARAMETER parent struct dwg object object* evalexpr Dwg EvalExpr name T, DXF 300

be_major BL, DXF 98
be_minor BL, DXF 99

eed1071 BL, DXF 1071

show_properties

B, DXF 280

chain_actions

B, DXF 281

def_basept

3BD, DXF 1010

def_endpt

3BD, DXF 1011

prop1 Dwg BLOCKPARAMETER PropInfo prop2 Dwg BLOCKPARAMETER PropInfo prop3 Dwg BLOCKPARAMETER PropInfo prop4 Dwg BLOCKPARAMETER PropInfo

prop_states

BL*, DXF 91

parameter_base_location

BS, DXF 177

upd_basept

3BD

basept  3BD
upd_endpt

3BD

endpt  3BD

x_label T, DXF 305

x_label_desc

T, DXF 306

y_label T, DXF 307

y_label_desc

T, DXF 308

x_value BD, DXF 142 y_value BD, DXF 141

x_value_set

Dwg BLOCKPARAMVALUESET

y_value_set

Dwg BLOCKPARAMVALUESET

BLOCK CONTROL BLOCK CONTROL is a table control object.

parent struct dwg object object*

num_entries

BS, DXF 70

entries H*

flags_r11

RS

model_space

H

paper_space

H

BLOCK HEADER BLOCK HEADER is a table object.

parent struct dwg object object*

flag  RC
name  TV
used  RSd
is_xref_ref

B

is_xref_resolved

BS

is_xref_dep

B

xref  H
__iterator

BL

anonymous

B

hasattrs B
blkisxref

B

xrefoverlaid

B

loaded_bit

B

num_owned

BL

base_pt 3DPOINT, DXF 10

xref_pname

T, DXF 1

num_inserts

RL

description

T, DXF 4

preview_size

BL

preview TF, DXF 310

insert_units

BS, DXF 70

explodable

B, DXF 280

block_scaling

RC, DXF 281

block_entity

H

first_entity

H

last_entity

H

entities H*
endblk_entity

H

inserts H*, DXF 331 layout H, DXF 340

flag2  RS
unknown_r11

RC

block_offset_r11

RL

BREAKDATA parent struct dwg object object*

num_pointrefs

BL, DXF 90

pointrefs

H*, DXF 330

dimref H, DXF 331 BREAKPOINTREF

parent struct dwg object object* CELLSTYLEMAP

parent struct dwg object object*

num_cells

BL, DXF 90

cells Dwg TABLESTYLE CellStyle* CONTEXTDATAMANAGER

parent struct dwg object object*

objectcontext

H

num_submgrs

BL

submgrs Dwg CONTEXTDATA submgr* CSACDOCUMENTOPTIONS

parent struct dwg object object*

class_version

BS

CURVEPATH parent struct dwg object object*

class_version

BS, DXF 90

entity H, DXF 340 DATALINK

parent struct dwg object object*

class_version

BS

data_adapter

T, DXF 1

description

T, DXF 300

tooltip T, DXF 301

connection_string

T, DXF 302

option BL, DXF 90

update_option

BL, DXF 91

bl92 BL, DXF 92 year BS, DXF 170 month BS, DXF 171 day BS, DXF 172 hour BS, DXF 173 minute BS, DXF 174 seconds BS, DXF 175 msec BS, DXF 176

path_option

BS, DXF 177

bl93 BL, DXF 93

update_status

T, DXF 304

num_customdata

BL, DXF 94

customdata

Dwg DATALINK customdata*

hardowner

H, DXF 360

DATATABLE parent struct dwg object object* flags BS, DXF 70

num_cols BL, DXF 90
num_rows BL, DXF 91
table_name

T, DXF 1

cols Dwg DATATABLE column* DBCOLOR

parent struct dwg object object* color CMC, DXF 62

DETAILVIEWSTYLE parent struct dwg object object*

mdoc_class_version

BS, DXF 70

desc T, DXF 3

is_modified_for_recompute

B, DXF 290

display_name

T, DXF 300

viewstyle_flags

BL, DXF 90

class_version

BS, DXF 70

flags BL, DXF 90

identifier_style

H, DXF 340

identifier_color

CMC, DXF 62

identifier_height

BD, DXF 40

identifier_exclude_characters

T, DXF 300

identifier_offset

BD, DXF 40

identifier_placement

RC, DXF 280

arrow_symbol

H, DXF 340

arrow_symbol_color

CMC, DXF 62

arrow_symbol_size

BD, DXF 40

boundary_ltype

H, DXF 340

boundary_linewt

BLd, DXF 90

boundary_line_color

CMC, DXF 62

viewlabel_text_style

H, DXF 340

viewlabel_text_color

CMC, DXF 62

viewlabel_text_height

BD, DXF 40

viewlabel_attachment

BL, DXF 90

viewlabel_offset

BD, DXF 40

viewlabel_alignment

BL, DXF 90

viewlabel_pattern

T, DXF 300

connection_ltype

H, DXF 340

connection_linewt

BLd, DXF 90

connection_line_color

CMC, DXF 62

borderline_ltype

H, DXF 340

borderline_linewt

BLd, DXF 90

borderline_color

CMC, DXF 62

model_edge

RC, DXF 280

DICTIONARY parent struct dwg object object*

numitems BL
is_hardowner

RC, DXF 280

cloning BS, DXF 281 texts T*, DXF 3

itemhandles

H*, DXF 350

DICTIONARYVAR parent struct dwg object object* schema RC, DXF 280

strvalue T, DXF 1

DICTIONARYWDFLT parent struct dwg object object*

numitems BL
is_hardowner

RC, DXF 280

cloning BS, DXF 281 texts T*, DXF 3

itemhandles

H*, DXF 350

defaultid

H, DXF 340

DIMASSOC parent struct dwg object object*

dimensionobj

H, DXF 330

associativity

BL, DXF 90

trans_space_flag

B, DXF 70

rotated_type

RC, DXF 71

ref Dwg DIMASSOC Ref* DIMSTYLE DIMSTYLE is a table object.

parent struct dwg object object* flag RC, DXF 70

name  T
used  RSd
is_xref_ref

B

is_xref_resolved

BS

is_xref_dep

B

xref  H

DIMTOL B, DXF 71 DIMLIM B, DXF 72 DIMTIH B, DXF 73 DIMTOH B, DXF 74 DIMSE1 B, DXF 75 DIMSE2 B, DXF 76 DIMALT B, DXF 170 DIMTOFL B, DXF 172 DIMSAH B, DXF 173 DIMTIX B, DXF 174 DIMSOXD B, DXF 175 DIMALTD BS, DXF 171 DIMZIN BS, DXF 78 DIMSD1 B, DXF 281 DIMSD2 B, DXF 282 DIMTOLJ BS, DXF 283 DIMJUST BS, DXF 280 DIMFIT BS, DXF 287 DIMUPT B, DXF 288 DIMTZIN BS, DXF 284 DIMALTZ BS, DXF 285 DIMALTTZ BS, DXF 286 DIMTAD BS, DXF 77 DIMUNIT BS, DXF 270 DIMAUNIT BS, DXF 275 DIMDEC BS, DXF 271 DIMTDEC BS, DXF 272 DIMALTU BS, DXF 273 DIMALTTD BS, DXF 274

DIMSCALE BD, DXF 40

DIMASZ BD, DXF 41 DIMEXO BD, DXF 42 DIMDLI BD, DXF 43 DIMEXE BD, DXF 44 DIMRND BD, DXF 45 DIMDLE BD, DXF 46 DIMTP BD, DXF 47 DIMTM BD, DXF 48 DIMFXL BD, DXF 49

DIMJOGANG

BD, DXF 50

DIMTFILL BS, DXF 69
DIMTFILLCLR

CMC, DXF 70

DIMAZIN BS, DXF 79

DIMARCSYM

BS, DXF 90

DIMTXT BD, DXF 140 DIMCEN BD, DXF 141 DIMTSZ BD, DXF 142 DIMALTF BD, DXF 143 DIMLFAC BD, DXF 144 DIMTVP BD, DXF 145 DIMTFAC BD, DXF 146 DIMGAP BD, DXF 147

DIMPOST T, DXF 3
DIMAPOST T, DXF 4
DIMBLK_T TV, DXF 5
DIMBLK1_T

TV, DXF 6

DIMBLK2_T

TV, DXF 7

DIMALTRND

BD, DXF 148

DIMCLRD_N

RS, DXF 176

DIMCLRE_N

RS, DXF 177

DIMCLRT_N

RS, DXF 178

DIMCLRD CMC, DXF 176 DIMCLRE CMC, DXF 177 DIMCLRT CMC, DXF 178 DIMADEC BS, DXF 179 DIMFRAC BS, DXF 276 DIMLUNIT BS, DXF 277 DIMDSEP BS, DXF 278 DIMTMOVE BS, DXF 279 DIMATFIT BS, DXF 289

DIMFXLON B, DXF 290
DIMTXTDIRECTION

B, DXF 294

DIMALTMZF

BD

DIMALTMZS

T

DIMMZF  BD
DIMMZS  T

DIMLWD BSd, DXF 371 DIMLWE BSd, DXF 372

flag0  B
DIMTXSTY H, DXF 340
DIMLDRBLK

H, DXF 341

DIMBLK H, DXF 342 DIMBLK1 H, DXF 343 DIMBLK2 H, DXF 344

DIMLTYPE H, DXF 345
DIMLTEX1 H, DXF 346
DIMLTEX2 H, DXF 347

DIMSTYLE CONTROL DIMSTYLE CONTROL is a table control object.

parent struct dwg object object*

num_entries

BS, DXF 70

entries H*
flags_r11

RS

num_morehandles

RC, DXF 71

morehandles

H*, DXF 340

DMDIMOBJECTCONTEXTDATA parent struct dwg object object*

class_version

BS, DXF 70

is_default

B, DXF 290

scale H, DXF 340

dimension

Dwg OCD Dimension

first_arc_pt

3BD, DXF 11

def_pt 3BD, DXF 12 DUMMY

parent struct dwg object object* DYNAMICBLOCKPROXYNODE

parent struct dwg object object* evalexpr Dwg EvalExpr

DYNAMICBLOCKPURGEPREVENTER parent struct dwg object object* flag BS, DXF 70

block  H

EVALUATION GRAPH parent struct dwg object object*

major  BL
minor  BL
first_nodeid

BLd, DXF 96

first_nodeid_copy

BLd, DXF 97

num_nodes

BL

nodes Dwg EVAL Node*

has_graph

B

num_edges

BL

edges Dwg EVAL Edge* FCFOBJECTCONTEXTDATA

parent struct dwg object object*

class_version

BS, DXF 70

is_default

B, DXF 290

scale H, DXF 340 location 3BD, DXF 10

horiz_dir

3BD, DXF 11

FIELD parent struct dwg object object* id T, DXF 1 code T, DXF 2

num_childs

BL, DXF 90

childs H*, DXF 360

num_objects

BL, DXF 97

objects H*, DXF 331 format TV, DXF 4

evaluation_option

BL, DXF 91

filing_option

BL, DXF 92

field_state

BL, DXF 94

evaluation_status

BL, DXF 95

evaluation_error_code

BL, DXF 96

evaluation_error_msg

T, DXF 300

value Dwg TABLE value

value_string

T, DXF 301

value_string_length

BL, DXF 98

num_childval

BL, DXF 93

childval Dwg FIELD ChildValue* FIELDLIST

parent struct dwg object object*

num_fields

BL, DXF 90

unknown B

fields H*, DXF 330 GEODATA

parent struct dwg object object*

class_version

BL, DXF 90

host_block

H, DXF 330

coord_type

BS, DXF 70

design_pt

3BD, DXF 10

ref_pt 3BD 1, DXF 11 obs_pt 3BD, DXF 11

scale_vec

3BD 1, DXF 43

unit_scale_horiz

BD, DXF 40

units_value_horiz

BL, DXF 91

unit_scale_vert

BD, DXF 41

units_value_vert

BL, DXF 92

up_dir 3BD, DXF 210

north_dir

2RD, DXF 12

scale_est

BL, DXF 95

user_scale_factor

BD, DXF 141

do_sea_level_corr

B, DXF 294

sea_level_elev

BD, DXF 142

coord_proj_radius

BD, DXF 143

coord_system_def

T, DXF 301

geo_rss_tag

T, DXF 302

coord_system_datum

T, DXF 303

coord_system_wkt

T, DXF 304

observation_from_tag

T, DXF 305

observation_to_tag

T, DXF 306

observation_coverage_tag

T, DXF 307

num_geomesh_pts

BL, DXF 93

geomesh_pts

Dwg GEODATA meshpt*

num_geomesh_faces

BL, DXF 96

geomesh_faces

Dwg GEODATA meshface*

has_civil_data

B

obsolete_false

B, DXF 292

ref_pt2d 2RD, DXF 15 zero1 3BD, DXF 16

unknown1 BL, DXF 93
unknown2 BL, DXF 94
unknown_b

B, DXF 293

north_dir_angle_deg

BD, DXF 54

north_dir_angle_rad

BD, DXF 140

GEOMAPIMAGE parent struct dwg object object*

class_version

BL, DXF 90

pt0 3BD, DXF 10 size 2RD, DXF 13

display_props

BS, DXF 70

clipping B, DXF 280
brightness

RC, DXF 281

contrast RC, DXF 282 fade RC, DXF 283

rotation BD
image_width

BD

image_height

BD

name  T
image_file

BD

image_visibility

BD

transparency

BS

height  BD
width  BD
show_rotation

B

scale_factor

BD

geoimage_brightness

BS

geoimage_contrast

BS

geoimage_fade

BS

geoimage_position

BS

geoimage_width

BS

geoimage_height

BS

GRADIENT BACKGROUND parent struct dwg object object*

class_version

BL, DXF 90

color_top

BLx, DXF 90

color_middle

BLx, DXF 91

color_bottom

BLx, DXF 92

horizon BD, DXF 140 height BD, DXF 141 rotation BD, DXF 142

GROUND PLANE BACKGROUND parent struct dwg object object*

class_version

BL, DXF 90

color_sky_zenith

BLx, DXF 90

color_sky_horizon

BLx, DXF 91

color_underground_horizon

BLx, DXF 92

color_underground_azimuth

BLx, DXF 93

color_near

BLx, DXF 94

color_far

BLx, DXF 95

GROUP parent struct dwg object object* name T, DXF 300 unnamed BS, DXF 70

selectable

BS, DXF 71

num_groups

BL

groups H*, DXF 340 IBL BACKGROUND

parent struct dwg object object*

class_version

BL, DXF 90

enable B, DXF 290 name T, DXF 1

rotation BD, DXF 40
display_image

B, DXF 290

secondary_background

H, DXF 340

IDBUFFER parent struct dwg object object*

unknown RC
num_obj_ids

BL

obj_ids H*, DXF 330 IMAGEDEF

parent struct dwg object object*

class_version

BL, DXF 90

image_size

2RD, DXF 10

file_path

T, DXF 1

is_loaded

B, DXF 280

resunits RC, DXF 281

pixel_size

2RD, DXF 11

IMAGEDEF REACTOR parent struct dwg object object*

class_version

BL, DXF 90

IMAGE BACKGROUND parent struct dwg object object*

class_version

BL, DXF 90

filename T, DXF 300
fit_to_screen

B, DXF 290

maintain_aspect_ratio

B, DXF 291

use_tiling

B, DXF 292

offset 2BD 1, DXF 140 scale 2BD 1, DXF 142

INDEX parent struct dwg object object*

last_updated

TIMEBLL, DXF 40 LAYER LAYER is a table object.

parent struct dwg object object*

flag  BS
name  T
used  RSd
is_xref_ref

B

is_xref_resolved

BS

is_xref_dep

B

xref  H
flag0  BS
frozen  B
on  B
frozen_in_new

B

locked  B
plotflag B, DXF 290

linewt RC, DXF 370 color CMC, DXF 62

plotstyle

H, DXF 390

material H, DXF 347

ltype H, DXF 6

visualstyle

H, DXF 348

LAYERFILTER parent struct dwg object object*

num_names

BL

names TV*, DXF 8 LAYER CONTROL LAYER CONTROL is a table control object.

parent struct dwg object object*

num_entries

BS, DXF 70

entries H*
flags_r11

RS

LAYER INDEX parent struct dwg object object*

last_updated

TIMEBLL, DXF 40

num_entries

BL

entries Dwg LAYER entry* LAYOUT

parent struct dwg object object*

plotsettings

Dwg Object PLOTSETTINGS

layout_name

T, DXF 1

tab_order

BS, DXF 71

layout_flags

BS, DXF 70

INSBASE 3DPOINT, DXF 12 LIMMIN 2DPOINT, DXF 10 LIMMAX 2DPOINT, DXF 11 UCSORG 3DPOINT, DXF 13 UCSXDIR 3DPOINT, DXF 16 UCSYDIR 3DPOINT, DXF 17

ucs_elevation

BD, DXF 146

UCSORTHOVIEW

BS, DXF 76

EXTMIN 3DPOINT, DXF 14 EXTMAX 3DPOINT, DXF 15

block_header

H, DXF 330

active_viewport

H, DXF 331

base_ucs H, DXF 346
named_ucs

H, DXF 345

num_viewports

BL

viewports

H*

LEADEROBJECTCONTEXTDATA parent struct dwg object object*

class_version

BS, DXF 70

is_default

B, DXF 290

scale H, DXF 340

num_points

BL, DXF 70

points 3DPOINT*, DXF 10 b290 B, DXF 290

x_direction

3DPOINT, DXF 11

inspt_offset

3DPOINT, DXF 12

endptproj

3DPOINT, DXF 13

LIGHTLIST parent struct dwg object object*

class_version

BL, DXF 90

num_lights

BL, DXF 90

lights Dwg LIGHTLIST light* LONG TRANSACTION

parent struct dwg object object* LTYPE LTYPE is a table object.

parent struct dwg object object*

flag  RC
name  TV
used  RSd
is_xref_ref

B

is_xref_resolved

BS

is_xref_dep

B

xref  H
description

T, DXF 3

pattern_len

BD, DXF 40

alignment

RC, DXF 72

numdashes

RC, DXF 73

dashes Dwg LTYPE dash*

dashes_r11

RD, DXF 49

has_strings_area

B

strings_area

TF

unknown_r11

RC

LTYPE CONTROL LTYPE CONTROL is a table control object.

parent struct dwg object object*

num_entries

BS, DXF 70

entries H*

flags_r11

RS

bylayer H
byblock H

MATERIAL parent struct dwg object object* name T, DXF 1

description

T, DXF 2

ambient_color

Dwg MATERIAL color

diffuse_color

Dwg MATERIAL color

diffusemap

Dwg MATERIAL mapper

specular_gloss_factor

BD, DXF 44

specular_color

Dwg MATERIAL color

specularmap

Dwg MATERIAL mapper

reflectionmap

Dwg MATERIAL mapper

opacity_percent

BD, DXF 140

opacitymap

Dwg MATERIAL mapper

bumpmap Dwg MATERIAL mapper

refraction_index

BD, DXF 145

refractionmap

Dwg MATERIAL mapper

color_bleed_scale

BD, DXF 460

indirect_bump_scale

BD, DXF 461

reflectance_scale

BD, DXF 462

transmittance_scale

BD, DXF 463

two_sided_material

B, DXF 290

luminance

BD, DXF 464

luminance_mode

BS, DXF 270

translucence

BD, DXF 148

self_illumination

BD, DXF 149

reflectivity

BD, DXF 468

illumination_model

BL, DXF 93

channel_flags

BL, DXF 94

mode BL, DXF 282

genprocname

T

genproctype

BS

genprocvalbool

B

genprocvalint

BS

genprocvalreal

BD

genprocvaltext

T

genprocvalcolor

CMC

genproctableend

B

num_gentextures

BS

gentextures

Dwg MATERIAL gentexture* MENTALRAYRENDERSETTINGS

parent struct dwg object object*

class_version

BL, DXF 90

name T, DXF 1

fog_enabled

B, DXF 290

fog_background_enabled

B, DXF 290

backfaces_enabled

B, DXF 290

environ_image_enabled

B, DXF 290

environ_image_filename

T, DXF 1

description

T, DXF 1

display_index

BL, DXF 90

has_predefined

B, DXF 290

mr_version

BL, DXF 90

sampling1

BL, DXF 90

sampling2

BL, DXF 90

sampling_mr_filter

BS, DXF 70

sampling_filter1

BD, DXF 40

sampling_filter2

BD, DXF 40

sampling_contrast_color1

BD, DXF 40

sampling_contrast_color2

BD, DXF 40

sampling_contrast_color3

BD, DXF 40

sampling_contrast_color4

BD, DXF 40

shadow_mode

BS, DXF 70

shadow_maps_enabled

B, DXF 290

ray_tracing_enabled

B, DXF 290

ray_trace_depth1

BL, DXF 90

ray_trace_depth2

BL, DXF 90

ray_trace_depth3

BL, DXF 90

global_illumination_enabled

B, DXF 290

gi_sample_count

BL, DXF 90

gi_sample_radius_enabled

B, DXF 290

gi_sample_radius

BD, DXF 40

gi_photons_per_light

BL, DXF 90

photon_trace_depth1

BL, DXF 90

photon_trace_depth2

BL, DXF 90

photon_trace_depth3

BL, DXF 90

final_gathering_enabled

B, DXF 290

fg_ray_count

BL, DXF 90

fg_sample_radius_state1

B, DXF 290

fg_sample_radius_state2

B, DXF 290

fg_sample_radius_state3

B, DXF 290

fg_sample_radius1

BD, DXF 40

fg_sample_radius2

BD, DXF 40

light_luminance_scale

BD, DXF 40

diagnostics_mode

BS, DXF 70

diagnostics_grid_mode

BS, DXF 70

diagnostics_grid_float

BD, DXF 40

diagnostics_photon_mode

BS, DXF 70

diagnostics_bsp_mode

BS, DXF 70

export_mi_enabled

B, DXF 290

mr_description

T, DXF 1

tile_size

BL, DXF 90

tile_order

BS, DXF 70

memory_limit

BL, DXF 90

diagnostics_samples_mode

B, DXF 290

energy_multiplier

BD, DXF 40

MLEADEROBJECTCONTEXTDATA parent struct dwg object object*

class_version

BS, DXF 70

is_default

B, DXF 290

scale H, DXF 340 MLEADERSTYLE

parent struct dwg object object*

class_version

BS, DXF 179

content_type

BS, DXF 170

mleader_order

BS, DXF 171

leader_order

BS, DXF 172

max_points

BL, DXF 90

first_seg_angle

BD, DXF 40

second_seg_angle

BD, DXF 41

type BS, DXF 173

line_color

CMC, DXF 91

line_type

H, DXF 340

linewt BLd, DXF 92

has_landing

B, DXF 290

has_dogleg

B, DXF 291

landing_gap

BD, DXF 42

landing_dist

BD, DXF 43

description

T, DXF 3

arrow_head

H, DXF 341

arrow_head_size

BD, DXF 44

text_default

T, DXF 300

text_style

H, DXF 342

attach_left

BS, DXF 174

attach_right

BS, DXF 178

text_angle_type

BS, DXF 175

text_align_type

BS, DXF 176

text_color

CMC, DXF 93

text_height

BD, DXF 45

has_text_frame

B, DXF 292

text_always_left

B, DXF 297

align_space

BD, DXF 46

block H, DXF 343

block_color

CMC, DXF 94

block_scale

3BD

use_block_scale

B, DXF 293

block_rotation

BD, DXF 141

use_block_rotation

B, DXF 294

block_connection

BS, DXF 177

scale BD, DXF 142

is_changed

B, DXF 295

is_annotative

B, DXF 296

break_size

BD, DXF 143

attach_dir

BS, DXF 271

attach_top

BS, DXF 273

attach_bottom

BS, DXF 272

text_extended

B, DXF 298

MLINESTYLE parent struct dwg object object* name T, DXF 2

description

T, DXF 3

flag BS, DXF 70

fill_color

CMC, DXF 62

start_angle

BD, DXF 51

end_angle

BD, DXF 52

num_lines

RC, DXF 71

lines Dwg MLINESTYLE line* MOTIONPATH

parent struct dwg object object*

class_version

BS, DXF 90

camera_path

H, DXF 340

target_path

H, DXF 340

viewtable

H, DXF 340 frames BS, DXF 90

frame_rate

BS, DXF 90

corner_decel

B, DXF 290

MTEXTATTRIBUTEOBJECTCONTEXTDATA parent struct dwg object object*

class_version

BS, DXF 70

is_default

B, DXF 290

scale H, DXF 340

horizontal_mode

BS, DXF 70

rotation BD, DXF 50

ins_pt 2RD, DXF 10

alignment_pt

2RD, DXF 11

enable_context

B, DXF 290

context Dwg Object SCALE MTEXTOBJECTCONTEXTDATA

parent struct dwg object object*

class_version

BS, DXF 70

is_default

B, DXF 290

scale H, DXF 340

attachment

BL, DXF 70

ins_pt 3BD, DXF 10

x_axis_dir

3BD, DXF 11

rect_height

BD, DXF 41

rect_width

BD, DXF 40

extents_width

BD, DXF 42

extents_height

BD, DXF 43

column_type

BL, DXF 71

column_width

BD, DXF 44

gutter BD, DXF 45

auto_height

B, DXF 73

flow_reversed

B, DXF 74

num_column_heights

BL, DXF 72

column_heights

BD*, DXF 46

NAVISWORKSMODELDEF parent struct dwg object object* flags BS, DXF 70 path T, DXF 1 status B, DXF 290

min_extent

3BD, DXF 10

max_extent

3BD, DXF 11

host_drawing_visibility

B, DXF 290

OBJECT PTR

parent struct dwg object object* ORDDIMOBJECTCONTEXTDATA

parent struct dwg object object*

class_version

BS, DXF 70

is_default

B, DXF 290

scale H, DXF 340

dimension

Dwg OCD Dimension

feature_location_pt

3BD, DXF 11

leader_endpt

3BD, DXF 12

PARTIAL VIEWING INDEX parent struct dwg object object*

num_entries

BL

has_entries

B

entries Dwg PARTIAL VIEWING INDEX Entry* PERSUBENTMGR

parent struct dwg object object*

class_version

BL, DXF 90

unknown_0

BL, DXF 90

unknown_2

BL, DXF 90

numassocsteps

BL, DXF 90

numassocsubents

BL, DXF 90

num_steps

BL, DXF 90

steps BL*, DXF 90

num_subents

BL, DXF 90

subents BL*, DXF 90 PLACEHOLDER

parent struct dwg object object* PLOTSETTINGS

parent struct dwg object object*

printer_cfg_file

T, DXF 1

paper_size

T, DXF 2

canonical_media_name

T, DXF 4

plot_flags

BS, DXF 70

plotview H, DXF 6
plotview_name

T, DXF 6

left_margin

BD, DXF 40

bottom_margin

BD, DXF 41

right_margin

BD, DXF 42

top_margin

BD, DXF 43

paper_width

BD, DXF 44

paper_height

BD, DXF 45

plot_origin

2BD 1, DXF 46

plot_window_ll

2BD 1, DXF 48

plot_window_ur

2BD 1, DXF 140

plot_paper_unit

BS, DXF 72

plot_rotation_mode

BS, DXF 73

plot_type

BS, DXF 74

paper_units

BD, DXF 142

drawing_units

BD, DXF 143

stylesheet

T, DXF 7

std_scale_type

BS, DXF 75

std_scale_factor

BD, DXF 147

paper_image_origin

2BD 1, DXF 148

shadeplot_type

BS, DXF 76

shadeplot_reslevel

BS, DXF 77

shadeplot_customdpi

BS, DXF 78

shadeplot

H, DXF 333

POINTCLOUDCOLORMAP parent struct dwg object object*

class_version

BS, DXF 70

def_intensity_colorscheme

T, DXF 1

def_elevation_colorscheme

T, DXF 1

def_classification_colorscheme

T, DXF 1

num_colorramps

BL, DXF 90

colorramps

Dwg POINTCLOUDCOLORMAP Ramp*

num_classification_colorramps

BL, DXF 90

classification_colorramps

Dwg POINTCLOUDCOLORMAP Ramp*

POINTCLOUDDEF parent struct dwg object object*

class_version

BL, DXF 90

source_filename

T, DXF 1

is_loaded

B, DXF 280

numpoints

RLL, DXF 160

extents_min

3BD, DXF 10

extents_max

3BD, DXF 11

POINTCLOUDDEFEX parent struct dwg object object*

class_version

BL, DXF 90

source_filename

T, DXF 1

is_loaded

B, DXF 280

numpoints

RLL, DXF 160

extents_min

3BD, DXF 10

extents_max

3BD, DXF 11

POINTCLOUDDEF REACTOR parent struct dwg object object*

class_version

BL, DXF 90

POINTCLOUDDEF REACTOR EX parent struct dwg object object*

class_version

BL, DXF 90

POINTPATH parent struct dwg object object*

class_version

BS, DXF 90

point 3BD, DXF 10 PROXY OBJECT

parent struct dwg object object*

class_id BL, DXF 90

version BL, DXF 71

maint_version

BL, DXF 97

from_dxf B, DXF 70
data_numbits

BL

data_size

BL, DXF 161

data TF, DXF 310

num_objids

BL

objids H*, DXF 340 RADIMLGOBJECTCONTEXTDATA

parent struct dwg object object*

class_version

BS, DXF 70

is_default

B, DXF 290

scale H, DXF 340

dimension

Dwg OCD Dimension

ovr_center

3BD, DXF 12

jog_point

3BD, DXF 13

RADIMOBJECTCONTEXTDATA parent struct dwg object object*

class_version

BS, DXF 70

is_default

B, DXF 290

scale H, DXF 340

dimension

Dwg OCD Dimension

first_arc_pt

3BD, DXF 11 RAPIDRTRENDERSETTINGS

parent struct dwg object object*

class_version

BL, DXF 90

name T, DXF 1

fog_enabled

B, DXF 290

fog_background_enabled

B, DXF 290

backfaces_enabled

B, DXF 290

environ_image_enabled

B, DXF 290

environ_image_filename

T, DXF 1

description

T, DXF 1

display_index

BL, DXF 90

has_predefined

B, DXF 290

rapidrt_version

BL, DXF 90

render_target

BL, DXF 70

render_level

BL, DXF 90

render_time

BL, DXF 90

lighting_model

BL, DXF 70

filter_type

BL, DXF 70

filter_width

BD, DXF 40

filter_height

BD, DXF 40 RASTERVARIABLES

parent struct dwg object object*

class_version

BL, DXF 90

image_frame

BS, DXF 70

image_quality

BS, DXF 71

units BS, DXF 72 RENDERENTRY

parent struct dwg object object*

class_version

BL, DXF 90

image_file_name

T, DXF 1

preset_name

T, DXF 1

view_name

T, DXF 1

dimension_x

BL, DXF 90

dimension_y

BL, DXF 90

start_year

BS, DXF 70

start_month

BS, DXF 70

start_day

BS, DXF 70

start_minute

BS, DXF 70

start_second

BS, DXF 70

start_msec

BS, DXF 70

render_time

BD, DXF 40

memory_amount

BL, DXF 90

material_count

BL, DXF 90

light_count

BL, DXF 90

triangle_count

BL, DXF 90

display_index

BL, DXF 90

RENDERENVIRONMENT parent struct dwg object object*

class_version

BL, DXF 90

fog_enabled

B, DXF 290

fog_background_enabled

B, DXF 290

fog_color_r

RC, DXF 280

fog_color_g

RC, DXF 280

fog_color_b

RC, DXF 280

fog_density_near

BD, DXF 40

fog_density_far

BD, DXF 40

fog_distance_near

BD, DXF 40

fog_distance_far

BD, DXF 40

environ_image_enabled

B, DXF 290

environ_image_filename

T, DXF 1

RENDERGLOBAL parent struct dwg object object*

class_version

BL, DXF 90

procedure

BL, DXF 90

destination

BL, DXF 90

save_enabled

B, DXF 290

save_filename

T, DXF 1

image_width

BL, DXF 90

image_height

BL, DXF 90

predef_presets_first

B, DXF 290

highlevel_info

B, DXF 290

RENDERSETTINGS parent struct dwg object object*

class_version

BL, DXF 90

name T, DXF 1

fog_enabled

B, DXF 290

fog_background_enabled

B, DXF 290

backfaces_enabled

B, DXF 290

environ_image_enabled

B, DXF 290

environ_image_filename

T, DXF 1

description

T, DXF 1

display_index

BL, DXF 90

has_predefined

B, DXF 290 SCALE

parent struct dwg object object* flag BS, DXF 70 name T, DXF 300

paper_units

BD, DXF 140

drawing_units

BD, DXF 141

is_unit_scale

B, DXF 290

SECTIONVIEWSTYLE parent struct dwg object object*

mdoc_class_version

BS, DXF 70

desc T, DXF 3

is_modified_for_recompute

B, DXF 290

display_name

T, DXF 300

viewstyle_flags

BL, DXF 90

class_version

BS, DXF 70

flags BL, DXF 90

identifier_style

H, DXF 340

identifier_color

CMC, DXF 62

identifier_height

BD, DXF 40

arrow_start_symbol

H, DXF 340

arrow_end_symbol

H, DXF 340

arrow_symbol_color

CMC, DXF 62

arrow_symbol_size

BD, DXF 40

identifier_exclude_characters

T, DXF 300

identifier_position

BLd, DXF 90

identifier_offset

BD, DXF 40

arrow_position

BLd, DXF 90

arrow_symbol_extension_length

BD, DXF 40

plane_ltype

H, DXF 340

plane_linewt

BLd, DXF 90

plane_line_color

CMC, DXF 62

bend_ltype

H, DXF 340

bend_linewt

BLd, DXF 90

bend_line_color

CMC, DXF 62

bend_line_length

BD, DXF 40

end_line_overshoot

BD, DXF 40

end_line_length

BD, DXF 40

viewlabel_text_style

H, DXF 340

viewlabel_text_color

CMC, DXF 62

viewlabel_text_height

BD, DXF 40

viewlabel_attachment

BL, DXF 90

viewlabel_offset

BD, DXF 40

viewlabel_alignment

BL, DXF 90

viewlabel_pattern

T, DXF 300

hatch_color

CMC, DXF 62

hatch_bg_color

CMC, DXF 62

hatch_pattern

T, DXF 300

hatch_scale

BD, DXF 40

hatch_transparency

BLd, DXF 90

unknown_b1

B, DXF 290

unknown_b2

B, DXF 290

num_hatch_angles

BL, DXF 90

hatch_angles

BD*, DXF 40

SECTION MANAGER parent struct dwg object object*

is_live B, DXF 70
num_sections

BS, DXF 90

sections H*, DXF 330 SECTION SETTINGS

parent struct dwg object object*

curr_type

BS, DXF 90

num_types

BL, DXF 91

types Dwg SECTION typesettings* SKYLIGHT BACKGROUND

parent struct dwg object object*

class_version

BL, DXF 90

sunid H, DXF 340 SOLID BACKGROUND

parent struct dwg object object*

class_version

BL, DXF 90

color BLx, DXF 90 SORTENTSTABLE

parent struct dwg object object*

num_ents BL
sort_ents

H*, DXF 5

block_owner

H

ents H*, DXF 331 SPATIAL FILTER

parent struct dwg object object*

num_clip_verts

BS, DXF 70

clip_verts

2RD*, DXF 10

extrusion

BE, DXF 210

origin 3BD, DXF 11

display_boundary_on

BS, DXF 71

front_clip_on

BS, DXF 72

front_clip_z

BD, DXF 40

back_clip_on

BS, DXF 73

back_clip_z

BD, DXF 41

inverse_transform

BD*, DXF 40

transform

BD*, DXF 40

SPATIAL INDEX parent struct dwg object object*

last_updated

TIMEBLL, DXF 40

num1 BD, DXF 40

num_hdls BL, DXF 90

hdls H*, DXF 330

bindata_size

BL, DXF 90

bindata TF, DXF 310 STYLE STYLE is a table object.

parent struct dwg object object*

flag  RC
name  TV
used  RSd
is_xref_ref

B

is_xref_resolved

BS

is_xref_dep

B

xref  H
is_shape B
is_vertical

B

text_size

BD, DXF 40

width_factor

BD, DXF 41

oblique_angle

BD, DXF 50

generation

RC, DXF 71

last_height

BD, DXF 42

font_file

T, DXF 3

bigfont_file

T, DXF 4

STYLE CONTROL STYLE CONTROL is a table control object.

parent struct dwg object object*

num_entries

BS, DXF 70

entries H*
flags_r11

RS

SUN parent struct dwg object object*

class_version

BL, DXF 90

is_on B, DXF 290 color CMC, DXF 63

intensity

BD, DXF 40

has_shadow

B, DXF 291

julian_day

BL, DXF 91

msecs BL, DXF 92 is_dst B, DXF 292

shadow_type

BL, DXF 70

shadow_mapsize

BS, DXF 71

shadow_softness

RC, DXF 280

SUNSTUDY parent struct dwg object object*

class_version

BL, DXF 90

setup_name

T, DXF 1

description

T, DXF 2

output_type

BL, DXF 70

sheet_set_name

T, DXF 3

use_subset

B, DXF 290

sheet_subset_name

T, DXF 4

select_dates_from_calendar

B, DXF 291

num_dates

BL, DXF 91

dates Dwg SUNSTUDY Dates*

select_range_of_dates

B, DXF 292

start_time

BL, DXF 93

end_time BL, DXF 94
interval BL, DXF 95
num_hours

BL, DXF 91

hours B*, DXF 290

shade_plot_type

BL, DXF 74

numvports

BL, DXF 75

numrows BL, DXF 76 numcols BL, DXF 77 spacing BD, DXF 40

lock_viewports

B, DXF 293

label_viewports

B, DXF 294

page_setup_wizard

H, DXF 340

view H, DXF 341

visualstyle

H, DXF 342

text_style

H, DXF 343

TABLECONTENT parent struct dwg object object* ldata Dwg LinkedData tdata Dwg LinkedTableData fdata Dwg FormattedTableData

tablestyle

H, DXF 340

TABLEGEOMETRY parent struct dwg object object* numrows BL, DXF 90 numcols BL, DXF 91

num_cells

BL, DXF 92

cells Dwg TABLEGEOMETRY Cell* TABLESTYLE

parent struct dwg object object*

class_version

BS

name T, DXF 3 flags BS, DXF 71

flow_direction

BS, DXF 70

horiz_cell_margin

BD, DXF 40

vert_cell_margin

BD, DXF 41

is_title_suppressed

B, DXF 280

is_header_suppressed

B, DXF 281

unknown_rc

RC, DXF 70

unknown_bl1

BL

unknown_bl2

BL

cellstyle

H

sty Dwg TABLESTYLE CellStyle

numoverrides

BL

unknown_bl3

BL

ovr Dwg TABLESTYLE CellStyle

num_rowstyles

BL

rowstyles

Dwg TABLESTYLE rowstyles*

TEXTOBJECTCONTEXTDATA parent struct dwg object object*

class_version

BS, DXF 70

is_default

B, DXF 290

scale H, DXF 340

horizontal_mode

BS, DXF 70

rotation BD, DXF 50

ins_pt 2RD, DXF 10

alignment_pt

2RD, DXF 11

TVDEVICEPROPERTIES parent struct dwg object object*

flags  BL
max_regen_threads

BS

use_lut_palette

BL

alt_hlt BLL
alt_hltcolor

BLL

geom_shader_usage

BLL

blending_mode

BL

antialiasing_level

BD

bd2  BD

UCS UCS is a table object.

parent struct dwg object object*

flag  RC
name  TV
used  RSd
is_xref_ref

B

is_xref_resolved

BS

is_xref_dep

B

xref  H

ucsorg 3BD, DXF 10 ucsxdir 3BD, DXF 11 ucsydir 3BD, DXF 12

ucs_elevation

BD, DXF 146

UCSORTHOVIEW

BS, DXF 79

base_ucs H, DXF 346

named_ucs

H

num_orthopts

BS

orthopts Dwg UCS orthopts* UCS CONTROL UCS CONTROL is a table control object.

parent struct dwg object object*

num_entries

BS, DXF 70

entries H*
flags_r11

RS

UNKNOWN OBJ

parent struct dwg object object* VBA PROJECT

parent struct dwg object object*

data_size

BL, DXF 90

data TF, DXF 310 VIEW VIEW is a table object.

parent struct dwg object object*

flag  RC
name  TV
used  RSd
is_xref_ref

B

is_xref_resolved

BS

is_xref_dep

B

xref  H
VIEWSIZE BD, DXF 40
view_width

BD, DXF 41

aspect_ratio

BD

VIEWCTR 2RD, DXF 10

view_target

3BD, DXF 12

VIEWDIR 3BD, DXF 11

twist_angle

BD, DXF 50

lens_length

BD, DXF 42

front_clip_z

BD, DXF 43

back_clip_z

BD, DXF 44

VIEWMODE 4BITS, DXF 71

render_mode

RC, DXF 281

use_default_lights

B, DXF 292

default_lightning_type

RC, DXF 282

brightness

BD, DXF 141

contrast BD, DXF 142

ambient_color

CMC, DXF 63

is_pspace

B

associated_ucs

B, DXF 72

ucsorg 3BD, DXF 110 ucsxdir 3BD, DXF 111 ucsydir 3BD, DXF 112

ucs_elevation

BD, DXF 146

UCSORTHOVIEW

BS, DXF 79

is_camera_plottable

B, DXF 73

background

H, DXF 332

visualstyle

H, DXF 348

sun H, DXF 361

base_ucs H, DXF 346
named_ucs

H, DXF 345

livesection

H, DXF 334

flag_3d RS
unknown_r2

RC

unknown_r11

RC

VIEW CONTROL VIEW CONTROL is a table control object.

parent struct dwg object object*

num_entries

BS, DXF 70

entries H*
flags_r11

RS

VISUALSTYLE parent struct dwg object object*

description

T, DXF 2

style_type

BL, DXF 70

ext_lighting_model

BS, DXF 177

internal_only

B, DXF 291

face_lighting_model

BL, DXF 71

face_lighting_model_int

BS, DXF 176

face_lighting_quality

BL, DXF 72

face_lighting_quality_int

BS, DXF 176

face_color_mode

BL, DXF 73

face_color_mode_int

BS, DXF 176

face_opacity

BD, DXF 40

face_opacity_int

BS, DXF 176

face_specular

BD, DXF 41

face_specular_int

BS, DXF 176

face_modifier

BL, DXF 90

face_modifier_int

BS, DXF 176

face_mono_color

CMC, DXF 63

face_mono_color_int

BS, DXF 176

edge_model

BS, DXF 74

edge_model_int

BS, DXF 176

edge_style

BL, DXF 91

edge_style_int

BS, DXF 176

edge_intersection_color

CMC, DXF 64

edge_intersection_color_int

BS, DXF 176

edge_obscured_color

CMC, DXF 65

edge_obscured_color_int

BS, DXF 176

edge_obscured_ltype

BL, DXF 75

edge_obscured_ltype_int

BS, DXF 176

edge_intersection_ltype

BL, DXF 175

edge_intersection_ltype_int

BS, DXF 176

edge_crease_angle

BD, DXF 42

edge_crease_angle_int

BS, DXF 176

edge_modifier

BL, DXF 92

edge_modifier_int

BS, DXF 176

edge_color

CMC, DXF 66

edge_color_int

BS, DXF 176

edge_opacity

BD, DXF 43

edge_opacity_int

BS, DXF 176

edge_width

BL, DXF 76

edge_width_int

BS, DXF 176

edge_overhang

BL, DXF 77

edge_overhang_int

BS, DXF 176

edge_jitter

BL, DXF 78

edge_jitter_int

BS, DXF 176

edge_silhouette_color

CMC, DXF 67

edge_silhouette_color_int

BS, DXF 176

edge_silhouette_width

BL, DXF 79

edge_silhouette_width_int

BS, DXF 176

edge_halo_gap

BL, DXF 170

edge_halo_gap_int

BS, DXF 176

edge_isolines

BL, DXF 171

edge_isolines_int

BS, DXF 176

edge_do_hide_precision

B, DXF 290

edge_do_hide_precision_int

BS, DXF 176

edge_style_apply

BL, DXF 174

edge_style_apply_int

BS

display_settings

BL, DXF 93

display_settings_int

BS, DXF 176

display_brightness_bl

BLd, DXF 44

display_brightness

BD, DXF 44

display_brightness_int

BS, DXF 176

display_shadow_type

BL, DXF 173

display_shadow_type_int

BS, DXF 176

bd2007_45

BD, DXF 45

num_props

BS, DXF 70

b_prop1c B, DXF 290
b_prop1c_int

BS, DXF 176

b_prop1d B, DXF 290
b_prop1d_int

BS, DXF 176

b_prop1e B, DXF 290
b_prop1e_int

BS, DXF 176

b_prop1f B, DXF 290
b_prop1f_int

BS, DXF 176

b_prop20 B, DXF 290
b_prop20_int

BS, DXF 176

b_prop21 B, DXF 290
b_prop21_int

BS, DXF 176

b_prop22 B, DXF 290
b_prop22_int

BS, DXF 176

b_prop23 B, DXF 290
b_prop23_int

BS, DXF 176

b_prop24 B, DXF 290
b_prop24_int

BS, DXF 176

bl_prop25

BL, DXF 90

bl_prop25_int

BS, DXF 176

bd_prop26

BD, DXF 40

bd_prop26_int

BS, DXF 176

bd_prop27

BD, DXF 40

bd_prop27_int

BS, DXF 176

bl_prop28

BL, DXF 90

bl_prop28_int

BS, DXF 176

c_prop29 CMC, DXF 62

c_prop29_int

BS, DXF 176

bl_prop2a

BL, DXF 90

bl_prop2a_int

BS, DXF 176

bl_prop2b

BL, DXF 90

bl_prop2b_int

BS, DXF 176

c_prop2c CMC, DXF 62

c_prop2c_int

BS, DXF 176

b_prop2d B, DXF 290
b_prop2d_int

BS, DXF 176

bl_prop2e

BL, DXF 90

bl_prop2e_int

BS, DXF 176

bl_prop2f

BL, DXF 90

bl_prop2f_int

BS, DXF 176

bl_prop30

BL, DXF 90

bl_prop30_int

BS, DXF 176

b_prop31 B, DXF 290
b_prop31_int

BS, DXF 176

bl_prop32

BL, DXF 90

bl_prop32_int

BS, DXF 176

c_prop33 CMC, DXF 62

c_prop33_int

BS, DXF 176

bd_prop34

BD, DXF 40

bd_prop34_int

BS, DXF 176

edge_wiggle

BL, DXF 90

edge_wiggle_int

BS, DXF 176

strokes T, DXF 1
strokes_int

BS, DXF 176

b_prop37 B, DXF 290
b_prop37_int

BS, DXF 176

bd_prop38

BD, DXF 40

bd_prop38_int

BS, DXF 176

bd_prop39

BD, DXF 40

bd_prop39_int

BS, DXF 176

VPORT VPORT is a table object.

parent struct dwg object object*

flag  RC

name  TV
used  RSd
is_xref_ref

B

is_xref_resolved

BS

is_xref_dep

B

xref  H
VIEWSIZE BD, DXF 40
view_width

BD

aspect_ratio

BD, DXF 41

VIEWCTR 2RD, DXF 12

view_target

3BD, DXF 17

VIEWDIR 3BD, DXF 16

view_twist

BD, DXF 51

lens_length

BD, DXF 42

front_clip_z

BD, DXF 43

back_clip_z

BD, DXF 44

VIEWMODE 4BITS, DXF 71

render_mode

RC, DXF 281

use_default_lights

B, DXF 292

default_lightning_type

RC, DXF 282

brightness

BD, DXF 141

contrast BD, DXF 142

ambient_color

CMC, DXF 63

lower_left

2RD, DXF 10

upper_right

2RD, DXF 11

UCSFOLLOW

B, DXF 71

circle_zoom

BS, DXF 72

FASTZOOM B, DXF 73

UCSICON RC, DXF 74

GRIDMODE B, DXF 76

GRIDUNIT 2RD, DXF 15

SNAPMODE B, DXF 75
SNAPSTYLE

B, DXF 77

SNAPISOPAIR

BS, DXF 78

SNAPANG BD, DXF 50 SNAPBASE 2RD, DXF 13 SNAPUNIT 2RD, DXF 14

ucs_at_origin

B

UCSVP B, DXF 71 ucsorg 3BD, DXF 110 ucsxdir 3BD, DXF 111 ucsydir 3BD, DXF 112

ucs_elevation

BD, DXF 146

UCSORTHOVIEW

BS, DXF 79

grid_flags

BS, DXF 60

grid_major

BS, DXF 61

background

H, DXF 332

visualstyle

H, DXF 348

sun H, DXF 361

named_ucs

H, DXF 345

base_ucs H, DXF 346

VPORT CONTROL VPORT CONTROL is a table control object.

parent struct dwg object object*

num_entries

BS, DXF 70

entries H*
flags_r11

RS

VX CONTROL VX CONTROL is a table control object.

parent struct dwg object object*

num_entries

BS, DXF 70

entries H*
flags_r11

RS

VX TABLE RECORD VX TABLE RECORD is a table object.

parent struct dwg object object*

flag  RC
name  TV
used  RSd
is_xref_ref

B

is_xref_resolved

BS

is_xref_dep

B

xref  H

is_on B, DXF 290

viewport H, DXF 338
prev_entry

H, DXF 340

vport_entity_address

RS

r11_viewport_index

RSd

r11_prev_entry_index

RSd

WIPEOUTVARIABLES parent struct dwg object object*

display_frame

BS, DXF 70

XRECORD parent struct dwg object object* cloning BS, DXF 280

xdata_size

BL

num_xdata

BL

xdata Dwg Resbuf*

num_objid_handles

BL

objid_handles

H*, DXF 340

PDFDEFINITION parent struct dwg object object*

filename T, DXF 1

name T, DXF 2 DGNDEFINITION

See [UNDERLAYDEFINITION], page 218, DWFDEFINITION

See [UNDERLAYDEFINITION], page 218, ASSOCARRAYMODIFYPARAMETERS

parent struct dwg object object*

aap_version

BL

num_items

BL

classname

TV

items Dwg ASSOCARRAYITEM*

numitems BL
numrows BL
numlevels

BL

ASSOCARRAYPATHPARAMETERS

See [ASSOCARRAYPARAMETERS], page 218, ASSOCARRAYPOLARPARAMETERS

See [ASSOCARRAYPARAMETERS], page 218, ASSOCARRAYRECTANGULARPARAMETERS

See [ASSOCARRAYPARAMETERS], page 218, Dwg 3DSOLID material

parent struct dwg entity 3DSOLID*

array_index

BL

mat_absref

BL

material_handle

H

Dwg 3DSOLID silhouette parent struct dwg entity 3DSOLID*

vp_id  BL
vp_target

3BD

vp_dir_from_target

3BD

vp_up_dir

3BD

vp_perspective

B

has_wires

B

num_wires

BL

wires Dwg 3DSOLID wire* Dwg 3DSOLID wire

parent struct dwg entity 3DSOLID*

type  RC
selection_marker

BLd

color  BL
acis_index

BLd

num_points

BL

points  3BD*
transform_present

B

axis_x  3BD
axis_y  3BD
axis_z  3BD
translation

3BD

scale  3BD
has_rotation

B

has_reflection

B

has_shear

B

Dwg ACSH HistoryNode

major  BL
minor  BL

trans BD*, DXF 40

color  CMC
step_id BL
material H

Dwg ACSH SubentColor

major  BL

minor  BL
transparency

BL

bl93  BL
is_face_variable

B

Dwg ACSH SubentMaterial

major  BL
minor  BL
reflectance

BL

displacement

BL

Dwg ACTIONBODY parent struct dwg object ASSOCNETWORK*

evaluatorid

T

expression

T

value  BL

Dwg ARRAYITEMLOCATOR parent struct dwg object ASSOCARRAYMODIFYACTIONBODY*

itemloc BL

Dwg ASSOCACTIONBODY action parent struct dwg object ASSOCMLEADERACTIONBODY*

depid  BL

dep H, DXF 330 Dwg ASSOCACTION Deps

parent struct dwg object ASSOCACTION*

is_owned B
dep  H

Dwg ASSOCARRAYITEM parent struct dwg abstractobject ASSOCARRAYPARAMETERS*

class_version

BL, DXF 90

itemloc BL
flags  BL
is_default_transmatrix

int

x_dir  3BD
transmatrix

BD*

rel_transform

BD*

has_h1  int
h1  H
h2  H

Dwg ASSOCPARAMBASEDACTIONBODY parent struct dwg object object*

version BL
minor  BL
num_deps BL, DXF 90
deps  H*
l4  BL
l5  BL
assocdep H
num_values

BL

values Dwg VALUEPARAM* Dwg ASSOCSURFACEACTIONBODY

parent struct dwg object object*

version BL
is_semi_assoc

B

l2  BL
is_semi_ovr

B

grip_status

BS

assocdep H

Dwg AcDs

file_signature

RL

file_header_size

RL

unknown_1

RL

version RL
unknown_2

RL

ds_version

RL

segidx_offset

RL

segidx_unknown

RL

num_segidx

RL

schidx_segidx

RL

datidx_segidx

RL

search_segidx

RL

prvsav_segidx

RL

file_size

RL

total_segments

BL

segidx Dwg AcDs SegmentIndex* datidx Dwg AcDs DataIndex data Dwg AcDs Data* blob01 Dwg AcDs DataBlob schidx Dwg AcDs SchemaIndex schdat Dwg AcDs SchemaData search Dwg AcDs Search

segments Dwg AcDs Segment* Dwg AcDs Data

record_hdrs

Dwg AcDs Data RecordHdr*

records Dwg AcDs Data Record* Dwg AcDs DataBlob

data_size

RLL

page_count

RL

record_size

RL

page_size

RL

unknown_1

RL

unknown_2

RL

ref Dwg AcDs DataBlobRef* Dwg AcDs DataBlob01

total_data_size

RLL

page_start_offset

RLL

page_index

int32 t

page_count

int32 t

page_data_size

RLL

page_data

RC*

Dwg AcDs DataBlobRef

total_data_size

RLL

num_pages

RL

record_size

RL

page_size

RL

unknown_1

RL

unknown_2

RL

pages Dwg AcDs DataBlobRef Page* Dwg AcDs DataBlobRef Page

segidx  RL
size  RL

Dwg AcDs DataIndex

num_entries

RL

di_unknown

RL

entries Dwg AcDs DataIndex Entry* Dwg AcDs DataIndex Entry

segidx  RL
offset  RL
schidx  RL

Dwg AcDs Data Record

data_size

RL

blob  RC*

Dwg AcDs Data RecordHdr

entry_size

RL

unknown RL
handle  RLL
offset  RL

Dwg AcDs Schema

num_index

RS

index  RLL*
num_props

RS

props Dwg AcDs Schema Prop* Dwg AcDs SchemaData

num_uprops

RL

uprops Dwg AcDs SchemaData UProp*

num_schemas

RL

schemas Dwg AcDs Schema*

num_propnames

RL

propnames

TV*

Dwg AcDs SchemaData UProp

size  RL
flags  RL

Dwg AcDs SchemaIndex

num_props

RL

si_unknown_1

RL

props Dwg AcDs SchemaIndex Prop*

si_tag  RLL
num_prop_entries

RL

si_unknown_2

RL

prop_entries

Dwg AcDs SchemaIndex Prop*

Dwg AcDs SchemaIndex Prop

index  RL
segidx  RL
offset  RL

Dwg AcDs Schema Prop

flags  RL
namidx  RL
type  RL
type_size

RL

unknown_1

RL

unknown_2

RL

num_values

RS

values  RC*

Dwg AcDs Search

num_search

RL

search Dwg AcDs Search Data* Dwg AcDs Search Data

schema_namidx

RL

num_sortedidx

RL

sortedidx

RLLd*

num_ididxs

RL

unknown RL

ididxs Dwg AcDs Search IdIdxs* Dwg AcDs Search IdIdx

handle  RLL
num_ididx

RL

ididx  RLL*

Dwg AcDs Search IdIdxs

num_ididx

RL

ididx Dwg AcDs Search IdIdx* Dwg AcDs Segment

signature

RS

name  RC
type  RCd
segment_idx

RL

is_blob01

RL

segsize RL
unknown_2

RL

ds_version

RL

unknown_3

RL

data_algn_offset

RL

objdata_algn_offset

RL

padding RC

Dwg AcDs SegmentIndex

offset  RLL
size  RL

Dwg BLOCKACTION connectionpts

code  BL
name  TV

Dwg BLOCKLOOKUPACTION lut parent struct dwg object BLOCKLOOKUPACTION* conn_pts Dwg BLOCKACTION connectionpts b282 B, DXF 282 b281 B, DXF 281

Dwg BLOCKPARAMETER PropInfo

num_connections

BL

connections

Dwg BLOCKPARAMETER connection*

Dwg BLOCKPARAMETER connection

code  BL
name  T

Dwg BLOCKPARAMVALUESET

desc  TV
flags  BL
minimum BD
maximum BD
increment

BD

num_valuelist

BS

valuelist

BD*

Dwg BLOCKSTRETCHACTION codes parent struct dwg object BLOCKSTRETCHACTION* bl95 BL, DXF 95 bs76 BS, DXF 76 bl94 BL, DXF 94

Dwg BLOCKSTRETCHACTION handles parent struct dwg object BLOCKSTRETCHACTION* hdl H, DXF 331 shrt BS, DXF 74 long1 BL, DXF 94 long2 BL, DXF 94

Dwg BLOCKVISIBILITYPARAMETER state parent struct dwg object BLOCKVISIBILITYPARAMETER* name T, DXF 303

num_blocks

BL, DXF 94

blocks H*, DXF 332

num_params

BL, DXF 95

params H*, DXF 333 Dwg COMPOUNDOBJECTID

parent struct dwg object object*

has_object

B

name  T
object  H

Dwg CONSTRAINTGROUPNODE parent struct dwg object ASSOC2DCONSTRAINTGROUP*

nodeid  BL
status  RC
num_connections

BL

connections

BL*

Dwg CONTEXTDATA dict parent struct dwg CONTEXTDATA submgr* text T, DXF 3

itemhandle

H, DXF 350

Dwg CONTEXTDATA submgr parent struct dwg object CONTEXTDATAMANAGER*

handle  H
num_entries

BL, DXF 90

entries Dwg CONTEXTDATA dict* Dwg CellContentGeometry

dist_top_left

3BD, DXF 10

dist_center

3BD, DXF 11

content_width

BD, DXF 43

content_height

BD, DXF 44 width BD, DXF 45 height BD, DXF 46 unknown BL, DXF 95

cell_parent

struct dwg TableCell*

geom_parent

struct dwg TABLEGEOMETRY Cell*

Dwg CellStyle type BL, DXF 90

data_flags

BS, DXF 170

property_override_flags

BL, DXF 91

merge_flags

BL, DXF 92

bg_color CMC, DXF 62

content_layout

BL, DXF 93

content_format

Dwg ContentFormat

margin_override_flags

BS, DXF 171

vert_margin

BD, DXF 40

horiz_margin

BD, DXF 40

bottom_margin

BD, DXF 40

right_margin

BD, DXF 40

margin_horiz_spacing

BD, DXF 40

margin_vert_spacing

BD, DXF 40

num_borders

BL, DXF 94

borders Dwg GridFormat*

tablerow_parent

struct dwg TableRow*

tabledatacolumn_parent

struct dwg TableDataColumn*

Dwg ColorRamp parent struct dwg POINTCLOUDCOLORMAP Ramp*

colorscheme

T, DXF 1

unknown_bl

BL, DXF 91

unknown_b

B, DXF 290

Dwg ContentFormat

property_override_flags

BL, DXF 90

property_flags

BL, DXF 91

value_data_type

BL, DXF 92

value_unit_type

BL, DXF 93

value_format_string

T, DXF 300

rotation BD, DXF 40
block_scale

BD, DXF 140

cell_alignment

BL, DXF 94

content_color

CMC, DXF 62

text_style

H

text_height

BD, DXF 144

Dwg DATALINK customdata parent struct dwg object DATALINK*

target  H

text T, DXF 304 Dwg DATATABLE column

parent struct dwg object DATATABLE* type BL, DXF 92 text T, DXF 2 rows Dwg DATATABLE row*

Dwg DATATABLE row parent struct dwg DATATABLE column* value Dwg TABLE value

Dwg DIMASSOC Ref parent struct dwg object DIMASSOC*

classname

T, DXF 1

osnap_type

RC, DXF 72

osnap_dist

BD, DXF 40

osnap_pt 3BD, DXF 10

num_xrefs

BS

xrefs H*, DXF 331

main_subent_type

BS, DXF 73

main_gsmarker

BL, DXF 91

num_xrefpaths

BS

xrefpaths

TV*, DXF 301

has_lastpt_ref

B, DXF 75

lastpt_ref

3BD

num_intsectobj

BL, DXF 74

intsectobj

H*, DXF 332 Dwg DIMENSION common

parent struct dwg object entity*

class_version

RC, DXF 280

extrusion

BE, DXF 210

def_pt 3BD, DXF 10

text_midpt

2RD, DXF 11

elevation

BD, DXF 31

flag RC, DXF 70

flag1  RC
user_text

TV, DXF 1

text_rotation

BD, DXF 53

horiz_dir

BD, DXF 51

ins_scale

3BD

ins_rotation

BD, DXF 54

attachment

BS, DXF 71

lspace_style

BS, DXF 72

lspace_factor

BD, DXF 41

act_measurement

BD, DXF 42

unknown B, DXF 73
flip_arrow1

B, DXF 74

flip_arrow2

B, DXF 75

clone_ins_pt

2RD, DXF 12

dimstyle H, DXF 3

block H, DXF 2 Dwg EVAL Edge

parent struct dwg object EVALUATION GRAPH* id BL, DXF 92 nextid BLd, DXF 93 e1 BLd, DXF 94 e2 BLd, DXF 91 e3 BLd, DXF 91

out_edge BLd

Dwg EVAL Node parent struct dwg object EVALUATION GRAPH* id BL, DXF 91

edge_flags

BL, DXF 93

nextid BLd, DXF 95

evalexpr H, DXF 360
node  BLd
active_cycles

B

Dwg EvalExpr

parentid BLd
major  BL
minor  BL
value_code

BSd

value.num40

BD

value.pt2d

2RD

value.pt3d

3BD

value.text1

TV

value.long90

BL

value.handle91

H

value.short70

BS

nodeid  BL

Dwg EvalVariant

code  BS
u.bd  BD
u.bl  BL
u.bs  BS
u.rc  RC
u.text  TV
u.handle H

Dwg FIELD ChildValue parent struct dwg object FIELD* key TV, DXF 6 value Dwg TABLE value

Dwg FileDepList Files

filename T32
filepath T32
fingerprint

T32

version T32
feature_index

RL

timestamp

RL

filesize RL
affects_graphics

RS

refcount RL

Dwg FormattedTableData parent struct dwg object TABLECONTENT*

cellstyle

Dwg CellStyle

num_merged_cells

BL, DXF 90

merged_cells

Dwg FormattedTableMerged*

Dwg FormattedTableMerged parent struct dwg FormattedTableData* top_row BL, DXF 91

left_col BL, DXF 92
bottom_row

BL, DXF 93

right_col

BL, DXF 94

Dwg GEODATA meshface

face1  BL
face2  BL
face3  BL

Dwg GEODATA meshpt

source_pt

2RD

dest_pt 2RD

Dwg GridFormat parent struct dwg CellStyle*

index_mask

BL, DXF 95

border_overrides

BL, DXF 90

border_type

BL, DXF 91

color CMC, DXF 62 linewt BLd, DXF 92 ltype H, DXF 340

visible B, DXF 93
double_line_spacing

BD, DXF 40 Dwg HATCH Color

parent struct dwg entity HATCH*

shift_value

BD, DXF 463

color CMC, DXF 63 Dwg HATCH ControlPoint

parent struct dwg HATCH PathSeg* point 2RD, DXF 10 weight BD, DXF 40

Dwg HATCH DefLine parent struct dwg entity HATCH* angle BD, DXF 53 pt0 2BD, DXF 43 offset 2BD, DXF 45

num_dashes

BS, DXF 79

dashes  BD*

Dwg HATCH Path parent struct dwg entity HATCH* flag BL, DXF 92

num_segs_or_paths

BL, DXF 93

segs Dwg HATCH PathSeg*

bulges_present

B, DXF 72

closed B, DXF 73

polyline_paths

Dwg HATCH PolylinePath*

num_boundary_handles

BL, DXF 97

boundary_handles

H*, DXF 330

Dwg HATCH PathSeg parent struct dwg HATCH Path*

curve_type

RC, DXF 72

first_endpoint

2RD, DXF 10

second_endpoint

2RD, DXF 11

center 2RD, DXF 10 radius BD, DXF 40

start_angle

BD, DXF 50

end_angle

BD, DXF 51

is_ccw B, DXF 73 endpoint 2RD, DXF 11

minor_major_ratio

BD, DXF 40

degree BL, DXF 94

is_rational

B, DXF 73

is_periodic

B, DXF 74

num_knots

BL, DXF 95

num_control_points

BL, DXF 96

knots  BD*
control_points

Dwg HATCH ControlPoint*

num_fitpts

BL, DXF 97

fitpts  2RD*
start_tangent

2RD

end_tangent

2RD

Dwg HATCH PolylinePath parent struct dwg HATCH Path* point 2RD, DXF 10 bulge BD, DXF 42

Dwg LAYER entry parent struct dwg object LAYER INDEX*

numlayers

BL, DXF 90

name T, DXF 8 handle H, DXF 360

Dwg LEADER ArrowHead parent struct dwg entity MULTILEADER*

is_default

B, DXF 94

arrowhead

H, DXF 345

Dwg LEADER BlockLabel parent struct dwg entity MULTILEADER* attdef H, DXF 330

label_text

TV, DXF 302

ui_index BS, DXF 177 width BD, DXF 44

Dwg LEADER Break parent struct dwg LEADER Line* start 3BD, DXF 11 end 3BD, DXF 12

Dwg LEADER Line parent struct dwg LEADER Node*

num_points

BL

points 3DPOINT*

num_breaks

BL

breaks Dwg LEADER Break*

line_index

BL, DXF 91 type BS, DXF 170 color CMC, DXF 92 ltype H, DXF 340 linewt BLd, DXF 171

arrow_size

BD, DXF 40

arrow_handle

H, DXF 341

flags BL, DXF 93 Dwg LEADER Node

parent struct dwg entity MULTILEADER*

has_lastleaderlinepoint

B, DXF 290

has_dogleg

B, DXF 291

lastleaderlinepoint

3BD, DXF 10

dogleg_vector

3BD, DXF 11

branch_index

BL, DXF 90

dogleg_length

BD, DXF 40

num_lines

BL

lines Dwg LEADER Line*

num_breaks

BL

breaks Dwg LEADER Break*

attach_dir

BS, DXF 271

Dwg LIGHTLIST light parent struct dwg object LIGHTLIST* name T, DXF 1

handle  H, DXF 5

Dwg LTYPE dash parent struct dwg object LTYPE* length BD, DXF 49

complex_shapecode

BS, DXF 75

style H, DXF 340

x_offset RD, DXF 44
y_offset RD, DXF 45

scale BD, DXF 46

rotation BD, DXF 50
shape_flag

BS, DXF 74

text T, DXF 9 Dwg LWPOLYLINE width

start BD, DXF 40 end BD, DXF 41

Dwg LinkedData name T, DXF 1

description

T, DXF 300

Dwg LinkedTableData

num_cols BL, DXF 90

cols Dwg TableDataColumn*

num_rows BL, DXF 90

rows Dwg TableRow*

num_field_refs

BL

field_refs

H*

Dwg MATERIAL color parent struct dwg object object*

flag  RC
factor  BD
rgb  BL

Dwg MATERIAL gentexture parent struct dwg object MATERIAL*

genprocname

T

material struct dwg object MATERIAL* Dwg MATERIAL mapper

parent struct dwg object object*

blendfactor

BD

transmatrix

BD*

filename T

color1 Dwg MATERIAL color color2 Dwg MATERIAL color

source  RC
projection

RC

tiling  RC
autotransform

RC

texturemode

BS

Dwg MESH edge parent struct dwg entity MESH* idxfrom BL, DXF 90 idxto BL, DXF 90

Dwg MLEADER AnnotContext

num_leaders

BL

leaders Dwg LEADER Node*

attach_dir

BS

scale_factor

BD, DXF 40

content_base

3BD, DXF 10

text_height

BD, DXF 41

arrow_size

BD, DXF 140

landing_gap

BD, DXF 145

text_left

BS, DXF 174

text_right

BS, DXF 175

text_angletype

BS, DXF 176

text_alignment

BS, DXF 177

has_content_txt

B, DXF 290

has_content_blk

B, DXF 296

content Dwg MLEADER Content base 3BD, DXF 110 base_dir 3BD, DXF 111

base_vert

3BD, DXF 112

is_normal_reversed

B, DXF 297

text_top BS, DXF 273

text_bottom

BS, DXF 272

Dwg MLEADER Content Block

type  RC
normal  3BD
location 3BD
rotation BD
block_table

H

scale  3BD
color  CMC
transform

BD* Dwg MLEADER Content MText

type  RC
normal  3BD
location 3BD
rotation BD
style  H
direction

3BD

color  CMC
width  BD
height  BD
line_spacing_factor

BD

default_text

T

line_spacing_style

BS

alignment

BS

flow  BS
bg_color CMC
bg_scale BD
bg_transparency

BL

is_bg_fill

B

is_bg_mask_fill

B

col_type BS
is_height_auto

B

col_width

BD

col_gutter

BD

is_col_flow_reversed

B

num_col_sizes

BL

col_sizes

BD*

word_break

B

unknown B

Dwg MLINESTYLE line parent struct dwg object MLINESTYLE* offset BD, DXF 49 color CMC, DXF 62

lt.index BSd, DXF 6
lt.ltype H, DXF 6

Dwg MLINE line parent struct dwg MLINE vertex*

num_segparms

BS, DXF 74

segparms BD*
num_areafillparms

BS, DXF 75

areafillparms

BD*

Dwg MLINE vertex parent struct dwg entity MLINE* vertex 3BD, DXF 11

vertex_direction

3BD, DXF 12

miter_direction

3BD, DXF 13

num_lines

RC

lines Dwg MLINE line* Dwg OCD Dimension

b293 B, DXF 293 def_pt 2RD, DXF 10

is_def_textloc

B, DXF 294

text_rotation

BD, DXF 140

block H, DXF 2 dimtofl B, DXF 298 dimosxd B, DXF 291

dimatfit B, DXF 70

dimtix B, DXF 292

dimtmove B, DXF 71
override_code

RC, DXF 280

has_arrow2

B, DXF 295

flip_arrow2

B, DXF 296

flip_arrow1

B, DXF 297

Dwg PARTIAL VIEWING INDEX Entry parent struct dwg object PARTIAL VIEWING INDEX*

extents_min

3BD

extents_max

3BD

object  H

Dwg POINTCLOUDCOLORMAP Ramp parent struct dwg object POINTCLOUDCOLORMAP*

class_version

BS, DXF 70

num_ramps

BL, DXF 90

ramps Dwg ColorRamp* Dwg POINTCLOUDEX Croppings

parent struct dwg entity POINTCLOUDEX* type BS, DXF 280

is_inside

B, DXF 290

is_inverted

B, DXF 290

crop_plane

3BD, DXF 13

crop_x_dir

3BD, DXF 213

crop_y_dir

3BD, DXF 213

num_pts BL, DXF 93 pts 3BD*

Dwg POINTCLOUD Clippings parent struct dwg entity POINTCLOUD*

is_inverted

B

type  BS
num_vertices

BL

vertices 2RD*
z_min  BD
z_max  BD

Dwg POINTCLOUD IntensityStyle parent struct dwg entity POINTCLOUD*

min_intensity

BD

max_intensity

BD

intensity_low_treshold

BD

intensity_high_treshold

BD

Dwg PROXY LWPOLYLINE parent struct dwg entity PROXY ENTITY*

size  RL
flags  BS

const_width

BD

elevation

BD

thickness

BD

extrusion

BE

num_points

BL

points  2RD*
num_bulges

BL

bulges  BD*
num_widths

BL

widths Dwg LWPOLYLINE width*

unknown_1

RC

unknown_2

RC

unknown_3

RC

Dwg R2004 Header

file_ID_string

RC

header_address

RLx

header_size

RL

x04  RL
root_tree_node_gap

RLd

lowermost_left_tree_node_gap

RLd

lowermost_right_tree_node_gap

RLd

unknown_long

RL

last_section_id

RL

last_section_address

RLL

secondheader_address

RLL

numgaps RL
numsections

RL

x20  RL
x80  RL
x40  RL
section_map_id

RL

section_map_address

RLL

section_info_id

RLd

section_array_size

RLd

gap_array_size

RL

crc32  RLx
padding RC
section_type

RL

decomp_data_size

RL

comp_data_size

RL

compression_type

RL

checksum RLx

Dwg SECTION geometrysettings parent struct dwg SECTION typesettings*

num_geoms

BL, DXF 90

hexindex BL, DXF 91

flags BL, DXF 92 color CMC, DXF 62 layer T, DXF 8 ltype T, DXF 6

ltype_scale

BD, DXF 40

plotstyle

T, DXF 1

linewt BLd, DXF 370

face_transparency

BS, DXF 70

edge_transparency

BS, DXF 71

hatch_type

BS, DXF 72

hatch_pattern

T, DXF 2

hatch_angle

BD, DXF 41

hatch_spacing

BD, DXF 42

hatch_scale

BD, DXF 43

Dwg SECTION typesettings parent struct dwg object SECTION SETTINGS* type BS, DXF 90

generation

BS, DXF 91

num_sources

BL, DXF 92

sources H*, DXF 330

destblock

H, DXF 331

destfile T, DXF 1

num_geom BL, DXF 93

geom Dwg SECTION geometrysettings* Dwg SPLINE control point

parent struct dwg entity SPLINE* x BD y BD z BD w BD, DXF 41

Dwg SUNSTUDY Dates

julian_day

BL, DXF 90

msecs BL, DXF 90 Dwg SummaryInfo Property

tag  T16
value  T16

Dwg TABLEGEOMETRY Cell parent struct dwg object TABLEGEOMETRY*

geom_data_flag

BL, DXF 93

width_w_gap

BD, DXF 40

height_w_gap

BD, DXF 41

tablegeometry

H, DXF 330

num_geometry

BL, DXF 94

geometry Dwg CellContentGeometry* Dwg TABLESTYLE CellStyle

parent struct dwg object TABLESTYLE* id BL, DXF 90 type BL, DXF 91 name T, DXF 300

cellstyle.type

BL

cellstyle.data_flags

BS

cellstyle.property_override_flags

BL

cellstyle.merge_flags

BL

cellstyle.bg_color

CMC

cellstyle.content_layout

BL

cellstyle.content_format

Dwg ContentFormat

cellstyle.margin_override_flags

BS

cellstyle.vert_margin

BD

cellstyle.horiz_margin

BD

cellstyle.bottom_margin

BD

cellstyle.right_margin

BD

cellstyle.margin_horiz_spacing

BD

cellstyle.margin_vert_spacing

BD

cellstyle.num_borders

BL

cellstyle.borders

Dwg GridFormat*

cellstyle.tablerow_parent

struct dwg TableRow*

cellstyle.tabledatacolumn_parent

struct dwg TableDataColumn*

cellstyle

struct dwg CellStyle

Dwg TABLESTYLE border

linewt  BSd

visible B
color  CMC

Dwg TABLESTYLE rowstyles parent struct dwg object TABLESTYLE*

text_style

H, DXF 7

text_height

BD, DXF 140

text_alignment

BS, DXF 170

text_color

CMC, DXF 62

fill_color

CMC, DXF 63

has_bgcolor

B, DXF 283

num_borders

BL

borders Dwg TABLESTYLE border*

data_type

BL

unit_type

BL

format_string

TU

Dwg TABLE AttrDef parent struct dwg TABLE Cell* attdef H, DXF 331 index BS, DXF 179 text T, DXF 300

Dwg TABLE BreakHeight parent struct dwg entity TABLE*

position 3BD
height  BD
flag  BL

Dwg TABLE BreakRow parent struct dwg entity TABLE*

position 3BD
start  BL
end  BL

Dwg TABLE Cell parent struct dwg entity TABLE* type BS, DXF 171 flags RC, DXF 172

is_merged_value

B, DXF 173

is_autofit_flag

B, DXF 174

merged_width_flag

BL, DXF 175

merged_height_flag

BL, DXF 176

rotation BD, DXF 145

text_value

T, DXF 1

text_style

H, DXF 7

block_handle

H, DXF 340

block_scale

BD, DXF 144

additional_data_flag

B

cell_flag_override

BL, DXF 177

virtual_edge_flag

RC, DXF 178

cell_alignment

RS, DXF 170

bg_fill_none

B, DXF 283

bg_color CMC, DXF 63

content_color

CMC, DXF 64

text_height

BD, DXF 140

top_grid_color

CMC, DXF 69

top_grid_linewt

BS, DXF 279

top_visibility

BS, DXF 289

right_grid_color

CMC, DXF 65

right_grid_linewt

BS, DXF 275

right_visibility

BS, DXF 285

bottom_grid_color

CMC, DXF 66

bottom_grid_linewt

BS, DXF 276

bottom_visibility

BS, DXF 286

left_grid_color

CMC, DXF 68

left_grid_linewt

BS, DXF 278

left_visibility

BS, DXF 288

unknown BL

value Dwg TABLE value

num_attr_defs

BL

attr_defs

Dwg TABLE AttrDef*

Dwg TABLE CustomDataItem name T, DXF 300 value Dwg TABLE value

cell_parent

struct dwg TableCell*

row_parent

struct dwg TableRow*

Dwg TABLE value

flags  BL
format_flags

BL

data_type

BL, DXF 90

data_size

BL, DXF 92

data_long

BL, DXF 91

data_double

BD, DXF 140

data_string

T, DXF 1

data_date

TF

data_point

2RD, DXF 11

data_3dpoint

3RD, DXF 11

data_handle

H

unit_type

BL, DXF 94

format_string

T, DXF 300

value_string

T, DXF 302

Dwg TableCell

flag BL, DXF 90

tooltip TV, DXF 300

customdata

BL, DXF 91

num_customdata_items

BL, DXF 90

customdata_items

Dwg TABLE CustomDataItem*

has_linked_data

BL

data_link

H

num_rows BL
num_cols BL
unknown BL
num_cell_contents

BL

cell_contents

Dwg TableCellContent*

style_id BL
has_geom_data

BL

geom_data_flag

BL

width_w_gap

BD

height_w_gap

BD

tablegeometry

H

num_geometry

BL

geometry Dwg CellContentGeometry*

style_parent

struct dwg CellStyle*

row_parent

struct dwg TableRow*

Dwg TableCellContent parent struct dwg TableCell* type BL, DXF 90 value Dwg TABLE value handle H, DXF 340

num_attrs

BL, DXF 91

attrs Dwg TableCellContent Attr*

has_content_format_overrides

BS

content_format

Dwg ContentFormat

Dwg TableCellContent Attr parent struct dwg TableCellContent* attdef H, DXF 330 value TV, DXF 301 index BL, DXF 92

Dwg TableDataColumn parent struct dwg LinkedTableData* name T, DXF 300

custom_data

BL, DXF 91

cellstyle

Dwg CellStyle

cellstyle_id

BL

width  BL

Dwg TableRow parent struct dwg LinkedTableData*

num_cells

BL

cells Dwg TableCell*

custom_data

BL

num_customdata_items

BL

customdata_items

Dwg TABLE CustomDataItem*

cellstyle

Dwg CellStyle

style_id BL
height  BL

Dwg UCS orthopts parent struct dwg object UCS* type BS, DXF 71 pt 3BD, DXF 13

Dwg VALUEPARAM parent struct dwg object object*

class_version

BL

name  T
unit_type

BL

num_vars BL

vars Dwg VALUEPARAM vars*

controlled_objdep

H

Dwg VALUEPARAM vars value Dwg EvalVariant

handle  H

Dwg MLEADER Content txt Dwg MLEADER Content MText blk Dwg MLEADER Content Block

Common Entity fields

__iterator

BL

color CMC, DXF 62

color_r11

RCd, DXF 62

dwg struct dwg struct*

edge_visualstyle

H, DXF 348

eed Dwg Eed*

elevation_r11

RD, DXF 38 entmode BB, DXF 67

extra_r11

RS

face_visualstyle

H, DXF 348

flag_r11 RC
full_visualstyle

H, DXF 348

has_ds_data

B

has_edge_visualstyle

B

has_face_visualstyle

B

has_full_visualstyle

B

invisible

BS, DXF 60

is_xdic_missing

B

isbylayerlt

B

layer H, DXF 8 linewt RC, DXF 370 ltype H, DXF 6

ltype_flags

BB

ltype_scale

BD, DXF 48

material H, DXF 347
material_flags

BB

next_entity

H

nolinks B
num_eed BL
num_reactors

BL

objid  BL
opts_r11 RS
ownerhandle

H, DXF 330

plotstyle

H, DXF 390

plotstyle_flags

BB

prev_entity

H

preview TF, DXF 310

preview_exists

B

preview_is_proxy

B

preview_size

BLL, DXF 160

reactors H*, DXF 330

shadow  H
shadow_flags

RC, DXF 284

thickness_r11

RD, DXF 39

viewport H
xdicobjhandle

H, DXF 360

Common Object fields dwg struct dwg struct* eed Dwg Eed*

handleref

Dwg Handle*

has_ds_data

B

is_xdic_missing

B

num_eed BL

num_reactors

BL

objid  BL
ownerhandle

H, DXF 330

reactors H*, DXF 330

xdicobjhandle

H, DXF 360

5 Sections

The r2000 format (used for r13-r2000) knows the following 6 sections:

HEADER CLASSES HANDLES OBJFREESPACE + 2NDHEADER MEASURE-MENT/TEMPLATE AUXHEADER (only r2000)

The r2004 and r2007 format (used for r2004-r2018) knows the following sections:

R2004 Header UNKNOWN SUMMARYINFO PREVIEW VBAPROJECT APPINFO APPINFOHISTORY FILEDEPLIST ACDS REVHISTORY SECURITY OBJECTS OBJFREESPACE TEMPLATE HANDLES CLASSES AUXHEADER HEADER SIGNATURE INFO SYSTEM MAP

The old pre-R13 formats (from r1.1 to r11) have those sections, with the tables interleaved into the HEADER.

HEADER ENTITIES BLOCKS entities EXTRAS entities

With (from r1.1 to r10) those 5 Section 5.22 [Tables], page 267:

BLOCKS LAYER STYLE LTYPE VIEW

With r11 came the additional tables:

UCS VPORT APPID DIMSTYLE VX

But we convert them internally to r2004 table record and table control objects. See Section 5.22 [Tables], page 267.

5.1 HEADER Section

See Chapter 4 [Objects], page 8.

5.2 OBJECTS Section

The OBJECTS Section is usually split up into multiple pages (separate sections of type AcDbObjects) and contains all entities and objects. It is indexed by Section 5.4 [HAN-DLES], page 265.

See Chapter 4 [Objects], page 8.

5.3 CLASSES Section

The Classes Section contains the basic info for all dynamically loaded types for entities and objects. Its types start with 500, and are variable. An entity which has no class loaded is displayed as proxy.

LibreDWG contains support for many classes, but not all. See src/classes.inc and src/classes.c. We define a stability for each class, one of stable, unstable, debugging and unhandled.

Objects in stable classes are treated as the fixed-type objects with full support. Changes are treated as API breaking.

Objects in unstable classes are sometimes written to DXF or JSON, but not to DWG. Changes are not treated as API breaking. Usually such objects are converted to UN-KNOWN OBJ or UNKNOWN ENT objects, and when written to DWG converted to PLACEHOLDER, DUMMY or POINT objects with EED pointing to the original class and content. Only when rewriting from-to the very same version with the full known unknown bits blob (e.g. dwgrewrite or json) such classes can persist as such.

Objects in debugging classes are only handled with the developer configure --enable-debug flag, otherwise ignored. See unstable above.

Objects in undhandled classes are always ignored. There are no fields known, only its type.

5.4 HANDLES Section

The Handles section contains a sorted list of all object handles and its position in the Objects stream. All values are stored relatively, as offsets. Handles only increase and can contain holews when an object is deleted, offsets can jump back also.

5.5 R2004 Header

The R2004 Header section at fixed position 0x100 in the DWG contains some meta-data for r2004 sections to find the two important sections INFO and SYSTEM MAP.

5.6 UNKNOWN Section

The content of the UNKNOWN section with type 0 is unknown and does not always exist.

5.7 SummaryInfo

All Section SummaryInfo fields:

TITLE TU16, DXF 1 SUBJECT TU16, DXF 1 AUTHOR TU16, DXF 1 KEYWORDS TU16, DXF 1 COMMENTS TU16, DXF 1

LASTSAVEDBY

TU16, DXF 1

REVISIONNUMBER

TU16, DXF 1

HYPERLINKBASE

TU16, DXF 1

TDINDWG TIMERLL
TDCREATE TIMERLL
TDUPDATE TIMERLL
num_props

RS

props Dwg SummaryInfo Property*

     unknown1 RL
     unknown2 RL

See [Dwg SummaryInfo Property], page 252,

5.8 Preview

The optional Preview section contains the thumbnail stream of BMP, WMF or PNG data of the drawing. Note that blocks or proxy objects can also contain its own preview fields. The program dwgbmp can extract the preview image from this section.

5.9 VBAProject 5.10 AppInfo

Which product and version exactly created that DWG.

5.11 AppInfoHistory 5.12 FileDepList

Features and File Dependencies. Image files, fonts, xrefs, plotconfigs.

5.13 AcDS

The AcDsPrototype 1b DataStorage, used mostly for binary ACIS blobs, embedded fonts, ...

5.14 RevHistory

Revision History

5.15 Security

Password Info

5.16 ObjFreeSpace

Some Objects meta-data

5.17 Template

Contains one Measurement Header variable.

5.18 AuxHeader

In case the original Header gets lost.

5.19 Signature 5.20 INFO

The info of all used sections.

5.21 SYSTEM MAP

The map of all used sections and its chunked pages.

5.22 Tables

The old pre-R13 formats (from r1.1 to r11) have no objects and no sections, just tables, which we store in the sections indexed by the enum Dwg Section Type r11.

BLOCKS LAYER STYLE LTYPE VIEW

With r11 came the additional tables:

UCS VPORT APPID DIMSTYLE VX

Since r13 all those tables are stored as table control objects and tablerecord objects. From pre-r13 DWG’s these tables are imported as old r11 sections and as new CONTROL

objects, so that all entities are accessible via the single BLOCK CONTROL.model space -> BLOCK HEADER.entities iterator, all layers via the LAYER CONTROL.entries -> LAYER objects, and so on. All blocks are accessed via all other model space

BLOCK HEADER’s, get_first_owned_block (BLOCK_HEADER). Each CONTROL object

holds a list of all table records, i.e. entries. Each table record entry has a name and other common table fields.

5.23 EXTRAS entities section

Before R13 we had no objects, just the 5-10 tables, and the entities divided into 3 sections. The entities, the blocks and the extras. Blocks just contains the entities from each BLOCK to the ENDBLK entity.

Extras contain entities which had no room in the original section, e.g. when closing a polyline, which needs one additional byte. Thus the original entity is replaced by an undocumented JUMP entity, which gives the offset into the EXTRAS section, until a JUMP in the EXTRAS jumps back to the next original entity.

   See e.g. r10/entities.dwg

The JUMP [31] replaces the POLYLINE 2D entity which got later closed. An open POLYLINE 2D needs size 8, but closed 9 bytes. Hence they added the new closed replacement to the extras section at offset 0, index [66], and added another JUMP [67] back to the next original entity. Here it jumps back to the VERTEX 2D [32] at offset 0x84f.

       type: 18 [RCd]
       Add entity JUMP [31] Decode entity JUMP
       ===========================
       Entity number: 31, Type: 18, Addr: 847
       flag_r11: 0x0 [RC 0]
       size: 8 [RS]
       jump_address_raw: 0x80000000 [RLx 0]
       jump_entity_section: DWG_EXTRA_SECTION

jump_address: 0x0
type: 20 [RCd]
Add entity VERTEX_2D [32] Decode entity VERTEX_2D
===========================
Entity number: 32, Type: 20, Addr: 84f
flag_r11: 0x0 [RC 0]
size: 24 [RS]
layer: 0 [H(RSd) 8]
opts_r11: 0x0 [RSx 0]
point: (5.500000, 7.500000) [2RD 10]
....
extras entities: (0x10c0-0x10d1 (0), size 17)
==========================================
type: 19 [RCd]
Add entity POLYLINE_2D [66] Decode entity POLYLINE_2D
===========================
Entity number: 66, Type: 19, Addr: 10c0
flag_r11: 0x80 [RC 0]
          HAS_ATTRIBS(0x80)
size: 9 [RS]
layer: 0 [H(RSd) 8]
opts_r11: 0x1 [RSx 0]
          HAS_FLAG(0x1)
flag: 0x1 [RC 70]
      CLOSED(0x1)
type: 18 [RCd]
Add entity JUMP [67] Decode entity JUMP
===========================
Entity number: 67, Type: 18, Addr: 10c9
flag_r11: 0x0 [RC 0]
size: 8 [RS]
jump_address_raw: 0x84f [RLx 0]
jump_entity_section: DWG_ENTITY_SECTION
jump_address: 0x84f
==========================================
extras entities: end

The entity iterator knows about these jumps. The DXF structure resolves those jumps, and inserts the replaced entities.

6 Structures

6.1 EED

“Extended Entity Data” (EED) may be optionally attached to each object. They consist of a handle to the registered APPID, and a list of typed data. Each block is preceded with a size, the processing stops with size 0.

Internally libredwg stores each eed line as an array of num eed structs. If the size > 0, then new block starts with a handle, an optional raw string (when reading from a DWG), and a number of typed data entries. Only the first eed struct of each block has a size, all subsequent eed structs have size 0.

Example:

EED[0] size: 109 [BS]
EED[0] handle: 5.2.762
EED[0] code: 70 [RC] short: 2 [RS]
EED[1] code: 70 [RC] short: 0 [RS]
EED[2] code: 70 [RC] short: 0 [RS]
EED[3] code: 11 [RC] 3dpoint: (0.000000, 0.000000, 0.000000) [3RD]
EED[4] code: 11 [RC] 3dpoint: (1.000000, 0.000000, 0.000000) [3RD]
EED[5] code: 11 [RC] 3dpoint: (0.000000, 1.000000, 0.000000) [3RD]
EED[6] code: 11 [RC] 3dpoint: (0.000000, 0.000000, 1.000000) [3RD]
EED[7] size: 6 [BS]
EED[7] handle: 5.2.763
EED[7] code: 70 [RC] short: 0 [RS]
EED[8] code: 70 [RC] short: 0 [RS]
EED[9] size: 23 [BS]
EED[9] handle: 5.1.12
EED[9] code: 0 [RC] string: "RTMaterial" len=10 cp=30
EED[10] code: 5 [RC] entity: 0x6507000000000000 [RLL]
- size: 0 [BS]

These 10 num eed structs consist of 3 blocks with 3 size and handle entries. EED[0] starts with size 109, the handle pointing to object 762, 3 shorts and 4 points. The next block at EED[7] has size 6, the handle pointing to object 763 and 2 shorts. The last block at EED[9] has size 9, the handle pointing to object 12 (the APPID.ACAD application) and a string and an entity reference. The size is calculated by the needed room for all data code + values, without the handle. E.g. EED[7] size: 6 is 1 + 2 for EED[7] RC + RS, and 1 + 2 for EED[8] RC + RS.

Each data block consists of a RC code, and a variable value. A string may be a an old pre-r2007 ASCII string with a RC length (max 255 chars), a codepage and the string. Or a r2007+ wide string with a RS length (max 32767 chars) and a UCS-2 wide string.

decode stores both, the raw data, and the structured data. in dxf just the data. encode prefers raw over the data.

6.2 XDATA

XRECORD XDATA are very similar to the EED array, but internally it is a single linkedlist, consisting of something like the EED data code + value pairs. There’s only one size, xdata size, and only one handle to the APPID, which handles this XRECORD XDATA.

7 Functions

You can use LibreDWG immediately upon loading, without any particular initialization. Only when using some see Section 7.4 [dynapi], page 274, functions you might need to initialize the version via dwg_api_init_version(&dwg), when you need other formats than r2000 and you call an API function which does not store the version internally. Most do. This limitation will soon be fixed.

You usually use one set of functions - either decoding or encoding - at a time. All functions use the common data types (see Chapter 3 [Types], page 6). All functions return an error code, and the high-level functions for multiple objects add the error bitmask, which is sorted by severity. When the error exceeds DWG ERR CRITICAL, processing is stopped.

The new see Section 7.4 [dynapi], page 274, has dynamic get and set functions for all objects and its fields. You can get and set a property value from any object pointer by the object name and the field name.

7.1 Decoding

The highest level function for decoding a file is dwg_read_file.

[Function] int dwg_read_file (char *filename, Dwg Data *dwg)

Open filename and decode it, saving information into dwg. Return 0 if successful.

You can then iterate over the entities in model space or paper space via two ways:

1. by using the dwg.h data structures. Via dwg->object[0], which is of type Dwg_

Object_BLOCK_CONTROL, and a custom void process_BLOCK_HEADER(Dwg_Object_Ref*
ref):
  Dwg_Object_BLOCK_CONTROL* block_control = dwg->block_control;
  // first all entities in the model space
  process_BLOCK_HEADER(dwg->header_vars.BLOCK_RECORD_MSPACE);
  // then all entities in the blocks
  for (i=0; i < block_control->num_entries; i++)
    {
      process_BLOCK_HEADER(block_control->block_headers[i]);
    }
  // and last all entities in the paper space
  process_BLOCK_HEADER(dwg->header_vars.BLOCK_RECORD_PSPACE);

or 2. by using the API functions from dwg_api.h:

  Dwg_Object_BLOCK_CONTROL* block_control = dwg_block_control(dwg);
  process_BLOCK_HEADER(dwg_model_space_ref(dwg));
  for (i=0; i < block_control->num_entries; i++)
    {
      process_BLOCK_HEADER(block_control->block_headers[i]);
    }
  process_BLOCK_HEADER(dwg_paper_space_ref(dwg));

and inside the process_BLOCK_HEADER function, you iterate over the entities from the block header via:

  Dwg_Object* obj = get_first_owned_entity(ref->obj);
  while (obj)
    {
      process_object(obj);
      obj = get_next_owned_entity(ref->obj, obj);
    }

where process_object checks the type of each entity under the Dwg Object* obj. For each entity or object type (i.e. a non-graphical dwg object, also tables) there also

exist the simple and expensive dwg_getall_ENTITY and dwg_getall_OBJECT functions:

[Function] int dwg_getall_ENTITY (Dwg Object Ref *block_header_ref)

Return a malloc’ed NULL-terminated array of all such entities for Model Space, Paper Space or an individual block.

[Function] int dwg_getall_OBJECT (Dwg Data *dwg)

Return a malloc’ed NULL-terminated array of all such DWG objects.

The decoder is driven by the fields definition in the src/dwg.spec, which adds each field

to the object. This is done in the src/decode.c or src/decode_r2007.c.

[Function] int dwg_decode_OBJECT (Bit Chain *dat, Dwg Object *obj)

Sets the fields for the object from the DWG bitstream.

Note: Pre-R13 DWG’s do contain all deleted entities, which e.g. where moved into a BLOCK. Those entities do have a type > 127. You need to filter them out by yourself, when processing the DWG.

7.2 Encoding

Encoding DWG files, i.e. DWG write support, can be disabled via ./configure --disable-write. The default formats and only useful ones are currently r1.1 - r2000. Experimentally work is ongoing for the r2004 format, which is also used for r2010, r2013, and r2018. The r2007 format version is not covered yet. The pre-r13 formats are much simpler and can be written, but need some hand-holding and manual conversions when converting from newer formats still.

See src/in_dxf.c for a high-level usage example. The default codepage is Latin-1, 30. The highest level function for encoding a bitstream to a file is dwg_write_file, which

dumps the dwg to a file.

[Function] int dwg_write_file (char *filename, Dwg Data *dwg)

Open filename and write the dwg to it. Return 0 if successful.

See Section 7.3 [add api], page 273, for:

[Function] Dwg_Data* dwg_add_Document (const Dwg Version Type version,

const int imperial, const int loglevel) Creates an initial template dwg structure in memory, suitable to be written to a DWG or DXF file, without any additional table records or entities. Creates ModelSpace, PaperSpace and most Tables and basic Dictionaries. and how to add entities and objects from scratch. Low level-functions:

[Function] int dwg_add_object (Dwg Data *dwg)

Adds a new uninitialized object to the dwg->object[] array. Return 0 or -1 if successful, otherwise DWG ERR OUTOFMEM. -1 is the array was re-allocated.

Then for each object or entity type there is a

[Function] int dwg_setup_<OBJECT> (Dwg Object *obj)

Initializes an object for the given OBJECT or ENTITY type, with all fields being zero’ed. This does not initialize the obj size, type, address, handlestream size, bitsize fields.

The encoder is driven by the fields definition in the src/dwg.spec and the generated src/dynapi.c, which adds each field to the object. This is done by src/encode.c or any src/in_*.c import module.

[Function] int dwg_encode_<OBJECT> (Bit Chain *dat, Dwg Object *obj)

Encodes the DWG bitstream from the fields of the object.

The iterator is similar to above, but you want to encode all data structures, not just the entities. But note that you need many helper functions, such as the Section 7.4 [dynapi], page 274, to create all needed sections to store a DWG if you didn’t read a DWG into the right a Dwg Data* dwg struct already. This is especially important when importing from DXF or from an earlier or later DWG version.

7.3 add api

The add api functions are useful for CAD programs which want to write DWG. All the other API’s are mostly to convert from and to DWG, so the main structures and links already do exist. With the add api you can easily create an empty DWG from scratch, add table entries (into fixed Tables or variables Dictionaries), and add entities. To set more entity fields use the Section 7.4 [dynapi], page 274.

For each almost each entity and table exists a function at to add it, with arguments to initialize some fields as in the VBA object model. The other objects are either created automatically, or handled separately.

All BITCODE T strings are encoded as UTF-8, as with the dynapi. See Section 7.5 [strings], page 276. Most names are copied, since most names are considered to be constant. If not, you need to free them by yourself. Exceptions are dxfname (there exists a separate dxfname u variant), the VX name, which does not exists anymore since r2000.

A very simple example using the add API is the example program See [dwgadd], page 281.

[Function] Dwg_Data dwg_add_Document (const Dwg Version Type version,

const int imperial, const int loglevel))

Creates an initial template dwg structure in memory, suitable to be written to a DWG or DXF file, without any additional table records or entities. Creates ModelSpace, PaperSpace and most Tables and basic Dictionaries. When writing DWG, a version of R 2000 is recommended, only R 1 2 - R 2000 are supported yet. For DXF you can try all versions R 13 - R 2018.

For each OBJECT and ENTITY type there exists a specific dwg_add_<OBJECT> function, which takes the owner and some default arguments. Entities are normally added to a block header, like modelspace, paperspace or any block. Objects are normally added to the dwg, or to some other object or entity. E.g.

[Function] Dwg_Entity_LINE *line = dwg add LINE

(Dwg Object BLOCK HEADER *modelspace,

dwg point 3d *start pt, dwg point 3d *end pt)

[Function] Dwg_Entity_TEXT* dwg_add_TEXT (Dwg Object BLOCK HEADER *restrict blkhdr, const char* restrict text_value, const dwg point 3d *restrict ins_pt, const double height)

Adds a TEXT entity to the ModelSpace, PaperSpace or a Block. Entity specific arguments are here the text, the point (as pointer to the struct of 3 doubles), and the text height.

[Function] Dwg_Object_LAYER *layer = dwg add LAYER (Dwg Data *dwg,

const char *name)

Adds a new layer the Layer Table, i.e. creates the new LAYER object, and adds it to LAYER CONTROL object, the list of layers.

Names and strings are encoded as UTF-8 and will be translated to type BITCODE T (i.e. versions specific TU or TV types, either UCS-2 unicode or single-byte codepage) internally, as with the dynapi. Only internally you will have to deal with 2 different DWG text representations: UCS-2 since r2007, single-byte before. see Section 7.5 [strings], page 276.

To understand the object model for the add API see some VBA Object model documen-

tation, such as e.g. http://entercad.ru/acadauto.en/.

The new add API mostly handles the direct Dwg_Entity_ENTITY structs, not all the generic Dwg_Object structs. Thus you can access the object specific fields directly, the common fields, not so easily.

The DWG Document consists of 3 basic entity containers ModelSpace, PaperSpace and

Blocks, plus Tables (Layers, Linetypes, ...), Dictionaries as generic replacements of

Tables with a root Dictionary, the NOD ("Named Object Dictionary"), and more support objects and complex entity groups.

Helper functions:

[Function] dwg_add_u8_input (Dwg_Data *restrict dwg, const char *restrict

        u8str)

Convert UTF-8 strings to BITCODE T fields. Returns a copy of the string. All external API’s only deal with UTF-8 strings.

7.4 dynapi

The new dynapi replaced the old dwg api functions to access each object field. The old dwg api functions were deprecated, and need to be re-enabled by defining CFLAGS="-DUSE_ DEPRECATED_API". See see Chapter 4 [Objects], page 8, for an description of each object and its fields..

For each of header, entity, common or subclass there is a function to get and set the value of any type, or converted utf8 string.

[Function] bool dwg_dynapi_entity_value (void *entity, const char

*dxfname, const char *fieldname, void *out, Dwg DYNAPI field *fp)

entity is of type dwg_ent_generic, that is the pointer to the object specific struct. dxfname is the dxfname of the object, fieldname is the field or property name of the field to be read from, *out the result pointer and the optiona *fp is filled by the information for this field.

[Function] bool dwg_dynapi_common_value (void *entity, const char

*fieldname, void *out, Dwg DYNAPI field *fp)

This accesses the common Dwg_Object_Object* or Dwg_Object_Entity* fields.
[Function] bool dwg_dynapi_header_value (void *dwg, const char *fieldname,

void *out, Dwg DYNAPI field *fp)

This accesses the Header (or sometimes also called Database) fields.

[Function] bool dwg_dynapi_subclass_value (void *ptr, const char

*subclass, const char *fieldname, void *out, Dwg DYNAPI field *fp)

This accesses a subclass, a structure within the object.

The utf8text functions convert version-specific text strings to UTF-8 strings. Internally the dwg stores strings as TU (unicode) or TV (single-byte codepage). The API treats all strings as UTF-8, as with JSON, DXF or the add API.

[Function] bool dwg_dynapi_entity_utf8text (void *entity, const char
   *dxfname, const char *fieldname, char *textp, int *isnewp,

Dwg DYNAPI field *fp)

isnewp is set to 1 f the string is a fresh copy, for unicode strings.

[Function] bool dwg_dynapi_common_utf8text (void *entity, const char

*fieldname, char *textp, int *isnewp, Dwg DYNAPI field *fp)

[Function] bool dwg_dynapi_header_utf8text (void *dwg, const char

*fieldname, char *textp, int *isnewp, Dwg DYNAPI field *fp)

This accesses the Header (or sometimes also called Database) fields.

[Function] bool dwg_dynapi_subclass_utf8text (void *ptr, const char
   *subclass, const char *fieldname, char *textp, int *isnewp,

Dwg DYNAPI field *fp)

This accesses a subclass, a structure within the object.

The setters don’t differentiate between common values and strings.

[Function] bool dwg_dynapi_entity_set_value (dwg ent generic *_obj, const

char *fieldname, const void *value, const bool is_utf8)

Sets the ENTITY.fieldname to a value. A malloc’ed struct is passed by ptr, not by the content. A non-malloc’ed struct is set by content. Arrays or strings must be malloced before. We just set the new pointer, the old value will be freed. If is utf8 is set, the given value is a UTF-8 string, and will be converted to TV or TU

[Function] bool dwg_dynapi_header_set_value (Dwg Data *dwg, const char

*fieldname, const void *value, const bool is_utf8)

[Function] bool dwg_dynapi_common_set_value (dwg ent generic *_obj, const

char *fieldname, const void *value, const bool is_utf8)

See the sourcecode of the importers or programs for the usage of the API’s.

7.5 strings

Internally the DWG consists of multiple different string formats, see Chapter 3 [Types], page 6. The most important are BITCODE_TV (i.e. char*) encoded according to

dwg->header.codepage, and BITCODE_TU (i.e. wchar_t on Windows, UCS-2).

Externally most functions get and set strings as UTF-8, as in DXF or JSON.

Before r2007 DWG’s the TV and T strings are encoded in its codepage, and converted from and to their proper codepage to UTF-8 or \U+XXXX.

To encode unicode characters special \U+XXXX sequences are used, and pre-r2007 DXF MIF \M+nXXXX sequences, where n is one of the asian wide-character codepages 932 (Japanese), 950 (trad. Chinese), 949 (Korean Wansung), 1361 (Johab), and 936 (simplified Chinese).

On DWG’s r2007 and later most strings (T and TU) are encoded in the Microsoft specific two-byte UCS-2 Unicode encoding, without proper support for surrogate pairs and the upper planes (i.e. emojis).

Fixed TF strings are not encoded and have a length stored also. Normal strings are all zero-delimited. EED and XDATA strings do have a length though, but have length limitations.

Strings in DXF and JSON also have quoting rules for special characters, like \r, \n, \" and so on.

Transformations:

DWG to DWG: decode reads the T and TU strings in its natural format into the field. encode translate it to TV or TU. encode needs header.from_version and how it was read, from DWG or from an importer (in dxf or in json) or the Section 7.3 [add api], page 273, (DWG OPTS IN).

DXF/JSON to DWG: in dxf/json keeps the T and TU strings as TV. encode to <r2007 keeps it as TV, r2007+ translates it to TU. Unicode is encoded as \U+XXXX. It sets DWG OPTS IN.

DWG to DXF/JSON: decode keeps the T and TU strings as TV or TU. out dxf/json translate them to TV or UTF-8 and quotes them via \U+XXXX.

add api to DWG/DXF: add reads strings as UTF-8, and encodes it from UTF-8 to TV or TU. (TU not yet, as we don’t encode r2004+ yet). add sets DWG OPTS IN.

7.6 Other Formats

7.6.1 DXF

We can write ASCII DXF files in various versions, with much more data than other free DWG libraries, but not as stable as the unfree Teigha library yet. AutoCAD ◯^{R} fails to import some of our files still (~10% failure rate).

Options: --minimal (see dwg2dxf or dwgread) creates only a short header with a ACAD-VER and HANDSEED element, and the entities, without any subclass markers, reactors or handles.

Support for the different r1.x DXF format is not planned.

Reading DXF works for most objects. Converting a minimal DXF to DWG may fail, and needs some fixes.

7.6.2 DXFB

We can almost write Binary DXF files in various versions.

Reading DXFB is under construction and un-tested.

7.6.3 JSON

We write to and read back from our own JSON format, which is a readable 1:1 mapping of the DWG structures, and carries much more information than the DXF format. The idea is to dump a DWG to JSON and filter/query or postprocess it with more powerful JSON query tools such as jq (https://stedolan.github.io/jq/), and optionally import it back in. See Chapter 9 [Programs: dwgfilter], page 279. JSON is much better structured than DXF.

The current first level objects are all the section names, like “HEADER”, “CLASSES”, “OBJECTS”. For more see the specs.

Note that for the versions before R13 we convert all tables to table control and record objects internally, so that we can use our iterators needed for DXF support. This does not reflect the internal DWG structure. See Section 5.22 [Tables], page 267.

7.6.4 GeoJSON

dwgread supports writing to the GeoJSON format as specified at http://geojson.org/

geojson-spec.html. See dwgread with the --fmt GeoJSON option.

We write in the RFC7946 format, the new GeoJSON format since 2016, which means smaller, less precision, and normalized polygons with proper right-hande rule orientation.

We write all coordinates as [x, y], not [y, x]. z-coordinates are optional, and only written if not 0.0. Colors are either written as palette index as integer if not 256 (ByLayer), or as TrueColor RGB hex string values for all r2004+ DWG’s.

Missing entities: No 3D entities, HATCH by definition. ELLIPSE, polyline bulges would need segmention into line segments. MLINE, SPLINE, MINSERT, SOLID, TRACE, RAY(?), XLINE(?)

Due to implementation quirks with ending commas in JSON, we mostly add an empty dummy feature at the very end, with null properties and null geometry.

8 Errors

LibreDWG is mostly a library, and as such collects error codes from the highest level function down to the lowest level functions. The error codes are sorted by severity, and only if the error exceeds DWG ERR CRITICAL, i.e. DWG_ERR_CLASSESNOTFOUND, processing is stopped.

All error bitmasks are collected during read or write and returned at the end.

DWG_ERR_WRONGCRC

1

DWG_ERR_NOTYETSUPPORTED

2

DWG_ERR_UNHANDLEDCLASS

4

DWG_ERR_INVALIDTYPE

8

DWG_ERR_INVALIDHANDLE

16

DWG_ERR_INVALIDEED

32

DWG_ERR_VALUEOUTOFBOUNDS

64

DWG_ERR_CLASSESNOTFOUND
          128 = DWG_ERR_CRITICAL
DWG_ERR_SECTIONNOTFOUND

256

DWG_ERR_PAGENOTFOUND

512

DWG_ERR_INTERNALERROR

1024

DWG_ERR_INVALIDDWG

2048

DWG_ERR_IOERROR

4096

DWG_ERR_OUTOFMEM

8192

Additionally, verbose warning and error messages are printed to stderr.

Unhandled class and Invalid type errors of objects are not severe. A DWG format can store a serialization of many third party classes and objects, and thus we will never be able read all possible types. Unknown types are just stored as binary blob without any DXF codes.

9 Programs

LibreDWG installs some binary programs to read or write DWG files.

dwgread

This reads a DWG file, and optionally converts its content to some output formats: JSON, Minimal JSON, GeoJSON, DXF, DXFB (i.e. Binary DXF), SVG.

dwgread [OPTION]... DWGFILE

Options: -v[0-9], –verbose [0-9] verbosity -O fmt, –format fmt fmt: JSON, DXF, DXFB, minJSON, GeoJSON. More planned formats: YAML, XML/OGR, GPX, SVG, PS. -o outfile, –file outfile also defines the output fmt. Default: stdout –help display this help and exit –version output version information and exit minJSON is minimized JSON, without whitespace.

dwgwrite

Create a DWG from a given input file (see Section 7.6.1 [DXF], page 276, see Section 7.6.2 [DXFB], page 277, see Section 7.6.3 [JSON], page 277), optionally via --as=rNNNN as another version. For now can only create r1.2-r2000 DWG files.

dxfwrite

Create a DXF from a given input file (DWG, see Section 7.6.1 [DXF], page 276, see Section 7.6.2 [DXFB], page 277, see Section 7.6.3 [JSON], page 277), optionally via --as=rNNNN as another version. Experimental. Supports the same options

as dwg2dxf.
dwg2dxf

Converts DWG files to DXF, optionally via --as=rNNNN as another version, an earlier or later version, or via -m or --minimal as a minimal DXF version, skipping most headers vars, classes, tables and objects. --binary as a binary DXF file, with full precision, under construction. The DXF files are created in the current directory and not overwritten, unless the option --overwrite or -y is given.

dxf2dwg

Converts DXF (or Binary DXF) files to DWG, optionally via --as=rVER as another version, an earlier or later version. The DWG files are created in the current directory and not overwritten, unless the option --overwrite or -y is given. This program is experimental and AutoCAD ◯^{R} may fail to import it. For now can only create r1.2-r2000 DWG. The default is writing as r2000.

dwgrewrite

Read and write the DWG, optionally via --as=rNNNN as another version, an earlier or later version. The default is writing as r2000. For now can only create r1.2-r2000 DWG.

dwglayers

Prints all layers in a DWG. With -x or --extnames prints the extended displayed layer name with spaces, not the internally stored old-style name with _ instead. Only relevant with old r13 and r14 DWGs, after that layers are always stored in the extended format. With -f or --flags also the status of frown, on/off and locked. With --on only the visible layers, which are on and not frozen. You can get the same effect via this json filter:

  dwgfilter '.OBJECTS[]' example.dwg | \
    grep -A22 '"object": "LAYER"' | grep name
dwggrep

Search regex pattern in all text values in a list of DWGs. dwggrep uses PCRE. With -i searches case-insensitive. With -c prints only the count of found texts. With -h or --no-filename does not print the filename. With --type NAME search only NAME entities or objects. With --dxf NUM search only in DXF group NUM fields. With --text searches only TEXT-like entities: TEXT, MTEXT, ATTRIB, ATTDEF.

dwgfilter

Search and modify a single DWG file via jq, using the powerful JQ query expression language on a temporary json file. See man jq. With -i replaces the DWG in-place. This only makes sense with an JQ expression which changes values.

dwg2SVG

Convert a DWG to a limited SVG. All paperspace or modelspace enties of type: TEXT, LINE, CIRCLE, ARC, POLYLINE 2D, LWPOLYLINE, INSERT, EL-LIPSE (unrotated), SOLID, 3DFACE, RAY, XLINE. With -m or --mspace all paper-space entities are ignored, and only model-space is printed. The default is to print all paper-space entities. But if there are none, print all model-space entities instead. Limitations: Many other graphical entities and some properties are still missing.

dwg2ps

Convert a DWG to a very limited Postscript file. All paperspace and modelspace entities of type LINE, POLYLINE 2D, LWPOLYLINE, ARC and CIRCLE.

This requires installation of pslib http://pslib.sourceforge.net/doc/
pslib.html.

Note that the graphical representation for PS and SVG output is severely lacking, block references (insert entities) are not yet exploded, UCS and paper space transformations per entity are not yet done.

Planned is dwgplot, via GNU Plotutils https://www.gnu.org/software/plotutils/,

to replace dwg2SVG and dwg2ps. This supports much more bitmap and vector formats. There are also some more examples in the source distribution:

load_dwg loads a DWG and adds some entities.

dwg2svg2 converts a DWG to SVG similar to dwg2SVG, but via the dwg_api.h only. The graphical representation for PS and SVG output is severely lacking, block references (insert entities) are not yet exploded , UCS and paper space transformations per entity are not yet done.

unknown lists the not yet reverse-engineered blobs from our examples files, and is the framework to guess the field layout for these. It is optionally using picat (http://picat-lang.org/) to solve some of the field-packing problems.

dwgfuzz afl++ fuzzing frontend, to test and debug various fast shared-memory options for afl-clang-fast, with the following runtime options: -indxf, -injson, -rw, -dwg, -dxf, -dxfb, -json, -geojson. All other output formats, like BMP, SVG, PS need to be fuzzed via their programs, which is the recommended way. The now default and fastest method INMEM does not need the 2nd file argument @@, the 2nd method STDIN neither.

See also https://github.com/LibreDWG/libredwg-fuzz for our fuzzing setup

to test new fuzzing campaigns automatically and find regressions.

dwgadd is the easiest way to create DWG’s (or DXF, JSON) from scratch or add entities to an existing DWG. It accepts a very simple file with commands to create entities or objects and set its properties. See man 1 dwgadd and man 5 dwgadd.

10 Bindings

LibreDWG generates library bindings to python and perl5 via swig. These can be quite huge, and it is recommended to use ccache. You can easily add bindings to other swigsupported languages, like Go, C#, ruby, php, D, lua, tcl, common lisp, ocaml, or others by yourself. Patches accepted.

Bindings for gambas (which looks very close to VBA) are at GitHub (https://github. com/LibreDWG/gambas3-bindings) and will soon be added to gambas3 as gb.dwg component. This is in development and about 80% finished.

11 Reference API

See the separate refman (https://www.gnu.org/software/libredwg/refman/) manual

(in pdf or html format, the pdf has ~1800 pages) for a detailed API description, or see the relevant dwg.h, dwg_api.h or the *.spec files.

For reference you might also want to check the public DXF reference manuals, the VBA

object model and the ODA OpenDesign_Specification_for_dwg_files.pdf.

12 Reporting bugs

To report bugs or suggest enhancements for GNU LibreDWG, please “submit a bug”

at Savannah (http://savannah.gnu.org/projects/libredwg) or send electronic mail to

libredwg@gnu.org. (If you use the web interface, you don’t need to also send email, since that is done automatically.) Issues and pull requests at the github mirror (https://github.

com/LibreDWG/libredwg) are also accepted.

For bug reports, please include enough information for the maintainers to reproduce the problem. Generally speaking, that means:

• The version numbers of LibreDWG and any other program(s) or manual(s) involved.

• Hardware and operating system names and versions.

• The contents of any input files necessary to reproduce the bug.

• The expected behavior and/or output.

• A description of the problem and samples of any erroneous output.

• Options you gave to configure other than specifying installation directories.

• Anything else that you think would be helpful. Usually that’s the failing part of the object processed with dwgread -v5, but only the failing part, not the whole output.

When in doubt whether something is needed or not, include it. It’s better to include too much than to leave out something important.

Patches are welcome; if possible, please make them with ‘git format-patch’ and include ChangeLog entries (see Section “Change Log” in The GNU Emacs Manual). Please follow the existing GNU coding conventions. For patches longer than 15 lines we need your copyright assignment to the FSF clerk. See CONTRIBUTING in the source distribution.

Appendix A GNU Free Documentation License

Version 1.3, 3 November 2008

Copyright \begin{matrix}c \\ ◯ \end{matrix} 2000, 2001, 2002, 2007, 2008 Free Software Foundation, Inc.

 http://fsf.org/

Everyone is permitted to copy and distribute verbatim copies of this license document, but changing it is not allowed.

0. PREAMBLE The purpose of this License is to make a manual, textbook, or other functional and useful document free in the sense of freedom: to assure everyone the effective freedom to copy and redistribute it, with or without modifying it, either commercially or noncommercially. Secondarily, this License preserves for the author and publisher a way to get credit for their work, while not being considered responsible for modifications made by others. This License is a kind of “copyleft”, which means that derivative works of the document must themselves be free in the same sense. It complements the GNU General Public License, which is a copyleft license designed for free software. We have designed this License in order to use it for manuals for free software, because free software needs free documentation: a free program should come with manuals providing the same freedoms that the software does. But this License is not limited to software manuals; it can be used for any textual work, regardless of subject matter or whether it is published as a printed book. We recommend this License principally for works whose purpose is instruction or reference.

1. APPLICABILITY AND DEFINITIONS This License applies to any manual or other work, in any medium, that contains a notice placed by the copyright holder saying it can be distributed under the terms of this License. Such a notice grants a world-wide, royalty-free license, unlimited in duration, to use that work under the conditions stated herein. The “Document”, below, refers to any such manual or work. Any member of the public is a licensee, and is addressed as “you”. You accept the license if you copy, modify or distribute the work in a way requiring permission under copyright law. A “Modified Version” of the Document means any work containing the Document or a portion of it, either copied verbatim, or with modifications and/or translated into another language. A “Secondary Section” is a named appendix or a front-matter section of the Document that deals exclusively with the relationship of the publishers or authors of the Document to the Document’s overall subject (or to related matters) and contains nothing that could fall directly within that overall subject. (Thus, if the Document is in part a textbook of mathematics, a Secondary Section may not explain any mathematics.) The relationship could be a matter of historical connection with the subject or with related matters, or of legal, commercial, philosophical, ethical or political position regarding them. The “Invariant Sections” are certain Secondary Sections whose titles are designated, as being those of Invariant Sections, in the notice that says that the Document is released under this License. If a section does not fit the above definition of Secondary then it is not allowed to be designated as Invariant. The Document may contain zero Invariant Sections. If the Document does not identify any Invariant Sections then there are none.

The “Cover Texts” are certain short passages of text that are listed, as Front-Cover Texts or Back-Cover Texts, in the notice that says that the Document is released under this License. A Front-Cover Text may be at most 5 words, and a Back-Cover Text may be at most 25 words.

A “Transparent” copy of the Document means a machine-readable copy, represented in a format whose specification is available to the general public, that is suitable for revising the document straightforwardly with generic text editors or (for images composed of pixels) generic paint programs or (for drawings) some widely available drawing editor, and that is suitable for input to text formatters or for automatic translation to a variety of formats suitable for input to text formatters. A copy made in an otherwise Transparent file format whose markup, or absence of markup, has been arranged to thwart or discourage subsequent modification by readers is not Transparent. An image format is not Transparent if used for any substantial amount of text. A copy that is not “Transparent” is called “Opaque”.

Examples of suitable formats for Transparent copies include plain ascii without markup, Texinfo input format, LaTEX input format, SGML or XML using a publicly available DTD, and standard-conforming simple HTML, PostScript or PDF designed for human modification. Examples of transparent image formats include PNG, XCF and JPG. Opaque formats include proprietary formats that can be read and edited only by proprietary word processors, SGML or XML for which the DTD and/or processing tools are not generally available, and the machine-generated HTML, PostScript or PDF produced by some word processors for output purposes only.

The “Title Page” means, for a printed book, the title page itself, plus such following pages as are needed to hold, legibly, the material this License requires to appear in the title page. For works in formats which do not have any title page as such, “Title Page” means the text near the most prominent appearance of the work’s title, preceding the beginning of the body of the text.

The “publisher” means any person or entity that distributes copies of the Document to the public.

A section “Entitled XYZ” means a named subunit of the Document whose title either is precisely XYZ or contains XYZ in parentheses following text that translates XYZ in another language. (Here XYZ stands for a specific section name mentioned below, such as “Acknowledgements”, “Dedications”, “Endorsements”, or “History”.) To “Preserve the Title” of such a section when you modify the Document means that it remains a section “Entitled XYZ” according to this definition.

The Document may include Warranty Disclaimers next to the notice which states that this License applies to the Document. These Warranty Disclaimers are considered to be included by reference in this License, but only as regards disclaiming warranties: any other implication that these Warranty Disclaimers may have is void and has no effect on the meaning of this License.

2. VERBATIM COPYING You may copy and distribute the Document in any medium, either commercially or noncommercially, provided that this License, the copyright notices, and the license notice saying this License applies to the Document are reproduced in all copies, and that you add no other conditions whatsoever to those of this License. You may not use technical measures to obstruct or control the reading or further copying of the copies you make or distribute. However, you may accept compensation in exchange for copies. If you distribute a large enough number of copies you must also follow the conditions in section 3. You may also lend copies, under the same conditions stated above, and you may publicly display copies.

3. COPYING IN QUANTITY If you publish printed copies (or copies in media that commonly have printed covers) of the Document, numbering more than 100, and the Document’s license notice requires Cover Texts, you must enclose the copies in covers that carry, clearly and legibly, all these Cover Texts: Front-Cover Texts on the front cover, and Back-Cover Texts on the back cover. Both covers must also clearly and legibly identify you as the publisher of these copies. The front cover must present the full title with all words of the title equally prominent and visible. You may add other material on the covers in addition. Copying with changes limited to the covers, as long as they preserve the title of the Document and satisfy these conditions, can be treated as verbatim copying in other respects. If the required texts for either cover are too voluminous to fit legibly, you should put the first ones listed (as many as fit reasonably) on the actual cover, and continue the rest onto adjacent pages. If you publish or distribute Opaque copies of the Document numbering more than 100, you must either include a machine-readable Transparent copy along with each Opaque copy, or state in or with each Opaque copy a computer-network location from which the general network-using public has access to download using public-standard network protocols a complete Transparent copy of the Document, free of added material. If you use the latter option, you must take reasonably prudent steps, when you begin distribution of Opaque copies in quantity, to ensure that this Transparent copy will remain thus accessible at the stated location until at least one year after the last time you distribute an Opaque copy (directly or through your agents or retailers) of that edition to the public. It is requested, but not required, that you contact the authors of the Document well before redistributing any large number of copies, to give them a chance to provide you with an updated version of the Document.

4. MODIFICATIONS You may copy and distribute a Modified Version of the Document under the conditions of sections 2 and 3 above, provided that you release the Modified Version under precisely this License, with the Modified Version filling the role of the Document, thus licensing distribution and modification of the Modified Version to whoever possesses a copy of it. In addition, you must do these things in the Modified Version: A. Use in the Title Page (and on the covers, if any) a title distinct from that of the

Document, and from those of previous versions (which should, if there were any, be listed in the History section of the Document). You may use the same title as a previous version if the original publisher of that version gives permission.

B. List on the Title Page, as authors, one or more persons or entities responsible for authorship of the modifications in the Modified Version, together with at least five of the principal authors of the Document (all of its principal authors, if it has fewer than five), unless they release you from this requirement.

C. State on the Title page the name of the publisher of the Modified Version, as the

publisher. D. Preserve all the copyright notices of the Document. E. Add an appropriate copyright notice for your modifications adjacent to the other

copyright notices. F. Include, immediately after the copyright notices, a license notice giving the public

permission to use the Modified Version under the terms of this License, in the form shown in the Addendum below.

G. Preserve in that license notice the full lists of Invariant Sections and required Cover

Texts given in the Document’s license notice. H. Include an unaltered copy of this License. I. Preserve the section Entitled “History”, Preserve its Title, and add to it an item

stating at least the title, year, new authors, and publisher of the Modified Version as given on the Title Page. If there is no section Entitled “History” in the Document, create one stating the title, year, authors, and publisher of the Document as given on its Title Page, then add an item describing the Modified Version as stated in the previous sentence.

J. Preserve the network location, if any, given in the Document for public access to a Transparent copy of the Document, and likewise the network locations given in the Document for previous versions it was based on. These may be placed in the “History” section. You may omit a network location for a work that was published at least four years before the Document itself, or if the original publisher of the version it refers to gives permission.

K. For any section Entitled “Acknowledgements” or “Dedications”, Preserve the Title of the section, and preserve in the section all the substance and tone of each of the contributor acknowledgements and/or dedications given therein.

L. Preserve all the Invariant Sections of the Document, unaltered in their text and in their titles. Section numbers or the equivalent are not considered part of the section titles.

M. Delete any section Entitled “Endorsements”. Such a section may not be included

in the Modified Version. N. Do not retitle any existing section to be Entitled “Endorsements” or to conflict in

title with any Invariant Section. O. Preserve any Warranty Disclaimers.

If the Modified Version includes new front-matter sections or appendices that qualify as Secondary Sections and contain no material copied from the Document, you may at your option designate some or all of these sections as invariant. To do this, add their titles to the list of Invariant Sections in the Modified Version’s license notice. These titles must be distinct from any other section titles. You may add a section Entitled “Endorsements”, provided it contains nothing but endorsements of your Modified Version by various parties—for example, statements of peer review or that the text has been approved by an organization as the authoritative definition of a standard. You may add a passage of up to five words as a Front-Cover Text, and a passage of up to 25 words as a Back-Cover Text, to the end of the list of Cover Texts in the Modified Version. Only one passage of Front-Cover Text and one of Back-Cover Text may be added by (or through arrangements made by) any one entity. If the Document already includes a cover text for the same cover, previously added by you or by arrangement made by the same entity you are acting on behalf of, you may not add another; but you may replace the old one, on explicit permission from the previous publisher that added the old one. The author(s) and publisher(s) of the Document do not by this License give permission to use their names for publicity for or to assert or imply endorsement of any Modified Version.

5. COMBINING DOCUMENTS You may combine the Document with other documents released under this License, under the terms defined in section 4 above for modified versions, provided that you include in the combination all of the Invariant Sections of all of the original documents, unmodified, and list them all as Invariant Sections of your combined work in its license notice, and that you preserve all their Warranty Disclaimers. The combined work need only contain one copy of this License, and multiple identical Invariant Sections may be replaced with a single copy. If there are multiple Invariant Sections with the same name but different contents, make the title of each such section unique by adding at the end of it, in parentheses, the name of the original author or publisher of that section if known, or else a unique number. Make the same adjustment to the section titles in the list of Invariant Sections in the license notice of the combined work. In the combination, you must combine any sections Entitled “History” in the various original documents, forming one section Entitled “History”; likewise combine any sections Entitled “Acknowledgements”, and any sections Entitled “Dedications”. You must delete all sections Entitled “Endorsements.”

6. COLLECTIONS OF DOCUMENTS You may make a collection consisting of the Document and other documents released under this License, and replace the individual copies of this License in the various documents with a single copy that is included in the collection, provided that you follow the rules of this License for verbatim copying of each of the documents in all other respects. You may extract a single document from such a collection, and distribute it individually under this License, provided you insert a copy of this License into the extracted document, and follow this License in all other respects regarding verbatim copying of that document.

7. AGGREGATION WITH INDEPENDENT WORKS A compilation of the Document or its derivatives with other separate and independent documents or works, in or on a volume of a storage or distribution medium, is called an “aggregate” if the copyright resulting from the compilation is not used to limit the legal rights of the compilation’s users beyond what the individual works permit. When the Document is included in an aggregate, this License does not apply to the other works in the aggregate which are not themselves derivative works of the Document. If the Cover Text requirement of section 3 is applicable to these copies of the Document, then if the Document is less than one half of the entire aggregate, the Document’s Cover Texts may be placed on covers that bracket the Document within the aggregate, or the electronic equivalent of covers if the Document is in electronic form. Otherwise they must appear on printed covers that bracket the whole aggregate.

8. TRANSLATION Translation is considered a kind of modification, so you may distribute translations of the Document under the terms of section 4. Replacing Invariant Sections with translations requires special permission from their copyright holders, but you may include translations of some or all Invariant Sections in addition to the original versions of these Invariant Sections. You may include a translation of this License, and all the license notices in the Document, and any Warranty Disclaimers, provided that you also include the original English version of this License and the original versions of those notices and disclaimers. In case of a disagreement between the translation and the original version of this License or a notice or disclaimer, the original version will prevail. If a section in the Document is Entitled “Acknowledgements”, “Dedications”, or “History”, the requirement (section 4) to Preserve its Title (section 1) will typically require changing the actual title.

9. TERMINATION You may not copy, modify, sublicense, or distribute the Document except as expressly provided under this License. Any attempt otherwise to copy, modify, sublicense, or distribute it is void, and will automatically terminate your rights under this License. However, if you cease all violation of this License, then your license from a particular copyright holder is reinstated (a) provisionally, unless and until the copyright holder explicitly and finally terminates your license, and (b) permanently, if the copyright holder fails to notify you of the violation by some reasonable means prior to 60 days after the cessation. Moreover, your license from a particular copyright holder is reinstated permanently if the copyright holder notifies you of the violation by some reasonable means, this is the first time you have received notice of violation of this License (for any work) from that copyright holder, and you cure the violation prior to 30 days after your receipt of the notice. Termination of your rights under this section does not terminate the licenses of parties who have received copies or rights from you under this License. If your rights have been terminated and not permanently reinstated, receipt of a copy of some or all of the same material does not give you any rights to use it.

10. FUTURE REVISIONS OF THIS LICENSE The Free Software Foundation may publish new, revised versions of the GNU Free Documentation License from time to time. Such new versions will be similar in spirit to the present version, but may differ in detail to address new problems or concerns.

See http://www.gnu.org/copyleft/.

Each version of the License is given a distinguishing version number. If the Document specifies that a particular numbered version of this License “or any later version” applies to it, you have the option of following the terms and conditions either of that specified version or of any later version that has been published (not as a draft) by the Free Software Foundation. If the Document does not specify a version number of this License, you may choose any version ever published (not as a draft) by the Free Software Foundation. If the Document specifies that a proxy can decide which future versions of this License can be used, that proxy’s public statement of acceptance of a version permanently authorizes you to choose that version for the Document.

11. RELICENSING “Massive Multiauthor Collaboration Site” (or “MMC Site”) means any World Wide Web server that publishes copyrightable works and also provides prominent facilities for anybody to edit those works. A public wiki that anybody can edit is an example of such a server. A “Massive Multiauthor Collaboration” (or “MMC”) contained in the site means any set of copyrightable works thus published on the MMC site. “CC-BY-SA” means the Creative Commons Attribution-Share Alike 3.0 license published by Creative Commons Corporation, a not-for-profit corporation with a principal place of business in San Francisco, California, as well as future copyleft versions of that license published by that same organization. “Incorporate” means to publish or republish a Document, in whole or in part, as part of another Document. An MMC is “eligible for relicensing” if it is licensed under this License, and if all works that were first published under this License somewhere other than this MMC, and subsequently incorporated in whole or in part into the MMC, (1) had no cover texts or invariant sections, and (2) were thus incorporated prior to November 1, 2008. The operator of an MMC Site may republish an MMC contained in the site under CC-BY-SA on the same site at any time before August 1, 2009, provided the MMC is eligible for relicensing.

ADDENDUM: How to use this License for your documents

To use this License in a document you have written, include a copy of the License in the document and put the following copyright and license notices just after the title page:

        Copyright (C)  year  your name.
        Permission is granted to copy, distribute and/or modify this document
        under the terms of the GNU Free Documentation License, Version 1.3
        or any later version published by the Free Software Foundation;
        with no Invariant Sections, no Front-Cover Texts, and no Back-Cover
        Texts.  A copy of the license is included in the section entitled ``GNU
        Free Documentation License''.

If you have Invariant Sections, Front-Cover Texts and Back-Cover Texts, replace the “with…Texts.” line with this:

          with the Invariant Sections being list their titles, with
          the Front-Cover Texts being list, and with the Back-Cover Texts
          being list.

If you have Invariant Sections without Cover Texts, or some other combination of the three, merge those two alternatives to suit the situation.

If your document contains nontrivial examples of program code, we recommend releasing these examples in parallel under your choice of free software license, such as the GNU General Public License, to permit their use in free software.

13 Index

13.1 General Index

(

(Dwg_Data . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 274

*

*layer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 274 *line. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 274

B

bug reporting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 284

C

checklist for bug reports . . . . . . . . . . . . . . . . . . . . . . 284 code, error . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 code, return. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 Common Entity fields. . . . . . . . . . . . . . . . . . . . . . . . . 260 Common Object fields . . . . . . . . . . . . . . . . . . . . . . . . 262 compilation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 coverage . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

D

data types . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 dwg file format. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 dwg_add_Document . . . . . . . . . . . . . . . . . . . . . . . 272, 273 dwg_add_object . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 273 dwg_add_TEXT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 274 dwg_decode_OBJECT . . . . . . . . . . . . . . . . . . . . . . . . . . . 272

dwg_dynapi_common_set_value . . . . . . . . . . . . . . . 276
dwg_dynapi_common_utf8text . . . . . . . . . . . . . . . . 275
dwg_dynapi_common_value . . . . . . . . . . . . . . . . . . . . 275
dwg_dynapi_entity_set_value . . . . . . . . . . . . . . . 275
dwg_dynapi_entity_utf8text . . . . . . . . . . . . . . . . 275
dwg_dynapi_entity_value . . . . . . . . . . . . . . . . . . . . 275
dwg_dynapi_header_set_value . . . . . . . . . . . . . . . 276
dwg_dynapi_header_utf8text . . . . . . . . . . . . . . . . 275
dwg_dynapi_header_value . . . . . . . . . . . . . . . . . . . . 275
dwg_dynapi_subclass_utf8text . . . . . . . . . . . . . . 275
dwg_dynapi_subclass_value. . . . . . . . . . . . . . . . . . 275

dwg_encode_<OBJECT> . . . . . . . . . . . . . . . . . . . . . . . . 273 dwg_getall_ENTITY . . . . . . . . . . . . . . . . . . . . . . . . . . . 272 dwg_getall_OBJECT . . . . . . . . . . . . . . . . . . . . . . . . . . . 272 dwg_read_file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 271 dwg_setup_<OBJECT>. . . . . . . . . . . . . . . . . . . . . . . . . . 273 dwg_write_file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 272 dwg2dxf . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 279 dwg2ps . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 280 dwg2svg2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 281 dwg2SVG. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 280 dwgadd. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 281 dwgfilter. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 280 dwgfuzz . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 281 dwggrep . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 280 dwglayers. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 280 dwgplot . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 281 dwgread . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 279 dwgrewrite . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 280 dwgwrite . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 279 dxf2dwg . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 279 DXF, ASCII DXF . . . . . . . . . . . . . . . . . . . . . . . . . . . . 276 DXFB, Binary DXF . . . . . . . . . . . . . . . . . . . . . . . . . . 277 dxfwrite . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 279

E

Chapter 13: Index 294

EED . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 269 ENTITIES. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 entity, 3DFACE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 entity, 3DLINE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 entity, 3DSOLID . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 entity, ALIGNMENTPARAMETERENTITY… 26 entity, ARC. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 entity, ARC DIMENSION . . . . . . . . . . . . . . . . . . . . . 28 entity, ARCALIGNEDTEXT. . . . . . . . . . . . . . . . . . . 26 entity, ATTDEF . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 entity, ATTRIB . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 entity, BASEPOINTPARAMETERENTITY … 32 entity, BLOCK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 entity, BODY . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 entity, CAMERA . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 entity, CIRCLE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 entity, DGNUNDERLAY . . . . . . . . . . . . . . . . . . . . . . 32 entity, DIMENSION ALIGNED. . . . . . . . . . . . . . . . 33 entity, DIMENSION ANG2LN . . . . . . . . . . . . . . . . . 34 entity, DIMENSION ANG3PT . . . . . . . . . . . . . . . . . 35 entity, DIMENSION DIAMETER. . . . . . . . . . . . . . 37 entity, DIMENSION LINEAR . . . . . . . . . . . . . . . . . 38 entity, DIMENSION ORDINATE . . . . . . . . . . . . . . 39 entity, DIMENSION RADIUS . . . . . . . . . . . . . . . . . 40 entity, DWFUNDERLAY . . . . . . . . . . . . . . . . . . . . . . 42 entity, ELLIPSE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 entity, ENDBLK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 entity, ENDREP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 entity, EXTRUDEDSURFACE . . . . . . . . . . . . . . . . . 42 entity, FLIPGRIPENTITY. . . . . . . . . . . . . . . . . . . . . 45 entity, FLIPPARAMETERENTITY. . . . . . . . . . . . 45 entity, GEOPOSITIONMARKER . . . . . . . . . . . . . . 45 entity, HATCH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 entity, HELIX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 entity, IMAGE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 entity, INSERT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 entity, JUMP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 entity, LARGE RADIAL DIMENSION . . . . . . . . 50 entity, LAYOUTPRINTCONFIG . . . . . . . . . . . . . . 51 entity, LEADER . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 entity, LIGHT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 entity, LINE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 entity, LINEARGRIPENTITY . . . . . . . . . . . . . . . . . 55 entity, LINEARPARAMETERENTITY . . . . . . . . 55 entity, LOAD . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 entity, LOFTEDSURFACE . . . . . . . . . . . . . . . . . . . . 55 entity, LWPOLYLINE . . . . . . . . . . . . . . . . . . . . . . . . . 58 entity, MESH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 entity, MINSERT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 entity, MLINE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 entity, MPOLYGON . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 entity, MTEXT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 entity, MULTILEADER. . . . . . . . . . . . . . . . . . . . . . . . 63 entity, NAVISWORKSMODEL. . . . . . . . . . . . . . . . . 65 entity, NURBSURFACE . . . . . . . . . . . . . . . . . . . . . . . 66 entity, OLE2FRAME . . . . . . . . . . . . . . . . . . . . . . . . . . 67 entity, OLEFRAME . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 entity, PDFUNDERLAY . . . . . . . . . . . . . . . . . . . . . . . 68 entity, PLANESURFACE . . . . . . . . . . . . . . . . . . . . . . 68 entity, POINT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 entity, POINTCLOUD . . . . . . . . . . . . . . . . . . . . . . . . . 70 entity, POINTCLOUDEX . . . . . . . . . . . . . . . . . . . . . . 71 entity, POINTPARAMETERENTITY . . . . . . . . . 73 entity, POLARGRIPENTITY . . . . . . . . . . . . . . . . . . 73 entity, POLYLINE 2D . . . . . . . . . . . . . . . . . . . . . . . . . 73 entity, POLYLINE 3D . . . . . . . . . . . . . . . . . . . . . . . . . 74 entity, POLYLINE MESH . . . . . . . . . . . . . . . . . . . . . 74 entity, POLYLINE PFACE . . . . . . . . . . . . . . . . . . . . 75 entity, PROXY ENTITY . . . . . . . . . . . . . . . . . . . . . . 75 entity, RAY. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 entity, REGION . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 entity, REPEAT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 entity, REVOLVEDSURFACE . . . . . . . . . . . . . . . . . 76 entity, ROTATIONGRIPENTITY. . . . . . . . . . . . . . 78 entity, ROTATIONPARAMETERENTITY. . . . . 78 entity, RTEXT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 entity, SECTIONOBJECT . . . . . . . . . . . . . . . . . . . . . 79 entity, SEQEND . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 entity, SHAPE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 entity, SOLID. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 entity, SPLINE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 entity, SWEPTSURFACE. . . . . . . . . . . . . . . . . . . . . . 81 entity, TABLE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 entity, TEXT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 entity, TOLERANCE . . . . . . . . . . . . . . . . . . . . . . . . . . 91 entity, TRACE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 entity, UNKNOWN ENT . . . . . . . . . . . . . . . . . . . . . . 92 entity, VERTEX 2D . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 entity, VERTEX 3D . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 entity, VERTEX MESH . . . . . . . . . . . . . . . . . . . . . . . 93 entity, VERTEX PFACE . . . . . . . . . . . . . . . . . . . . . . 93 entity, VERTEX PFACE FACE . . . . . . . . . . . . . . . 93 entity, VIEWPORT . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 entity, VISIBILITYGRIPENTITY . . . . . . . . . . . . . 95 entity, VISIBILITYPARAMETERENTITY …. 95 entity, WIPEOUT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 entity, XLINE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 entity, XYGRIPENTITY . . . . . . . . . . . . . . . . . . . . . . 96 entity, XYPARAMETERENTITY . . . . . . . . . . . . . 96 enums . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 error. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 278 error code. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

F

features, still missing . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 271 functions, create . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 273 functions, decoding . . . . . . . . . . . . . . . . . . . . . . . . . . . 271 functions, dynamic field access . . . . . . . . . . . . . . . . 274 functions, encoding . . . . . . . . . . . . . . . . . . . . . . . . . . . 272 functions, other formats. . . . . . . . . . . . . . . . . . . . . . . 276 functions, read path . . . . . . . . . . . . . . . . . . . . . . . . . . 271 functions, write path. . . . . . . . . . . . . . . . . . . . . . . . . . 272

G

gambas . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 282 GeoJSON. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 277

H

header . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 HEADER . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

J

JSON . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 277

L

license . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 linking . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 load dwg . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 281

O

object, ACMECOMMANDHISTORY . . . . . . . . . . 96 object, ACMESCOPE . . . . . . . . . . . . . . . . . . . . . . . . . 96 object, ACMESTATEMGR . . . . . . . . . . . . . . . . . . . . 96 object, ACSH BOOLEAN CLASS . . . . . . . . . . . . . 96 object, ACSH BOX CLASS. . . . . . . . . . . . . . . . . . . . 97 object, ACSH BREP CLASS . . . . . . . . . . . . . . . . . . 97 object, ACSH CHAMFER CLASS . . . . . . . . . . . . . 99 object, ACSH CONE CLASS . . . . . . . . . . . . . . . . . . 99 object, ACSH CYLINDER CLASS . . . . . . . . . . . . 99 object, ACSH EXTRUSION CLASS . . . . . . . . . . 100 object, ACSH FILLET CLASS . . . . . . . . . . . . . . . 101 object, ACSH HISTORY CLASS . . . . . . . . . . . . . 102 object, ACSH LOFT CLASS . . . . . . . . . . . . . . . . . 102 object, ACSH PYRAMID CLASS . . . . . . . . . . . . 102 object, ACSH REVOLVE CLASS. . . . . . . . . . . . . 103 object, ACSH SPHERE CLASS . . . . . . . . . . . . . . 104 object, ACSH SWEEP CLASS . . . . . . . . . . . . . . . 104 object, ACSH TORUS CLASS. . . . . . . . . . . . . . . . 105 object, ACSH WEDGE CLASS. . . . . . . . . . . . . . . 105 object, ALDIMOBJECTCONTEXTDATA …. 106 object, ANGDIMOBJECTCONTEXTDATA . . 106 object,

ANNOTSCALEOBJECTCONTEXTDATA . . 106 object, APPID. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 107 object, APPID CONTROL . . . . . . . . . . . . . . . . . . . 107 object, ASSOC2DCONSTRAINTGROUP . . . . . 107 object,

ASSOC3POINTANGULARDIMACTIONBODY . . 108 object, ASSOCACTION . . . . . . . . . . . . . . . . . . . . . . 109 object, ASSOCACTIONPARAM . . . . . . . . . . . . . . 109 object,

ASSOCALIGNEDDIMACTIONBODY . . . . . 109 object, ASSOCARRAYACTIONBODY . . . . . . . 110 object,

ASSOCARRAYMODIFYACTIONBODY … 110 object,

ASSOCARRAYMODIFYPARAMETERS … 218 object,

ASSOCARRAYPATHPARAMETERS . . . . . . 219 object,

ASSOCARRAYPOLARPARAMETERS …. 219 object,

ASSOCARRAYRECTANGULARPARAMETERS . . 219 object, ASSOCASMBODYACTIONPARAM . . 110 object,

ASSOCBLENDSURFACEACTIONBODY . . 112 object,

ASSOCCOMPOUNDACTIONPARAM . . . . . 113 object, ASSOCDEPENDENCY . . . . . . . . . . . . . . . 113 object, ASSOCDIMDEPENDENCYBODY …. 114 object, ASSOCEDGEACTIONPARAM . . . . . . . 114 object,

ASSOCEDGECHAMFERACTIONBODY . . 115 object,

ASSOCEDGEFILLETACTIONBODY . . . . . . 115 object,

ASSOCEXTENDSURFACEACTIONBODY . . 115 object,

ASSOCEXTRUDEDSURFACEACTIONBODY . . 115 object, ASSOCFACEACTIONPARAM . . . . . . . . 116 object,

ASSOCFILLETSURFACEACTIONBODY . . 116 object, ASSOCGEOMDEPENDENCY . . . . . . . . 116 object,

ASSOCLOFTEDSURFACEACTIONBODY . . 117 object, ASSOCMLEADERACTIONBODY …. 117 object, ASSOCNETWORK . . . . . . . . . . . . . . . . . . . 117 object,

ASSOCNETWORKSURFACEACTIONBODY . . 118 object, ASSOCOBJECTACTIONPARAM. . . . . 119 object,

ASSOCOFFSETSURFACEACTIONBODY . . 119 object,

ASSOCORDINATEDIMACTIONBODY …. 119 object,

ASSOCOSNAPPOINTREFACTIONPARAM . . 120 object,

ASSOCPATCHSURFACEACTIONBODY . . 120 object, ASSOCPATHACTIONPARAM. . . . . . . . 121 object, ASSOCPERSSUBENTMANAGER …. 121 object,

ASSOCPLANESURFACEACTIONBODY . . 123 object, ASSOCPOINTREFACTIONPARAM . . 124 object,

ASSOCRESTOREENTITYSTATEACTIONBODY . . 124 object,

ASSOCREVOLVEDSURFACEACTIONBODY . . 125 object,

ASSOCROTATEDDIMACTIONBODY . . . . . 125 object,

ASSOCSWEPTSURFACEACTIONBODY . . 125 object,

ASSOCTRIMSURFACEACTIONBODY …. 125 object, ASSOCVALUEDEPENDENCY . . . . . . . 126 object, ASSOCVARIABLE . . . . . . . . . . . . . . . . . . . 126 object, ASSOCVERTEXACTIONPARAM …. 127 object, BLKREFOBJECTCONTEXTDATA . . 127 object, BLOCK CONTROL . . . . . . . . . . . . . . . . . . 154 object, BLOCK HEADER . . . . . . . . . . . . . . . . . . . . 155 object,

BLOCKALIGNEDCONSTRAINTPARAMETER . . 128 object, BLOCKALIGNMENTGRIP. . . . . . . . . . . 129 object,

BLOCKALIGNMENTPARAMETER . . . . . . . 129 object,

BLOCKANGULARCONSTRAINTPARAMETER . . 130 object, BLOCKARRAYACTION. . . . . . . . . . . . . . 131 object, BLOCKBASEPOINTPARAMETER . . 132 object,

BLOCKDIAMETRICCONSTRAINTPARAMETER . . 132 object, BLOCKFLIPACTION. . . . . . . . . . . . . . . . . 133 object, BLOCKFLIPGRIP . . . . . . . . . . . . . . . . . . . . 134 object, BLOCKFLIPPARAMETER . . . . . . . . . . . 135 object,

BLOCKGRIPLOCATIONCOMPONENT … 136 object,

BLOCKHORIZONTALCONSTRAINTPARAMETER . . 136 object,

BLOCKLINEARCONSTRAINTPARAMETER . . 137 object, BLOCKLINEARGRIP . . . . . . . . . . . . . . . . 138 object, BLOCKLINEARPARAMETER . . . . . . . 138 object, BLOCKLOOKUPACTION . . . . . . . . . . . . 139 object, BLOCKLOOKUPGRIP . . . . . . . . . . . . . . . 140 object, BLOCKLOOKUPPARAMETER . . . . . . 140 object, BLOCKMOVEACTION. . . . . . . . . . . . . . . 141 object,

BLOCKPARAMDEPENDENCYBODY. . . . . 142 object, BLOCKPOINTPARAMETER. . . . . . . . . 142 object, BLOCKPOLARGRIP . . . . . . . . . . . . . . . . . 143 object, BLOCKPOLARPARAMETER . . . . . . . . 143 object, BLOCKPOLARSTRETCHACTION… 144 object, BLOCKPROPERTIESTABLE . . . . . . . . 145 object, BLOCKPROPERTIESTABLEGRIP… 145 object,

BLOCKRADIALCONSTRAINTPARAMETER . . 145 object, BLOCKREPRESENTATION. . . . . . . . . . 146 object, BLOCKROTATEACTION . . . . . . . . . . . . 146 object, BLOCKROTATIONGRIP . . . . . . . . . . . . . 147 object, BLOCKROTATIONPARAMETER …. 147 object, BLOCKSCALEACTION . . . . . . . . . . . . . . 148 object, BLOCKSTRETCHACTION. . . . . . . . . . . 149 object, BLOCKUSERPARAMETER . . . . . . . . . . 150 object,

BLOCKVERTICALCONSTRAINTPARAMETER . . 150 object, BLOCKVISIBILITYGRIP . . . . . . . . . . . . 152 object, BLOCKVISIBILITYPARAMETER … 152 object, BLOCKXYGRIP. . . . . . . . . . . . . . . . . . . . . . 153 object, BLOCKXYPARAMETER. . . . . . . . . . . . . 153 object, BREAKDATA . . . . . . . . . . . . . . . . . . . . . . . . 156 object, BREAKPOINTREF. . . . . . . . . . . . . . . . . . . 156 object, CELLSTYLEMAP . . . . . . . . . . . . . . . . . . . . 157 object, CONTEXTDATAMANAGER . . . . . . . . . 157 object, CSACDOCUMENTOPTIONS. . . . . . . . . 157 object, CURVEPATH . . . . . . . . . . . . . . . . . . . . . . . . . 157 object, DATALINK . . . . . . . . . . . . . . . . . . . . . . . . . . . 157 object, DATATABLE . . . . . . . . . . . . . . . . . . . . . . . . . 158 object, DBCOLOR . . . . . . . . . . . . . . . . . . . . . . . . . . . 158 object, DETAILVIEWSTYLE. . . . . . . . . . . . . . . . . 158 object, DGNDEFINITION . . . . . . . . . . . . . . . . . . . . 218 object, DICTIONARY . . . . . . . . . . . . . . . . . . . . . . . . 160 object, DICTIONARYVAR . . . . . . . . . . . . . . . . . . . 160 object, DICTIONARYWDFLT. . . . . . . . . . . . . . . . 160 object, DIMASSOC. . . . . . . . . . . . . . . . . . . . . . . . . . . 161 object, DIMSTYLE . . . . . . . . . . . . . . . . . . . . . . . . . . . 161 object, DIMSTYLE CONTROL . . . . . . . . . . . . . . 164 object, DMDIMOBJECTCONTEXTDATA … 165 object, DUMMY. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 165 object, DWFDEFINITION . . . . . . . . . . . . . . . . . . . 218 object, DYNAMICBLOCKPROXYNODE. . . . . 165 object,

DYNAMICBLOCKPURGEPREVENTER . . 165 object, EVALUATION GRAPH . . . . . . . . . . . . . . 165 object, FCFOBJECTCONTEXTDATA . . . . . . . 166 object, FIELD . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 166 object, FIELDLIST . . . . . . . . . . . . . . . . . . . . . . . . . . . 167 object, GEODATA. . . . . . . . . . . . . . . . . . . . . . . . . . . . 167 object, GEOMAPIMAGE . . . . . . . . . . . . . . . . . . . . . 169 object, GRADIENT BACKGROUND. . . . . . . . . 170 object, GROUND PLANE BACKGROUND . . 170 object, GROUP. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 object, IBL BACKGROUND . . . . . . . . . . . . . . . . . 171 object, IDBUFFER . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 object, IMAGE BACKGROUND . . . . . . . . . . . . . 172 object, IMAGEDEF . . . . . . . . . . . . . . . . . . . . . . . . . . 172 object, IMAGEDEF REACTOR . . . . . . . . . . . . . . 172 object, INDEX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 172 object, LAYER . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173 object, LAYER CONTROL. . . . . . . . . . . . . . . . . . . 173 object, LAYER INDEX. . . . . . . . . . . . . . . . . . . . . . . 174 object, LAYERFILTER . . . . . . . . . . . . . . . . . . . . . . . 173 object, LAYOUT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 174 object, LEADEROBJECTCONTEXTDATA . . 175 object, LIGHTLIST . . . . . . . . . . . . . . . . . . . . . . . . . . 175 object, LONG TRANSACTION . . . . . . . . . . . . . . 175 object, LTYPE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 176 object, LTYPE CONTROL . . . . . . . . . . . . . . . . . . . 176 object, MATERIAL. . . . . . . . . . . . . . . . . . . . . . . . . . . 177 object, MENTALRAYRENDERSETTINGS … 179 object,

MLEADEROBJECTCONTEXTDATA . . . . . 181 object, MLEADERSTYLE . . . . . . . . . . . . . . . . . . . . 182 object, MLINESTYLE . . . . . . . . . . . . . . . . . . . . . . . . 184 object, MOTIONPATH . . . . . . . . . . . . . . . . . . . . . . . 184 object,

MTEXTATTRIBUTEOBJECTCONTEXTDATA . . 185 object, MTEXTOBJECTCONTEXTDATA … 185 object, NAVISWORKSMODELDEF . . . . . . . . . . 186 object, OBJECT PTR . . . . . . . . . . . . . . . . . . . . . . . . 186 object, ORDDIMOBJECTCONTEXTDATA . . 186 object, PARTIAL VIEWING INDEX . . . . . . . . . 187 object, PDFDEFINITION . . . . . . . . . . . . . . . . . . . . 218 object, PERSUBENTMGR . . . . . . . . . . . . . . . . . . . 187 object, PLACEHOLDER . . . . . . . . . . . . . . . . . . . . . 187 object, PLOTSETTINGS . . . . . . . . . . . . . . . . . . . . . 187 object, POINTCLOUDCOLORMAP . . . . . . . . . . 189 object, POINTCLOUDDEF. . . . . . . . . . . . . . . . . . . 189 object, POINTCLOUDDEF REACTOR . . . . . . 190 object, POINTCLOUDDEF REACTOR EX . . 190 object, POINTCLOUDDEFEX . . . . . . . . . . . . . . . 190 object, POINTPATH . . . . . . . . . . . . . . . . . . . . . . . . . 190 object, PROXY OBJECT . . . . . . . . . . . . . . . . . . . . 190 object,

RADIMLGOBJECTCONTEXTDATA . . . . . . 191 object, RADIMOBJECTCONTEXTDATA …. 191 object, RAPIDRTRENDERSETTINGS . . . . . . . 192 object, RASTERVARIABLES. . . . . . . . . . . . . . . . . 193 object, RENDERENTRY . . . . . . . . . . . . . . . . . . . . . 193 object, RENDERENVIRONMENT . . . . . . . . . . . 194 object, RENDERGLOBAL. . . . . . . . . . . . . . . . . . . . 194 object, RENDERSETTINGS. . . . . . . . . . . . . . . . . . 195 object, SCALE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 196 object, SECTION MANAGER. . . . . . . . . . . . . . . . 198 object, SECTION SETTINGS . . . . . . . . . . . . . . . . 198 object, SECTIONVIEWSTYLE . . . . . . . . . . . . . . . 196 object, SKYLIGHT BACKGROUND . . . . . . . . . 198 object, SOLID BACKGROUND . . . . . . . . . . . . . . 199 object, SORTENTSTABLE . . . . . . . . . . . . . . . . . . . 199 object, SPATIAL FILTER . . . . . . . . . . . . . . . . . . . . 199 object, SPATIAL INDEX . . . . . . . . . . . . . . . . . . . . . 200 object, STYLE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 200 object, STYLE CONTROL . . . . . . . . . . . . . . . . . . . 201 object, SUN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 201 object, SUNSTUDY . . . . . . . . . . . . . . . . . . . . . . . . . . 201 object, TABLECONTENT. . . . . . . . . . . . . . . . . . . . 203 object, TABLEGEOMETRY . . . . . . . . . . . . . . . . . . 203 object, TABLESTYLE. . . . . . . . . . . . . . . . . . . . . . . . 203 object, TEXTOBJECTCONTEXTDATA . . . . . 204 object, TVDEVICEPROPERTIES . . . . . . . . . . . . 204 object, UCS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 205 object, UCS CONTROL . . . . . . . . . . . . . . . . . . . . . . 206 object, UNKNOWN OBJ . . . . . . . . . . . . . . . . . . . . . 206 object, VBA PROJECT . . . . . . . . . . . . . . . . . . . . . . 206 object, VIEW . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 206 object, VIEW CONTROL . . . . . . . . . . . . . . . . . . . . 208 object, VISUALSTYLE . . . . . . . . . . . . . . . . . . . . . . . 208 object, VPORT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 214 object, VPORT CONTROL . . . . . . . . . . . . . . . . . . 217 object, VX CONTROL . . . . . . . . . . . . . . . . . . . . . . . 217 object, VX TABLE RECORD . . . . . . . . . . . . . . . . 217 object, WIPEOUTVARIABLES. . . . . . . . . . . . . . . 218 object, XRECORD . . . . . . . . . . . . . . . . . . . . . . . . . . . 218 OBJECTS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 OCS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

P

patches, contributing. . . . . . . . . . . . . . . . . . . . . . . . . . 284 perl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 282 problems . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 284 programs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 279 projects, related. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 python . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 282

R

Reference API . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 283 reporting bugs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 284 return code . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

S

Sections . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 264 strings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 276 structs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 structures . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 269 SummaryInfo . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 265

T

table, APPID . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 107 table, BLOCK HEADER . . . . . . . . . . . . . . . . . . . . . 155 table, DIMSTYLE . . . . . . . . . . . . . . . . . . . . . . . . . . . . 161 table, LAYER . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173 table, LTYPE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 176 table, STYLE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 200 table, UCS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 205 table, VIEW. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 206 table, VPORT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 214 table, VX TABLE RECORD . . . . . . . . . . . . . . . . . 217 table control, APPID CONTROL. . . . . . . . . . . . . 107 table control, BLOCK CONTROL. . . . . . . . . . . . 154 table control, DIMSTYLE CONTROL . . . . . . . . 164 table control, LAYER CONTROL . . . . . . . . . . . . 173 table control, LTYPE CONTROL . . . . . . . . . . . . 176 table control, STYLE CONTROL . . . . . . . . . . . . 201 table control, UCS CONTROL . . . . . . . . . . . . . . . 206 table control, VIEW CONTROL . . . . . . . . . . . . . 208 table control, VPORT CONTROL . . . . . . . . . . . . 217 table control, VX CONTROL . . . . . . . . . . . . . . . . 217 Tables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 267

U

unknown . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 281

V

version, API/ABI . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

X

XDATA . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 270

13.2 Object and Field Index

__iterator . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 155, 260 _3DDWFPREC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 _dxf_sab_converted …. 25, 43, 56, 66, 69, 77, 82,

98, 111

3

3DFACE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 3DLINE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 3DSOLID . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24

A

aaab_version . . . . . . . . . 108, 109, 110, 117, 119, 125 aab_version. . . . . 108, 109, 110, 112, 115, 116, 117,

118, 119, 120, 123, 124, 125, 126

aap_version. . . . . 109, 111, 113, 114, 116, 119, 120,

121, 124, 127, 218

ACADMAINTVER . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 acis_data . . . . . . . 25, 43, 56, 66, 69, 76, 82, 97, 111 acis_empty . . . . . . 24, 42, 55, 66, 68, 76, 81, 97, 111 acis_empty_bit . . 26, 44, 57, 67, 70, 78, 83, 98, 112 acis_empty2. . . . . 25, 43, 56, 66, 69, 77, 82, 98, 112 acis_index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 220 ACMECOMMANDHISTORY . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 ACMESCOPE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 ACMESTATEMGR. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 ACSH_BOOLEAN_CLASS . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 ACSH_BOX_CLASS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 ACSH_BREP_CLASS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 ACSH_CHAMFER_CLASS . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 ACSH_CONE_CLASS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 ACSH_CYLINDER_CLASS. . . . . . . . . . . . . . . . . . . . . . . . . . 99 ACSH_EXTRUSION_CLASS . . . . . . . . . . . . . . . . . . . . . . . 100 ACSH_FILLET_CLASS . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 ACSH_HISTORY_CLASS. . . . . . . . . . . . . . . . . . . . . . . . . . 102 ACSH_LOFT_CLASS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102 ACSH_PYRAMID_CLASS. . . . . . . . . . . . . . . . . . . . . . . . . . 102 ACSH_REVOLVE_CLASS. . . . . . . . . . . . . . . . . . . . . . . . . . 103 ACSH_SPHERE_CLASS . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 ACSH_SWEEP_CLASS . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 ACSH_TORUS_CLASS . . . . . . . . . . . . . . . . . . . . . . . . . . . . 105 ACSH_WEDGE_CLASS . . . . . . . . . . . . . . . . . . . . . . . . . . . . 105 act_measurement …. 28, 34, 35, 36, 37, 39, 40, 41,

51, 234

action_index . . . . . . . . . . . . . . . . . . 107, 109, 118, 126 action_offset_x. . . . . . . . . . . . . . . . . . . . 134, 142, 150 action_offset_y. . . . . . . . . . . . . . . . . . . . 134, 142, 150 action_type. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 115 actionbody. . . . . . 107, 108, 109, 110, 117, 118, 119,

125, 126

actions …. 108, 117, 118, 131, 134, 140, 141, 144,

147, 149

active_cycles . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 235 active_viewport . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 175 adb_version . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114, 142 additional_data_flag . . . . . . . . . . . . . . . . . . . . . . . 255 affects_graphics . . . . . . . . . . . . . . . . . . . . . . . . . . . . 236 ALDIMOBJECTCONTEXTDATA . . . . . . . . . . . . . . . . . . . . . 106 align_angle. . . . . . . . . . . . . . . . . . . . . . 44, 83, 101, 105 align_direction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 align_option . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101, 105 align_perpendicular . . . . . . . . . . . . . . . . . . . . . . . . 130 align_space. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 183 align_start . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45, 84 alignment. . . . . . . . . . . . . . . . . . . . . . . . . . . . 27, 176, 245 alignment_pt . . . . . . . . . . . . . . . . . 29, 31, 91, 185, 204

ALIGNMENTPARAMETERENTITY . . . . . . . . . . . . . . . . . . . . 26

alt_hlt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 205 alt_hltcolor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 205 ambient_color . . . . . . . . . . . . . . . . . . 94, 177, 207, 215 ANGBASE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

ANGDIMOBJECTCONTEXTDATA . . . . . . . . . . . . . . . . . . . . 106

ANGDIR. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 angle . . . . . . . . . . . . . . . . . . . . 33, 47, 61, 131, 148, 237 angle_desc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 144, 148 angle_name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 144, 148 angle_offset . . . . . . . . . . . . . . . . . . . . . . . 134, 142, 150 angle_value_set . . . . . . . . . . . . . . . . . . . . . . . . 144, 148 annot_type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 annotative_app. . . . . . . . . . . . . . . . . . . . . . . . . . . . 30, 32 annotative_data_bytes. . . . . . . . . . . . . . . . . . . . 30, 32 annotative_data_size . . . . . . . . . . . . . . . . . . . . . 30, 32 annotative_short . . . . . . . . . . . . . . . . . . . . . . . . . 30, 32

ANNOTSCALEOBJECTCONTEXTDATA . . . . . . . . . . . . . . . 106

anonymous . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 155 antialiasing_level. . . . . . . . . . . . . . . . . . . . . . . . . . 205 appid. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 APPID. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 107 APPID_CONTROL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 107 APPID_CONTROL_OBJECT. . . . . . . . . . . . . . . . . . . . . . . . . . 8 arc_end_param . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 arc_handle . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28

arc_length_parameterization . . . . . . . . . . . . . . . . 57

arc_pt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106 arc_start_param . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 ARC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 ARC_DIMENSION . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 ARCALIGNEDTEXT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 areafillparms . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 246 array_index. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 219 arrow_end_symbol . . . . . . . . . . . . . . . . . . . . . . . . . . . . 196 arrow_handle . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64, 240 arrow_head . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 182 arrow_head_size . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 182 arrow_position . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 197 arrow_size . . . . . . . . . . . . . . . . . . . . . . . . . . 64, 240, 243 arrow_start_symbol. . . . . . . . . . . . . . . . . . . . . . . . . . 196 arrow_symbol . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 159 arrow_symbol_color . . . . . . . . . . . . . . . . . . . . . 159, 196

arrow_symbol_extension_length . . . . . . . . . . . . . 197

arrow_symbol_size . . . . . . . . . . . . . . . . . . . . . . 159, 196 arrowhead . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 239 arrowhead_on. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 arrowhead_type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 arrowheads . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 asdap_class_version . . . . . 111, 114, 116, 119, 127 aspect_ratio . . . . . . . . . . . . . . . . . . . . . . . . 21, 207, 215 assoc_dep . . . . . . . . . . . . . . . . . 108, 109, 117, 119, 125 ASSOC2DCONSTRAINTGROUP . . . . . . . . . . . . . . . . . . . . . 107

ASSOC3POINTANGULARDIMACTIONBODY. . . . . . . . . . . 108

ASSOCACTION. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 ASSOCACTIONPARAM . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109

ASSOCALIGNEDDIMACTIONBODY. . . . . . . . . . . . . . . . . . 109

ASSOCARRAYACTIONBODY . . . . . . . . . . . . . . . . . . . . . . . 110

ASSOCARRAYMODIFYACTIONBODY . . . . . . . . . . . . . . . . 110
ASSOCARRAYMODIFYPARAMETERS . . . . . . . . . . . . . . . . 218

ASSOCARRAYPARAMETERS . . . . . . . . . . . . . . . . . . . . . . . 218

ASSOCARRAYPATHPARAMETERS. . . . . . . . . . . . . . . . . . . 219
ASSOCARRAYPOLARPARAMETERS. . . . . . . . . . . . . . . . . . 219
ASSOCARRAYRECTANGULARPARAMETERS. . . . . . . . . . . 219
ASSOCASMBODYACTIONPARAM . . . . . . . . . . . . . . . . . . . . 110
ASSOCBLENDSURFACEACTIONBODY . . . . . . . . . . . . . . . 112
ASSOCCOMPOUNDACTIONPARAM. . . . . . . . . . . . . . . . . . . 113

assocdep . . . . . . . . . . . . . . . . . . . . . . . 108, 116, 126, 222 ASSOCDEPENDENCY . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 ASSOCDIMDEPENDENCYBODY . . . . . . . . . . . . . . . . . . . . . 114 ASSOCEDGEACTIONPARAM . . . . . . . . . . . . . . . . . . . . . . . 114

ASSOCEDGECHAMFERACTIONBODY . . . . . . . . . . . . . . . . 115
ASSOCEDGEFILLETACTIONBODY. . . . . . . . . . . . . . . . . . 115

ASSOCEXTENDSURFACEACTIONBODY . . . . . . . . . . . . . . 115
ASSOCEXTRUDEDSURFACEACTIONBODY . . . . . . . . . . . . 115

ASSOCFACEACTIONPARAM . . . . . . . . . . . . . . . . . . . . . . . 116

ASSOCFILLETSURFACEACTIONBODY . . . . . . . . . . . . . . 116

ASSOCGEOMDEPENDENCY . . . . . . . . . . . . . . . . . . . . . . . . 116 associated_annotation . . . . . . . . . . . . . . . . . . . . . . . 53 associated_ucs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 207 associativity . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 161

ASSOCLOFTEDSURFACEACTIONBODY . . . . . . . . . . . . . . 117

ASSOCMLEADERACTIONBODY . . . . . . . . . . . . . . . . . . . . . 117 ASSOCNETWORK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 117

ASSOCNETWORKSURFACEACTIONBODY . . . . . . . . . . . . . 118

ASSOCOBJECTACTIONPARAM . . . . . . . . . . . . . . . . . . . . . 119

ASSOCOFFSETSURFACEACTIONBODY . . . . . . . . . . . . . . 119
ASSOCORDINATEDIMACTIONBODY . . . . . . . . . . . . . . . . 119
ASSOCOSNAPPOINTREFACTIONPARAM . . . . . . . . . . . . . 120
ASSOCPATCHSURFACEACTIONBODY . . . . . . . . . . . . . . . 120

ASSOCPATHACTIONPARAM . . . . . . . . . . . . . . . . . . . . . . . 121 ASSOCPERSSUBENTMANAGER . . . . . . . . . . . . . . . . . . . . . 121

ASSOCPLANESURFACEACTIONBODY . . . . . . . . . . . . . . . 123
ASSOCPOINTREFACTIONPARAM. . . . . . . . . . . . . . . . . . . 124
ASSOCRESTOREENTITYSTATEACTIONBODY . . . . . . . . 124
ASSOCREVOLVEDSURFACEACTIONBODY . . . . . . . . . . . . 125
ASSOCROTATEDDIMACTIONBODY. . . . . . . . . . . . . . . . . . 125
ASSOCSWEPTSURFACEACTIONBODY . . . . . . . . . . . . . . . 125
ASSOCTRIMSURFACEACTIONBODY . . . . . . . . . . . . . . . . 125

ASSOCVALUEDEPENDENCY . . . . . . . . . . . . . . . . . . . . . . . 126 assocvariable . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 ASSOCVARIABLE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 126 ASSOCVERTEXACTIONPARAM . . . . . . . . . . . . . . . . . . . . . 127 attach_bottom . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65, 184 attach_dir. . . . . . . . . . . . . . . . . . . . . . 65, 184, 241, 243 attach_left. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 183 attach_right . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 183 attach_top . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65, 184 attachment… 28, 34, 35, 36, 37, 38, 40, 41, 51, 62,

185, 234

attdef . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 240, 254, 258 attdef_class_version . . . . . . . . . . . . . . . . . . . . . . . . 30 ATTDEF. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 ATTDIA. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 attenuation_end_limit . . . . . . . . . . . . . . . . . . . . . . . 53

attenuation_start_limit . . . . . . . . . . . . . . . . . . . . . 53

attenuation_type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 ATTMODE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 attr_defs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 256 ATTREQ. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 attribs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49, 60, 90 ATTRIB. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 attrs. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 258 AUNITS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 AUPREC. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 AUTHOR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 264 auto_height . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63, 186 autotransform . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 243 av_class_version . . . . . . . . . . . . . . . . . . . . . . . . . . . . 126 axis_base_pt. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 axis_point . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 axis_pt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 axis_ratio . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 axis_vector . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48, 78 axis_x . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 220 axis_y . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 220 axis_z . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 220 AXISMODE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 AXISUNIT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

B

b_prop1c . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 212 b_prop1c_int . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 212 b_prop1d . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 212 b_prop1d_int . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 212 b_prop1e . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 212 b_prop1e_int . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 212 b_prop1f . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 212 b_prop1f_int . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 212 b_prop20 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 212 b_prop20_int . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 212 b_prop21 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 212 b_prop21_int . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 212 b_prop22 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 212 b_prop22_int . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 212 b_prop23 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 212 b_prop23_int . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 212 b_prop24 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 212 b_prop24_int . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 212 b_prop2d . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 213 b_prop2d_int . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 213 b_prop31 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 214 b_prop31_int . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 214 b_prop37 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 214 b_prop37_int . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 214 b1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108, 112, 119, 126 b2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112, 126 b280. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 140 b281. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 228 b282. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 228 b290 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103, 127, 175 b293. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 246 b3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 b4 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 b5 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 back_clip_on . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 199 back_clip_z. . . . . . . . . . . . . . . . . . . . . 93, 199, 207, 215 backfaces_enabled . . . . . . . . . . . . . . . . . 179, 192, 195 background . . . . . . . . . . . . . . . . . . . . . . . . . . 95, 208, 216 BACKZ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 bank. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45, 84, 101, 105 base. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 244 base_dir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 244 base_dist . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 base_point . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 base_point_set. . . . . . . . . . . . . . . . . . . . . . . . . . . . 45, 84 base_pt . . . . . . . . . . . . . . . . . . . . 32, 132, 147, 149, 155 base_state_label . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135 base_ucs . . . . . . . . . . . . . . . . . . . 95, 175, 205, 208, 217 base_vert . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 244

BASEPOINTPARAMETERENTITY . . . . . . . . . . . . . . . . . . . . 32

basept. . . . . . 128, 130, 131, 133, 135, 137, 138, 139,

144, 146, 148, 151, 154

bd_prop26 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 213 bd_prop26_int . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 213 bd_prop27 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 213 bd_prop27_int . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 213 bd_prop34 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 214 bd_prop34_int . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 214 bd_prop38 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 214 bd_prop38_int . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 214 bd_prop39 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 214 bd_prop39_int . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 214 bd2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 205 bd2007_45 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 212 bd44. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 bd45. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 be_major … 128, 129, 130, 131, 132, 133, 134, 135, 136, 137, 138, 139, 140, 141, 142, 143, 144, 145, 147,

148, 149, 150, 151, 152, 153

be_minor … 128, 129, 130, 131, 132, 133, 134, 135, 136, 137, 138, 139, 140, 141, 142, 143, 144, 145, 147,

148, 149, 150, 151, 152, 153

beg_tan_vec . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47, 81 begin_addr_r11 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 bend_line_color . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 197 bend_line_length . . . . . . . . . . . . . . . . . . . . . . . . . . . . 197 bend_linewt. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 197 bend_ltype . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 197 bg_bl91 . . 129, 134, 138, 140, 143, 145, 147, 152, 153 bg_bl92 . . 129, 134, 138, 140, 143, 145, 147, 152, 153 bg_color. . . . . . . . . . . . . . . . . . . . . . . . . . . . 231, 245, 255 bg_fill_color . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 bg_fill_flag. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 bg_fill_none . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 255 bg_fill_scale . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 bg_fill_trans . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 bg_insert_cycling . . 129, 134, 138, 140, 143, 145,

147, 152, 153

bg_insert_cycling_weight . . . . 129, 134, 138, 140,

143, 145, 147, 152, 153

bg_location. . . . . 129, 134, 138, 140, 143, 145, 147,

152, 153

bg_scale . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 245 bg_transparency . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 245 bigfont_file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 201 bindata. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44, 200 bindata_size . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44, 200 bitsize . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 bitsize_hi . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 bl_prop25 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 212 bl_prop25_int . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 212 bl_prop28 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 213 bl_prop28_int . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 213 bl_prop2a . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 213 bl_prop2a_int . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 213 bl_prop2b . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 213 bl_prop2b_int . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 213 bl_prop2e . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 213 bl_prop2e_int . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 213 bl_prop2f . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 213 bl_prop2f_int . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 213 bl_prop30 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 213 bl_prop30_int . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 214 bl_prop32 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 214 bl_prop32_int . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 214 bl2. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113, 120, 121, 124 bl92 . . . . . . . . . . . . . . . . . . . . . . . . 99, 100, 101, 104, 157 bl93 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 158, 221 bl95 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 bl96. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 136 blend_options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 blendfactor. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 242 blending_mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 205 BLIPMODE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 blk . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 260 blkisxref . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 155

BLKREFOBJECTCONTEXTDATA . . . . . . . . . . . . . . . . . . . . 127

blob. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 225 blob01 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 223 block… 29, 34, 35, 36, 38, 39, 40, 41, 51, 146, 165,

183, 234, 246

block_color . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65, 183 block_connection . . . . . . . . . . . . . . . . . . . . . . . . . . . . 183 block_entity . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 156 block_handle . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 255 block_header. . . . . . . . . . . . . . . . . . . . . . 49, 60, 90, 174 block_name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 block_offset_r11 . . . . . . . . . . . . . . . . . . . . . . . . . . . . 156 block_owner. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 199 block_rotation. . . . . . . . . . . . . . . . . . . . . . . . . . . 65, 183 block_scale. . . . . . . . . . . . . . . . . . . . . 65, 183, 232, 255 block_scaling . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 156 block_size . . . . . . 25, 42, 56, 66, 68, 76, 81, 97, 111 block_style . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 block_table. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 244 BLOCK. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 BLOCK_CONTROL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 154 BLOCK_CONTROL_OBJECT. . . . . . . . . . . . . . . . . . . . . . . . . . 8 BLOCK_HEADER . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 155 BLOCK_RECORD_MSPACE. . . . . . . . . . . . . . . . . . . . . . . . . . . 8 BLOCK_RECORD_PSPACE. . . . . . . . . . . . . . . . . . . . . . . . . . . 8

BLOCKALIGNEDCONSTRAINTPARAMETER. . . . . . . . . . . 128

BLOCKALIGNMENTGRIP. . . . . . . . . . . . . . . . . . . . . . . . . . 129

BLOCKALIGNMENTPARAMETER . . . . . . . . . . . . . . . . . . . . 129
BLOCKANGULARCONSTRAINTPARAMETER. . . . . . . . . . . 130

BLOCKARRAYACTION . . . . . . . . . . . . . . . . . . . . . . . . . . . . 131

BLOCKBASEPOINTPARAMETER . . . . . . . . . . . . . . . . . . . . 132
BLOCKDIAMETRICCONSTRAINTPARAMETER . . . . . . . . 132

BLOCKFLIPACTION . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 133 BLOCKFLIPGRIP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 134 BLOCKFLIPPARAMETER. . . . . . . . . . . . . . . . . . . . . . . . . . 135

BLOCKGRIPLOCATIONCOMPONENT . . . . . . . . . . . . . . . . 136
BLOCKHORIZONTALCONSTRAINTPARAMETER . . . . . . . 136

blocklabels . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65

BLOCKLINEARCONSTRAINTPARAMETER . . . . . . . . . . . . 137

BLOCKLINEARGRIP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 138 BLOCKLINEARPARAMETER . . . . . . . . . . . . . . . . . . . . . . . 138 BLOCKLOOKUPACTION . . . . . . . . . . . . . . . . . . . . . . . . . . . 139 BLOCKLOOKUPGRIP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 140 BLOCKLOOKUPPARAMETER . . . . . . . . . . . . . . . . . . . . . . . 140 BLOCKMOVEACTION . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 141

BLOCKPARAMDEPENDENCYBODY. . . . . . . . . . . . . . . . . . . 142

BLOCKPOINTPARAMETER . . . . . . . . . . . . . . . . . . . . . . . . 142 BLOCKPOLARGRIP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 143 BLOCKPOLARPARAMETER . . . . . . . . . . . . . . . . . . . . . . . . 143

BLOCKPOLARSTRETCHACTION . . . . . . . . . . . . . . . . . . . . 144

BLOCKPROPERTIESTABLE . . . . . . . . . . . . . . . . . . . . . . . 145

BLOCKPROPERTIESTABLEGRIP. . . . . . . . . . . . . . . . . . . 145
BLOCKRADIALCONSTRAINTPARAMETER . . . . . . . . . . . . 145

BLOCKREPRESENTATION . . . . . . . . . . . . . . . . . . . . . . . . 146 BLOCKROTATEACTION . . . . . . . . . . . . . . . . . . . . . . . . . . . 146 BLOCKROTATIONGRIP . . . . . . . . . . . . . . . . . . . . . . . . . . . 147 BLOCKROTATIONPARAMETER . . . . . . . . . . . . . . . . . . . . . 147 blocks. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153, 229 BLOCKSCALEACTION . . . . . . . . . . . . . . . . . . . . . . . . . . . . 148 BLOCKSTRETCHACTION. . . . . . . . . . . . . . . . . . . . . . . . . . 149 BLOCKUSERPARAMETER. . . . . . . . . . . . . . . . . . . . . . . . . . 150

BLOCKVERTICALCONSTRAINTPARAMETER . . . . . . . . . 150

blockvisi_desc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153 blockvisi_name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153 BLOCKVISIBILITYGRIP . . . . . . . . . . . . . . . . . . . . . . . . 152

BLOCKVISIBILITYPARAMETER. . . . . . . . . . . . . . . . . . . 152

BLOCKXYGRIP. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153 BLOCKXYPARAMETER . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153 blverts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 BODY . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32

border_color_overrides_flag . . . . . . . . . . . . . . . . 87
border_lineweight_overrides_flag. . . . . . . . . . . 88

border_overrides . . . . . . . . . . . . . . . . . . . . . . . . . . . . 237 border_type. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 237

border_visibility_overrides_flag. . . . . . . . . . . 89

borderline_color . . . . . . . . . . . . . . . . . . . . . . . . . . . . 160 borderline_linewt . . . . . . . . . . . . . . . . . . . . . . . . . . . 160 borderline_ltype . . . . . . . . . . . . . . . . . . . . . . . . . . . . 160 borders . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 231, 253 bottom_grid_color . . . . . . . . . . . . . . . . . . . . . . . . . . . 255 bottom_grid_linewt. . . . . . . . . . . . . . . . . . . . . . . . . . 255 bottom_height . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 bottom_margin. . . . . . . . . . . . . . . . . . . . . . . . . . . 188, 231 bottom_row . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 236 bottom_visibility . . . . . . . . . . . . . . . . . . . . . . . . . . . 256 boundary_handles . . . . . . . . . . . . . . . . . . . . . . . . . . . . 238 boundary_line_color . . . . . . . . . . . . . . . . . . . . . . . . 159 boundary_linewt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 159 boundary_ltype . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 159 box_height . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 box_width . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 branch_index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 241 break_flag . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 break_flow_direction . . . . . . . . . . . . . . . . . . . . . . . . 90 break_heights . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 break_rows . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 break_size . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 184 break_spacing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 break_unknown1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 break_unknown2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 BREAKDATA . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 156 BREAKPOINTREF . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 156 breaks. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 240, 241 brightness . . . . . . . . . . . . . . 49, 94, 95, 169, 207, 215 bs1 . . . . . . . . . . . . . . . . . . . . . . . . . 27, 113, 120, 121, 124 bs2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27, 113 bulge . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92, 239 bulges . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58, 248 bulges_present . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 238 bumpmap . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 177 byblock . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 177 byblock_color . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 bylayer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 177 byte . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

C

c_prop29 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 213 c_prop29_int . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 213 c_prop2c . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 213 c_prop2c_int . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 213 c_prop33 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 214 c_prop33_int . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 214 camera_path. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 184 CAMERA. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 CAMERADISPLAY. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 CAMERAHEIGHT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 canonical_media_name . . . . . . . . . . . . . . . . . . . . . . . 188 cast_shadows. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 CECOLOR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 cell_alignment . . . . . . . . . . . . . . . . . . . . . . . . . 232, 255 cell_contents . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 257 cell_flag_override. . . . . . . . . . . . . . . . . . . . . . . . . . 255 cell_parent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 230, 256 cells . . . . . . . . . . . . . . . . . . . . . . . . . . . 85, 157, 203, 259 cellstyle. . . . . . . . . . . . . . . . . . . . . . 204, 236, 253, 259 cellstyle.bg_color. . . . . . . . . . . . . . . . . . . . . . . . . . 252 cellstyle.borders . . . . . . . . . . . . . . . . . . . . . . . . . . . 253

                                                cellstyle.bottom_margin . . . . . . . . . . . . . . . . . . . . 252
                                                cellstyle.content_format. . . . . . . . . . . . . . . . . . . 252
                                                cellstyle.content_layout. . . . . . . . . . . . . . . . . . . 252

cellstyle.data_flags . . . . . . . . . . . . . . . . . . . . . . . 252 cellstyle.horiz_margin . . . . . . . . . . . . . . . . . . . . . 252

                                                cellstyle.margin_horiz_spacing . . . . . . . . . . . . 253
                                                cellstyle.margin_override_flags. . . . . . . . . . . 252
                                                cellstyle.margin_vert_spacing . . . . . . . . . . . . . 253

cellstyle.merge_flags . . . . . . . . . . . . . . . . . . . . . . 252 cellstyle.num_borders . . . . . . . . . . . . . . . . . . . . . . 253

                                                cellstyle.property_override_flags . . . . . . . . 252

cellstyle.right_margin . . . . . . . . . . . . . . . . . . . . . 253

                                                cellstyle.tabledatacolumn_parent . . . . . . . . . 253
                                                cellstyle.tablerow_parent. . . . . . . . . . . . . . . . . . 253

cellstyle.type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 252 cellstyle.vert_margin . . . . . . . . . . . . . . . . . . . . . . 252 cellstyle_id . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 259 CELLSTYLEMAP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 157 CELTSCALE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 CELTYPE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 CELWEIGHT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 center . . . . . . . . . . . . . . . . . . . . . 26, 27, 32, 42, 93, 238 center_pt. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29, 37, 131 CEPSNTYPE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 chain_actions . . 128, 129, 130, 132, 135, 136, 137,

139, 141, 142, 143, 146, 148, 150, 151, 152, 153 CHAMFERA . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 CHAMFERB . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 CHAMFERC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 CHAMFERD . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 channel_flags . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 178 char . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 char* . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 char_spacing. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 check_intersections. . . . . . . . . . . . . . . . . . . . 101, 105 checksum . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 250 child_id . . . . . . . . . . . . . . . . . . . . . . . 113, 120, 121, 124 child_param . . . . . . . . . . . . . . . . . . . 113, 120, 121, 124 child_status . . . . . . . . . . . . . . . . . . 113, 120, 121, 124 childs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 166 childval . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 167 circle_zoom . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94, 216 circle_zoom_percent. . . . . . . . . . . . . . . . . . . . . . . . . . 21 CIRCLE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 class_id . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76, 191 class_version … 28, 30, 31, 33, 34, 35, 37, 38, 39, 40, 44, 45, 48, 50, 51, 53, 63, 70, 71, 78, 83, 95, 96, 106, 107, 108, 109, 110, 111, 112, 113, 114, 115, 116, 117, 118, 119, 120, 121, 124, 125, 126, 127, 142, 157, 159, 165, 166, 167, 169, 170, 171, 172, 175, 179, 181, 182, 184, 185, 186, 187, 189, 190, 191, 192, 193, 194, 195, 196, 199, 201, 202, 203, 204, 221, 233, 247, 259

classification_colorramps. . . . . . . . . . . . . . . . . . 189
classification_colorscheme. . . . . . . . . . . . . . . . . . 72

classname . . . . . . . . . . . . . . . . . . . . . . . . . . 117, 219, 232 CLAYER. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 clip_boundary . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 clip_boundary_type . . . . . . . . . . . . . . . . . . . . . . . 49, 95 clip_inverts. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 clip_mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49, 95 clip_verts . . . . . . . . . . . . . . . . . . . . . . . . 33, 49, 96, 199 clipping . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49, 95, 169 clippings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 clone_ins_pt . . 29, 34, 35, 36, 38, 39, 40, 41, 51, 234 cloning . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 160, 161, 218 close_to_axis . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 closed . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 238 closed_b. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47, 81 closed_surfaces . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 CMATERIAL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 CMLJUST . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 CMLSCALE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 CMLSTYLE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

code . . . . . . . . . . . . . . . . . . . . . . . . . . . 166, 228, 229, 235
codepage . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21
codes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 145, 150
col_gutter . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 245
col_sizes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 245
col_spacing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50, 60
col_type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 245
col_width . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 245
col_widths . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85
color … 27, 158, 173, 199, 201, 220, 237, 240, 244,

245, 250, 253, 260

color_bleed_scale . . . . . . . . . . . . . . . . . . . . . . . . . . . 177 color_bottom . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 170 color_far . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 color_middle . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 170 color_near . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 color_r11 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 260 color_sky_horizon . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 color_sky_zenith . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 color_top . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 170

                                                color_underground_azimuth. . . . . . . . . . . . . . . . . . 171
                                                color_underground_horizon. . . . . . . . . . . . . . . . . . 171

color1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 242 color2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 242 colorramps . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 189 colors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46, 61 colorscheme. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 231 cols . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 158, 242 colspacing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 column_heights. . . . . . . . . . . . . . . . . . . . . . . . . . . 63, 186 column_offset . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 132 column_type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63, 186 column_width . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63, 186 combined_state . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 134 COMMENTS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 264 comp_data_size . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 250 complex_shapecode . . . . . . . . . . . . . . . . . . . . . . . . . . . 241 compression_type . . . . . . . . . . . . . . . . . . . . . . . . . . . . 250 conn_pts. . . . . . . . . 132, 134, 141, 145, 147, 149, 228 connection_line_color . . . . . . . . . . . . . . . . . . . . . . 160 connection_linewt . . . . . . . . . . . . . . . . . . . . . . . . . . . 160 connection_ltype . . . . . . . . . . . . . . . . . . . . . . . . . . . . 160 connection_string . . . . . . . . . . . . . . . . . . . . . . . . . . . 157 connections . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 229, 230 const_width . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58, 248 constraint_type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 content . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 243 content_base . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 243 content_color. . . . . . . . . . . . . . . . . . . . . . . . . . . 232, 255 content_format . . . . . . . . . . . . . . . . . . . . . . . . . 231, 258 content_height . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 230 content_layout . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 231 content_type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 182 content_width . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 230 context . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 185 CONTEXTDATAMANAGER. . . . . . . . . . . . . . . . . . . . . . . . . . 157 contrast . . . . . . . . . . . . . 33, 49, 94, 95, 169, 207, 215 control_points . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 239 controlled_objdep . . . . . . . . . . . . . . . . . . . . . . . . . . . 259 coord_proj_radius . . . . . . . . . . . . . . . . . . . . . . . . . . . 168 coord_system_datum. . . . . . . . . . . . . . . . . . . . . . . . . . 168 coord_system_def . . . . . . . . . . . . . . . . . . . . . . . . . . . . 168 coord_system_wkt . . . . . . . . . . . . . . . . . . . . . . . . . . . . 168 coord_type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 167 COORDS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 corner_decel . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 185 corner1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24, 80, 92 corner2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24, 80, 92 corner3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24, 80, 92 corner4 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24, 80, 92 CPSNID. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 crc32. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 250 crease. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 crop_plane . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 247 crop_x_dir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 247 crop_y_dir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 247 croppings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 cross_sections . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 crosssects . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102 CSACDOCUMENTOPTIONS . . . . . . . . . . . . . . . . . . . . . . . . 157 CSHADOW . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 ctrl_pts. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48, 81 ctrl_tol. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47, 81 ctx . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 cur_colorscheme . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 curr_type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 198 curve_type . . . . . . . . . . . . . . . . . . . . . . . . 73, 74, 75, 238 CURVEPATH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 157 custom_data . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 258, 259 customdata . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 158, 257 customdata_items . . . . . . . . . . . . . . . . . . . . . . . 257, 259 cv_hull_display . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67

D

d_node . . . . . . . . . . . . . . . . . . . . . . . . . 108, 110, 120, 125 dashes. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 176, 238 dashes_r11 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 176 data . . . . . . . . . . . . . . . . . . . . . . . . . 68, 76, 191, 206, 223 data_3dpoint . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 257 data_adapter . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 157 data_algn_offset . . . . . . . . . . . . . . . . . . . . . . . . . . . . 228 data_date . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 256 data_double. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 256 data_flags . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 230 data_handle. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 257

data_horiz_bottom_color . . . . . . . . . . . . . . . . . . . . . 87
data_horiz_bottom_linewt . . . . . . . . . . . . . . . . . . . . 88
data_horiz_bottom_visibility . . . . . . . . . . . . . . . 89

data_horiz_ins_color . . . . . . . . . . . . . . . . . . . . . . . . 87 data_horiz_ins_linewt . . . . . . . . . . . . . . . . . . . . . . . 88

data_horiz_ins_visibility. . . . . . . . . . . . . . . . . . . 89

data_horiz_top_color . . . . . . . . . . . . . . . . . . . . . . . . 87 data_horiz_top_linewt . . . . . . . . . . . . . . . . . . . . . . . 88

data_horiz_top_visibility. . . . . . . . . . . . . . . . . . . 89

data_link . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 257

data_long . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 256
data_numbits . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76, 191
data_point . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 257
data_row_alignment . . . . . . . . . . . . . . . . . . . . . . . . . . . 86
data_row_color . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86
data_row_fill_color. . . . . . . . . . . . . . . . . . . . . . . . . . 86
data_row_fill_none . . . . . . . . . . . . . . . . . . . . . . . . . . . 86
data_row_height . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86

                                                data_row_style_override . . . . . . . . . . . . . . . . . . . . . 90

data_size. . . . . . . . . . 68, 76, 191, 206, 224, 225, 256
data_string. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 256
data_text_style . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86
data_type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 254, 256
data_vert_ins_color. . . . . . . . . . . . . . . . . . . . . . . . . . 87
data_vert_ins_linewt . . . . . . . . . . . . . . . . . . . . . . . . 89

                                                data_vert_ins_visibility . . . . . . . . . . . . . . . . . . . . 90

data_vert_left_color . . . . . . . . . . . . . . . . . . . . . . . . 87
data_vert_left_linewt . . . . . . . . . . . . . . . . . . . . . . . 88

                                                data_vert_left_visibility. . . . . . . . . . . . . . . . . . . 90

data_vert_right_color . . . . . . . . . . . . . . . . . . . . . . . 88
data_vert_right_linewt . . . . . . . . . . . . . . . . . . . . . . 89

                                                data_vert_right_visibility. . . . . . . . . . . . . . . . . . 90

dataflags . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30, 31, 91
DATALINK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 157
DATATABLE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 158
dates. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 202
datidx . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 223
datidx_segidx . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 223
day . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 158
DBCOLOR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 158
decomp_data_size . . . . . . . . . . . . . . . . . . . . . . . . . . . . 250
def_base_angle_pt . . . . . . . . . . . . . . . . . . . . . . . . . . . 148
def_basept. . . . . . 128, 129, 130, 133, 135, 136, 137,

139, 143, 146, 148, 151, 154

                                                def_classification_colorscheme . . . . . . . . . . . . 189
                                                def_elevation_colorscheme. . . . . . . . . . . . . . . . . . 189

def_endpt . . 128, 129, 130, 133, 135, 136, 137, 139,

143, 146, 148, 151, 154

                                                def_intensity_colorscheme. . . . . . . . . . . . . . . . . . 189

def_label_pt . . . . . . . . . . . . . . . . . . . . . . . . . . . . 136, 142 def_pt . . 28, 33, 34, 36, 37, 38, 39, 40, 50, 132, 141,

142, 150, 152, 165, 233, 246

default_flag. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 default_lighting_type . . . . . . . . . . . . . . . . . . . . . . . 94 default_lightning_type . . . . . . . . . . . . . . . . 207, 215 default_text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 244 default_value . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 defaultid . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 161 definition . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 definition_id . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 deflines. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47, 62 degree . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47, 80, 239 DELOBJ. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 dep . . . . . . . . . . . . . . . . . . . 111, 114, 116, 119, 127, 221 dep_body . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 dep_on . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 depbodyid . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 dependency . . . . . . 128, 131, 133, 137, 138, 146, 151 dependent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 147, 149

dependent_on_compound_object . . . . . . . . . . . . . . 117

depid. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 221 deps … 108, 109, 118, 126, 131, 134, 140, 141, 145,

147, 149, 222

desc . . . . . . . . . . . . . . . . . . . . . . . . . . . 127, 158, 196, 229 description. . . . . 156, 157, 176, 177, 179, 182, 184,

192, 195, 202, 208, 242

design_pt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 167 dest_pt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 237 destblock . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 251 destfile . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 251 destination. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 195 DETAILVIEWSTYLE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 158 DGNDEFINITION . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 218 DGNFRAME . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 DGNUNDERLAY . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 di_unknown . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 225 diagnostics_bsp_mode . . . . . . . . . . . . . . . . . . . . . . . 181 diagnostics_grid_float . . . . . . . . . . . . . . . . . . . . . 181 diagnostics_grid_mode . . . . . . . . . . . . . . . . . . . . . . 181 diagnostics_mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . 181

diagnostics_photon_mode . . . . . . . . . . . . . . . . . . . . 181
diagnostics_samples_mode. . . . . . . . . . . . . . . . . . . 181

DICTIONARY . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 160 DICTIONARY_ACAD_GROUP . . . . . . . . . . . . . . . . . . . . . . . . 9

DICTIONARY_ACAD_MLINESTYLE. . . . . . . . . . . . . . . . . . . 9

DICTIONARY_COLOR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 DICTIONARY_LAYOUT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 DICTIONARY_LIGHTLIST. . . . . . . . . . . . . . . . . . . . . . . . . . 9 DICTIONARY_MATERIAL. . . . . . . . . . . . . . . . . . . . . . . . . . . 9

DICTIONARY_NAMED_OBJECT . . . . . . . . . . . . . . . . . . . . . . 9
DICTIONARY_PLOTSETTINGS . . . . . . . . . . . . . . . . . . . . . . 9
DICTIONARY_PLOTSTYLENAME . . . . . . . . . . . . . . . . . . . . . 9

DICTIONARY_VISUALSTYLE . . . . . . . . . . . . . . . . . . . . . . . 9 DICTIONARYVAR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 160 DICTIONARYWDFLT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 160 diffuse_color . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 177 diffusemap . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 177 dim_rotation. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 DIMADEC. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10, 164 DIMALT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10, 162 DIMALTD. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10, 162 DIMALTF. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10, 163 DIMALTMZF . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10, 164 DIMALTMZS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10, 164 DIMALTRND . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10, 163 DIMALTTD . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10, 162 DIMALTTZ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10, 162 DIMALTU. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10, 162 DIMALTZ. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10, 162 DIMAPOST . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10, 163 DIMAPOST_T . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 DIMARCSYM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10, 163 DIMARROW . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 DIMASO. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 DIMASSOC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10, 161 dimasz. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 DIMASZ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10, 162 dimatfit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 246 DIMATFIT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10, 164 DIMAUNIT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10, 162 DIMAZIN. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10, 163 dimbase_version . . . . . . . . . . . . . . . . . . . . . . . . 114, 142 DIMBLK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10, 164 DIMBLK_T . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10, 163 DIMBLK1. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10, 164 DIMBLK1_T . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10, 163 DIMBLK2. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10, 164 DIMBLK2_T . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10, 163 DIMCEN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10, 163 DIMCLRD. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11, 163 DIMCLRD_C . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 DIMCLRD_N . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 163 DIMCLRE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11, 164 DIMCLRE_C . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 DIMCLRE_N . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 163 DIMCLRT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11, 164 DIMCLRT_C . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 DIMCLRT_N . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 163 DIMDEC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11, 162 DIMDLE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11, 162 DIMDLI . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11, 162 DIMDSEP. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11, 164 dimension. . . . . . . . . . . . . . . . . . . . . . 106, 165, 187, 191 dimension_x. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 193 dimension_y. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 193 DIMENSION_ALIGNED . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 DIMENSION_ANG2LN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 DIMENSION_ANG3PT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 DIMENSION_DIAMETER . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 DIMENSION_LINEAR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 DIMENSION_ORDINATE . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 DIMENSION_RADIUS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 dimensionobj . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 161 DIMEXE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11, 162 DIMEXO . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11, 162 DIMFIT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11, 162 DIMFRAC. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11, 164 DIMFXL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11, 163 DIMFXLON . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11, 164 dimgap . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52, 92 DIMGAP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11, 163 DIMJOGANG . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11, 163 DIMJUST. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11, 162 DIMLDRBLK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11, 164 DIMLFAC. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11, 163 DIMLIM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11, 161 dimline_pt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106 DIMLTEX1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11, 164 DIMLTEX2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11, 164 DIMLTYPE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11, 164 DIMLUNIT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11, 164 DIMLWD . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11, 164 DIMLWE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11, 164 DIMMZF . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11, 164 DIMMZS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12, 164 dimosxd . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 246 DIMPOST. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12, 163 DIMPOST_T . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 dimref . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 156 DIMRND . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12, 162 DIMSAH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12, 162 DIMSAV. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 DIMSCALE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12, 162 DIMSD1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12, 162 DIMSD2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12, 162 DIMSE1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12, 162 DIMSE2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12, 162 DIMSHO. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 DIMSOXD. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12, 162 dimstyle . . 29, 34, 35, 36, 38, 39, 40, 41, 51, 53, 92,

234

DIMSTYLE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12, 161 DIMSTYLE_CONTROL . . . . . . . . . . . . . . . . . . . . . . . . . . . . 164

DIMSTYLE_CONTROL_OBJECT . . . . . . . . . . . . . . . . . . . . . 12

DIMTAD . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12, 162 DIMTDEC. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12, 162 DIMTFAC. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12, 163 DIMTFILL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12, 163 DIMTFILLCLR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12, 163 DIMTIH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12, 161 dimtix . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 246 DIMTIX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12, 162 dimtmove . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 246 DIMTM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12, 163 DIMTMOVE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12, 164 dimtofl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 246 DIMTOFL. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12, 162 DIMTOH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12, 162 DIMTOL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12, 161 DIMTOLJ. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12, 162 DIMTP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12, 163 DIMTSZ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12, 163 DIMTVP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12, 163 DIMTXSTY . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13, 164 DIMTXT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13, 163 DIMTXTDIRECTION . . . . . . . . . . . . . . . . . . . . . . . . . 13, 164 DIMTZIN. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13, 162 DIMUNIT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13, 162 DIMUPT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13, 162 DIMZIN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13, 162 direction. . . . . . . . . . . . . . . . . . . . . . 100, 103, 104, 244 displacement . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 221 display_boundary_on . . . . . . . . . . . . . . . . . . . . . . . . 199 display_brightness. . . . . . . . . . . . . . . . . . . . . . . . . . 211 display_brightness_bl . . . . . . . . . . . . . . . . . . . . . . 211 display_brightness_int . . . . . . . . . . . . . . . . . . . . . 211 display_frame . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 218 display_image . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 display_index . . . . . . . . . . . . . . . . . 179, 192, 194, 195 display_location. … 131, 134, 140, 141, 144, 147,

149

display_name . . . . . . . . . . . . . . . . . . . . . . . . . . . . 159, 196 display_props . . . . . . . . . . . . . . . . . . . . . . . . 48, 95, 169 display_settings . . . . . . . . . . . . . . . . . . . . . . . . . . . . 211 display_settings_int . . . . . . . . . . . . . . . . . . . . . . . 211 display_shadow_type . . . . . . . . . . . . . . . . . . . . . . . . 211

display_shadow_type_int . . . . . . . . . . . . . . . . . . . . 212

DISPSILH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 dist_center. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 230 dist_top_left . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 230 distance . . . . . . . . . . . . . . . . . . . . . . . 126, 133, 139, 146 distance_desc. . . . . . . . . . . . . . . . . . . . . . . . . . . 139, 144 distance_name. . . . . . . . . . . . . . . . . . . . . . . . . . . 139, 144 distance_value_set. . . . . . . . . . . . . . . . . . . . . . . . . . 144 dlevel. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 DMDIMOBJECTCONTEXTDATA . . . . . . . . . . . . . . . . . . . . . 165 do_sea_level_corr . . . . . . . . . . . . . . . . . . . . . . . . . . . 168 dogleg_length . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 241 dogleg_vector . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 241 double . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6, 7 double[3] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 double_flag . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47, 62 double_line_spacing . . . . . . . . . . . . . . . . . . . . . . . . 237 draft_angle . . . . . . . . . . . . . 44, 78, 83, 100, 103, 104 draft_end_distance . . . . . . . . . . . . . . . . . . . 44, 78, 83 draft_start_distance . . . . . . . . . . . . . . . . . 44, 78, 83 DRAGMODE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 DRAGVS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 drawing_units. . . . . . . . . . . . . . . . . . . . . . . . . . . 188, 196 ds_version . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 223, 228 DUMMY. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 165 DWFDEFINITION . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 218 DWFFRAME . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 DWFUNDERLAY . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 dwg . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 260, 262 Dwg_3DSOLID_material . . . . . . . . . . . . . . . . . . . . . . . 219 Dwg_3DSOLID_silhouette . . . . . . . . . . . . . . . . . . . . . 219 Dwg_3DSOLID_wire . . . . . . . . . . . . . . . . . . . . . . . . . . . . 220 Dwg_AcDs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 223 Dwg_AcDs_Data . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 224 Dwg_AcDs_Data_Record . . . . . . . . . . . . . . . . . . . . . . . 225

Dwg_AcDs_Data_RecordHdr . . . . . . . . . . . . . . . . . . . . 225

Dwg_AcDs_DataBlob . . . . . . . . . . . . . . . . . . . . . . . . . . . 224 Dwg_AcDs_DataBlob01 . . . . . . . . . . . . . . . . . . . . . . . . 224 Dwg_AcDs_DataBlobRef . . . . . . . . . . . . . . . . . . . . . . . 224

Dwg_AcDs_DataBlobRef_Page. . . . . . . . . . . . . . . . . . 225

Dwg_AcDs_DataIndex. . . . . . . . . . . . . . . . . . . . . . . . . . 225

Dwg_AcDs_DataIndex_Entry. . . . . . . . . . . . . . . . . . . 225

Dwg_AcDs_Schema . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 225 Dwg_AcDs_Schema_Prop . . . . . . . . . . . . . . . . . . . . . . . 227 Dwg_AcDs_SchemaData . . . . . . . . . . . . . . . . . . . . . . . . 226

Dwg_AcDs_SchemaData_UProp. . . . . . . . . . . . . . . . . . 226

Dwg_AcDs_SchemaIndex . . . . . . . . . . . . . . . . . . . . . . . 226

Dwg_AcDs_SchemaIndex_Prop. . . . . . . . . . . . . . . . . . 226

Dwg_AcDs_Search . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 227 Dwg_AcDs_Search_Data . . . . . . . . . . . . . . . . . . . . . . . 227 Dwg_AcDs_Search_IdIdx . . . . . . . . . . . . . . . . . . . . . . 227 Dwg_AcDs_Search_IdIdxs . . . . . . . . . . . . . . . . . . . . . 227 Dwg_AcDs_Segment . . . . . . . . . . . . . . . . . . . . . . . . . . . . 228 Dwg_AcDs_SegmentIndex . . . . . . . . . . . . . . . . . . . . . . 228 Dwg_ACSH_HistoryNode . . . . . . . . . . . . . . . . . . . . . . . 220 Dwg_ACSH_SubentColor . . . . . . . . . . . . . . . . . . . . . . . 220

Dwg_ACSH_SubentMaterial . . . . . . . . . . . . . . . . . . . . 221

Dwg_ACTIONBODY . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 221 Dwg_ARRAYITEMLOCATOR . . . . . . . . . . . . . . . . . . . . . . . 221 Dwg_ASSOCACTION_Deps . . . . . . . . . . . . . . . . . . . . . . . 221

Dwg_ASSOCACTIONBODY_action . . . . . . . . . . . . . . . . 221

Dwg_ASSOCARRAYITEM. . . . . . . . . . . . . . . . . . . . . . . . . . 221

Dwg_ASSOCPARAMBASEDACTIONBODY . . . . . . . . . . . . . 222
Dwg_ASSOCSURFACEACTIONBODY . . . . . . . . . . . . . . . . 222
Dwg_BLOCKACTION_connectionpts . . . . . . . . . . . . . 228
Dwg_BLOCKLOOKUPACTION_lut. . . . . . . . . . . . . . . . . . 228
Dwg_BLOCKPARAMETER_connection . . . . . . . . . . . . . 229
Dwg_BLOCKPARAMETER_PropInfo . . . . . . . . . . . . . . . 228

Dwg_BLOCKPARAMVALUESET . . . . . . . . . . . . . . . . . . . . . 229

Dwg_BLOCKVISIBILITYPARAMETER_state . . . . . . . 229
Dwg_CellContentGeometry . . . . . . . . . . . . . . . . . . . . 230

Dwg_CellStyle . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 230 Dwg_ColorRamp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 231 Dwg_COMPOUNDOBJECTID . . . . . . . . . . . . . . . . . . . . . . . 229

Dwg_CONSTRAINTGROUPNODE . . . . . . . . . . . . . . . . . . . . 229

Dwg_ContentFormat . . . . . . . . . . . . . . . . . . . . . . . . . . . 231 Dwg_CONTEXTDATA_dict . . . . . . . . . . . . . . . . . . . . . . . 230 Dwg_CONTEXTDATA_submgr . . . . . . . . . . . . . . . . . . . . . 230

Dwg_DATALINK_customdata . . . . . . . . . . . . . . . . . . . . 232

Dwg_DATATABLE_column . . . . . . . . . . . . . . . . . . . . . . . 232 Dwg_DATATABLE_row . . . . . . . . . . . . . . . . . . . . . . . . . . . 232 Dwg_DIMASSOC_Ref . . . . . . . . . . . . . . . . . . . . . . . . . . . . 232 Dwg_DIMENSION_common . . . . . . . . . . . . . . . . . . . . . . . 233 Dwg_EVAL_Edge . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 234 Dwg_EVAL_Node . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 234 Dwg_EvalExpr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 235 Dwg_EvalVariant . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 235 Dwg_FIELD_ChildValue . . . . . . . . . . . . . . . . . . . . . . . 236 Dwg_FileDepList_Files . . . . . . . . . . . . . . . . . . . . . . 236 Dwg_FormattedTableData . . . . . . . . . . . . . . . . . . . . . 236

Dwg_FormattedTableMerged. . . . . . . . . . . . . . . . . . . 236

Dwg_GEODATA_meshface . . . . . . . . . . . . . . . . . . . . . . . 237 Dwg_GEODATA_meshpt. . . . . . . . . . . . . . . . . . . . . . . . . . 237 Dwg_GridFormat . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 237 Dwg_HATCH_Color . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 237 Dwg_HATCH_ControlPoint . . . . . . . . . . . . . . . . . . . . . 237 Dwg_HATCH_DefLine . . . . . . . . . . . . . . . . . . . . . . . . . . . 237 Dwg_HATCH_Path . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 238 Dwg_HATCH_PathSeg . . . . . . . . . . . . . . . . . . . . . . . . . . . 238 Dwg_HATCH_PolylinePath . . . . . . . . . . . . . . . . . . . . . 239 Dwg_LAYER_entry . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 239 Dwg_LEADER_ArrowHead . . . . . . . . . . . . . . . . . . . . . . . 239 Dwg_LEADER_BlockLabel . . . . . . . . . . . . . . . . . . . . . . 239 Dwg_LEADER_Break . . . . . . . . . . . . . . . . . . . . . . . . . . . . 240 Dwg_LEADER_Line . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 240 Dwg_LEADER_Node . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 240 Dwg_LIGHTLIST_light . . . . . . . . . . . . . . . . . . . . . . . . 241 Dwg_LinkedData . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 241 Dwg_LinkedTableData . . . . . . . . . . . . . . . . . . . . . . . . 242 Dwg_LTYPE_dash . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 241 Dwg_LWPOLYLINE_width . . . . . . . . . . . . . . . . . . . . . . . 241 Dwg_MATERIAL_color. . . . . . . . . . . . . . . . . . . . . . . . . . 242

                                                Dwg_MATERIAL_gentexture . . . . . . . . . . . . . . . . . . . . 242

Dwg_MATERIAL_mapper . . . . . . . . . . . . . . . . . . . . . . . . 242 Dwg_MESH_edge . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 243

                                                Dwg_MLEADER_AnnotContext. . . . . . . . . . . . . . . . . . . 243

Dwg_MLEADER_Content . . . . . . . . . . . . . . . . . . . . . . . . 260

                                                Dwg_MLEADER_Content_Block. . . . . . . . . . . . . . . . . . 244
                                                Dwg_MLEADER_Content_MText. . . . . . . . . . . . . . . . . . 244

Dwg_MLINE_line . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 245 Dwg_MLINE_vertex . . . . . . . . . . . . . . . . . . . . . . . . . . . . 246 Dwg_MLINESTYLE_line . . . . . . . . . . . . . . . . . . . . . . . . 245 Dwg_OCD_Dimension . . . . . . . . . . . . . . . . . . . . . . . . . . . 246

                                                Dwg_PARTIAL_VIEWING_INDEX_Entry. . . . . . . . . . . 247
                                                Dwg_POINTCLOUD_Clippings. . . . . . . . . . . . . . . . . . . 247
                                                Dwg_POINTCLOUD_IntensityStyle . . . . . . . . . . . . . 248
                                                Dwg_POINTCLOUDCOLORMAP_Ramp . . . . . . . . . . . . . . . 247
                                                Dwg_POINTCLOUDEX_Croppings . . . . . . . . . . . . . . . . 247

Dwg_PROXY_LWPOLYLINE . . . . . . . . . . . . . . . . . . . . . . . 248 Dwg_R2004_Header . . . . . . . . . . . . . . . . . . . . . . . . . . . . 249

                                                Dwg_SECTION_geometrysettings . . . . . . . . . . . . . . 250
                                                Dwg_SECTION_typesettings. . . . . . . . . . . . . . . . . . . 251

dwg_size . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21

                                                Dwg_SPLINE_control_point. . . . . . . . . . . . . . . . . . . 251
                                                Dwg_SummaryInfo_Property. . . . . . . . . . . . . . . . . . . 251

Dwg_SUNSTUDY_Dates. . . . . . . . . . . . . . . . . . . . . . . . . . 251 Dwg_TABLE_AttrDef . . . . . . . . . . . . . . . . . . . . . . . . . . . 254 Dwg_TABLE_BreakHeight . . . . . . . . . . . . . . . . . . . . . . 254 Dwg_TABLE_BreakRow. . . . . . . . . . . . . . . . . . . . . . . . . . 254 Dwg_TABLE_Cell . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 254

                                                Dwg_TABLE_CustomDataItem. . . . . . . . . . . . . . . . . . . 256

Dwg_TABLE_value . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 256 Dwg_TableCell . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 257 Dwg_TableCellContent . . . . . . . . . . . . . . . . . . . . . . . 258

                                                Dwg_TableCellContent_Attr. . . . . . . . . . . . . . . . . . 258

Dwg_TableDataColumn . . . . . . . . . . . . . . . . . . . . . . . . 258 Dwg_TABLEGEOMETRY_Cell . . . . . . . . . . . . . . . . . . . . . 251 Dwg_TableRow . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 259 Dwg_TABLESTYLE_border . . . . . . . . . . . . . . . . . . . . . . 253

                                                Dwg_TABLESTYLE_CellStyle. . . . . . . . . . . . . . . . . . . 252
                                                Dwg_TABLESTYLE_rowstyles. . . . . . . . . . . . . . . . . . . 253

Dwg_UCS_orthopts . . . . . . . . . . . . . . . . . . . . . . . . . . . . 259 Dwg_VALUEPARAM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 259 Dwg_VALUEPARAM_vars . . . . . . . . . . . . . . . . . . . . . . . . 260

                                                DWG_ERR_CLASSESNOTFOUND . . . . . . . . . . . . . . . . . . . . 277

DWG_ERR_INTERNALERROR . . . . . . . . . . . . . . . . . . . . . . 277 DWG_ERR_INVALIDDWG. . . . . . . . . . . . . . . . . . . . . . . . . . 277 DWG_ERR_INVALIDEED. . . . . . . . . . . . . . . . . . . . . . . . . . 277 DWG_ERR_INVALIDHANDLE . . . . . . . . . . . . . . . . . . . . . . 277 DWG_ERR_INVALIDTYPE . . . . . . . . . . . . . . . . . . . . . . . . 277 DWG_ERR_IOERROR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 277

                                                DWG_ERR_NOTYETSUPPORTED . . . . . . . . . . . . . . . . . . . . 277

DWG_ERR_OUTOFMEM . . . . . . . . . . . . . . . . . . . . . . . . . . . . 277 DWG_ERR_PAGENOTFOUND . . . . . . . . . . . . . . . . . . . . . . . 277

                                                DWG_ERR_SECTIONNOTFOUND . . . . . . . . . . . . . . . . . . . . 277

DWG_ERR_UNHANDLEDCLASS . . . . . . . . . . . . . . . . . . . . . 277

                                                DWG_ERR_VALUEOUTOFBOUNDS. . . . . . . . . . . . . . . . . . . 277

DWG_ERR_WRONGCRC . . . . . . . . . . . . . . . . . . . . . . . . . . . . 277 DWGCODEPAGE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 DYNAMICBLOCKPROXYNODE . . . . . . . . . . . . . . . . . . . . . . 165

DYNAMICBLOCKPURGEPREVENTER . . . . . . . . . . . . . . . . 165

E

e1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 234 e2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 234 e3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 234 edge_color . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 210 edge_color_int . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 210 edge_crease_angle . . . . . . . . . . . . . . . . . . . . . . . . . . . 210 edge_crease_angle_int . . . . . . . . . . . . . . . . . . . . . . 210 edge_do_hide_precision . . . . . . . . . . . . . . . . . . . . . 211

edge_do_hide_precision_int . . . . . . . . . . . . . . . . 211

edge_flags . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 235 edge_halo_gap . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 211 edge_halo_gap_int . . . . . . . . . . . . . . . . . . . . . . . . . . . 211

edge_intersection_color . . . . . . . . . . . . . . . . . . . . 209
edge_intersection_color_int . . . . . . . . . . . . . . . 209
edge_intersection_ltype . . . . . . . . . . . . . . . . . . . . 210
edge_intersection_ltype_int . . . . . . . . . . . . . . . 210

edge_isolines . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 211 edge_isolines_int . . . . . . . . . . . . . . . . . . . . . . . . . . . 211 edge_jitter. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 210 edge_jitter_int . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 211 edge_model . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 209 edge_model_int . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 209 edge_modifier . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 210 edge_modifier_int . . . . . . . . . . . . . . . . . . . . . . . . . . . 210 edge_obscured_color . . . . . . . . . . . . . . . . . . . . . . . . 210

edge_obscured_color_int . . . . . . . . . . . . . . . . . . . . 210

edge_obscured_ltype . . . . . . . . . . . . . . . . . . . . . . . . 210

edge_obscured_ltype_int . . . . . . . . . . . . . . . . . . . . 210

edge_opacity . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 210 edge_opacity_int . . . . . . . . . . . . . . . . . . . . . . . . . . . . 210 edge_overhang . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 210 edge_overhang_int . . . . . . . . . . . . . . . . . . . . . . . . . . . 210 edge_silhouette_color . . . . . . . . . . . . . . . . . . . . . . 211

edge_silhouette_color_int. . . . . . . . . . . . . . . . . . 211

edge_silhouette_width . . . . . . . . . . . . . . . . . . . . . . 211

edge_silhouette_width_int. . . . . . . . . . . . . . . . . . 211

edge_style . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 209 edge_style_apply . . . . . . . . . . . . . . . . . . . . . . . . . . . . 211 edge_style_apply_int . . . . . . . . . . . . . . . . . . . . . . . 211 edge_style_int . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 209 edge_transparency . . . . . . . . . . . . . . . . . . . . . . . . . . . 250 edge_visualstyle . . . . . . . . . . . . . . . . . . . . . . . . . . . . 260 edge_width . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 210 edge_width_int . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 210 edge_wiggle. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 214 edge_wiggle_int . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 214 edges. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59, 99, 101, 166 eed . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 260, 262 eed1071 …. 128, 129, 130, 131, 132, 133, 134, 135, 136, 137, 138, 139, 140, 141, 142, 143, 144, 145, 147,

148, 149, 150, 151, 152, 153

elevation …. 28, 29, 31, 33, 34, 36, 37, 38, 39, 41,

46, 50, 58, 61, 74, 80, 91, 92, 233, 248

elevation_apply_to_fixed_range . . . . . . . . . . . . . 72

elevation_as_gradient . . . . . . . . . . . . . . . . . . . . . . . 73 elevation_max . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 elevation_min . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72

elevation_out_of_range_behavior . . . . . . . . . . . . 72

elevation_r11 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 260 ELEVATION . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 ELLIPSE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 enable . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 enable_context . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 185 enable_frame_text . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 enabled . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 117 encr_sat_data . . 25, 42, 56, 66, 68, 76, 81, 97, 111 end . . . . . . . . . . . . . . . . . . . . . . . . . . 24, 55, 240, 241, 254 end_angle . . . . . . . . . . . . . . . . . . . . 26, 27, 42, 184, 238 end_draft_angle . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 end_draft_dist . . . . . . . . . . . . . . . . . . . . . . . . . 100, 104 end_draft_magnitude. . . . . . . . . . . . . . . . . . . . . . . . . . 57 end_line_length . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 197 end_line_overshoot. . . . . . . . . . . . . . . . . . . . . . . . . . 197 end_marker . . . . . . 26, 44, 57, 67, 69, 77, 82, 98, 112 end_pt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 131 end_tan_vec . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47, 81 end_tangent. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 239 end_time . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 202 end_width . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73, 74, 92 endblk_entity . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 156 ENDBLK. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 ENDCAPS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 endpoint . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 238 endpt . . 128, 130, 131, 133, 135, 137, 138, 139, 144,

146, 148, 151, 154

endptproj . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52, 175 ENDREP. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 endsetbacks. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102 energy_multiplier . . . . . . . . . . . . . . . . . . . . . . . . . . . 181 entities . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 156 entity. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 124, 157 entmode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 260 entries …. 107, 154, 165, 174, 176, 187, 201, 206,

208, 217, 225, 230

entry_size . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 225 ents. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 199 environ_image_enabled . . . . . . . . 179, 192, 194, 195

environ_image_filename . . . . . . . 179, 192, 194, 195

evalexpr … 96, 97, 98, 99, 100, 101, 102, 103, 104, 105, 106, 128, 129, 130, 131, 132, 133, 134, 135, 136, 137, 138, 139, 140, 141, 142, 143, 144, 145, 147, 149,

150, 151, 152, 153, 165, 235

evaluation_error_code . . . . . . . . . . . . . . . . . . . . . . 167 evaluation_error_msg . . . . . . . . . . . . . . . . . . . . . . . 167 evaluation_option . . . . . . . . . . . . . . . . . . . . . . . . . . . 166 evaluation_status . . . . . . . . . . . . . . . . . . . . . . . . . . . 166 EVALUATION_GRAPH . . . . . . . . . . . . . . . . . . . . . . . . . . . . 165 evaluator . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 127 evaluatorid. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 221 explodable . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 156 export_mi_enabled . . . . . . . . . . . . . . . . . . . . . . . . . . . 181 expr. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 expr_description. … 128, 131, 133, 137, 138, 146,

151

expr_name . . . . . . . 128, 131, 133, 137, 138, 146, 151 expression . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 221 exprs. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 140 ext_lighting_model. . . . . . . . . . . . . . . . . . . . . . . . . . 208 extents_height. . . . . . . . . . . . . . . . . . . . . . . . . . . 62, 186 extents_max . . . . . . . . . . . . . . . . . . . . . . . . . 71, 190, 247 extents_min. . . . . . . . . . . . . . . . . . . . . . 70, 71, 190, 247 extents_width . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62, 186 extlight_length . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 extlight_radius . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 extlight_shape . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 extlight_width . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 EXTMAX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13, 174 EXTMIN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13, 174 EXTNAMES . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 extra_acis_data . . 25, 43, 56, 67, 69, 77, 82, 98, 112 extra_r11 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 260 extra_r11_size . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 extra_r11_text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 EXTRUDEDSURFACE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 extrusion …. 24, 26, 27, 28, 29, 31, 32, 33, 34, 36, 37, 38, 39, 40, 42, 46, 49, 50, 52, 55, 58, 60, 61, 62,

70, 74, 79, 80, 85, 91, 92, 199, 233, 248

F

face_color_mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 209 face_color_mode_int . . . . . . . . . . . . . . . . . . . . . . . . 209 face_lighting_model . . . . . . . . . . . . . . . . . . . . . . . . 208

face_lighting_model_int . . . . . . . . . . . . . . . . . . . . 209

face_lighting_quality . . . . . . . . . . . . . . . . . . . . . . 209

face_lighting_quality_int. . . . . . . . . . . . . . . . . . 209

face_modifier . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 209 face_modifier_int . . . . . . . . . . . . . . . . . . . . . . . . . . . 209 face_mono_color . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 209 face_mono_color_int . . . . . . . . . . . . . . . . . . . . . . . . 209 face_opacity . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 209 face_opacity_int . . . . . . . . . . . . . . . . . . . . . . . . . . . . 209 face_specular . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 209 face_specular_int . . . . . . . . . . . . . . . . . . . . . . . . . . . 209 face_transparency . . . . . . . . . . . . . . . . . . . . . . . . . . . 250 face_visualstyle . . . . . . . . . . . . . . . . . . . . . . . . . . . . 260 face1. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 237 face2. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 237 face3. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 237 faces. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 FACETRES . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 factor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 242 fade . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33, 49, 95, 169 falloff_angle . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 FASTZOOM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13, 216 FCFOBJECTCONTEXTDATA . . . . . . . . . . . . . . . . . . . . . . . 166 fdata . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84, 203 feature_index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 236 feature_location_pt . . . . . . . . . . . . . . . . . . . . . 40, 187 fg_ray_count . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 180

fg_sample_radius_state1 . . . . . . . . . . . . . . . . . . . . 180
fg_sample_radius_state2 . . . . . . . . . . . . . . . . . . . . 180
fg_sample_radius_state3 . . . . . . . . . . . . . . . . . . . . 181

fg_sample_radius1 . . . . . . . . . . . . . . . . . . . . . . . . . . . 181 fg_sample_radius2 . . . . . . . . . . . . . . . . . . . . . . . . . . . 181 field_length . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30, 31 field_refs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 242 field_state. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 166 FIELDLIST . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 167 fields . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 167 FIELD. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 166 file_header_size . . . . . . . . . . . . . . . . . . . . . . . . . . . . 223 file_ID_string . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 249 file_name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 file_path . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 172 file_signature . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 223 file_size . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 223 filename . . . . . . . . . . . . . . . . . . . . . . . 172, 218, 236, 242 filepath . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 236 filesize . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 236 filing_option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 166 fill_color . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 184, 253 FILLETRAD . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 FILLMODE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 filter_height . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 192 filter_type. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 192 filter_width . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 192

final_gathering_enabled . . . . . . . . . . . . . . . . . . . . 180

fingerprint. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 236 FINGERPRINTGUID . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 first_arc_pt . . . . . . . . . . . . . . . . . 38, 41, 51, 165, 191 first_attrib . . . . . . . . . . . . . . . . . . . . . . . . . . 49, 60, 90 first_endpoint . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 238 first_entity . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 156 first_nodeid . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 165 first_nodeid_copy . . . . . . . . . . . . . . . . . . . . . . . . . . . 165 first_seg_angle . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 182 first_vertex . . . . . . . . . . . . . . . . . . . . . . . . . . 73, 74, 75 fit_pts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48, 81 fit_to_screen . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 172 fit_tol . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47, 81 fitpts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 239 flag . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28, 33, 34, 36, 37, 38, 39, 41, 47, 50, 51, 58, 68, 73, 74, 75, 80, 92, 93, 107, 146, 150, 155, 161, 165, 173, 176, 184, 196,

200, 205, 206, 214, 217, 233, 238, 242, 254, 257 flag_3d. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21, 208 flag_for_table_value . . . . . . . . . . . . . . . . . . . . . . . . 85 flag_r11 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 260 flag0 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 164, 173 flag1 . . . . . . . . 28, 33, 34, 36, 37, 38, 39, 41, 50, 233 flag2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40, 156 flags …. 30, 31, 60, 64, 65, 79, 158, 159, 186, 196, 203, 205, 222, 226, 227, 229, 240, 248, 250, 254, 256 flags_r11 . . 107, 155, 165, 174, 177, 201, 206, 208,

217

FLAGS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 FLATLAND . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 flip_arrow1 . . . . . 28, 34, 35, 36, 38, 39, 40, 41, 51,

234, 247

flip_arrow2 . . . . . 28, 34, 35, 36, 38, 39, 40, 41, 51,

234, 246

flip_label . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135 flip_label_desc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135 FLIPGRIPENTITY . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 FLIPPARAMETERENTITY. . . . . . . . . . . . . . . . . . . . . . . . . . 45 flipped_state_label . . . . . . . . . . . . . . . . . . . . . . . . 135 flow. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 245 flow_dir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 flow_direction. . . . . . . . . . . . . . . . . . . . . . . . . . . 85, 203 flow_reversed . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63, 186

fog_background_enabled . . . . . . . 179, 192, 194, 195

fog_color_b. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 194 fog_color_g. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 194 fog_color_r. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 194 fog_density_far . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 194 fog_density_near . . . . . . . . . . . . . . . . . . . . . . . . . . . . 194 fog_distance_far . . . . . . . . . . . . . . . . . . . . . . . . . . . . 194 fog_distance_near . . . . . . . . . . . . . . . . . . . . . . . . . . . 194 fog_enabled . . . . . . . . . . . . . . . . . . . 179, 192, 194, 195 font . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 font_19 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 font_file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 201 format . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 166 format_flags . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 256 format_string. . . . . . . . . . . . . . . . . . . . . . . . . . . 254, 257 frame_rate . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 185 frames . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 185 from_dxf . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76, 191 front_clip_on . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 199 front_clip_z . . . . . . . . . . . . . . . . . . . 93, 199, 207, 215 FRONTZ. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 frozen . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173 frozen_in_new . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173 frozen_layers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 full_visualstyle . . . . . . . . . . . . . . . . . . . . . . . . . . . . 260

G

gap_array_size . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 250 generation . . . . . . . . . . . . . . . . . . . 30, 31, 91, 200, 251 genprocname . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 178, 242 genproctableend . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 178 genproctype. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 178 genprocvalbool . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 178 genprocvalcolor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 178 genprocvalint . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 178 genprocvalreal . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 178 genprocvaltext . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 178 gentextures. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 178 geo_rss_tag. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 168 GEODATA . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 167 geoimage_brightness . . . . . . . . . . . . . . . . . . . . . . . . 170 geoimage_contrast . . . . . . . . . . . . . . . . . . . . . . . . . . . 170 geoimage_fade . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 170 geoimage_height . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 170 geoimage_position . . . . . . . . . . . . . . . . . . . . . . . . . . . 170 geoimage_width . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 170 geom. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 251 geom_data_flag . . . . . . . . . . . . . . . . . . . . . . . . . 252, 257 geom_parent. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 230 geom_shader_usage . . . . . . . . . . . . . . . . . . . . . . . . . . . 205 GEOMAPIMAGE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 169 geomesh_faces . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 168 geomesh_pts. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 168 geometry . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 252, 258 geometry_status . . . . . . . . . . . . . . . 107, 109, 117, 126 GEOPOSITIONMARKER . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 gi_photons_per_light . . . . . . . . . . . . . . . . . . . . . . . 180 gi_sample_count . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 180 gi_sample_radius . . . . . . . . . . . . . . . . . . . . . . . . . . . . 180

gi_sample_radius_enabled. . . . . . . . . . . . . . . . . . . 180
global_illumination_enabled . . . . . . . . . . . . . . . 180

glyph_display_type . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 gradient_angle. . . . . . . . . . . . . . . . . . . . . . . . . . . . 46, 61 gradient_name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46, 61 gradient_shift. . . . . . . . . . . . . . . . . . . . . . . . . . . . 46, 61 gradient_tint . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46, 61 GRADIENT_BACKGROUND . . . . . . . . . . . . . . . . . . . . . . . . 170 grid_flags . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 216 grid_major . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94, 216 GRIDMODE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13, 216 GRIDUNIT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13, 94, 216 grip_expr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 136 grip_status. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 222 grip_type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 136

GROUND_PLANE_BACKGROUND . . . . . . . . . . . . . . . . . . . . 170

groups . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 GROUP. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 guide_curves. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 guides . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102 gutter . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63, 186

H

h_nodeid . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102 h1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108, 222 h2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 222 h330_2 . . . . . . . . . . . . . . . . . . . . . . . . . 113, 120, 121, 124 h330_3 . . . . . . . . . . . . . . . . . . . . . . . . . 113, 120, 121, 124 HALOGAP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 handedness . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 handle . . . . . . . . . . . 225, 227, 230, 239, 241, 258, 260 handleref . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 262 HANDLING . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 HANDSEED . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 hardowner . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 158 has_action . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 115 has_align_start . . . . . . . . . . . . . . . . . . . . . . . . 101, 105 has_arrow2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 246 has_attribs. . . . . . . . . . . . . . . . . . . . . . . . . . . . 49, 60, 85 has_bgcolor. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 253

has_border_color_overrides. . . . . . . . . . . . . . . . . . 87
has_border_lineweight_overrides . . . . . . . . . . . . 88

has_border_visibility_overrides . . . . . . . . . . . . 89

has_break_data . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 has_child_param . . . . . . . . . . . . . . . 113, 120, 121, 124 has_civil_data . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 169 has_content_blk . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 243

has_content_format_overrides . . . . . . . . . . . . . . 258

has_content_txt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 243 has_derived . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 has_dogleg . . . . . . . . . . . . . . . . . . . . . . . . . . 64, 182, 240 has_ds_data . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 260, 262 has_edge_visualstyle . . . . . . . . . . . . . . . . . . . . . . . 260 has_entries. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 187 has_face_visualstyle . . . . . . . . . . . . . . . . . . . . . . . 260 has_full_visualstyle . . . . . . . . . . . . . . . . . . . . . . . 260 has_geom_data . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 257 has_graph . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 166 has_h1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 222 has_landing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64, 182

has_lastleaderlinepoint . . . . . . . . . . . . . . . . . . . . 240

has_lastpt_ref . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 233 has_leader . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 has_linked_data . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 257 has_name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 has_no_flags. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 has_object . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 229 has_photometric_data . . . . . . . . . . . . . . . . . . . . . . . . 54 has_predefined. . . . . . . . . . . . . . . . . . . . . 179, 192, 195 has_reflection . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 220 has_revision_guid . . 26, 44, 57, 67, 70, 78, 83, 98,

112

has_rotation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 220 has_shadow . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 201 has_shear . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 220 has_strings_area . . . . . . . . . . . . . . . . . . . . . . . . . . . . 176 has_t78 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 127 has_table_overrides. . . . . . . . . . . . . . . . . . . . . . . . . . 85 has_target_grip . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 has_text_frame. . . . . . . . . . . . . . . . . . . . . . . . . . . 64, 183 has_vertex. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73, 74, 75 has_webfile . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 has_wires . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 219 hasattrs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 155 hatch_angle. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 251 hatch_angles . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 198 hatch_bg_color . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 198 hatch_color . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62, 198 hatch_pattern. . . . . . . . . . . . . . . . . . . . . . . . . . . 198, 251 hatch_scale . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 198, 251 hatch_spacing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 251 hatch_transparency. . . . . . . . . . . . . . . . . . . . . . . . . . 198 hatch_type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 250 HATCH. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 hdls . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 145, 150, 200 header_address . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 249

header_horiz_bottom_color. . . . . . . . . . . . . . . . . . . 87
header_horiz_bottom_linewt. . . . . . . . . . . . . . . . . . 88
header_horiz_bottom_visibility . . . . . . . . . . . . . 89

header_horiz_ins_color . . . . . . . . . . . . . . . . . . . . . . 87

                                                header_horiz_ins_linewt . . . . . . . . . . . . . . . . . . . . . 88
                                                header_horiz_ins_visibility . . . . . . . . . . . . . . . . 89

header_horiz_top_color . . . . . . . . . . . . . . . . . . . . . . 87

                                                header_horiz_top_linewt . . . . . . . . . . . . . . . . . . . . . 88
                                                header_horiz_top_visibility . . . . . . . . . . . . . . . . 89

header_row_alignment . . . . . . . . . . . . . . . . . . . . . . . . 86
header_row_color . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86
header_row_fill_color . . . . . . . . . . . . . . . . . . . . . . . 86
header_row_fill_none . . . . . . . . . . . . . . . . . . . . . . . . 86
header_row_height . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86

                                                header_row_style_override. . . . . . . . . . . . . . . . . . . 90

header_size. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 249
header_suppressed . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85
header_text_style . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86
header_vert_ins_color . . . . . . . . . . . . . . . . . . . . . . . 87
header_vert_ins_linewt . . . . . . . . . . . . . . . . . . . . . . 88

                                                header_vert_ins_visibility. . . . . . . . . . . . . . . . . . 89

header_vert_left_color . . . . . . . . . . . . . . . . . . . . . . 87

                                                header_vert_left_linewt . . . . . . . . . . . . . . . . . . . . . 88
                                                header_vert_left_visibility . . . . . . . . . . . . . . . . 89
                                                header_vert_right_color . . . . . . . . . . . . . . . . . . . . . 87
                                                header_vert_right_linewt . . . . . . . . . . . . . . . . . . . . 88
                                                header_vert_right_visibility . . . . . . . . . . . . . . . 89

height …. 29, 31, 79, 91, 93, 97, 99, 100, 103, 106,

170, 230, 244, 254, 259

height_w_gap . . . . . . . . . . . . . . . . . . . . . . . . . . . . 252, 258 HELIX. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 hexindex . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 250 HIDETEXT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 highlevel_info . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 195 history_id . . . . . . 26, 44, 57, 67, 70, 77, 83, 98, 112 history_node . . 96, 97, 99, 100, 101, 102, 103, 104,

105, 106

hookline_dir. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 hookline_on . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 hor_dir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 horiz_alignment . . . . . . . . . . . . . . . . . . . . . . . 30, 31, 91 horiz_cell_margin . . . . . . . . . . . . . . . . . . . . . . . 85, 203 horiz_dir… 28, 33, 35, 36, 37, 38, 40, 41, 50, 166,

234

horiz_direction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 horiz_margin . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 231 horizon . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 170 horizontal_mode . . . . . . . . . . . . . . . . . . . . . . . . 185, 204 host_block . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 167

                                                host_drawing_visibility . . . . . . . . . . . . . . . . . . . . 186

hotspot_angle . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 hour. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 158 hours. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 202 HYPERLINKBASE . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14, 264

I

IBL_BACKGROUND . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 id . . . . . . . . . . . . . . . . . . . 78, 92, 93, 166, 234, 235, 252 IDBUFFER . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 identifier_color . . . . . . . . . . . . . . . . . . . . . . . 159, 196

identifier_exclude_characters . . . . . . . . 159, 197

identifier_height . . . . . . . . . . . . . . . . . . . . . . 159, 196 identifier_offset . . . . . . . . . . . . . . . . . . . . . . 159, 197 identifier_placement . . . . . . . . . . . . . . . . . . . . . . . 159 identifier_position . . . . . . . . . . . . . . . . . . . . . . . . 197 identifier_style . . . . . . . . . . . . . . . . . . . . . . . 159, 196 ididx . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 227, 228 ididxs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 227 idxfrom . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 243 idxto. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 243 ignore_attachment . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 illuminance_dist . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 illumination_model. . . . . . . . . . . . . . . . . . . . . . . . . . 178 image_file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 169 image_file_name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 193 image_frame. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 193 image_height . . . . . . . . . . . . . . . . . . . . . . . . . . . . 169, 195 image_quality . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 193 image_size . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 172 image_visibility . . . . . . . . . . . . . . . . . . . . . . . . . . . . 169 image_width . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 169, 195 imagedef. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49, 96 imagedefreactor. . . . . . . . . . . . . . . . . . . . . . . . . . . 49, 96 IMAGE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 IMAGE_BACKGROUND . . . . . . . . . . . . . . . . . . . . . . . . . . . . 172 IMAGEDEF . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 172 IMAGEDEF_REACTOR . . . . . . . . . . . . . . . . . . . . . . . . . . . . 172 increment . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 229 index. . . . . . . . . . . . . . . . . . . . . . 116, 141, 226, 254, 258 index_mask . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 237 INDEX. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 172 INDEXCTL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 indicator_alpha . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 indicator_color . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 indirect_bump_scale . . . . . . . . . . . . . . . . . . . . . . . . 177 ins_pt… 29, 31, 32, 49, 59, 62, 80, 85, 91, 92, 127,

185, 204

ins_rotation . . 28, 33, 35, 36, 37, 38, 40, 41, 51, 234 ins_scale … 28, 33, 35, 36, 37, 38, 40, 41, 50, 234 INSBASE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14, 174 insert_units . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 156 inserts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 156 INSERT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 inspt_offset . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52, 175 INSUNITS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 intensity . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53, 201 intensity_as_gradient . . . . . . . . . . . . . . . . . . . . . . . 72 intensity_colorscheme . . . . . . . . . . . . . . . . . . . . . . . 72

intensity_high_treshold . . . . . . . . . . . . . . . . . . . . 248

intensity_low_treshold . . . . . . . . . . . . . . . . . . . . . 248 intensity_max . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 intensity_min . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72

intensity_out_of_range_behavior . . . . . . . . . . . . 72

intensity_scheme . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71
intensity_style . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71
INTERFERECOLOR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14
INTERFEREOBJVS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14
INTERFEREVPVS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14
internal_only . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 208
INTERSECTIONCOLOR . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14
INTERSECTIONDISPLAY. . . . . . . . . . . . . . . . . . . . . . . . . . 14
interval . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 202
intsectobj . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 233
inverse_transform . . . . . . . . . . . . . . . . . . . . . . . . . . . 199
invis_flags . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24
invisible . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 260
ipe_alignment . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65
is_annotative . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65, 184
is_associative. . . . . . . . . . . . . . . . . . . . . . . . . . . . 46, 61
is_attached_to_object . . . . . . . . . . . . . . . . . . . . . . 114
is_autofit_flag . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 254
is_bg_fill . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 245
is_bg_mask_fill . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 245
is_blob01 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 228
is_camera_plottable . . . . . . . . . . . . . . . . . . . . . . . . 208
is_ccw . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 238
is_changed . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 183
is_close_to_axis . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103
is_col_flow_reversed . . . . . . . . . . . . . . . . . . . . . . . 245
is_def_textloc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 246
is_default. . . . . . 106, 127, 165, 166, 175, 181, 185,

186, 191, 204, 239

is_default_transmatrix . . . . . . . . . . . . . . . . . . . . . 222

                                                is_delegating_to_owning_action . . . . . . . . . . . . 114

is_dst . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 201 is_face_variable . . . . . . . . . . . . . . . . . . . . . . . . . . . . 221 is_gradient_fill . . . . . . . . . . . . . . . . . . . . . . . . . 46, 61 is_hardowner . . . . . . . . . . . . . . . . . . . . . . . . . . . . 160, 161 is_header_suppressed . . . . . . . . . . . . . . . . . . . . . . . 204 is_height_auto . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 245 is_initialized . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 152 is_inside . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 247 is_inverted. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 247 is_live . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 198 is_loaded . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 172, 190 is_locked . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 is_merged_value . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 254

                                                is_modified_for_recompute . . . . . . . . . . . . . 159, 196

is_neg_textdir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 is_normal_reversed. . . . . . . . . . . . . . . . . . . . . . . . . . 244 is_not_annotative . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 is_on . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 201, 217 is_owned . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 221 is_partial . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 is_periodic. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 239 is_photometric . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 is_pspace . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 207 is_r2013 … 109, 110, 113, 114, 116, 119, 120, 121,

124, 127

is_rational. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 239 is_read_dep. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 is_reverse . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 is_semi_assoc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 222 is_semi_ovr. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 222 is_shape . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 200 is_shx. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 is_solid. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44, 84 is_solid_fill . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46, 61 is_text_extended . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 is_title_suppressed . . . . . . . . . . . . . . . . . . . . . . . . 204 is_underlined . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 is_unit_scale . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 196 is_vertical. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 200 is_watertight . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 is_write_dep . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 is_xdic_missing . . . . . . . . . . . . . . . . . . . . . . . . 261, 262 is_xref_dep. . . . . 107, 155, 161, 173, 176, 200, 205,

206, 215, 217

is_xref_ref. . . . . 107, 155, 161, 173, 176, 200, 205,

206, 215, 217

is_xref_resolved. … 107, 155, 161, 173, 176, 200,

205, 206, 215, 217

isbylayerlt. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 261 isoline_present . . 25, 43, 56, 66, 69, 77, 82, 98, 111 isolines . . . . . . . . 25, 43, 56, 66, 69, 77, 82, 97, 111 ISOLINES . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 itemhandle . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 230 itemhandles . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 160, 161 itemloc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 221, 222 items . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110, 219

J

jog_point . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51, 191 JOINSTYLE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 julian_day . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 201, 251 jump_address. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 jump_address_raw . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 jump_entity_section. . . . . . . . . . . . . . . . . . . . . . . . . . 50 JUMP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 justification . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60, 65

K

key . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 236 KEYWORDS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 264 knot_tol. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47, 81 knotparam . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47, 81 knots . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48, 81, 239

L

l2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 222 l4 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 222 l5 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 222 label_text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 240 label_viewports . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 203 lamp_color_preset . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 lamp_color_rgb . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 lamp_color_temp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 lamp_color_type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 landing_dist . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64, 182 landing_gap . . . . . . . . . . . . . . . . . . . . . . . . . 45, 182, 243 LARGE_RADIAL_DIMENSION . . . . . . . . . . . . . . . . . . . . . . 50 last_attrib. . . . . . . . . . . . . . . . . . . . . . . . . . . . 49, 60, 90 last_entity. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 156 last_height. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 201 last_section_address . . . . . . . . . . . . . . . . . . . . . . . 249 last_section_id . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 249 last_updated . . . . . . . . . . . . . . . . . . . . . . . 172, 174, 200 last_vertex. . . . . . . . . . . . . . . . . . . . . . . . . . . . 73, 74, 75 lastleaderlinepoint . . . . . . . . . . . . . . . . . . . . . . . . 240 lastpt_ref . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 233 LASTSAVEDBY. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 264 LATITUDE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 layer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 250, 261 layer_colors. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 LAYER. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173 LAYER_CONTROL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173 LAYER_CONTROL_OBJECT . . . . . . . . . . . . . . . . . . . . . . . . 14 LAYER_INDEX. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 174 LAYERFILTER. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173 layout . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 156 layout_flags . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 174 layout_name. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 174 LAYOUT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 174 LAYOUTPRINTCONFIG . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 ldata . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84, 203 leader_endpt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40, 187 leader_len. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38, 41, 51 leader_order . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 182 leader1_pt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 leader2_pt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 leaders . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 243 LEADER. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51

LEADEROBJECTCONTEXTDATA . . . . . . . . . . . . . . . . . . . . 175

left_col . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 236 left_grid_color . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 256 left_grid_linewt . . . . . . . . . . . . . . . . . . . . . . . . . . . . 256 left_margin. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 188 left_offset . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 left_visibility . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 256 length . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97, 106, 241 lens_length . . . . . . . . . . . . . . . . . . . . . . . . . 93, 207, 215 LENSLENGTH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 light_color . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 light_count. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 194 light_luminance_scale . . . . . . . . . . . . . . . . . . . . . . 181 lighting_model . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 192 lights . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 175 LIGHT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 LIGHTGLYPHDISPLAY . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 LIGHTLIST . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 175 LIMCHECK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 LIMMAX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14, 174 LIMMIN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14, 174 line_color . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64, 182 line_index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 240 line_linewt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 line_ltype . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 line_spacing_factor . . . . . . . . . . . . . . . . . . . . . . . . 244 line_spacing_style. . . . . . . . . . . . . . . . . . . . . . . . . . 245 line_type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 182 LINE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 LINEARGRIPENTITY . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 LINEARPARAMETERENTITY . . . . . . . . . . . . . . . . . . . . . . . 55 lines . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 184, 241, 246 linespace_factor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 linespace_style . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 linewt . . . . . . . . . . . 173, 182, 237, 240, 250, 253, 261 livesection. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 208 loaded_bit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 155 LOAD . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 location . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 166, 244 lock_aspect . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 lock_position_flag . . . . . . . . . . . . . . . . . . . . . . . 30, 31 lock_viewports . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 202 locked . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173

loft_entity_transmatrix . . . . . . . . . . . . . . . . . . . . . 57

LOFTANG1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 LOFTANG2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 LOFTEDSURFACE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 LOFTMAG1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 LOFTMAG2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 LOFTNORMALS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 LOFTPARAM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 long . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 LONG_TRANSACTION . . . . . . . . . . . . . . . . . . . . . . . . . . . . 175 LONGITUDE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 lookup_desc. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 141 lookup_name. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 141 lower_left . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 216

lowermost_left_tree_node_gap . . . . . . . . . . . . . . 249
lowermost_right_tree_node_gap . . . . . . . . . . . . . 249

lspace_factor … 28, 34, 35, 36, 37, 39, 40, 41, 51,

234

lspace_style . . 28, 34, 35, 36, 37, 39, 40, 41, 51, 234 lt.index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 245 lt.ltype . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 245 LTSCALE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 ltype. . . . . . . . . . . . . . . . . . . . . . 173, 237, 240, 250, 261 ltype_flags. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 261 ltype_scale . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 250, 261 LTYPE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 176 LTYPE_BYBLOCK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 LTYPE_BYLAYER . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 LTYPE_CONTINUOUS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 LTYPE_CONTROL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 176 LTYPE_CONTROL_OBJECT . . . . . . . . . . . . . . . . . . . . . . . . 15 luminance . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 178 luminance_mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 178 LUNITS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 LUPREC. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 lut . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 140 LWDISPLAY . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 LWPOLYLINE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58

M

m_density . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 main_gsmarker . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 233 main_subent_type . . . . . . . . . . . . . . . . . . . . . . . . . . . . 233 maint_version . . . . . . . . . . . . . . . . . . . . . . . . 48, 76, 191 maintain_aspect_ratio . . . . . . . . . . . . . . . . . . . . . . 172 major . . . . . 96, 97, 99, 100, 101, 102, 103, 104, 105,

106, 165, 220, 221, 235

major_radius . . . . . . . . . . . . . . . . . . . . . . . . 99, 100, 105 major_version . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 margin_horiz_spacing . . . . . . . . . . . . . . . . . . . . . . . 231 margin_override_flags . . . . . . . . . . . . . . . . . . . . . . 231 margin_vert_spacing . . . . . . . . . . . . . . . . . . . . . . . . 231 mat_absref . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 219 material . . . . . . . . . . . . . . . . . . . . . . . 173, 220, 242, 261 material_count . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 194 material_flags . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 261 material_handle . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 219 materials . . . . . . . 25, 43, 56, 67, 69, 77, 82, 98, 112 MATERIAL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 177 max_assoc_dep_index . . . . . . . . . . 107, 109, 118, 126 max_extent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 186 max_intensity . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 248 max_points . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 182 max_regen_threads . . . . . . . . . . . . . . . . . . . . . . . . . . . 205 MAXACTVP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 maximum . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 229 mdoc_class_version . . . . . . . . . . . . . . . . . . . . . 158, 196 MEASUREMENT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 memory_amount . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 194 memory_limit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 181

                                                MENTALRAYRENDERSETTINGS . . . . . . . . . . . . . . . . . . . . 179

MENU . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 MENUEXT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 merge_flags. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 231 merged_cells . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 236 merged_height_flag. . . . . . . . . . . . . . . . . . . . . . . . . . 254 merged_width_flag . . . . . . . . . . . . . . . . . . . . . . . . . . . 254 MESH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 min_extent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 186 min_intensity . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 248 minimum . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 229 minor . . . . . 96, 97, 99, 100, 101, 102, 103, 104, 105,

106, 165, 220, 221, 222, 235

minor_major_ratio . . . . . . . . . . . . . . . . . . . . . . . . . . . 238 minor_radius . . . . . . . . . . . . . . . . . . . . . . . . 99, 100, 105 MINSERT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 minute . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 158 MIRRTEXT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 miter_direction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 246 miter_option . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101, 105 mleader_order . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 182

MLEADEROBJECTCONTEXTDATA. . . . . . . . . . . . . . . . . . . 181

mleaderstyle. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 MLEADERSTYLE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 182 mlinestyle . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 MLINE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 MLINESTYLE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 184 mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68, 178 model_edge . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 160 model_space. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 155

modeler_format_version. . . . . . . . 44, 57, 70, 78, 83

month. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 158 morehandles. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 165 MOTIONPATH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 184 MPOLYGON . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 mr_description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 181 mr_version . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 179 msec. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 158 msecs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 201, 251 mtext. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 mtext_style . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30, 31 mtext_visible . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 MTEXT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62

MTEXTATTRIBUTEOBJECTCONTEXTDATA. . . . . . . . . . . 185

MTEXTOBJECTCONTEXTDATA . . . . . . . . . . . . . . . . . . . . . 185 MULTILEADER . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63

N

n_density . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 name. . . . . . . . . . . . . . . . . . . . . . . . . . . . 32, 46, 53, 61, 72, 79, 107, 109, 111, 113, 114, 116, 119, 120, 121, 124, 127, 128, 129, 130, 131, 132, 133, 134, 135, 136, 137, 138, 139, 140, 141, 142, 143, 144, 145, 147, 148, 149, 150, 151, 152, 153, 155, 161, 169, 171, 173, 176, 177, 179, 184, 192, 195, 196, 200, 203, 205, 206, 215, 217,

218, 228, 229, 239, 241, 252, 256, 258, 259

named_ucs . . . . . . . . . . . . . . . . . . 95, 175, 206, 208, 217 names. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173 namidx . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 227 NAVISWORKSMODEL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 NAVISWORKSMODELDEF. . . . . . . . . . . . . . . . . . . . . . . . . . 186 network_action_index . . . . . . . . . . . . . . . . . . . . . . . 118 network_version . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 118 next_entity. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 261 nextid. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 234, 235 no_twist . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 node . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114, 235 nodeid. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 229, 235 nodes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108, 166 nolinks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 261 normal . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 244 north_dir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 168 north_dir_angle_deg . . . . . . . . . . . . . . . . . . . . . . . . 169 north_dir_angle_rad . . . . . . . . . . . . . . . . . . . . . . . . 169 NORTHDIRECTION . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 notes. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 num_actions. . . . . 108, 117, 118, 131, 134, 140, 141,

144, 147, 149

num_areafillparms . . . . . . . . . . . . . . . . . . . . . . . . . . . 246 num_arrowheads . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 num_attr_defs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 256 num_attrs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 258 num_blocklabels . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 num_blocks . . . . . 25, 42, 56, 66, 68, 76, 81, 97, 111,

153, 229

num_blverts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 num_borders . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 231, 253 num_boundary_handles. . . . . . . . . . . . . . . . . . . . 62, 238 num_break_heights . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 num_break_rows . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 num_breaks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 240, 241 num_bulges . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58, 248 num_cell_contents . . . . . . . . . . . . . . . . . . . . . . . . . . . 257 num_cells . . . . . . . . . . . . . . . . . . . . . . . 85, 157, 203, 259 num_childs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 166 num_childval . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 167

                                                num_classification_colorramps . . . . . . . . . . . . . 189

num_clip_inverts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 num_clip_verts . . . . . . . . . . . . . . . . . . . 33, 49, 96, 199 num_clippings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 num_codes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 145, 150 num_col_sizes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 245 num_colorramps . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 189 num_colors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46, 61 num_cols . . . . . . . . . . . . . . . . . 50, 60, 85, 158, 242, 257 num_column_heights . . . . . . . . . . . . . . . . . . . . . . 63, 186 num_connections . . . . . . . . . . . . . . . . . . . . . . . . 228, 230 num_control_points. . . . . . . . . . . . . . . . . . . . . . . . . . 239 num_crease . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 num_croppings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 num_cross_sections . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 num_crosssects . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102 num_ctrl_pts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48, 81 num_customdata . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 158 num_customdata_items. . . . . . . . . . . . . . . . . . . 257, 259 num_dashes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 238 num_dates . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 202 num_deflines . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47, 62 num_deps … 108, 109, 118, 126, 131, 134, 140, 141,

144, 147, 149, 222

num_edges . . . . . . . . . . . . . . . . . . . . . . . . 59, 99, 101, 166 num_eed . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 261, 262 num_endsetbacks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102 num_entries. . . . . 107, 154, 164, 174, 176, 187, 201,

206, 208, 217, 225, 230

num_ents . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 199 num_faces . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 num_field_refs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 242 num_fields . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 167 num_fit_pts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48, 81 num_fitpts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 239 num_frozen_layers . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 num_gentextures . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 178 num_geom . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 251 num_geomesh_faces . . . . . . . . . . . . . . . . . . . . . . . . . . . 168 num_geomesh_pts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 168 num_geometry . . . . . . . . . . . . . . . . . . . . . . . . . . . . 252, 258 num_geoms . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 250 num_groups . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 num_guide_curves . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 num_guides . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102 num_hatch_angles . . . . . . . . . . . . . . . . . . . . . . . . . . . . 198 num_hdls. . . . . . . . . . . . . . . . . . . . . . . . . . . . 145, 150, 200 num_hours . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 202 num_ididx . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 227 num_ididxs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 227 num_index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 225 num_inserts. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 155 num_intsectobj . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 233 num_items . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110, 219 num_knots. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48, 81, 239 num_leaders. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 243 num_lights . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 175 num_lines . . . . . . . . . . . . . . . . . . . . . . . 60, 184, 241, 246 num_m_verts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74, 75 num_materials . . 25, 43, 56, 67, 69, 77, 82, 98, 112 num_merged_cells . . . . . . . . . . . . . . . . . . . . . . . . . . . . 236 num_morehandles . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 165 num_n_verts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74, 75 num_names . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173 num_nodes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108, 166 num_obj_ids. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 num_objects. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 166 num_objid_handles . . . . . . . . . . . . . . . . . . . . . . . . . . . 218 num_objids . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76, 191 num_orthopts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 206 num_owned . . . . . . . . . . . . . . 49, 60, 73, 74, 75, 85, 155 num_owned_actions . . . . . . . . . . . . . . . . . . . . . . . . . . . 118 num_owned_params. . . . . . . . . . . . . . 108, 109, 118, 126 num_pages . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 224 num_params . . . . . . . . . . . . . . . . 113, 120, 121, 124, 229 num_paths . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46, 61 num_pointrefs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 156 num_points . . . . . . . . . . . . . 52, 58, 175, 220, 240, 248 num_prop_entries . . . . . . . . . . . . . . . . . . . . . . . . . . . . 226 num_propinfos . . . . . . . . . . . . 132, 141, 142, 150, 152 num_propnames . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 226 num_props . . . . . . . . . . . . . . . . . . . . . . . . . . 212, 226, 264 num_pts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 145, 149, 247 num_radiuses . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102 num_ramps . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 247 num_reactors . . . . . . . . . . . . . . . . . . . . . . . . . . . . 261, 262 num_rows . . . . . . . . . . . . . . . . . 50, 60, 85, 158, 242, 257 num_rowstyles . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 204 num_schemas. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 226 num_search . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 227 num_sections . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 198 num_seeds . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 num_segidx . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 223 num_segparms . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 246 num_segs_or_paths . . . . . . . . . . . . . . . . . . . . . . . . . . . 238 num_silhouettes . . 25, 43, 56, 66, 69, 77, 82, 98, 111 num_sortedidx . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 227 num_source_files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 num_sources. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 251 num_startsetbacks . . . . . . . . . . . . . . . . . . . . . . . . . . . 102 num_states . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153 num_steps . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 121, 187 num_subdiv_vertex . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 num_subents . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 122, 187 num_submgrs. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 157 num_types . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 198 num_uprops . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 226 num_valuelist . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 229 num_values . . . . . . . . . . . 108, 109, 118, 127, 222, 227 num_vars . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 259 num_vertex . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 num_vertexids . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 num_vertices . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 248 num_verts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60, 79 num_viewports . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 175 num_widths . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59, 248 num_wires . . 25, 43, 56, 66, 69, 77, 82, 98, 111, 219 num_xdata . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 218 num_xrefpaths . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 233 num_xrefs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 233 num1. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 200 numassocsteps . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 187 numassocsubents . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 187 numcols . . . . . . . . . . . . . . . . . . . . . . . . . 42, 140, 202, 203 numdashes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 176 numelems . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 140 numentities . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 numfaces . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 numfragments. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 numgaps . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 249 numitems. . . . . . . . . . . . . . . . . . . . . . . . . . . . 160, 161, 219 numlayers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 239 numlevels . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 219 numoverrides . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 204 numpoints . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71, 190 numrows . . . . . . . . . . . . . . . . . . . . 42, 140, 202, 203, 219 numsections. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 249 numverts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 numvports . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 202 NURBSURFACE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66

O

obj_ids . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 172 objdata_algn_offset . . . . . . . . . . . . . . . . . . . . . . . . 228 object. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 229, 247 OBJECT_PTR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 186 objectcontext . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 157 objects . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 166 objid . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 261, 262 objid_handles . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 218 objids . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76, 191 oblique_angle. . . . . . . . . . 29, 31, 34, 39, 80, 91, 200 obs_pt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 167 OBSCOLOR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

observation_coverage_tag. . . . . . . . . . . . . . . . . . . 168

observation_from_tag . . . . . . . . . . . . . . . . . . . . . . . 168 observation_to_tag. . . . . . . . . . . . . . . . . . . . . . . . . . 168 OBSLTYPE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 obsolete_false . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 169 offset . . 144, 147, 149, 172, 225, 226, 228, 238, 245 offset_from_arc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 oldCECOLOR_hi . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 oldCECOLOR_lo . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 OLE2FRAME . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 oleclient . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 OLEFRAME . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 OLESTARTUP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 oleversion . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 on . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173 on_off. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 opacity_percent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 177 opacitymap . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 177 operand1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 operand2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 operation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 option. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 115, 157 opts_r11 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 261

ORDDIMOBJECTCONTEXTDATA . . . . . . . . . . . . . . . . . . . . 186

order. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 orientation . . . . . . . . . . . . . . . . . . . . . . . . 129, 134, 138

orientation_on_both_grips . . . . . . . . . . . . . 131, 133

origin . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52, 70, 199 ORTHOMODE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 orthopts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 206 OSMODE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 osnap_dist . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 233 osnap_mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 120 osnap_pt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 233 osnap_type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 233 other_dist . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 out_edge . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 234 output_type. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 202 override_code . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 246 ovr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 204 ovr_center . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51, 191 owned_actions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 118 owned_params . . . . . . . . . . . . . . . . . . 108, 109, 118, 126 owner. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102 ownerhandle . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 261, 262 owningnetwork . . . . . . . . . . . . . . . . . 107, 109, 118, 126

P

pab …. 108, 110, 112, 115, 116, 117, 118, 119, 120,

124, 125, 126

padding . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 228, 250 page_count . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 224 page_data . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 224 page_data_size . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 224 page_index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 224 page_setup_wizard . . . . . . . . . . . . . . . . . . . . . . . . . . . 203 page_size . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 224, 225 page_start_offset . . . . . . . . . . . . . . . . . . . . . . . . . . . 224 pages. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 225 paper_height . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 188 paper_image_origin. . . . . . . . . . . . . . . . . . . . . . . . . . 189 paper_size . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 188 paper_space. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 155 paper_units . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 188, 196 paper_width. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 188 param . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114, 120 paramblock . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110

parameter_base_location . . . . . 128, 130, 131, 133,

135, 136, 137, 139, 144, 146, 148, 151, 154

params . . . . . . . . . . . . . . . . . . . . 113, 120, 121, 124, 229 parent. . . . . . . . . . . . . . . 24, 26, 28, 29, 30, 32, 33, 34, 35, 37, 38, 39, 40, 42, 45, 46, 47, 48, 49, 50, 51, 53, 55, 58, 59, 60, 61, 62, 63, 65, 66, 67, 68, 70, 71, 73, 74, 75, 76, 78, 79, 80, 81, 84, 91, 92, 93, 95, 96, 97, 99, 100, 101, 102, 103, 104, 105, 106, 107, 108, 109, 110, 112, 113, 114, 115, 116, 117, 118, 119, 120, 121, 123, 124, 125, 126, 127, 128, 129, 130, 131, 132, 133, 134, 135, 136, 137, 138, 139, 140, 141, 142, 143, 144, 145, 146, 147, 148, 149, 150, 152, 153, 154, 155, 156, 157, 158, 160, 161, 164, 165, 166, 167, 169, 170, 171, 172, 173, 174, 175, 176, 177, 179, 181, 182, 184, 185, 186, 187, 189, 190, 191, 192, 193, 194, 195, 196, 198, 199, 200, 201, 203, 204, 205, 206, 208, 214, 217, 218, 219, 220, 221, 222, 228, 229, 230, 231, 232, 233, 234, 236, 237, 238, 239, 240, 241, 242, 243, 245, 246, 247,

248, 250, 251, 252, 253, 254, 258, 259

parentid . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 235 PARTIAL_VIEWING_INDEX . . . . . . . . . . . . . . . . . . . . . . 187 path. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 186 path_curve . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 path_entity . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45, 84 path_entity_id . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83

path_entity_transform_computed . . . . . . . . . 45, 84
path_entity_transmatrix . . . . . . . . . . . . . . . . . 44, 84

path_flags . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45, 84 path_option. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 158 path_type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 pathdata . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 pathdata_size . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 pathentity_transform. . . . . . . . . . . . . . . . . . . 101, 105 paths . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46, 61 pattern_len. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 176 pattern_type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46, 61 pbsab_status … 112, 115, 116, 117, 118, 119, 120,

124, 125, 126 PDFDEFINITION . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 218 PDFUNDERLAY . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 PDMODE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 PDSIZE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 PELEVATION . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 PELLIPSE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 periodic. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47, 81 PERSUBENTMGR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 187 PEXTMAX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 PEXTMIN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 photon_trace_depth1 . . . . . . . . . . . . . . . . . . . . . . . . 180 photon_trace_depth2 . . . . . . . . . . . . . . . . . . . . . . . . 180 photon_trace_depth3 . . . . . . . . . . . . . . . . . . . . . . . . 180 physical_intensity . . . . . . . . . . . . . . . . . . . . . . . . . . . 54

physical_intensity_method. . . . . . . . . . . . . . . . . . . 54

PICKSTYLE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 PINSBASE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 pixel_size . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47, 172 PLACEHOLDER. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 187 plane_line_color . . . . . . . . . . . . . . . . . . . . . . . . . . . . 197 plane_linewt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 197 plane_ltype. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 197

plane_normal_lofting_type. . . . . . . . . . . . . . . . . . . 57

PLANESURFACE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 PLIMCHECK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 PLIMMAX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 PLIMMIN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 PLINEGEN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 PLINEWID . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 plot_flags . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 188 plot_glyph . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 plot_origin. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 188 plot_paper_unit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 188 plot_rotation_mode. . . . . . . . . . . . . . . . . . . . . . . . . . 188 plot_type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 188 plot_window_ll . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 188 plot_window_ur . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 188 plotflag . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173 plotsettings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 174 PLOTSETTINGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 187 plotstyle . . . . . . . . . . . . . . . . . . . . . . . . . . 173, 250, 261 plotstyle_flags . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 261 plotview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 188 plotview_name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 188 point . . . . . 25, 43, 56, 66, 69, 76, 77, 82, 92, 93, 97,

111, 190, 237, 239

point_present . . 25, 43, 56, 66, 69, 77, 82, 97, 111 POINT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 POINTCLOUD . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 POINTCLOUDCOLORMAP. . . . . . . . . . . . . . . . . . . . . . . . . . 189 pointclouddef . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 POINTCLOUDDEF_REACTOR . . . . . . . . . . . . . . . . . . . . . . 190

POINTCLOUDDEF_REACTOR_EX. . . . . . . . . . . . . . . . . . . 190

pointclouddefex . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 POINTCLOUDDEF . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 189 POINTCLOUDDEFEX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 190 POINTCLOUDEX. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 POINTPARAMETERENTITY . . . . . . . . . . . . . . . . . . . . . . . . 73 POINTPATH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 190 pointrefs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 156 points . . . . . . . . . . . . . . . . . . 52, 58, 175, 220, 240, 248 POLARGRIPENTITY . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 polyline_paths . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 238 POLYLINE_2D . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 POLYLINE_3D . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 POLYLINE_MESH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 POLYLINE_PFACE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 position . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45, 53, 254 position_desc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 142 position_name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 142 predef_presets_first . . . . . . . . . . . . . . . . . . . . . . . 195 preset_name. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 193 prev_entity. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 261 prev_entry . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 218 preview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 156, 261 preview_exists . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 261 preview_is_proxy . . . . . . . . . . . . . . . . . . . . . . . . . . . . 261 preview_size . . . . . . . . . . . . . . . . . . . . . . . . . . . . 156, 261 printer_cfg_file . . . . . . . . . . . . . . . . . . . . . . . . . . . . 187 procedure . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 195 projection . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 242 PROJECTNAME . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 prompt. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 prop_entries . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 226 prop_states. . . . . 128, 130, 133, 135, 136, 137, 139,

144, 146, 148, 151, 154

prop1 . . 128, 129, 130, 132, 133, 135, 136, 137, 139,

141, 142, 143, 146, 148, 150, 151, 152, 154

prop2 . . 128, 130, 132, 133, 135, 136, 137, 139, 141,

142, 143, 146, 148, 150, 151, 152, 154

prop3 . . 128, 130, 133, 135, 136, 137, 139, 143, 146,

148, 151, 154

prop4 . . 128, 130, 133, 135, 136, 137, 139, 143, 146,

148, 151, 154

property_flags . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 232

                                                property_override_flags . . . . . . . . . . . . . . . 230, 231

propnames . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 226 props . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 226, 264 PROXY_ENTITY. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 PROXY_OBJECT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 190 PROXYGRAPHICS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 prvsav_segidx . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 223 PSLTSCALE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 PSOLHEIGHT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 PSOLWIDTH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 PSTYLEMODE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 PSVPSCALE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 pt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79, 127, 132, 259 pt0 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48, 95, 169, 237 pt1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68, 116 pt2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68, 101, 105, 116 pts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 145, 149, 247 PUCSBASE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 PUCSNAME . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 PUCSORG . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 PUCSORGBACK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 PUCSORGBOTTOM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 PUCSORGFRONT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 PUCSORGLEFT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 PUCSORGRIGHT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 PUCSORGTOP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 PUCSORTHOREF. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 PUCSORTHOVIEW . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 PUCSXDIR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 PUCSYDIR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17

Q

QTEXTMODE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17

R

r_node . . . . . . . . . . . . . . . . . . . . . . . . . 108, 110, 119, 125 r11_prev_entry_index . . . . . . . . . . . . . . . . . . . . . . . 218 r11_viewport_index. . . . . . . . . . . . . . . . . . . . . . . . . . 218

RADIMLGOBJECTCONTEXTDATA. . . . . . . . . . . . . . . . . . . 191

RADIMOBJECTCONTEXTDATA . . . . . . . . . . . . . . . . . . . . . 191 radius . . . . . . . . . . . . 26, 27, 32, 45, 48, 103, 104, 238 radiuses . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102 ramps. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 247 rapidrt_version . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 192 RAPIDRTRENDERSETTINGS . . . . . . . . . . . . . . . . . . . . . . 192 RASTERVARIABLES . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 193 rational. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47, 81 ray_trace_depth1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . 180 ray_trace_depth2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . 180 ray_trace_depth3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . 180 ray_tracing_enabled . . . . . . . . . . . . . . . . . . . . . . . . 180 RAY . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 reactor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71, 72 reactors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 262 readdep . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 REALWORLDSCALE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 record_hdrs. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 224 record_history . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102 record_size . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 224, 225 records . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 224 rect_height . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62, 185 rect_width . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62, 185 ref . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 161, 224 ref_pt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 167 ref_pt2d . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 169 refcount . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 236

reference_vector_for_

controlling_twist . . . . . . . . . . . . . . . . . . . . . . 45, 84 reflectance. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 221 reflectance_scale . . . . . . . . . . . . . . . . . . . . . . . . . . . 177 reflectionmap . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 177 reflectivity . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 178 refraction_index . . . . . . . . . . . . . . . . . . . . . . . . . . . . 177 refractionmap . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 177 REGENMODE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 REGION. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 rel_transform . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 222 render_level . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 192 render_mode . . . . . . . . . . . . . . . . . . . . . . . . . 94, 207, 215 render_target . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 192 render_time . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 192, 193 RENDERENTRY. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 193 RENDERENVIRONMENT . . . . . . . . . . . . . . . . . . . . . . . . . . . 194 RENDERGLOBAL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 194 RENDERSETTINGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 195 REPEAT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 REQUIREDVERSIONS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 reserved. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46, 61 resunits . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 172 revision_bytes . . 26, 43, 57, 67, 69, 77, 82, 98, 112 revision_guid . . 25, 43, 56, 67, 69, 77, 82, 98, 112 revision_major . . 26, 43, 57, 67, 69, 77, 82, 98, 112 revision_minor1 . . 26, 43, 57, 67, 69, 77, 82, 98, 112 revision_minor2 . . 26, 43, 57, 67, 69, 77, 82, 98, 112 REVISIONNUMBER . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 264 revolve_angle . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78, 103

revolved_entity_transmatrix . . . . . . . . . . . . . . . . 78

REVOLVEDSURFACE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 rgb . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 242 right_col . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 236 right_grid_color . . . . . . . . . . . . . . . . . . . . . . . . . . . . 255 right_grid_linewt . . . . . . . . . . . . . . . . . . . . . . . . . . . 255 right_margin . . . . . . . . . . . . . . . . . . . . . . . . . . . . 188, 231 right_offset. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 right_visibility . . . . . . . . . . . . . . . . . . . . . . . . . . . . 255 root_tree_node_gap. . . . . . . . . . . . . . . . . . . . . . . . . . 249 rotated_type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 161 rotation … 29, 31, 49, 60, 79, 80, 85, 91, 127, 169,

170, 171, 185, 204, 232, 241, 244, 254

ROTATIONGRIPENTITY . . . . . . . . . . . . . . . . . . . . . . . . . . . 78

ROTATIONPARAMETERENTITY . . . . . . . . . . . . . . . . . . . . . 78

row_heights . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 row_offset . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 132 row_parent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 256, 258 row_spacing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50, 60 rows . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 232, 242 rowspacing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 rowstyles . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 204 RTEXT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 ruled_surface . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58

S

sab . . 112, 115, 116, 117, 118, 119, 120, 124, 125, 126 sab_size . . . . . . . . 25, 43, 56, 66, 68, 76, 81, 97, 111

sampling_contrast_color1. . . . . . . . . . . . . . . . . . . 179
sampling_contrast_color2. . . . . . . . . . . . . . . . . . . 179
sampling_contrast_color3. . . . . . . . . . . . . . . . . . . 179
sampling_contrast_color4. . . . . . . . . . . . . . . . . . . 180

sampling_filter1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . 179 sampling_filter2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . 179 sampling_mr_filter. . . . . . . . . . . . . . . . . . . . . . . . . . 179 sampling1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 179 sampling2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 179 save_enabled . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 195 save_filename . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 195 saved_filename . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 SAVEIMAGES . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 scale …. 32, 49, 59, 60, 80, 85, 106, 127, 165, 166, 172, 175, 182, 183, 185, 186, 191, 204, 220, 241, 244 scale_est . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 168 scale_factor… 44, 65, 83, 101, 104, 127, 170, 243 scale_flag. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49, 59, 85 scale_spacing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47, 61 scale_vec . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 167 SCALE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 196 scenario. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47, 80 schdat . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 223 schema . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 160 schema_namidx . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 227 schemas . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 226 schidx. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 223, 225 schidx_segidx . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 223 sea_level_elev . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 168 search. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 223, 227 search_segidx . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 223 second_endpoint . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 238 second_seg_angle . . . . . . . . . . . . . . . . . . . . . . . . . . . . 182 secondary_background . . . . . . . . . . . . . . . . . . . . . . . 171 secondheader_address . . . . . . . . . . . . . . . . . . . . . . . 249 seconds . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 158 section_array_size. . . . . . . . . . . . . . . . . . . . . . . . . . 250 section_info_id . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 250 section_map_address . . . . . . . . . . . . . . . . . . . . . . . . 249 section_map_id . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 249 section_settings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 section_type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 250 SECTION_MANAGER . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 198 SECTION_SETTINGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . 198 SECTIONOBJECT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 sections . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 198 SECTIONVIEWSTYLE . . . . . . . . . . . . . . . . . . . . . . . . . . . . 196 seeds. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 segidx . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 223, 225, 226 segidx_offset . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 223 segidx_unknown . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 223 segment_idx. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 228 segments . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 224 segparms . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 246 segs. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 238 segsize . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 228

select_dates_from_calendar . . . . . . . . . . . . . . . . 202

select_range_of_dates . . . . . . . . . . . . . . . . . . . . . . 202 selectable . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 selection_marker . . . . . . . . . . . . . . . . . . . . . . . . . . . . 220 self_illumination . . . . . . . . . . . . . . . . . . . . . . . . . . . 178 seqend . . . . . . . . . . . . . . . . . . . . . . 49, 60, 73, 74, 75, 90 SEQEND. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 setup_name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 202 shade_plot_type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 202 SHADEDGE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 SHADEDIF . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 shadeplot . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95, 189 shadeplot_customdpi . . . . . . . . . . . . . . . . . . . . . . . . 189 shadeplot_mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 shadeplot_reslevel. . . . . . . . . . . . . . . . . . . . . . . . . . 189 shadeplot_type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 189 shadow . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 262 shadow_flags . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 262 shadow_map_size . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 shadow_map_softness. . . . . . . . . . . . . . . . . . . . . . . . . . 54 shadow_maps_enabled . . . . . . . . . . . . . . . . . . . . . . . . 180 shadow_mapsize . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 201 shadow_mode. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 180 shadow_softness . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 201 shadow_type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53, 201 SHADOWPLANELOCATION. . . . . . . . . . . . . . . . . . . . . . . . . . 17 shape_flag . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 241 SHAPE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 sheet_set_name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 202 sheet_subset_name . . . . . . . . . . . . . . . . . . . . . . . . . . . 202 shift_value. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 237 short . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 short170 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 shorts. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 145, 150 show_clipping . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 show_cropping . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 show_history . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102 show_intensity. . . . . . . . . . . . . . . . . . . . . . . . . . . . 71, 72 show_properties. . . . . 128, 129, 130, 132, 135, 136, 137, 139, 141, 142, 143, 146, 148, 150, 151, 152, 153 show_rotation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 170 SHOWHIST . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 shsw_b294 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101, 105 shsw_b295 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101, 105 shsw_b296 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101, 105 shsw_bl93 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100, 104 shsw_text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100, 104 shsw_text_size . . . . . . . . . . . . . . . . . . . . . . . . . 100, 104 shsw_text2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100, 104 shsw_text2_size . . . . . . . . . . . . . . . . . . . . . . . . 100, 104 si_tag . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 226 si_unknown_1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 226 si_unknown_2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 226 sides. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 signature . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 228 silhouettes. . . . . 25, 43, 56, 66, 69, 77, 82, 98, 111 simple_surfaces . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 single_color_gradient. . . . . . . . . . . . . . . . . . . . 46, 61 size . . . . . . . . . . . . 21, 48, 95, 169, 225, 226, 228, 248 SKETCHINC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 SKPOLY. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 SKYLIGHT_BACKGROUND . . . . . . . . . . . . . . . . . . . . . . . . 198 sm_axis . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 SNAPANG . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17, 93, 216 SNAPBASE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17, 93, 216 SNAPISOPAIR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17, 216 SNAPMODE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17, 216 SNAPSTYLE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17, 216 SNAPUNIT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17, 93, 216 solid . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58, 78 SOLID. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 SOLID_BACKGROUND . . . . . . . . . . . . . . . . . . . . . . . . . . . . 199 SOLIDHIST . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 sort_ents . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 199 sortedidx . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 227 SORTENTS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 SORTENTSTABLE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 199 source . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 242 source_filename . . . . . . . . . . . . . . . . . . . . . . . . 189, 190 source_files. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 source_pt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 237 sources . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 251 spacing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 202 SPATIAL_FILTER . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 199 SPATIAL_INDEX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 200 specular_color . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 177 specular_gloss_factor . . . . . . . . . . . . . . . . . . . . . . 177 specularmap. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 177 SPLFRAME . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 splineflags . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47, 80 SPLINE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 SPLINESEGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 SPLINETYPE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 start . . . . . . . . . . . . . . . . . . . . . . . . 24, 55, 240, 241, 254 start_angle. . . . . . . . . . 26, 27, 42, 78, 103, 184, 238 start_day . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 193 start_draft_angle . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 start_draft_dist . . . . . . . . . . . . . . . . . . . . . . . 100, 104 start_draft_magnitude . . . . . . . . . . . . . . . . . . . . . . . 57 start_minute . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 193 start_month. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 193 start_msec . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 193 start_pt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 start_second . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 193 start_tangent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 239 start_time . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 202 start_width. . . . . . . . . . . . . . . . . . . . . . . . . . . . 73, 74, 92 start_year . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 193 startsetbacks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102 state . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79, 134 states . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153 status . . . . . . . . . . . . 53, 110, 113, 116, 120, 186, 229 status_flag . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 std_scale_factor . . . . . . . . . . . . . . . . . . . . . . . . . . . . 189 std_scale_type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 189 step_id . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 220 steps . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 122, 187 STEPSIZE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 STEPSPERSEC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 strings_area . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 176 strokes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 214 strokes_int. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 214 struct. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 strvalue . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 160 sty . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 204 style … 27, 30, 31, 46, 61, 62, 79, 80, 91, 241, 244 style_attachment . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 style_content . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 style_id . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80, 257, 259 style_parent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 258 style_sheet . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 style_type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 208 STYLE_CONTROL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 201 STYLE_CONTROL_OBJECT . . . . . . . . . . . . . . . . . . . . . . . . 18 stylesheet . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 188 STYLE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 200 STYLESHEET . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 stylization_type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 subdiv_vertex . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 subent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 115 subents . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 122, 187 SUBJECT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 264 submgrs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 157 sun . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95, 208, 217 sunid. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 199 SUN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 201 SUNSTUDY . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 201 SURFTAB1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 SURFTAB2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 SURFTYPE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 SURFU. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 SURFV. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 sweep_alignment_flags. . . . . . . . . . . . . . . . . . . . 44, 84 sweep_entity . . . . . . . . . . . . . . . . . . . . . . . . . 45, 84, 103 sweep_entity_id . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83

                                                sweep_entity_transform_computed . . . . . . . . 45, 84
                                                sweep_entity_transmatrix . . . . . . . . . . . . . . . . 44, 83

sweep_transmatrix . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 sweep_vector. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 sweepdata . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 sweepdata_size . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 sweepentity_transform . . . . . . . . . . . . . . . . . 101, 105 SWEPTSURFACE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81

T

t2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 t3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 t58 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 127 t78 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 127 tab_order . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 174 table_flag_override. . . . . . . . . . . . . . . . . . . . . . . . . . 85 table_name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 158 TABLECONTENT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 203 tabledatacolumn_parent . . . . . . . . . . . . . . . . . . . . . 231 tablegeometry. . . . . . . . . . . . . . . . . . . . . . . . . . . 252, 258 TABLEGEOMETRY . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 203 tablerow_parent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 231 tablestyle . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84, 203 TABLE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 TABLESTYLE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 203 tag. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30, 31, 251 tangent_dir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 target . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53, 232 target_path. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 184 TARGET. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 tdata . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84, 203 TDCREATE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18, 264 TDINDWG. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18, 264 TDUCREATE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 TDUPDATE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18, 264 TDUSRTIMER . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 TDUUPDATE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 text . . . . . . . . . . . . . . . . . . . . . . . . 62, 230, 232, 241, 254 text_align_type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 183 text_alignment . . . . . . . . . . . . . . . . . . 45, 64, 243, 253 text_always_left . . . . . . . . . . . . . . . . . . . . . . . . . . . . 183 text_angle_type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 183 text_angletype. . . . . . . . . . . . . . . . . . . . . . . . . . . 64, 243 text_bottom. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 244 text_color . . . . . . . . . . . . . . . . . . . . . . . . . . 64, 183, 253 text_default . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 183 text_direction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 text_extended . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 184 text_height . . . . . . . . . . . 62, 183, 232, 243, 253, 255 text_left . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64, 243 text_midpt . . 28, 33, 34, 36, 37, 38, 39, 41, 50, 233 text_position . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 text_right . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64, 243 text_rotation … 28, 33, 35, 36, 37, 38, 40, 41, 50,

234, 246

text_size . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26, 200 text_style . . . . . . . . . . . . 64, 183, 203, 232, 253, 255 text_top . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 244 text_value. . . . . . . . . . . . . . . . . 27, 31, 79, 91, 92, 255 texts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 160, 161 TEXT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 TEXTOBJECTCONTEXTDATA . . . . . . . . . . . . . . . . . . . . . . 204 TEXTQLTY . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 TEXTSIZE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 TEXTSTYLE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 texturemode. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 243 thickness …. 24, 26, 29, 31, 32, 55, 58, 70, 73, 80,

91, 92, 248

thickness_r11 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 262 THICKNESS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 tile_order . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 181 tile_size . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 181 TILEMODE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 TILEMODELIGHTSYNCH . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 tiling . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 242 timestamp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 236 TIMEZONE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19

title_horiz_bottom_color . . . . . . . . . . . . . . . . . . . . 87
title_horiz_bottom_linewt. . . . . . . . . . . . . . . . . . . 88
title_horiz_bottom_visibility . . . . . . . . . . . . . . 89

title_horiz_ins_color . . . . . . . . . . . . . . . . . . . . . . . 87 title_horiz_ins_linewt . . . . . . . . . . . . . . . . . . . . . . 88

title_horiz_ins_visibility. . . . . . . . . . . . . . . . . . 89

title_horiz_top_color . . . . . . . . . . . . . . . . . . . . . . . 87 title_horiz_top_linewt . . . . . . . . . . . . . . . . . . . . . . 88

title_horiz_top_visibility. . . . . . . . . . . . . . . . . . 89

title_row_alignment. . . . . . . . . . . . . . . . . . . . . . . . . . 86 title_row_color . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 title_row_fill_color . . . . . . . . . . . . . . . . . . . . . . . . 86 title_row_fill_none. . . . . . . . . . . . . . . . . . . . . . . . . . 86 title_row_height . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86

                                                title_row_style_override . . . . . . . . . . . . . . . . . . . . 90

title_suppressed . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 title_text_style . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 title_vert_ins_color . . . . . . . . . . . . . . . . . . . . . . . . 87 title_vert_ins_linewt . . . . . . . . . . . . . . . . . . . . . . . 88

                                                title_vert_ins_visibility. . . . . . . . . . . . . . . . . . . 89

title_vert_left_color . . . . . . . . . . . . . . . . . . . . . . . 87 title_vert_left_linewt . . . . . . . . . . . . . . . . . . . . . . 88

                                                title_vert_left_visibility. . . . . . . . . . . . . . . . . . 89

title_vert_right_color . . . . . . . . . . . . . . . . . . . . . . 87

                                                title_vert_right_linewt . . . . . . . . . . . . . . . . . . . . . 88
                                                title_vert_right_visibility . . . . . . . . . . . . . . . . 89

TITLE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 264 TOLERANCE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 tooltip . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 136, 157, 257 top_grid_color . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 255 top_grid_linewt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 255 top_height . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 top_margin . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 188 top_row . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 236 top_visibility . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 255 topradius . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 total_data_size . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 224 total_segments . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 223 TRACE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 TRACEWID . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 trans. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 220 trans_space_flag . . . . . . . . . . . . . . . . . . . . . . . . . . . . 161 transform . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 200, 244 transform_present . . . . . . . . . . . . . . . . . . . . . . . . . . . 220 translation. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 220 translucence . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 178 transmatrix. . . . . . . . . . . . . . . . . . . . . 66, 110, 222, 242 transmittance_scale . . . . . . . . . . . . . . . . . . . . . . . . 178 transparency . . . . . . . . . . . . . . . . . . . . . . . . . . . . 170, 221 TREEDEPTH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 triangle_count . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 194 TSTACKALIGN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 TSTACKSIZE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 turn_height . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 turns. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 TVDEVICEPROPERTIES. . . . . . . . . . . . . . . . . . . . . . . . . . 204 twist_angle . . . . . 44, 78, 83, 93, 101, 103, 105, 207 two_sided_material. . . . . . . . . . . . . . . . . . . . . . . . . . 178 txt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 260 type …. 30, 31, 53, 64, 68, 150, 182, 220, 227, 228, 230, 232, 240, 244, 247, 251, 252, 254, 258, 259 type_size . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 227 types. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 198

U

u.bd. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 235 u.bl. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 235 u.bs. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 235 u.handle . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 236 u.rc. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 235 u.text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 235 u_isolines . . . . . . . . . . . . . . . . . . 44, 57, 67, 70, 78, 83 ucs_at_origin . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94, 216 ucs_elevation . . . . . . . . . . . . . 94, 174, 205, 207, 216 ucs_name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 ucs_origin . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 ucs_x_dir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71, 72 ucs_y_dir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71, 72 ucs_z_dir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71, 72 UCS_CONTROL. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 206 UCS_CONTROL_OBJECT . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 UCSBASE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 UCSFOLLOW . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 216 UCSICON. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19, 216 UCSNAME . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 ucsorg . . . . . . . . . . . . . . . . . . . . . . . . . . 94, 205, 207, 216 UCS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 205 UCSORG . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19, 174 UCSORGBACK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 UCSORGBOTTOM. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 UCSORGFRONT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 UCSORGLEFT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 UCSORGRIGHT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 UCSORGTOP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 UCSORTHOREF . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 UCSORTHOVIEW . . . . . . . . . . . 19, 94, 174, 205, 207, 216 UCSVP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94, 216 ucsxdir . . . . . . . . . . . . . . . . . . . . . . . . . 94, 205, 207, 216 UCSXDIR. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19, 174 ucsydir . . . . . . . . . . . . . . . . . . . . . . . . . 94, 205, 207, 216 UCSYDIR. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19, 174 ui_index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 240 uint64_t . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 UNDERLAY . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 UNDERLAYDEFINITION. . . . . . . . . . . . . . . . . . . . . . . . . . 218 unit_scale_horiz . . . . . . . . . . . . . . . . . . . . . . . . . . . . 167 unit_scale_vert . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 168 unit_type . . . . . . . . . . . . . . . . . . . . . . . . . . 254, 257, 259 unit1_name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 unit1_ratio . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 unit2_name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 unit2_ratio . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 unit3_name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 unit3_ratio . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 unit4_name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 unit4_ratio . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 unitfactor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 UNITMODE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 units. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 193 units_value_horiz . . . . . . . . . . . . . . . . . . . . . . . . . . . 167 units_value_vert . . . . . . . . . . . . . . . . . . . . . . . . . . . . 168 unknown . . . . . . . . . . . . . 25, 28, 34, 35, 36, 37, 39, 40, 41, 42, 51, 56, 66, 68, 76, 81, 97, 107, 111, 167, 171,

225, 227, 230, 234, 245, 256, 257

unknown_0 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 121, 187 unknown_1 . . . . . . . . . . . . . . . . . 223, 224, 225, 227, 249 unknown_10 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 unknown_11 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 unknown_12 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 unknown_13 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 unknown_14 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 unknown_14b . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 unknown_15 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 unknown_16 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 unknown_17 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 unknown_2… 121, 187, 223, 224, 225, 227, 228, 249 unknown_20 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 unknown_21 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 unknown_22 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 unknown_23 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 unknown_3 . . . . . . . . . . . . . . . . . . . . . . . . . . 121, 228, 249 unknown_4f2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 unknown_5 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 unknown_51e . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 unknown_520 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 unknown_52c . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 unknown_52e . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 unknown_530 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 unknown_54 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 unknown_55 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 unknown_56 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 unknown_57 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 unknown_59 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 unknown_6 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 unknown_6a . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 unknown_6b . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 unknown_6c . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 unknown_8 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 unknown_9 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 unknown_b. . . . . . . . . . . . . . . . . . . . . . . . . . . . 84, 169, 231 unknown_b0 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 unknown_b1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59, 198 unknown_b2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59, 198 unknown_b37. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123 unknown_bit_1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 unknown_bit_2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 unknown_bit_3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 unknown_bit_4 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 unknown_bit_5 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 unknown_bl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84, 231 unknown_bl0 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 unknown_bl1 . . . . . . . . . . . . . . . . . . . . . . . . . . 73, 85, 204 unknown_bl10 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 122 unknown_bl11 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 122 unknown_bl12 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 122 unknown_bl13 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 122 unknown_bl14 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 122 unknown_bl15 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 122 unknown_bl16 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 122 unknown_bl17 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 122 unknown_bl18 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 122 unknown_bl19 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 122 unknown_bl2. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 204 unknown_bl20 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 122 unknown_bl21 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123 unknown_bl22 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123 unknown_bl23 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123 unknown_bl24 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123 unknown_bl25 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123 unknown_bl26 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123 unknown_bl27 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123 unknown_bl28 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123 unknown_bl29 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123 unknown_bl3. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 204 unknown_bl30 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123 unknown_bl31 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123 unknown_bl32 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123 unknown_bl33 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123 unknown_bl34 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123 unknown_bl35 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123 unknown_bl36 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123 unknown_bl6. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 122 unknown_bl6a . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 122 unknown_bl7. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 122 unknown_bl7a . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 122 unknown_bl8. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 122 unknown_bl9. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 122 unknown_bool . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 152 unknown_bs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 unknown_day . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 unknown_h . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 unknown_hour. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 unknown_long . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 249 unknown_min . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 unknown_mon . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 unknown_ms . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 unknown_r11. . . . . . . . . . . . . . . . . . . . . 55, 156, 176, 208 unknown_r2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 208 unknown_rc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84, 204 unknown_sec . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 unknown_short . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 unknown_short_1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 unknown_string . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 unknown_t . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 141 unknown_year. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 unknown1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 169, 265 unknown2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 169, 265 UNKNOWN_ENT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 UNKNOWN_OBJ. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 206 unnamed . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 up_dir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 168 upd_basept. . . . . . 128, 130, 131, 133, 135, 136, 138,

139, 144, 146, 148, 151, 154

upd_endpt . . 128, 130, 131, 133, 135, 137, 138, 139,

144, 146, 148, 151, 154

upd_state . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 134 update_option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 157

update_status . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 158
upper_right. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 216
uprops . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 226
use_attenuation_limits . . . . . . . . . . . . . . . . . . . . . . 53
use_block_rotation. . . . . . . . . . . . . . . . . . . . . . . . . . 183
use_block_scale . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 183
use_default_lights . . . . . . . . . . . . . . . . . 94, 207, 215
use_lut_palette . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 205
use_subset . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 202
use_tiling . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 172
used … 107, 155, 161, 173, 176, 200, 205, 206, 215,

217

user_scale_factor . . . . . . . . . . . . . . . . . . . . . . . . . . . 168 user_text … 28, 33, 34, 36, 37, 38, 39, 41, 50, 234 USERI1. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 USERI2. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 USERI3. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 USERI4. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 USERI5. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 USERR1. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 USERR2. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 USERR3. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 USERR4. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 USERR5. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 USRTIMER . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 uvec . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48, 95 uvec1. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 uvec2. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67

V

v_isolines . . . . . . . . . . . . . . . . . . 44, 57, 67, 70, 78, 83 value . . 127, 128, 137, 138, 150, 151, 167, 221, 232,

236, 251, 256, 258, 260

value.handle91 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 235 value.long90 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 235 value.num40. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 235 value.pt2d . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 235 value.pt3d . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 235 value.short70 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 235 value.text1. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 235 value_code . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 235 value_data_type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 232 value_format_string . . . . . . . . . . . . . . . . . . . . . . . . 232 value_set… 129, 131, 133, 137, 138, 139, 146, 151 value_string . . . . . . . . . . . . . . . . . . . . . . . . . . . . 167, 257 value_string_length . . . . . . . . . . . . . . . . . . . . . . . . 167 value_unit_type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 232 valuelist . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 229 values. . . . . . . . . . . . . . . . 108, 109, 118, 127, 222, 227 vars. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 259 VBA_PROJECT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 206 vector. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 version …. 25, 42, 56, 66, 68, 76, 81, 97, 108, 111,

121, 191, 222, 223, 236

VERSIONGUID . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 vert_alignment . . . . . . . . . . . . . . . . . . . . . . . . 30, 31, 91 vert_cell_margin . . . . . . . . . . . . . . . . . . . . . . . . 86, 203 vert_dir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 vert_margin. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 231 vertex . . . . . . . . . . . . . . . . . . . . . . . . . 59, 73, 74, 75, 246 vertex_direction . . . . . . . . . . . . . . . . . . . . . . . . . . . . 246 VERTEX_2D . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 VERTEX_3D . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 VERTEX_MESH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 VERTEX_PFACE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 VERTEX_PFACE_FACE . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 vertexids . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 vertices . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 248 vertind . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 verts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60, 79 view . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32, 203 view_name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 193 view_target . . . . . . . . . . . . . . . . . . . . . . . . . 93, 207, 215 view_twist . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 215 view_width . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 206, 215 VIEW. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 206 VIEW_CONTROL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 208 VIEW_CONTROL_OBJECT. . . . . . . . . . . . . . . . . . . . . . . . . . 20 VIEWCTR . . . . . . . . . . . . . . . . . . . . . . . . . . 20, 93, 207, 215 VIEWDIR . . . . . . . . . . . . . . . . . . . . . . . . . . 20, 93, 207, 215 viewlabel_alignment. . . . . . . . . . . . . . . . . . . . 160, 198 viewlabel_attachment. . . . . . . . . . . . . . . . . . . 160, 197 viewlabel_offset . . . . . . . . . . . . . . . . . . . . . . . 160, 197 viewlabel_pattern . . . . . . . . . . . . . . . . . . . . . . 160, 198 viewlabel_text_color. . . . . . . . . . . . . . . . . . . 159, 197 viewlabel_text_height . . . . . . . . . . . . . . . . . 159, 197 viewlabel_text_style. . . . . . . . . . . . . . . . . . . 159, 197 VIEWMODE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20, 207, 215 viewport . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 218, 262 viewports . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 175 VIEWPORT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 VIEWSIZE . . . . . . . . . . . . . . . . . . . . . . . . . 20, 93, 206, 215 viewstyle_flags . . . . . . . . . . . . . . . . . . . . . . . . 159, 196 viewtable . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 184 VIEWTWIST . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 virtual_edge_flag . . . . . . . . . . . . . . . . . . . . . . . . . . . 255 virtual_guide . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 VISIBILITYGRIPENTITY . . . . . . . . . . . . . . . . . . . . . . . . 95

VISIBILITYPARAMETERENTITY. . . . . . . . . . . . . . . . . . . 95

visible . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 237, 253 VISRETAIN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 visualstyle . . . . . . . . . . . . . . . . 95, 173, 203, 208, 217 VISUALSTYLE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 208 void* . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 vp_dir_from_target. . . . . . . . . . . . . . . . . . . . . . . . . . 219 vp_id. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 219 vp_perspective . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 219 vp_target . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 219 vp_up_dir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 219 VPOINTX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 VPOINTXALT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 VPOINTY . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 VPOINTYALT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 VPOINTZ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 VPOINTZALT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 vport_entity_address . . . . . . . . . . . . . . . . . . . . . . . 218 vport_entity_header. . . . . . . . . . . . . . . . . . . . . . . . . . 94 VPORT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 214 VPORT_CONTROL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 217 VPORT_CONTROL_OBJECT . . . . . . . . . . . . . . . . . . . . . . . . 20 vvec . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48, 95 vvec1. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 vvec2. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 VX_CONTROL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 217 VX_CONTROL_OBJECT . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 VX_TABLE_RECORD . . . . . . . . . . . . . . . . . . . . . . . . . 21, 217

W

w . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 251 wchar*. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 web_angle1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 web_angle2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 web_angle3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 web_angle4 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 web_angle5 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 web_flux . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 web_rotation. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 web_symetry . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 webfile . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 webfile_type. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 weight . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 237 weighted. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47, 81 width. . . . . . . . . . 93, 97, 106, 170, 230, 240, 244, 259 width_factor . . . . . . . . . . . . . . . . . . 30, 31, 80, 91, 200 width_w_gap . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 252, 258 widths . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59, 248 WIPEOUT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 WIPEOUTVARIABLES . . . . . . . . . . . . . . . . . . . . . . . . . . . . 218

                                                wireframe_data_present . . . 25, 43, 56, 66, 69, 77,

82, 97, 111

WIREFRAME . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 wires . . . . . . . 25, 43, 56, 66, 69, 77, 82, 98, 111, 220 wizard_flag . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 word_break . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 245 workplane . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 WORLDVIEW . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21

X

x . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70, 251 x_ang. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 x_axis_dir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62, 185 x_dir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62, 222 x_direction . . . . . . . . . . . . . . . . . . . . . . . . . . 52, 92, 175 x_label . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 154 x_label_desc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 154 x_offset . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 241 x_radius . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99, 100 x_value . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 154 x_value_set. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 154 x04 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 249 x20 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 249 x40 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 249 x80 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 249 XCLIPFRAME . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 xdata. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 218 xdata_size . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 218 xdicobjhandle . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 262 XEDIT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 xline1_pt . . . . . . . . . . . . . . . . . . . . . . . . . . 29, 34, 36, 39 xline1end_pt. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 xline1start_pt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 xline2_pt . . . . . . . . . . . . . . . . . . . . . . . . . . 29, 34, 37, 39 xline2end_pt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35, 37 xline2start_pt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 XLINE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 XRECORD . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 218 xref … 107, 155, 161, 173, 176, 200, 205, 206, 215,

217

xref_pname . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32, 155 xrefoverlaid . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 155 xrefpaths . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 233 xrefs. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 233 xscale. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 XYGRIPENTITY. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 XYPARAMETERENTITY . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96

Y

y . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70, 251 y_label . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 154 y_label_desc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 154 y_offset . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 241 y_value . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 154 y_value_set. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 154 year. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 158

Z

z . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70, 251 z_is_zero . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24, 55 z_max. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 248 z_min. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 248 zero1. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 169